### v1

# Gravitational wave probes of particle dark matter: a review

Abstract
Various theories of dark matter predict distinctive astrophysical signatures in gravitational-wave sources that could be observed by ground- and space-based laser interferometers. Different candidates—including axions, dark photons, macroscopic dark matter, WIMPs, and dark-matter spikes—may appear in interferometer data via their coupling to gravity or the Standard Model, altering the measured gravitational-wave strain in distinct ways. Despite their differences, these candidates share two key features: (1) they can be probed through their effects on gravitational waves from inspiraling compact objects, isolated black holes, and neutron stars, or via direct interactions with detectors, and (2) their signatures likely persist far longer than the seconds-long mergers detected today, necessitating new data analysis methods beyond matched filtering. This review outlines these dark matter candidates, their observational signatures, and approaches for their detection.
I Introduction
Our knowledge of physics has been incomplete for decades, if not for over a century Bertone and Hooper (2018). Dark energy and dark matter (DM) could comprise 69% and 26% of the universe, and yet we have not been able to directly detect either. The experimental evidence for the existence of some “invisible” or DM is overwhelming: stars orbiting the center of the Milky Way would have flown out of their orbits if DM did not hold them in place Freeman (1970); Rogstad and Shostak (1972); Whitehurst and Roberts (1972); Roberts and Rots (1973); anisotropies in the cosmic microwave background could result from DM potential wells in the early universe Peebles (1982); Burles and Tytler (1998); DM could seed the formation of large-scale structures visible today, which would have been impossible with ordinary matter alone Pagels and Primack (1982); Davis et al. (1982); Blumenthal et al. (1984); Davis et al. (1985); Navarro et al. (1996); and gravitational lensing of light by galaxies in between us and the light source could be explained by invisible mass inside the galaxy Paczynski (1986); Alcock et al. (1993); Aubourg et al. (1993); Alcock et al. (2000); Lasserre (2000); Tisserand et al. (2007).
While the observational evidence indicates that DM may exist, it does not explain what it is made of. Could it be a single new particle, such as the axion Peccei and Quinn (1977a, b); Weinberg (1978) or dark photon Galison and Manohar (1984); Holdom (1986); Pierce et al. (2018); Filippi and De Napoli (2020), a collection of new particles, or macroscopic, in the form of PBHs formed within $\sim 1$ second of the Big Bang Green and Kavanagh (2021)? Or, particle DM may (or may not) exist alongside PBHs, which may comprise a portion or the totality of DM Hawking (1971); Carr et al. (2010); Clesse and García-Bellido (2015); Carr et al. (2016); Clesse and García-Bellido (2017); Clesse and Garcia-Bellido (2022); Miller (2024).
The numerous hypotheses regarding the existence of DM suggest that the mass of its constituents can vary across hundreds of orders of magnitude Belenchia et al. (2022). Significant time and resources, however, have been spent on searching a somewhat narrow mass range comprising so-called Weakly Interacting Massive Particles (WIMPs) Jungman et al. (1996). WIMPs have been probed through collider experiments Goodman et al. (2010) and DM/ nucleus interactions by experiments such as CDMS Akerib et al. (2003, 2006); Ahmed et al. (2009), SuperCDMS Agnese et al. (2019), XENON Angle et al. (2008, 2011); Aprile et al. (2012, 2016, 2024), and LUX Akerib et al. (2013, 2016). Other experiments have tried to detect annihilating DM indirectly through its gamma-ray emission in the Galactic Center, galaxy clusters, or other areas, e.g. MAGIC Aleksic et al. (2012); Aleksić et al. (2016), VERITAS Holder et al. (2009), HESS Aharonian et al. (2006) and Fermi Hooper and Goodenough (2011), or through cosmic rays, e.g. AMS Aguilar et al. (2007, 2013); Bergstrom et al. (2013) or Fermi-LAT Abdo et al. (2009); Ackermann et al. (2012). Despite these efforts, however, WIMPS remain undetected [^1] DAMA claims to see a periodic oscillation of a DM signal, though this result has not yet been reproduced Petriello and Zurek (2008), and so it is worthwhile to ask whether other models could better explain DM. Thus, the experimental development from the particle physics community has widened to include probing ultralight DM through torsion balance experiments, such as Eöt-Wash Su et al. (1994); Schlamminger et al. (2008) and the MICROSCOPE satellite Touboul et al. (2012); Bergé et al. (2018), whose results can be interpreted as constraints on DM, and resonant cavity experiments, such as ADMX Du et al. (2018), among others. It is therefore also worth asking whether high-precision GW interferometers, such as LIGO, Virgo and KAGRA Aasi et al. (2015); Acernese et al. (2015); Aso et al. (2013), can be used to search for DM, and in particular, whether continuous gravitational-wave (CWs) can be used as the probe. While not yet as extensive as particle physics experiments, GW probes of DM have already, in many cases, surpassed limits from existing experiments, or will beat them as the detectors, both in ground and in space, develop.
CWs are quasi-monochromatic, quasi-infinite-duration signals that canonically arise from asymmetrically rotating, lumpy neutron stars. Over the last few decades, much research has focused on developing ways to probe such neutron stars, both those that are known electromagnetically and those that may only be emitting GWs (“gravitars”) Sieniawska and Bejger (2019); Tenorio et al. (2021); Riles (2023); Piccinni (2022); Miller (2023); Wette (2023). Though the signal model is simple, the unknown sky positions and the uncertain neutron-star physics (e.g. the unknown equation of state, “spin wandering” Akmal et al. (1998); Cutler (2002), changes in the magnetosphere Lyne et al. (2010), etc.) complicates these searches, and thus methods that are robust not only to non-Gaussian noise disturbances but also to theoretical uncertainties had to be developed to look for such systems Krishnan et al. (2004); Dhurandhar et al. (2008); Astone et al. (2014); Suvorova et al. (2016); Bayley et al. (2019). Until now, such searches have only yielded constraints on the maximum possible deformation of the size of known Abbott et al. (2021a, 2022a); Ashok et al. (2021, 2024); Abac et al. (2025) and unknown neutron stars Abbott et al. (2022b); Steltner et al. (2023).
It turns out that with CWs, we can not only probe neutron stars, but also a wide range of DM models. If DM particles couple to standard model ones, we could see them in GW data via their interactions with detector components, which would give rise to a narrow-band, stochastic “correlated-noise” signal Pierce et al. (2018); Guo et al. (2019); Grote and Stadnik (2019); Ismail et al. (2023). Separately, if a new boson that may or may not be DM exists and couples to gravity, and is of just the right mass, it will form clouds around rotating BHs, extracting mass and spin from them as they grow, and emitting CWs as the DM particles annihilate Brito et al. (2015). Heavy, strongly-interacting DM could induce collapses of stars that form in binaries, and as these objects orbit each other, they will emit GWs at particular frequencies distinct from if they were ordinary stars of the same mass Singh et al. (2021); Bhattacharya et al. (2024). Such ultralight particles could also interact with the material inside neutron stars, changing their moments of inertia and thus their CW emission Khmelnitsky and Rubakov (2014). CWs can be used to probe the millisecond pulsar hypothesis to the GeV excess, which, as it is more constrained, lends more strength to the WIMP DM annihilation hypothesis for the GeV excess Calore et al. (2019); Miller and Zhao (2023); Bartel and Profumo (2024). Finally, axions can be constrained using the very slow orbit of the Hulse-Taylor binary and other known binaries to be below to be $f_{a}<\mathcal{O}(10^{11})$ GeV, depending on the system Kim and Yang (2024).
This review article is broken into different parts depending both on the type of DM considered and constraints on DM interactions with gravity or standard model particles. In Section II , we describe different ways that ultralight dark matter could couple to standard model particles in GW interferometers. Section III details the constraints that have been derived from searches of recent LIGO, Virgo and KAGRA data on the different ways that DM could couple to the interferometers. Section IV focuses on how macroscopic DM could transit through the interferometers, giving rise to a measurable differential acceleration of the mirrors. In Section V, we describe how scalar and vector bosons (that could be DM) form around rotating BHs, and the types of GW signals that could be emitted if these BHs are isolated or are in binary systems. Search results for annihilating scalar and vector boson cloud systems are explained in Section VI. Next, we discuss how soliton DM would impact GW signals from binary systems and neutron stars in Section VII. The collection of infamous WIMPs around BHs and in the Galactic Center are discussed in Section VIII in the context of how they can be probed with GWs. We explain how DM spikes around rotating BHs could alter GW signals in Section IX. In Section X, we describe a model of atomic DM that would give rise to “dark atoms” and “dark BHs”, and how it can be constrained with GW observations and non-observations of merging compact objects. We conclude in Section XI about the future for the burgeoning field of GW probes of particle DM.
II Ultralight dark matter interacting with the interferometers
II.1 Generic features of ultralight dark matter signal
If it exists, cold, ultralight DM could interact with standard-model particles in model-dependent ways. Such interactions would cause macroscopic differences in the materials, and, depending on the type of DM, would lead to an observable signal in different components of GW interferometers. While the physics behind each type of DM is different, there are some characteristics that are model-independent. First, the number of ultralight DM particles in a given region in space is gigantic, and can be calculated by attributing the DM energy density $\rho_{\text{DM}}$ to result from ultralight DM:
$N_{0}$ $=$ $\lambda_{\rm C}^{3}\frac{\rho_{\text{DM}}}{m_{\rm DM}c^{2}}=\left(\frac{2\pi\hbar}{m_{\rm DM}v_{0}}\right)^{3}\frac{\rho_{\text{DM}}}{{m_{\rm DM}c^{2}}},$ (1)
$\approx$ $1.69\times 10^{54}\left(\frac{10^{-12}\text{ eV}}{m_{\rm DM}}\right)^{4},$
where $v_{0}\simeq 220$ km/s is the virial velocity Smith et al. (2007), $\hbar$ is Planck’s reduced constant, $c$ is the speed of light, $m_{\rm DM}$ is the mass of the ultralight dark matter particle, $\lambda_{\rm C}$ is the Compton wavelength of DM, and $\rho_{\text{DM}}=0.3\,\mathrm{GeV}/$m3 is the DM energy density de Salas and Widmark (2021) (unless otherwise noted). The large occupation number $N_{0}$ implies that ultralight DM can be treated as a coherent region, where within that region, DM can be approximated as a plane wave that oscillates at a fixed frequency $f_{0}$, which is given by the DM mass:
$f_{0}=\frac{m_{\rm DM}c^{2}}{2\pi\hbar}\simeq 241\text{ Hz}\left(\frac{m_{\rm DM}}{10^{-12}\text{ eV}/c^{2}}\right),$ (2)
In other words, the observable effect is that the interferometer components will forever oscillate at a fixed frequency since they always sit in the DM field Hall et al. (2018). The DM wave also has a finite coherence time $T_{\text{coh}}$ that relates to the virial velocity of DM:
$T_{\text{coh}}=\frac{4\pi\hbar}{m_{\rm DM}v_{0}^{2}}=1.4\times 10^{4}\text{ s}\left(\frac{10^{-12}\text{ eV}}{m_{\rm DM}}\right),$ (3)
If we observe for a duration longer than the coherence time, the signal will no longer be monochromatic but have its power stochastically distributed at frequencies slightly higher than the characteristic frequency of the DM particle:
$\Delta f_{v}=\frac{1}{2}\left(\frac{v_{0}}{c}\right)^{2}f_{0}\approx 2.94\times 10^{-7}f_{0}.$ (4)
This frequency shift arises because in a given coherent region, the number of DM particles is huge, and all are traveling at different velocities that are Maxwell-Boltzmann distributed around the virial velocity. Furthermore, the DM signal has an associated coherence length $L_{\text{coh}}$:
$L_{\text{coh}}=\frac{2\pi\hbar}{m_{\rm DM}v_{0}}=1.6\times 10^{9}\text{ m}\left(\frac{10^{-12}\text{ eV}/c^{2}}{m_{\rm DM}}\right).$ (5)
If observing the interaction of DM with standard-model model particles in detectors separated by distances less than $L_{\text{coh}}$, which is the case for ground-based detectors, we expect to observe a correlated signal, again at an almost fixed frequency. In Fig. 2, we show an example of what a simulated vector DM signal interacting with one of the LIGO detectors looks like.
Figure 2: Taken from Miller et al. (2021). We show the strain time series $h(t)$ (left) and the Fourier transform of it (right) of a simulated dark photon DM signal. The structure in the amplitude spectrum density in the right-hand plot arises because we have taken the length of the Fourier transform to exceed that of the coherence time of the signal, and the signal is a superposition of $1000$ dark photons traveling with distinct Maxwell-Boltzmann-distributed velocities, which cause small frequency deviations away from the minimal frequency $f_{0}=740.436$ Hz ($m_{A}=3.06\times 10^{-12}$ eV$/c^{2}$). This plot shows why it is important to choose $T_{\text{FFT}}\sim T_{\text{coh}}$ for the analysis of DM signals. Though we have simulated a dark photon signal, the power spectrum will appear similarly for the other kinds of DM interaction signals. Here, $T_{\text{obs}}\sim 10^{5}$ s, meaning the frequency resolution is $\delta f=10^{-5}$ Hz. The coherence time and length of this signal are: $T_{\text{coh}}=4595$ s and $L_{\text{coh}}=5.28\times 10^{8}$ m; the coupling strength is $\epsilon=3\times 10^{-21}$. The signal is actually simulated for $\sim 233$ days, though we only show the first day of its time evolution.
We will now describe different ways in which DM could interact with particles in the interferometers. We will assume that we are working within $T_{\text{coh}}$, in order to avoid technical difficulties from working with particles of slightly different frequencies. However, we note that in practice, this consideration is necessary when modelling the signal and assessing the sensitivity of GW interferometers to DM/standard-model interactions.
II.2 Axions
Axions could couple to photons, given by the following Chern-Simons interaction Lagrangian:
$\mathcal{L}_{I}$ $=\frac{g_{\text{a}\gamma}}{4}a(t)F_{\mu\nu}\tilde{F}^{\mu\nu}$
[…]
where $a(t)$ is the axion field, $g_{\text{a}\gamma}$ is the axion-photon coupling constant, $F_{\mu\nu}$ is field strength of electromagnetic field, $\tilde{F}^{\mu\nu}\equiv\epsilon^{\mu\nu\rho\sigma}F_{\rho\sigma}/2$ is its Hodge dual with the Levi-Civita anti-symmetric tensor $\epsilon^{\mu\nu\rho\sigma}$, the dot denotes the time derivative, and $A_{\mu}$ is the electromagnetic vector potential. $F_{\mu\nu}\equiv\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$. All spatial derivatives here are being neglected.
In Fourier space, we can write $A_{i}$ in terms of two circular polarization modes:
$A_{i}(t,\bm{x})=\sum_{\lambda=\text{L},\text{R}}\int\dfrac{\mathrm{d}^{3}k}{(2\pi)^{3}}A_{\lambda}(t,\bm{k})\,e^{\lambda}_{i}(\hat{\bm{k}})\,e^{i\bm{k}\cdot\bm{x}}\ ,$ (7)
where $\bm{k}$ is the wave number vector, and “L” and “R” denote the left- and right-hand polarizations, and $\bm{x}$ is a spatial point at which the DM field is measured. By completely fixing gauge degrees of freedom ($\partial_{i}A_{i}=0$ and $A_{0}=0$, respectively), and noting how the polarization vectors $e^{\lambda}_{i}(\hat{\bm{k}})$ in different directions relate to each other Nagano et al. (2019), the equations of motion for $A_{\text{L/R}}(t,\bm{k})$ can be derived:
$\ddot{A}_{\text{L/R}}+\omega^{2}_{\text{L/R}}A_{\text{L/R}}$ $=0$ (8)
$\omega_{\text{L/R}}^{2}$ $=k^{2}\left(1\mp g_{\text{a}\gamma}\dot{a}/k\right),$ (9)
where $k$ is the momentum and $\omega_{\rm L/R}$ is the angular frequency of the light. Here, we can see a modified dispersion relationship based on the strength of the axion-photon coupling.
Noting that the axion field can be written as (assuming to work within one $T_{\text{coh}}$:
$a(t)=a_{0}\cos(m_{\rm DM}t)=\left(\frac{\hbar\sqrt{2\rho_{\text{DM}}}}{m_{\rm DM}c}\right)\cos(m_{\rm DM}t)$ (10)
where $a_{0}$ is the axion field amplitude normalized by the DM energy density, and $t$ is time. Using the ordinary photon dispersion relationship is $c_{0}=\frac{\omega}{k}$, we can derive how the velocities of different polarizations of light will differ:
$\frac{c_{\text{\text{L/R}}}(t)}{c}$ $\simeq 1\pm\delta c(t)\equiv 1\pm\delta c_{0}\sin(m_{\rm DM}t),$ (11)
where $\delta c_{0}=g_{\text{a}\gamma}a_{0}m/(2k)$ is the maximum change in the phase velocity of light induced by the axion, and $\delta c_{0}\ll 1$ is assumed. In physical quantities, $\delta c_{0}$ is the amplitude of the signal we expect, and is equal to:
$\delta c_{0}\simeq 1.3\times 10^{-24}\left(\frac{\lambda}{1550\text{ nm}}\right)\left(\frac{g_{\text{a}\gamma}}{10^{-12}\text{ GeV}^{-1}}\right),$ (12)
in which the laser light is assumed to have a wavelength that relates to $k$ as: $\lambda=2\pi/k$.
For our purposes, in the presence of an axion field, the phase velocities of circularly polarized photons in the interferometers are modulated Nagano et al. (2019). This signal can only be seen if additional but simplistic optical components are added to measure the optical path difference between two orthogonal polarizations of light ($p$- and $s$-) Nagano et al. (2021). In practice, GW detectors work with linearly polarized light ($p$-) that will be modulated by axions, making $s$- polarized light. Furthermore, such deviations in polarizations can also be measured in pulsar timing arrays Xue et al. (2024); Porayko et al. (2024).
II.3 Dilaton scalar dark matter
Dilaton-like DM Preskill et al. (1983); Abbott and Sikivie (1983); Dine and Fischler (1983); Cho and Keum (1998); Cho and Kim (2009); Arvanitaki et al. (2015a); Stadnik and Flambaum (2015a, b, 2016) could cause the rest mass of the electron and other physical constants to oscillate with time. In other words, atoms’ Bohr radii of different interferometer components would be modulated at the dilaton frequency Grote and Stadnik (2019), and would induce changes in the size and index of refraction of the beam splitter Vermeulen et al. (2021) and the arm mirrors Göttel et al. (2024). Since beam-splitter size oscillations would asymmetrically affect both rays, DM effects would lead to a differential phase shift (i.e. a measurable signal). In fact, at first the beam splitter effect was assumed to be the only one that mattered, but recently, the finite differences between the thicknesses of the mirrors in each arm of the interferometer lead to non-negligible differences in the path lengths of the light in the interferometer Göttel et al. (2024).
Similarly to the axion, the scalar ultralight dark-matter field $\phi$ can be written as Arvanitaki et al. (2015a); Derevianko (2018); Vermeulen et al. (2021):
$\phi(t,\vec{r})=\left(\frac{\hbar\sqrt{2\rho_{\rm{DM}}}}{m_{\rm DM}c}\right)\cos\left(\omega t-\vec{k}\cdot\vec{r}+\Psi\right),$ (13)
where $\vec{r}$ is a position vector, $\omega=(m_{\rm DM}c^{2})/\hbar$ is the angular Compton frequency, $\vec{k}=(m_{\rm DM}\vec{v}_{\text{obs}})/\hbar$ is the wave vector,$\Psi$ is a phase factor, and $\vec{v}_{\text{obs}}$ is the velocity of the DM relative to the observer.
The Lagrangian $\mathcal{L}_{\mathrm{int}}$ for this scalar field $\phi$ is Vermeulen et al. (2021):
$\mathcal{L}_{\mathrm{int}}\supset\frac{\phi}{\Lambda_{\gamma}}\frac{F_{\mu\nu}F^{\mu\nu}}{4}-\frac{\phi}{\Lambda_{e}}m_{e}\bar{\psi}_{e}\psi_{e},$ (14)
where $\psi_{e}$ and $\bar{\psi}_{e}$ are the standard-model electron field and its Dirac conjugate, $m_{e}$ is the mass of the electron, and $\Lambda_{\gamma}$ and $\Lambda_{e}$ denote the scalar DM coupling parameters to the photon and electron, respectively.
For comparison, we write a part of the standard model Lagrangian as:
$\mathcal{L}_{\textrm{SM}}\supset-\frac{F_{\mu\nu}F^{\mu\nu}}{4}-m_{e}\bar{\psi}_{e}\psi_{e}-eJ_{\mu}A^{\mu}$ (15)
and we can see that the rest mass of the electron effectively changes in the presence of DM:
$m_{e}^{\prime}=m_{e}\left(1+\frac{\phi}{\Lambda_{e}}\right)$ (16)
Additionally, the four-vector potential must be canonically normalized to reproduce Maxwell’s equations, so:
$A^{\prime}_{\mu}=A_{\mu}\sqrt{1-\frac{\phi}{\Lambda_{\gamma}}},$ (17)
which further implies that the charge of the electron $e$ changes as:
$e^{\prime}=\frac{e}{\sqrt{1-\frac{\phi}{\Lambda_{\gamma}}}}$ (18)
and thus the fine structure constant changes as:
$\alpha^{\prime}=\frac{e^{\prime 2}}{4\pi}=\alpha\left(1-\frac{\phi}{\Lambda_{\gamma}}\right)^{-1}\simeq\alpha\left(1+\frac{\phi}{\Lambda_{\gamma}}\right)$ (19)
II.3.1 Changes to size and indices of refraction of solids
The changes in fundamental constants will lead to a change in length of, or strain on, solid bodies. Noting that the length $l$ of an object is roughly $l\sim N_{A}a_{B}$, where $N_{A}$ is the number of atoms in the body and $a_{B}=(m_{e}\alpha)^{-1}$ is the Bohr radius, we can obtain the strain on a solid body in general:
$\delta l=-N_{A}(m_{e}\alpha)^{-2}(m_{e}\delta\alpha+\alpha\delta m_{e})$ (20)
$\frac{\delta l}{l}=-\left(\frac{\delta\alpha(t)}{\alpha}+\frac{\delta m_{e}(t)}{m_{e}}\right)$ (21)
$\delta l=l\phi\left(\frac{1}{\Lambda_{\gamma}}+\frac{1}{\Lambda_{e}}\right)$ (22)
where $\delta$ refers to the change in each parameter:
$\delta m_{e}$ $=m_{e}^{\prime}-m_{e}=m_{e}\frac{\phi}{\Lambda_{e}}$ (23)
$\delta\alpha$ $=\alpha^{\prime}-\alpha=\alpha\frac{\phi}{\Lambda_{\gamma}}$ (24)
Then, the differential displacement $\delta(L_{x}-L_{y})$ on the beam splitter in GW interferometers can be derived Grote and Stadnik (2019). Along the $x$-arm, assuming that light is incident on the beam splitter at 45 degrees, a reflectivity of the mirror of 50%, the change in the $x$-arm length, $\delta L_{x}$, is calculated as:
$\delta L_{x}\approx\delta[\sqrt{2}nl-l/(2\sqrt{2})-w/2],$ (25)
where $w$ is the thickness of the mirrors. Along the $y$-arm, the change in arm length is:
$\delta L_{y}=-\delta l/(2\sqrt{2})-\delta w/2,$ (26)
and finally, the differential length change is:
$\delta(L_{x}-L_{y})$ $\approx\sqrt{2}\left[\left(n-\frac{1}{2}\right)\delta l\right]$ (27)
$\approx\left(\frac{1}{\Lambda_{\gamma}}+\frac{1}{\Lambda_{e}}\right)\left(\frac{n\,l\,\hbar\,\sqrt{2\,\rho_{\mathrm{DM}}}}{m_{\rm DM}\,c}\right)\cos\left(\omega t\right),$ (28)
where $n$ and $l$ are the index of refraction and length of the beam splitter, respectively. We only consider the $\delta l$ contribution, not one from $\delta n$, because it is much larger than than the $\delta n$ one.
Additionally, dilatons could be visible in a different part of the interferometers, that is, in the reference cavity. Laser light locked to an optical cavity made out of a solid material will have its frequency modulated due to the presence of a dilaton DM field, compared to light locked to a free-space suspended cavity, whose frequency would remain stable Derevianko (2018); Geraci et al. (2019); Hall and Aggarwal (2022). The strain amplitude of such a signal can be written as:
${h}_{\text{DM}}(t)=\left[\frac{\hbar}{m_{\rm DM}c^{3}}\sqrt{8\pi\rho_{\text{DM}}G}(d_{e}+d_{m_{e}})\right]\cos\omega t$ (29)
where $d_{e}$ and $d_{m_{e}}$ relate to the other couplings as:
$d_{e,m_{e}}=\frac{M_{\mathrm{Pl}}}{(\sqrt{4\pi}\Lambda_{\gamma,e})},$ (30)
where $M_{\mathrm{Pl}}$ is the Planck mass, and $m_{e}$ is the electron mass.
II.3.2 Changes to neutron-star moment of inertia
Fluctuations in fundamental constants could also cause other effects, such as changing the moment of inertia of neutron stars Kaplan et al. (2022), or changing the times measured cesium clocks Kaplan et al. (2022). These effects will be relevant when we describe how ultralight dark matter could be detected in pulsar timing arrays. To study the former, we must also consider interactions in QCD Damour and Donoghue (2010); Kaplan et al. (2022):
$\mathcal{L}_{\phi,\rm QCD}\supset\frac{\phi}{\Lambda}\left(\frac{d_{g}\beta_{3}}{2g_{3}}G_{\mu\nu}^{A}G_{A}^{\mu\nu}-\sum_{q=u,d}(d_{q}+\gamma_{q}d_{g})m_{q}\bar{q}q\right)$ (31)
where $\beta_{3}$ is the QCD beta function, $g_{3}$ is the QCD gauge coupling, $G_{\mu\nu}^{A}$ is the gluon field strength tensor, $\gamma_{q}$ are the light quark anomalous dimensions, $\Lambda=M_{\text{Pl}}/\sqrt{4\pi}$, and $d_{g}$, and $d_{q}$ are the dilaton coupling constants to the gluon and the quarks, respectively. Reading off the Lagrangian, and noting the difference from the standard model, the change in masses of quarks is:
$\frac{\delta m_{q}}{m_{q}}$ $=\frac{d_{q}}{\Lambda}\phi\,,$ (32)
while in nucleon masses:
$\frac{\delta m_{p,n}}{m_{p,n}}\simeq\frac{1}{\Lambda}\Big{(}d_{g}+C_{n}\,d_{\hat{m}}\Big{)}\phi\,,$ (33)
where $C_{n}=0.048$ Damour and Donoghue (2010), $m_{p}$ is proton mass, $m_{n}$ is the neutron mass, and the symmetric combination of the quark mass couplings has been defined to be:
$d_{\hat{m}}\equiv\frac{d_{u}m_{u}+d_{d}m_{d}}{m_{u}+m_{d}}\,.$ (34)
Moving back to the neutron star, conservation of angular momentum requires that the fluctuations in the spin frequency $\delta\omega$ relate to changes in the pulsar’s moment of inertia $\delta I$ as:
$\frac{\delta\omega}{\omega_{0}}=-\frac{\delta I}{I_{0}}\,.$ (35)
where $I_{0}$ is the unperturbed moment of inertia and $\omega_{0}$ is the unperturbed angular frequency of the neutron star. Based on the above discussion, DM interactions can cause changes in the masses of fundamental particles, which, in turn, cause time-dependent oscillations in the moment of inertia. In the case of neutron stars, the mass of the neutron will oscillate, and thus, indirectly, the mass of the neutron star itself will change as well.
Assuming a simplistic model of a non-rotating, spherically symmetric neutron star composed only of neutrons with a polytropic equation-of-state, the radius and moment of inertia can be written as Shapiro and Teukolsky (1983):
$R$ $\propto M_{0}^{-1/3}m_{n}^{-8/3}$ (36)
$I_{0}$ $\propto M_{0}R^{2}=M^{1/3}m_{n}^{-16/3},$ (37)
where $R$ and $M_{0}$ are the radius and unperturbed mass of the neutron star. Solving explicitly the Tolman–Oppenheimer–Volkoff (TOV) equations for a realistic equation-of-state would cause $\mathcal{O}(1)$ deviations from the simple scalings. These “unknown” scalings can be parameterized in terms of $\eta$ and $\delta\eta$
$\frac{\delta I}{I_{0}}=\eta\frac{\delta M}{M_{0}}+\delta\eta\frac{\delta m_{n}}{m_{n}}\,.$ (38)
which are set to $\eta=1/3$ and $\delta\eta=-16/3$ in line with the simplistic model, and to illustrate the concept.
Because of the changes in the pulsar’s moment of inertia, the pulsar’s spin frequency will change. For pulsar timing array experiments, the measured quantities are the times of arrivals of pulses from each pulsar, which will change in response to the oscillating moment of inertia. In practice, the observable here is the residual $h$ of the times of arrivals of pulses from the neutron star, which is given by:
[…]
Now, the changes in the particle mass must be related to the macroscopic change in the pulsar mass
$\frac{\delta M}{M_{0}}$ $=\sum_{f\in\{e,\mu,p,n\}}Y_{f}\frac{m_{f}}{m_{n}}\frac{\delta m_{f}}{m_{f}}\,,$ (40)
where $\delta M$ is the change in the pulsar’s mass. The sum is over the different types of particles present in the pulsar: electrons, muons, protons and neutrons, $Y_{f}\equiv N_{f}/(N_{n}+N_{p})$ is a weighting function that gives the number of $f$ particles relative to the sum of protons and neutrons (which dominate the pulsar’s mass). For neutron star observed by pulsar timing arrays, the following is assumed for $Y_{f}$: $Y_{n}=0.9$, $Y_{p}=Y_{e}=0.1$ and $Y_{\mu}=0.05$ Bell et al. (2019); Kaplan et al. (2022).
By plugging in the relations in Eq. 19 and Eq. 16 to Eq. 40, and subsequently integrating over time in Eq. 39, the changes in the timing residuals due to the ultralight dark-matter particles can be obtained:
$h(t)=$ $\frac{\sqrt{2\rho_{\phi}}}{m_{\rm DM}^{2}\Lambda}\left(\vec{y}\cdot\vec{d}\,\right)\hat{\phi}_{P}\sin{\big{(}m_{\rm DM}t+\gamma(\vec{x}_{P})\big{)}}\,,$ (41)
where $\vec{x}_{P}$ is the location of the pulsar, $\hat{\phi}_{P}$ is the pulsar normalized signal amplitude, $\gamma$ is a phase, and $\vec{y}\cdot\vec{d}\equiv\sum_{i}y_{i}d_{i}$ where $i\in\{m_{e},m_{\mu},g,\hat{m}\}$ and:
[…]
This type of DM interaction would leave imprints on pulsar timing array experiments Porayko et al. (2018); Smarra et al. (2023); Afzal et al. (2023), which will be discussed later.
II.3.3 Changes to atomic clock times
Changes in fundamental constants would also affect the way in which atomic clocks measure time. Since pulsar timing array experiments reference their measurements of the times of arrivals of pulses to terrestrial time (TT) using atomic clocks composed primarily of cesium, any oscillation of the fundamental constants will affect the atomic clocks, and thus the times measured Flambaum et al. (2004). Likewise, a shift in the frequency of the atomic clock will give rise to an apparent shift in the measured pulsar spin frequency, given by Eq. 39.
From Flambaum et al. (2004); Kaplan et al. (2022), the frequency of an atomic clock is:
$f\propto\Big{(}m_{e}\alpha^{2}\Big{)}\Big{[}\alpha^{2}F_{\rm rel}(Z\alpha)\Big{]}\left(\mu\frac{m_{e}}{m_{p}}\right)^{\zeta}$ (43)
where $Z$ is the nuclear charge, $F_{\rm rel}(Z\alpha)$ is a relativistic correction to an atom’s energy levels, $\mu$ is the nuclear magnetic moment, and $\zeta=1$ if the clocks that have hyperfine transitions (the case of cesium), and $\zeta=0$ for optical transitions [^2] This equation arises from solving the non-relativistic Schrodinger equation for the valence electron wavefunction and evaluating it at the nucleus of the atom.. The Casmir effect $F_{\rm rel}(Z\alpha)$ is a correction that comes from solving the relativistic wave equation for an electron near the nucleus
The fluctuations in fundamental constants will cause fluctuations in the frequencies of atomic clocks composed of atoms $A$:
$\begin{split}\frac{\delta f_{A}}{f_{A}}\simeq&\Bigg{[}\frac{\delta m_{e}}{m_{e}}+(4+K_{A})\frac{\delta\alpha}{\alpha}\\ &+\zeta\Bigg{(}\frac{\delta m_{e}}{m_{e}}+C_{A}\sum_{q=u,d}\frac{\delta m_{q}}{m_{q}}-\frac{\delta m_{p}}{m_{p}}\Bigg{)}\Bigg{]}\,,\end{split}$ (44)
where the derivative of the Casmir effect term is $\delta F_{\rm rel}/F_{\rm rel}=K_{A}$, and the derivative of the nuclear magnetic moment is $\delta\mu/\mu=C_{A}\,\delta m_{q}/m_{q}$. For Cesium, $K_{A}=0.83$ depends nontrivially on $Z$ and $\alpha$, and $C_{A}=0.110$ Flambaum et al. (2004). $m_{q}$ is a quark mass, and the index $q$ runs over up and down quarks.
Finally, by substituting the aforementioned equations for fluctuations of the fundamental constants into Eq. 44, and then further plugging this result into Eq. 39, the timing residuals induced by the ultralight dark-matter are :
$h(t)=\frac{\sqrt{2\rho_{\text{DM}}}}{m_{\rm DM}^{2}\Lambda}\left(\vec{y}\cdot\vec{d}\right)\hat{\phi}_{E}\sin{\big{(}m_{\rm DM}t+\gamma\big{)}}\,.$ (45)
where $\vec{x}_{E}$ is the position of the Earth, and the sensitivity parameters of this search are given by:
$\left\{y_{g},\,y_{\gamma},\,y_{\hat{m}},\,y_{m_{e}}\right\}\simeq\left\{\zeta,\,\xi_{A},\,\zeta\left(C_{n}+\hat{C}_{A}\right),1+\zeta\right\}\,,$ (46)
where $\xi_{A}\equiv 4+K_{A}$, and $\hat{C}_{A}=C_{A}(m_{u}+m_{d})^{2}/2m_{u}m_{d}$. This type of DM interaction would leave imprints on pulsar timing array experiments Porayko et al. (2018); Smarra et al. (2023); Afzal et al. (2023), which will be discussed later.
II.4 Vector dark matter
DM could be spin-1, i.e. vectors, which we refer to as “dark photons”. Dark photons could explain the entirety of the relic abundance of DM, which could arise from the misalignment mechanism Nelson and Scholtz (2011); Arias et al. (2012); Graham et al. (2016), parametric resonance or the tachyonic instability of a scalar field Agrawal et al. (2020); Co et al. (2019); Bastero-Gil et al. (2019); Dror et al. (2019), or from cosmic string network decays Long and Wang (2019). They could couple to standard model particles, either to baryon or baryon-lepton number. In the interferometers, these interactions would occur everywhere, but would be detectable in the four interferometer mirrors. The dark photons [^3] These are different dark photons than those that kinetically mix with the ordinary photon, which will be discussed later in the context of luminous superradiance Siemonsen et al. (2023a). would cause a “dark” force on the mirrors, causing quasi-sinusoidal oscillations Pierce et al. (2018); Miller et al. (2021).
Dark photons can be described analogously to ordinary photons, i.e. with a four-vector potential, which can be written as:
[…]
where $(A_{0})_{\mu}$ is the four-amplitude of $A_{\mu}$ $\Upsilon$ is a random phase, and $\vec{x}$ is the position at which $A_{\mu}$ is measured. The index $\mu={0,1,2,3}$ refers to both time and spatial components.
As always, we have the freedom to choose a gauge, and usually, the Lorenz gauge is easiest to work in ($\partial^{\mu}A_{\mu}=0$). After making this choice, we can compare the time component to magnitude of the spatial components of the four-potential:
$\frac{(A_{0})_{0}}{|\vec{A}_{0}|}=\frac{v_{0}}{c}\simeq 7.667\times 10^{-4},$ (48)
where $|\vec{A}_{0}|=\left(\frac{\hbar\sqrt{2\rho_{\text{DM}}}}{m_{\rm DM}c^{2}}\frac{1}{\sqrt{\epsilon_{0}}}\right)$ is the magnitude of the spatial components of $A_{\mu}$, normalized by the present DM energy density of the universe, and $\epsilon_{0}$ is the permittivity of vacuum. In Eq. 48, we see that the dark scalar potential is suppressed by about three orders of magnitude compared to the the dark three-vector potential. Therefore, the time-component of the four-vector potential can be safely neglected, leaving only the three-vector potential:
$\vec{A}=\vec{A}_{0}\sin\left(\omega t-\vec{k}\cdot\vec{x}+\Upsilon\right),$ (49)
The Lagrangian $\mathcal{L}$ that characterizes the dark photon coupling to a number current density $J^{\mu}$ of baryons or baryons minus leptons is:
$\mathcal{L}=-\frac{1}{4\mu_{0}}F^{\mu\nu}F_{\mu\nu}+\frac{1}{2\mu_{0}}\left(\frac{m_{\rm DM}c}{\hbar}\right)^{2}A^{\mu}A_{\mu}-\epsilon eJ^{\mu}A_{\mu},\$ (50)
where $\mu_{0}$ is the magnetic permeability in vacuum and $\epsilon$ is the strength of the particle/dark photon coupling normalized by the electromagnetic coupling constant.
Using Eq. 47, the “dark” electric and magnetic fields can be derived:
$\vec{E}$ $=\partial_{0}\vec{A}-\vec{\nabla}A_{0}\simeq\omega\vec{A}_{0}\cos(\omega t-\vec{k}\cdot\vec{x}+\phi),$ (51)
$\vec{B}$ =×$=\vec{\nabla}\times\vec{A}=-\vec{k}\times\vec{A}_{0}\cos(\omega t-\vec{k}\cdot\vec{x}+\phi),$ (52)
and we note that the electric field is much stronger than the magnetic one:
$\frac{|\vec{E}|}{|c\vec{B}|}\sim\frac{\omega}{|\vec{k}|}=\frac{c^{2}}{|\vec{v}|}\sim 10^{3}c,$ (53)
which implies that we need to only consider the electric field. This dark electric field causes the test masses to oscillate at a quasi-fixed frequency given by the dark photon mass. The acceleration of a given mirror can thus be derived Guo et al. (2019); Pierce et al. (2018):
$\vec{a}_{j}(t,\vec{x}_{j})$ $=\frac{\vec{F}_{j}(t,\vec{x}_{j})}{M_{j}}\simeq\epsilon e\frac{Q_{D,j}}{M_{j}}\omega|\vec{A}_{0}|\hat{A}\cos(\omega t-\vec{k}\cdot\vec{x}_{j}+\phi),$
$\epsilon^{2}$ $=\frac{\alpha_{\text{DP}}}{\alpha}$ (54)
where $\alpha_{\text{DP}}$ is the dark photon fine structure constant, $\epsilon^{2}$ is the ratio between the dark photon fine structure constant and the electromagnetic fine structure constant $\alpha$, and $Q_{D,j}$ is the total charge in the $j$th mirror of mass $M_{j}$. $\hat{A}$ is a unit vector, and $\phi$ is a function of the spatial coordinates. If dark photons couple to the baryon number, $q_{j}$ is the number of protons and neutrons in each mirror; if they couple to the difference between the baryon and lepton numbers, $q_{j}$ is the number of neutrons in each mirror. Each mirror is positioned differently with respect to the incoming dark photon DM field, and thus is accelerated at a slightly different amount over time. Thus, the interferometer experiences a measurable differential acceleration, and thus a differential strain.
Assuming that all mirrors are the same, integrating Eq. 54 twice over time, and averaging over random polarization and propagation directions, the strain on the interferometers caused by a dark photon DM signal is Pierce et al. (2018):
$\sqrt{\langle h^{2}_{D}\rangle}$ $=C\frac{Q_{D}}{M}\frac{\hbar e}{c^{4}\sqrt{\epsilon_{0}}}\sqrt{2\rho_{\text{DM}}}v_{0}\frac{\epsilon}{f_{0}},$
$\simeq 6.56\times 10^{-26}\left(\frac{\epsilon}{10^{-22}}\right)\left(\frac{100\text{ Hz}}{f_{0}}\right),$ (55)
where $C=\sqrt{2}/3$ (in LIGO) is a geometrical factor obtained by averaging over all possible dark photon propagation and polarization directions, and different detector geometricians (the calculation for $C$ is shown in the appendix of Pierce et al. (2018) and in Morisaki et al. (2021)).
A second strain results due to the so-called “finite light travel time effect, in which the mirrors in the interferometers have moved in the time that it takes the light to reach them from the beamsplitter. This strain can actually be larger than that given in Eq. 55:
$\sqrt{\langle h_{C}^{2}\rangle}$ $=\frac{\sqrt{3}}{2}\sqrt{\langle h_{D}^{2}\rangle}\frac{2\pi f_{0}L}{v_{0}},$
$\simeq 6.58\times 10^{-26}\left(\frac{\epsilon}{10^{-23}}\right).$ (56)
The total strain is: $\langle h_{{\rm total}}^{2}\rangle=\langle h_{D}^{2}\rangle+\langle h_{C}^{2}\rangle$.
II.5 Tensor dark matter
Modifications to gravity due to an additional spin-2 particle could act as DM. Specifically, bimetric gravity Hassan and Rosen (2012); Schmidt-May and von Strauss (2016), a theory in which a massless and massive spin-2 field interact, provides a plausible candidate for ultralight dark matter Marzola et al. (2018), produced by the misalignment mechanism Preskill et al. (1983).
Here, we consider a massive spin-2 field $M_{\mu\nu}$ described by the Fierz-Pauli Lagrangian density Armaleo et al. (2020):
${\mathcal{L}}$ $\,:=\frac{1}{2}M_{\mu\nu}\mathcal{E}^{\mu\nu\rho\sigma}M_{\rho\sigma}-\frac{1}{4}m_{\rm DM}^{2}\left(M_{\mu\nu}M^{\mu\nu}-M^{2}\right)\,,$ (57)
where $M:=g^{\mu\nu}M_{\mu\nu}$ and $\mathcal{E}^{\mu\nu\rho\sigma}$ is the Lichnerowicz operator, defined in equation 2.2 of Armaleo et al. (2020).
For the Friedman-Lemaitre-Robertson-Walker (FLRW) background metric, the equations of motion for this field, assuming it is ultralight, can be derived at late times as in  Marzola et al. (2018); Aoki and Maeda (2018); Kun et al. (2018). The field spatial component can be written in a similar way as for dark photons and scalar bosons at given position $\mathbf{x}$:
$M_{ij}$ $=\frac{\sqrt{2\rho_{\text{DM}}(\vec{x})}}{m}\cos{\left(\frac{mc^{2}}{\hbar}t+\vec{k}\cdot\vec{x}+\Upsilon(\vec{x})\right)}\varepsilon_{ij}(\vec{x})\,,$ (58)
where $\Upsilon$ is a random phase, and $\varepsilon_{ij}$ encodes information about the five polarizations of the massive spin-2 field.
The strain on the GW detectors arises analogously to that from GWs: a stretching of space-time in the presence of the field, since the massive spin-2 metric and its coupling constant can be absorbed into the definition of the massless spin-2 metric in the linear regime of the coupling constant $\alpha_{\rm TB}$ [^4] In the literature, $\alpha_{\rm TB}=\alpha$ simply, but to disentangle it from the fine structure constants, we rename it. TB is tensor boson. Armaleo et al. (2020); thus, the perturbation $h_{ij}$ is:
$h_{ij}(t)$ $=\frac{\alpha_{\rm TB}}{M_{\text{Pl}}}M_{ij}(t)=\frac{\alpha_{\rm TB}\sqrt{2\rho_{\text{DM}}}}{mM_{\text{Pl}}}\cos{\left(\frac{DMc^{2}}{\hbar}t+\Upsilon\right)}\varepsilon_{ij}(\mathbf{x})\,,$ (59)
where $\alpha$ is a dimensionless constant that quantifies the difference between the strengths of each of the spin-2 fields, and $M_{\text{Pl}}$ is the Planck mass.
Contracting $h_{ij}$ with the GW detector response function $D^{ij}$ in the tensor case, the resulting strain is:
$h(t)=D^{ij}h_{ij}=\frac{\alpha_{\rm TB}\sqrt{\rho_{\text{DM}}}}{\sqrt{2}mM_{\text{Pl}}}\cos{\left(\frac{m_{\rm DM}c^{2}}{\hbar}t+\Upsilon\right)}\Delta\varepsilon,$ (60)
where $\Delta\varepsilon:=\,\varepsilon_{ij}(n^{i}n^{j}-m^{i}m^{j})$.
II.6 Gravitational interactions of dark matter
The previous sections assumed a coupling of DM to standard-model particles. However, in principle DM need not interact at all with the standard model; thus, models of DM have been devised in which only a minimal coupling to gravity is assumed. This is, in fact, the simplest assumption that can be made, since we only have evidence for the existence of DM through gravitational interactions.
One such hypothetical scalar DM particle has a Lagrangian of:
$\mathcal{L}=\sqrt{-g}\left[\frac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{2}m_{\rm DM}^{2}\phi^{2}\right]\,.$ (61)
and Khmelnitsky and Rubakov (2014):
$\phi(\vec{x},t)=\frac{\sqrt{2\rho_{\text{DM}}}}{m_{\rm DM}}\hat{\phi}(\vec{x})\cos\left(m_{\rm DM}t+\gamma(\vec{x})\right),$ (62)
where $\gamma(\vec{x})$ is a phase that depends on position $\vec{x}$, and $\hat{\phi}(\vec{x})$ describes the interference pattern caused by the ultralight dark matter. As before, the scalar field density is normalized by $\rho_{\text{DM}}$.
Scalar DM could cause the arrival times of pulses from millisecond pulsars on earth to change, in an analogous way that GWs do Khmelnitsky and Rubakov (2014). Such a difference in the arrival times can be written as Khmelnitsky and Rubakov (2014); Porayko et al. (2018):
$\delta t_{\text{DM}}=\frac{\Psi_{\text{c}}(\vec{x})}{2m_{\rm DM}}[\hat{\phi}^{2}_{\text{E}}\sin{(2m_{\rm DM}+\gamma_{\text{E}})}-\hat{\phi}^{2}_{\text{P}}\sin{(2m_{\rm DM}+\gamma_{\text{P}})}],$ (63)
where
$\Psi_{\text{c}}(\vec{x})\approx 6.52\times 10^{-18}\left(\frac{10^{-22~{}}\text{eV}}{m_{\rm DM}}\right)^{2}\left(\frac{\rho_{\phi}}{0.4~{}\text{GeV}/\text{cm}^{3}}\right),$ (64)
where $\gamma_{\text{P}}\equiv 2\gamma(\vec{x}_{\text{p}})-2m_{\rm DM}d_{p}/c$ ($\gamma_{\text{E}}\equiv 2\gamma(\vec{x}_{\text{e}})$) parameterize random phases evaluated at the pulsar (P) or Earth (E), and $d_{p}$ is the distance between the pulsar and the Earth. The DM energy density background is assumed to be constant when calculating $\Psi_{\text{c}}(\vec{x})$, though the possibility of deviations, due to the coherently oscilalting ultralight dark-matter field, are parametrized in terms of two phase factors: one for the pulsar $\hat{\phi}^{2}(\vec{x}_{\text{p}})\equiv\hat{\phi}^{2}_{\text{P}}$ and one for the Earth ($\hat{\phi}^{2}(\vec{x}_{\text{e}})\equiv\hat{\phi}^{2}_{\text{E}}$). Since all the pulsars are separated by at least a kpc, assuming a constant DM energy density is valid Porayko et al. (2018).
III Observational constraints on dark matter directly interacting with gravitational-wave interferometers
Section II focused on the currently-known types of DM that could cause a signal in GW interferometer data. While we have indicated the strain amplitudes of such signals for some of the DM models considered, we have not yet discussed whether the interferometers are actually sensitive to such weak signals. In this section, we will describe how searches for DM interactions with GW detectors are performed, show current constraints on various models, and explain what future constraints will be on others.
Each method discussed below can achieve a certain sensitivity, that is, a minimum detectable strain amplitude on the detector at a given confidence level. This strain amplitude, the left-hand side of many equations in Section II, is then mapped to constraints on the physical coupling constant of DM to the standard model. The methods themselves are agnostic towards the particular model of DM that is being searched for.
III.1 Methods
As explained in Section II.1, the expected signal frequency is fixed by the DM mass, but has some stochastic variations of $\mathcal{O}(10^{-6})$ Centers et al. (2021). Because the interferometers always exist in the DM field, the signal is always present. Thus, we can describe the expected signal as: quasi-monochromatic and quasi-infinite duration. Essentially, we are looking for a resonance at a particular frequency fixed by the DM mass.
If the signal were purely monochromatic, we could simply take a single fast Fourier transform of the data and look for peaks in the power spectrum. However, the stochastic frequency variations prevent us from doing that. If we observe for a duration $T_{\text{obs}}$ longer than $T_{\text{coh}}$, the signal will not be sinusoidal and thus its power will be spread among different frequency bins, which would inhibit a possible detection. Thus, the following methods have been designed to combine signal power across chunks of data $T_{\text{FFT}}\ll T_{\text{obs}}$ in a way that optimizes sensitivity towards particle DM interactions with the standard model and gravity.
III.1.1 Cross-correlation
Cross-correlation Pierce et al. (2018); Guo et al. (2019) requires at least two separate time-domain datasets that are Fourier transformed and multiplied together to compute the cross-power in each frequency bin. The cross power is then divided by the auto-power of each detector, and then summed over all the Fourier transforms to arrive at a measure of power at each frequency. It is then divided by the standard deviation of the noise to compute the signal-to-noise ratio.
The cross-correlated signal strength for detector pair $IJ$ is:
$S_{IJ,j}=\frac{1}{N_{\rm FFT}}\sum_{i=1}^{N_{\rm FFT}}\frac{z_{I,ij}z_{J,ij}^{*}}{P_{I,ij}P_{J,ij}},$ (65)
in the $j^{\rm th}$ frequency bin at the $i^{\rm th}$ time. “*” denotes the complex conjugate, and $z_{1,ij}$ and $z_{2,ij}$ is the Fourier transform of the time-domain data from detectors 1 and 2, respectively, and $N_{\text{FFT}}$ is the total number of FFTs taken over the observing run. If the observing run lasts $T_{\text{obs}}$ and the FFT length is $T_{\text{FFT}}$, then $N_{\text{FFT}}=[T_{\text{obs}}/T_{\text{FFT}}]$. $P_{1,ij}$ and $P_{2,ij}$ are the individual detector power spectral densities, which are typically estimated using a running median of each FFT. Typically $T_{\text{FFT}}=1800$ s, independently of the DM mass, which ensures a relatively easy analysis but implies the sensitivity towards DM interactions is not optimal at each mass analyzed.
To obtain a detection statistic, i.e. the signal-to-noise ratio, we have to divide the signal strength by the variance of the data, which is:
$\sigma_{IJ,j}^{2}=\frac{1}{N_{\rm FFT}}\left\langle\frac{1}{2P_{I,ij}P_{J,ij}}\right\rangle_{N_{\rm FFT}},$ (66)
where $\langle...\rangle_{N_{\rm FFT}}$ is the average over $N_{\rm FFT}$ time segments. We note that the average is over inverse noise-weighted power spectral densities, (analogous to adding resistors in parallel), which helps to suppress spurious power due to large noise disturbances.
The signal-to-noise ratio (SNR), the dection statistic, is then simply:
${\rm SNR}_{IJ,j}=\frac{S_{IJ,j}}{\sigma_{IJ,j}}.$ (67)
In the presence of pure Gaussian noise, the SNR will follow a normal distribution, with a mean of 0 and a standard deviation of 1. If the SNR exceeds a certain threshold, which is set both theoretically and by the trials factor, then a particular frequency is classified as being “significant”.
Depending on the type of DM searched for, the signal-to-noise ratio for spin-0 (labeled SD), spin-1 (DP) and spin-2 (DG) cross-correlation searches can be written as Manita et al. (2024):
${\rm SNR}^{\rm SD}_{IJ,j}$ $\approx\left|\frac{2\sqrt{2}\alpha_{A}^{2}f_{\rm SD}T_{\rm eff}({\cal B}^{\prime}_{IJ}\Gamma_{IJ}+{\cal C}^{\prime}\gamma_{IJ})}{\sqrt{S_{n,I}(\frac{m_{\rm DM}}{2\pi})S_{n,J}(\frac{m_{\rm DM}}{2\pi})}}\right|\,.$ (68)
${\rm SNR}^{\rm DP}_{IJ,j}$ $\approx\left|\frac{2\sqrt{2}\epsilon_{D}^{2}f_{\rm DP}T_{\rm eff}({\cal B}_{IJ}\Gamma_{IJ}+{\cal C}(1+\frac{1}{3}\Omega^{z}_{\rm DP})\gamma_{IJ})}{\sqrt{S_{n,I}(\frac{m_{\rm DM}}{2\pi})S_{n,J}(\frac{m_{\rm DM}}{2\pi})}}\right|\,,$ (69)
${\rm SNR}^{\rm DG}_{IJ,j}$ $\approx\left|\frac{2\sqrt{2}\alpha^{2}f_{\rm DG}T_{\rm eff}{\cal A}\gamma_{IJ}}{\sqrt{S_{n,I}(\frac{m_{\rm DM}}{2\pi})S_{n,J}(\frac{m_{\rm DM}}{2\pi})}}\right|\,,$ (70)
For spin-0: $f_{\rm SD}\equiv\rho_{\rm SD}/\rho_{\text{DM}}$ is the ratio of the energy density of the spin-0 DM to the total energy density of DM. $\alpha_{A}$ is the dimensionless coupling constant that depends on the coupling constants of the scalar DM particles to standard model ones, e.g. the electron and photon, and are given by Eq. 44-47 in Manita et al. (2024). $\gamma_{IJ}$ is the overlap-reduction function that quantities the sensitivity loss by cross correlating detectors that are not co-located and co-aligned, and is given in Tab. I in Manita et al. (2024) for different pairs of detectors. It specifically arises from the spatial displacement of mirrors by DM. $\Gamma_{IJ}$ is the overlap-reduction function that arises from the finite-light time travel effects of the signal on the mirrors Morisaki et al. (2021).
For spin-1: $f_{\rm DP}\equiv\rho_{\rm DP}/\rho_{\text{DM}}$ is the ratio of the energy density of the spin-1 DM to the total energy density of DM. $\Omega^{z}_{\rm DP}$ is a term that depends on the polarization of DM, and $\Omega^{z}_{\rm DP}=1/3$ is when the polarizations are randomized.
For spin-2: The SNR is independent of how the DM energy density is distributed in each polarization mode. $f_{\rm DG}\equiv\rho_{\rm DG}/\rho_{\text{DM}}$ is the ratio of the energy density of the spin-2 DM to the total energy density of DM.
In all the expressions $T_{\rm eff}=T_{\text{obs}}$ if $T_{\text{obs}}<T_{\text{coh}}$, and $T_{\rm eff}=\sqrt{T_{\text{obs}}T_{\text{FFT}}}$ if $T_{\text{obs}}>T_{\text{coh}}$. Furthermore, the definitions of the dimensionless coefficients ${\cal C}$, ${\cal B}_{IJ}$ and ${\cal A}$ are:
${\cal C}$ $\equiv\frac{\rho_{\rm DM}e^{2}v_{0}^{2}}{5\omega^{2}}\left(\frac{Q_{D}}{M}\right)^{2}\,,$ (71)
${\cal B}_{IJ}$ $\equiv\frac{8\rho_{\rm DM}e^{2}}{3\omega^{4}L_{I}L_{J}}\left(\frac{Q_{D}}{M}\right)^{2}\sin^{2}\left(\frac{\omega L_{I}}{2}\right)\sin^{2}\left(\frac{\omega L_{J}}{2}\right)\,,$ (72)
${\cal A}$ $\equiv\frac{\rho_{\rm DM}}{5\omega^{2}M_{\text{Pl}}^{2}}$ (73)
where $\omega=2\pi f$ is the angular frequency of the ultralight dark matter, $L_{I}$ and $L_{J}$ are the arm-lengths of each interferometer,
III.1.2 Excess power time-frequency method
Cross-correlation is able to capture the phase information in each FFT; however, at the moment the software has not yet been implemented to change $T_{\text{FFT}}$ to match $T_{\text{coh}}$, which would result in optimal sensitivity to each DM mass. Thus, another method Miller et al. (2021) was developed that can vary $T_{\text{FFT}}$ as a function of frequency, and that is more robust against noise disturbances. It relies on creating time-frequency spectrograms in which frequencies in each FFT are only kept if their powers are (1) above a given threshold and (2) local maxima. After applying these cuts, a time-frequency “peakmap” is created, which is a collection of ones that indicates particular time/frequency points at which the aforementioned two conditions are met. A peakmap can be created every Hz, and $T_{\text{FFT}}$ can be varied to match $T_{\text{coh}}$ in each 1 Hz window. Because $T_{\text{FFT}}\sim T_{\text{coh}}$, the signal is expected to be sinusoidal. Thus, we can sum, in each frequency bin, the ones that are present, in essence creating a histogram. On this histogram, at each frequency, we compute a detection statistic called the “critical ratio”:
$CR=\frac{n-\mu}{\sigma}$ (74)
where $n$ is the number of peaks at a given frequency, and $\mu$ and $\sigma$ are the mean and standard deviation of the number of peaks in the histogram. The CR, like the SNR, follows a normal distribution. The minimum detectable strain amplitude of this method is:
$h_{\text{0,min}}$ $\approx\frac{\mathcal{G}}{N_{\text{FFT}}^{1/4}\theta^{1/2}_{\rm thr}}\sqrt{\frac{S_{n}(f)}{T_{\text{FFT,max}}}}$
×$\times\left(\frac{p_{0}(1-p_{0})}{p_{1}^{2}}\right)^{1/4}\sqrt{CR_{\text{thr}}-\sqrt{2}\mathrm{erfc}^{-1}(2\Gamma)},$
$N_{\text{FFT}}$ $=\frac{T_{\text{obs}}}{T_{\text{FFT,max}}},$
$p_{0}$ $=e^{-\theta_{\text{thr}}}-e^{-2\theta_{\text{thr}}}+\frac{1}{3}e^{-3\theta_{\text{thr}}}=0.0755\text{ for $\theta_{\text{thr}}=2.5$,}$
$p_{1}$ $=e^{-\theta_{\text{thr}}}-2e^{-2\theta_{\text{thr}}}+e^{-3\theta_{\text{thr}}}\text{ }=0.0692\text{ for $\theta_{\text{thr}}=2.5$,}$ (75)
where $\mathcal{G}=[1.3,2.8,2.08]$ for scalar, vector or tensor DM interactions, respectively. The $\mathcal{O}(1)$ differences arise from averages over the spin-1 or spin-2 polarizations. $\theta_{\text{thr}}$ is a threshold on the equalized power in the time-frequency spectrogram, and $\Gamma$ is the confidence level.
III.1.3 Logarithmic power spectral density (LPSD) method
To truly maximize the sensitivity towards DM signals, one must, with infinite precision, analyze chunks of data with $T_{\text{coh}}$ as the DM mass changes. LPSD Tröbs and Heinzel (2006); Vermeulen et al. (2021); Göttel et al. (2024) is an attempt to do this, which creates spectra that are logarithmically spaced in frequency. It is extremely computationally intensive, since it requires an implementation separate from the discrete FFT. The benefit of using such an algorithm is its ability to estimate power spectral densities logarithmically, thereby capturing important features of the signals that could be missed when estimating the power spectral density across a linear frequency range. Given that the ultralight dark matter causes frequency variations of $\mathcal{O}(10^{-6})$, it is worthwhile to use LPSD to capture these tiny fluctuations in frequency.
Accounting for the strain on the test masses, the loss of sensitivity due to the number of round-trips of the laser light, as well, the induced strain can also be written in terms of transfer functions between scalar DM and the interferometers:
$h(\omega)\approx\left(\frac{1}{\Lambda_{\gamma}}+\frac{1}{\Lambda_{e}}\right)\cdot\left(\frac{\hbar\,\sqrt{2\,\rho_{\mathrm{DM}}}}{m_{\rm DM}\,c}\right)A_{\rm cal}^{-1}(\omega)$ (76)
where $A_{\rm cal}$ is a function given in Göttel et al. (2024) that encapsulates the transfer functions between the DM and the beam splitter and test masses, as well as the thicknesses and indices of refraction of the mirrors, the mirrors’ transmissivities, and the GW/ detector transfer function. Then, $h(\omega)$ can be estimated from the data using Eq. 11 and 12 in Göttel et al. (2024), and Eq. 16 in Cowan et al. (2011), which can be mapped to constraints on the coupling constants using Eq. 76.
III.1.4 Stochastic summing method
This method leverages the fact that the signal power will be stochastically spread at frequencies above the characteristic one if $T_{\text{FFT}}>T_{\text{coh}}$, and essentially sums the power at those frequencies. Afterwards, a detection statistic, like the signal-to-noise ratio, is calculated by dividing by the noise power spectral density:
$\rho(f_{c})\equiv\sum_{i}^{N_{\text{FFT}}}\sum_{f_{0}\leq f_{n}\leq f_{0}(1+\kappa^{2}v^{2}_{0}/c^{2})}\frac{4|\tilde{d}(f_{n};t_{i})|^{2}}{T_{\text{FFT}}S(f_{n};t_{i})},$ (77)
where $\tilde{d}(f_{n};t_{i})$ is the FFT of $i^{\rm th}$ data chunk, $S(f_{n};t_{i})$ is the one-sided noise power spectral density at $t=t_{i}$, and $\kappa$ = 3.17 to ensure that no more than 1% of signal power is lost by restricting the inner sum. The outer sum is over all chunks of length $T_{\text{FFT}}$; the inner sum is over the frequency spread of the signal $\delta f_{v}$.
The sensitivity of this method has been estimated in Nakatsuka et al. (2023) for both axions and dark photons, analytically and numerically in the case of $T_{\text{FFT}}=T_{\text{obs}}$. In the absence of a signal, assuming Gaussian noise, a detection threshold can be established as a function of the false alarm probability $\alpha$ (Eq. 47 in Nakatsuka et al. (2023)), and this threshold can be fed back into an integral over the likelihood function (Eq. 46 in Nakatsuka et al. (2023)), to compute, at a given confidence level, the minimum detectable signal amplitude $\bar{\lambda}_{\mathrm{up}}$:
$\bar{\lambda}_{\mathrm{up}}(T_{\text{obs}})=\begin{cases}(\Delta_{X,\mathrm{tot}})^{-1/2}\sqrt{\frac{\ln(\alpha)}{\ln(\beta)}-1},&\mathrm{for}~{}T_{\text{obs}}<\tau/\kappa,\\ (\Delta_{X,\mathrm{tot}})^{-1/2}\sqrt{M_{\alpha}+M_{1-\beta}}\left(\kappa T_{\text{obs}}/T_{\text{coh}}\right)^{1/4}&\mathrm{for}~{}T_{\text{obs}}\gg T_{\text{coh}}.\end{cases}$ (78)
$\Delta_{X,\text{tot}}$ encodes the deterministic part of the spectral shape of the signal, and depends on the relative motion of the DM velocity with respect to the interferometer arms. $\beta$ is the detection probability. In principle, Eq. 11 and 17 of Nakatsuka et al. (2023) give these relative motion terms; in practice, however, conservative ($\vec{v}=[0,0,v_{0}]$) and optimistic ($\vec{v}=[v_{0},0,0]]$) choices of the direction of the DM velocity with respect to the interferometers result in values of [0.38,0.81], respectively (see Eq. 52 in Nakatsuka et al. (2023)). $M_{\chi}$ represents the relation between the peak width and the area of the Gaussian distribution defined by
[…]
Depending on the type of DM signal assumed to be interacting with the interferometers, $\bar{\lambda}_{\mathrm{up}}$ can be equated to the following physical normalized (by the noise) DM signal amplitudes to obtain the upper limit on the corresponding coupling constant:
$\bar{\lambda}_{\mathrm{axion}}$ $=\frac{T_{\text{obs}}}{\sqrt{T_{\text{obs}}S_{n}}}g_{\alpha\gamma}\sqrt{\frac{\rho_{\rm DM}}{m^{2}}}\frac{m}{2k},\quad$
$\bar{\lambda}_{\mathrm{time}}$ $=\epsilon e\frac{2T_{\text{obs}}}{\sqrt{T_{\text{obs}}S_{n}}}\sqrt{\frac{2\rho_{\rm DM}}{3m^{2}}}\frac{(Q/M)_{\mathrm{in}}}{mL}\sin^{2}\left(\frac{mL}{2}\right),\quad$
$\bar{\lambda}_{\mathrm{space}}$ $=\epsilon e\frac{2T_{\text{obs}}}{\sqrt{T_{\text{obs}}S_{n}}}\sqrt{\frac{2\rho_{\rm DM}}{3m^{2}}}\frac{(Q/M)_{\mathrm{in}}\bar{v}}{2{\sqrt{2}}},\quad$
$\bar{\lambda}_{\mathrm{charge}}$ $=\epsilon e\frac{2T_{\text{obs}}}{\sqrt{T_{\text{obs}}S_{n}}}\sqrt{\frac{2\rho_{\rm DM}}{3m^{2}}}\frac{|(Q/M)_{e}-(Q/M)_{\mathrm{in}}|}{2Lm}.$ (80)
where $L$ is the length of the interferometers and $(Q/M)_{e},(Q/M)_{\mathrm{in}}$ are the charge-to-mass ratios of the end mirrors and input mirrors, respectively. The above approach works if $T_{\text{coh}}\neq T_{\text{obs}}$, however in practice, this does happen, since the data are typically broken in to smaller chunks with $T_{\text{FFT}}\sim T_{\text{coh}}$. The likelihood equation derived in Nakatsuka et al. (2023) is actually unstable in this region, and so a numerical approach is used to estimate sensitivity and derive upper limits as in Abac et al. (2024).
III.1.5 Distinguishing amongst dark-matter models
Only a couple of works Miller et al. (2022a); Manita et al. (2024) have attempted to find ways of determining whether a particular type of DM interacted with GW detectors. In Miller et al. (2022a), the authors employ the Wiener filter Wiener et al. (1964) to compute the cross power across different detector pairs in the follow-up stages of the analysis, using different $T_{\text{FFT}}$, to exploit the ways that types of DM will couple to different interferometer baselines. Not only was the Wiener filter shown to produce low residuals when the signal in the data matched the model, and high values when the data were filtered with the wrong signal but also it could, with high confidence, veto candidates returned in a real search of O3 data. Ref. Manita et al. (2024) uses standard cross correlation to argue that the overlap reduction function can be used as way to distinguish amongst different spins of DM particles, since the cross power will be different. In both cases, cross power is important, since the individual power spectra are more or less indistinguishable for all types of ultralight dark matter, which arises because the phase evolutions for the considered types of ultralight dark matter here are the same.
III.2 Ground-based interferometers
Searches for DM particles interacting with the LIGO, Virgo, KAGRA and GEO600 interferometers have been performed using data from the most recent observing run, O3, resulting in constraints on dilatons and dark photons Abbott et al. (2022c); Vermeulen et al. (2021); Göttel et al. (2024); Abac et al. (2024). We summarize the results that have arisen from each of the existing ground-based interferometers.
III.2.1 LIGO, Virgo and GEO600
In Abbott et al. (2022c), two methods were used to search for ultralight dark matter: cross-correlation and excess power. Cross-correlation benefits from utilizing the phase information of the signal, something that is lost when using the time-frequency excess power method. On the other hand, the excess power method matches its coherence time $T_{\text{FFT}}$ to the signal coherence time $T_{\text{coh}}$ in every Hz band analyzed, allowing an improved sensitivity across the frequency range compared to cross-corelation, which fixed $T_{\text{FFT}}=1800$ s.
In Fig. 3, we show the upper limits on the coupling constant $\epsilon^{2}$ for both methods, in comparison to existing experiments. We can see, across a broad frequency range, that results from GW interferometers surpass by orders of magnitude constraints from Eöt-Wash and the MICROSCOPE experiments. These upper limits were derived for cross-correlation and excess power by following the procedure outlined in Section III.1.1 and Section III.1.2, respectively. Both methods employed the Feldman-Cousins approach Feldman and Cousins (1998) to set upper limits, which is robust against non-Gaussian noise and has been shown to produce conservative upper limits with respect to those that would have been obtained through simulations Miller et al. (2021).
Figure 3: Adapted from Abbott et al. (2022c). Upper limits obtained from analyzing LIGO O3 data on the coupling of dark photons to baryons $U(1)_{\rm B}$ in the LIGO mirrors. MICROSCOPE Bergé et al. (2018) and Eot-wash torsion balance upper limits are plotted as a comparison to the results here Schlamminger et al. (2008). To produce limits on dark photon/baryon-lepton coupling, $U(1)_{\rm B-L}$ in the LIGO mirrors, these limits should be multiplied by four.
Additionally, a recent search for scalar bosons interacting with LIGO was performed Göttel et al. (2024) using LPSD, resulting in tight constraints on the coupling constants given in Eq. 27. In Fig. 4, we can see, in blue, the constraints that have been derived, which supersede, at least a low masses, those that were derived in a GEO600 search for the same kind of DM Vermeulen et al. (2021). The enhancement comes from the improved low-frequency sensitivity of LIGO relative to GEO600, as well as incorporating the interaction of scalar DM with the mirrors, as well as the beam splitter, into the search. These upper limits were derived for by following the procedure outlined in Section III.1.3 and Section III.1.2.
Figure 4: Taken from Göttel et al. (2024). 95% confidence-level upper limits on $\Lambda_{i}^{-1}$, the coupling of dilaton DM to electrons or photons, as a function of mass and frequency from LIGO O3 data, which would have cause time-dependent oscillations of the sizes and indices of refraction of the beamsplitter and the LIGO mirrors. a) and b) show these constraints compared to existing ones on $\Lambda_{e}$ and $\Lambda_{\gamma}$, respectively. The results from the LIGO O3 search are shown by the thick blue line, constraints from direct experimental searches for DM Vermeulen et al. (2021); Aiello et al. (2022); Aharony et al. (2021); Savalle et al. (2021); Kennedy et al. (2020); Antypas et al. (2019, 2021); Tretiak et al. (2022); Oswald et al. (2022); Campbell et al. (2021); Beloy et al. (2021); Zhang et al. (2023); Fukusumi et al. (2023); Sherrill et al. (2023) are shown in thin grey, and constraints from searches for “fifth forces” Bergé et al. (2022); Hees et al. (2018) are depicted by the dashed red lines.
III.2.2 KAGRA
KAGRA is another interferometer in Japan that, while not yet as sensitive to GWs as LIGO and Virgo, can be used to search for DM. In particular, the different materials used in the mirrors in KAGRA can be exploited to be sensitive to vector DM that couples to baryon-lepton number Michimura et al. (2020), since the signal-to-noise ratio is enhanced for objects with different charge-to-mass ratios, as indicated by Eq. 80. Furthermore, this kind of DM can actually be seen in non-GW channels, such as the auxiliary length channels PRCL and MICH. The raw outputs of PRCL and MICH are time-varying displacement sof the forms (we also write the GW channel for comparison, $\delta L_{\rm DARM}$) Abac et al. (2024):
$\delta L_{\rm DARM}$ $=$ $\delta(L_{\rm x}-L_{\rm y}),$ (81)
$\delta L_{\rm MICH}$ $=$ $\delta(l_{\rm x}-l_{\rm y}),$ (82)
$\delta L_{\rm PRCL}$ $=$ $\delta[(l_{\rm x}+l_{\rm y})/2+l_{\rm p}],$ (83)
In KAGRA, $L_{x}=L_{y}=3000$m is the arm length, $l_{x}=26.7$m is the distance between the beam splitter and the input $x$ mirror, $l_{y}=23.3$m is the distance between the beam splitter and the input $y$ mirror, and $l_{p}=41.6$m is the the length of the power recycling cavity. All of these lengths are much shorter than the DM coherence length $L_{\text{coh}}$. See Fig. 1 of Abac et al. (2024) for a schematic of the interferometer.
These two channels contain both fused silica mirrors used for the beam splitter and power recycling mirrors, and cryogenic sapphire test masses. The charge-to-mass ratios for baryon-lepton number differ by $\sim 9\times 10^{-3}$ for fused silica and sapphire, while for baryon number alone, these quantities differ by only $\sim 4\times 10^{-5}$, showing a large enhancement of the baryon-lepton DM interaction signal.
No significant signals were found in this analysis of KAGRA O3 data Abac et al. (2024), so upper limits, as shown in Fig. 5, were calculated following the procedure outlined in Section III.1.4. The upper limits are shown for each channel analyzed, and many narrow-band noise artifacts are visible, which tends to limit the sensitivity of this search at low freqeuncies. The constraints are not as strong as those that come from LIGO; however, future searches with KAGRA are expected to yield powerful constraints on this type of DM due to the difference in charge-to-mass ratios of the mirrors Michimura et al. (2020).
Figure 5: Taken from Abac et al. (2024). 95% confidence-level upper limit on the strength of the coupling between dark photons and baryon-lepton number using KAGRA O3 data, derived from two particular interferometer channels: MICH (blue line) and PRCL (orange line). Because the KAGRA input mirrors, end mirrors, and power recycling mirrors are made of different material, the coupling strength of dark photons to baryon-lepton number is enhanced, and is more visible in these particular channels than in the ordinary strain one. In the low-mass (low-frequency) range, many narrow-band noise disturbances are seen, which arise due to unknown sources.
III.3 Space-based interferometers
Spaced-based GW interferometers, such as LISA Danzmann and Team (2003); Babak et al. (2021), Taiji Hu and Wu (2017), Tianqin Luo et al. (2016) and DECIGO Hu and Wu (2017), will hopefully fly within the next 10-15 years. The exquisite sensitivity of GWs in the $\mu$Hz to mHz band also allows the prospect to probe ultralight dark matter with masses of [$10^{-19},10^{-14}$] eV, a few orders of magnitude lower than those currently searched for in ground-based interferometers Kim (2023). Until now, however, only a protoype for LISA, called LISA Pathfinder, has been flown as a “proof-of-concept” mission Armano et al. (2016); Vetrugno (2017); Armano et al. (2018). Nevertheless, pilot searches can be performed on LISA Pathfinder data as a way to prepare for when space-based detectors fly, which allows for the development of robust data analysis pipelines and for handling of the peculiarities of future data, e.g. glitches, downtimes, etc.
Such a search was preformed using LISA Pathfinder data Miller and Mendes (2023) using the excess power method described in Section III.1.2 for ultralight dark photons coupling to baryons, however, no physical constraints on the coupling strength were obtained, since the noise level was too high, the arm length was too short (only $40$ cm compared to $\mathcal{O}(10^{6})$ km), and only one arm existed. However, the search procedure designed will serve as a roadmap for future analyses of spaced-based detector data, since many problems, including limited sampling at low frequency, gaps, noise non-stationarities and glitches had to be handled in this analysis of data from a space-based detector.
After this search was performed, however, it was realized that the relative acceleration of the spacecraft and one of the test masses would produce a much stronger signal than that which was searched for previously, which was the relative acceleration of the two test masses Frerick et al. (2024). A conservative upper limit on the baryon-lepton coupling was set without performing a search, assuming no signal would have been detected in LISA Pathfinder data, as shown in Fig. 6, by using a reference acceleration power spectral density of the spacecraft with respect to one of the test masses. We can see that the upper limits from the relative acceleration of space-craft/ test mass are much better than that from Miller and Mendes (2023), labeled .(decoherence). The constraint on B-L coupling is stronger with respect to that of baryon coupling because the test mass and spacecraft are primarily made of gold, and gold/carbon, respectively.
Figure 6: Taken from Frerick et al. (2024). Upper limits on the coupling of dark photons to baryon-lepton number considering the relative acceleration of the spacecraft and one of the test masses in LISA Pathfinder. The constraints derived from a search for the induced acceleration of the two test masses by dark photons in LISA Pathfinder data is shown as “decoherence” . Using the relative acceleration caused by dark photon interaction with baryon-lepton number allows for stringent constraints to be placed on the coupling, since the spacecraft and test masses are made of different materials. Constraints from other experiments and future projections for LISA are shown as well.
Only B-L constraints are shown from Frerick et al. (2024) because those arising from coupling to baryons are not yet strong enough to probe a tighter constraint than provided by MICROSCOPE. However, many conservative assumptions were made in Frerick et al. (2024), including the composition of the spacecraft and the use of a reference acceleration power spectral density, both of which could, if studied in more detail, lead to more stringent constraints for both the baryon and B-L coupling scenarios.
III.4 Pulsar timing arrays
Because pulsars are extremely accurate clocks, they can also be used to search for GWs, which will distort the spacetime between earth and pulsars. Practically speaking, GWs will cause a change in the times of arrivals of pulses with respect to that in the absence of GWs. Of course, other effects between the pulsars and earth must be taken into account, e.g. interstellar medium dispersion and scattering, instabilities in the rotation of the pulsar itself, which makes these measurements particularly challenging. Nevertheless, pulsar timing arrays can be formed by cross correlating the times of arrivals from multiple pulsars in the galaxy, mitigating many of these effects.
Ultralight scalar DM will also cause deviations in the times of arrivals of pulses as given in Khmelnitsky and Rubakov (2014), though the mass of such particles are 10 orders of magnitude lower than those probed with ground-based interferometers.
In Section II, we discussed various models of scalar DM that can be constrained with pulsar timing arrays. We again divide this section in the different types of DM that can be probed with pulsar timing arrays.
III.4.1 Gravitationally-interacting ultralight dark matter
As discussed before, DM could interact only via gravity, and alter the distances between the pulsars and earth, analgous to GWs. Using European pulsar timing array data, a search was performed in the residuals for this kind of DM interaction. In Fig. 8, we show recent constraints on ultralight scalar DM in two ways: (1) the minimum detectable strain amplitude as a function of DM mass, and (2) the fraction of DM that ultralight dark matter could compose. We can see that the masses below $\sim 10^{-23.2}$ eV can be well constrained by pulsar timing arrays, and that ultralight dark matter cannot make up all of DM at these masses. Fig. 8 presents upper limits in three scenarios: (1) uncorrelated, (2) correlated and (3) pulsar-correlated.
  1. 1.
    “Uncorrelated” refers to the case when the ultralight dark-matter $L_{\text{coh}}$ is less than the average pulsar separation and the earth-pulsar distance, and means that the pulsars experience different phases of the ultralight dark-matter signal. Thus, $\hat{\phi}_{P}^{2}$ and $\hat{\phi}_{E}^{2}$ are independent.
  2. 2.
    “Correlated” means that the earth-pulsar distance and the average distance between pulsars is smaller than $L_{\text{coh}}$, and that the DM $L_{\text{coh}}$ comprises the inner 20 kpc region of the galaxy for which galaxy rotation curves have been used to test the DM hypothesis. This means that each pulsar experiences the same phase of the ultralight dark-matter signal. In this case, the same coherence patch and local DM energy density are measured.
  3. 3.
    “Pulsar-correlated” means that $L_{\text{coh}}$ is smaller than the radius of the galactic center sampled by experiments that derive galaxy rotation curves. However, estimates of the DM energy density are from different coherence patches and thus measurements averaged over different patches.
Each type of analysis optimally probes a different mass regime depending on the DM coherence length. Therefore, the “correlated” curve is valid for masses less than $\sim 2\times 10^{-24}$ eV; the “pulsar-correlated” curve can be applied for $2\times 10^{-24}\,\text{eV}\lesssim m_{\rm DM}\lesssim 5\times 10^{-23}\,\text{eV}$ and the “uncorrelated” curve holds for $m_{\rm DM}\gtrsim 5\times 10^{-23}~{}\text{eV}$. In both correlated cases, $\hat{\phi}_{P}^{2}=\hat{\phi}_{E}^{2}$.
Figure 8: Taken from Smarra et al. (2023). 95% confidence-level upper limits from analyses of European pulsar timing array data on ultralight dark matter interacting only gravitationally and changing the times of arrivals of pulses in an analogous way as GWs do. The panels show (a) the dimensionless strain amplitude $\Psi_{\text{c}}$ and (b) the fraction of the DM energy density that scalar DM could compose as a function of DM mass and frequency. $\rho_{\text{DM}}=0.4\text{GeV}/\text{cm}^{3}$. For comparison, upper limits on these quantities are shown from previous searches Porayko and Postnov (2014); Porayko et al. (2018). In the right panel, we zoom in on the excluded ultralight dark matter masses. The horizontal dotted line represents the scenario in which ultralight dark matter is responsible for all of the local DM energy density . Scalar ultralight dark-matter particles with mass $-24.0<\text{log}_{10}\text{~{}}(m_{\rm DM}/\text{eV})<-23.7$ can comprise at most $30-40$ of $\rho_{\text{DM}}$, while particles with masses $-23.7<\text{log}_{10}\text{~{}}(m_{\rm DM}/\text{eV})<-23.3$ could comprise up to $\sim 70$ of $\rho_{\text{DM}}$.
Additionally, NANOGrav data can be used to constrain this kind of ultralight dark matter. In Fig. 9, a similar constraint is shown on a quantity that relates to the fraction of the DM energy density that gravitationally-interacting DM could compose. $S_{\phi}$ is the spin of DM, which can be 0 or 1 in this model. NANOGrav writes the timing residuals for the $I^{\rm th}$ pulsar as:
$h_{I}(t)$ $=\sum_{i}A_{E,I}^{i}(\bm{x}_{E})\sin(\omega t+\gamma_{E}^{i})$ (84)
$+A_{P,I}^{i}(\bm{x}_{P,I})\sin(\omega t+\gamma_{P,I}^{i})\,,$ (85)
where $i$ indicates the polarization ($=0$ for spin-0 and $=1,2,3$ for spin-1). This expression is divided into “earth” and “pulsar” terms with amplitudes and arbitrary phases of $A_{E,I}^{i}(\bm{x}_{E})$ and phase $\gamma_{E}^{i}$, and $A_{P,I}^{i}(\bm{x}_{P,I})$ and time-independent phase $\gamma_{P,I}^{i}$, respectively. These phases depend on the positions of the earth and the pulsar with respect to the ultralight dark-matter field (see Eq. 60 and 61 of Afzal et al. (2023)).
Like the EPTA, NANOGrav assumes a model of ultralight dark matter that induces metric perturbations and therefore changes in the times of arrivals of pulses Khmelnitsky and Rubakov (2014). The amplitude of the signal that is constrained is similar to Eq. 64:
$A_{E,I}^{i}(\bm{x}_{E})=A_{P,I}^{i}(\bm{x}_{E})=A_{\rm grav}=\left(2S_{\phi}+1\right)\displaystyle\frac{\pi G\rho_{\text{DM}}}{2m_{\rm DM}^{2}}\hat{\phi}^{2}(\bm{x})$ (86)
where $S_{\phi}=0$ or 1 for scalar and vector DM, respectively.
Fig. 9 shows the NANOGrav constraints on this amplitude, which can be compared to the right-hand panel of Fig. 8. NANOGrav cannot yet probe a physical constraint for the fraction of DM energy density that this model of DM could compose, likely because their dataset only spans 15 years, while the European pulsar timing array collaboration analyzed almost 25 years of data.
Figure 9: Taken from Afzal et al. (2023). Upper limits on the local ULDM density for the correlated (solid line) and uncorrelated (dashed line) signals at the $95\%$ credible level using data from the NANOGrav pulsar timing array. The model of DM considered here is the same as that in Fig. 8. The gray dashed line indicates the predicted DM abundance.
III.4.2 Variation of fundamental constants constraints
From the most recent NANOGrav dataset, constraints on ultralight dark matter that causes fundamental constants to oscillate Afzal et al. (2023) can be set as well for masses below $10^{-22}$ eV. In the case that ultralight dark matter causes changes in the moment of inertia of the pulsar, Fig. 10 depicts constraints on the couplings of the scalar DM particle to up- and down- quarks weighted by their masses ($\hat{m})$, electrons $(d_{e})$, muons $(d_{\mu})$, photons $(d_{\gamma})$ and gluons $(d_{g})$. To derive these constraints, the simplest model of the equation-of-state of the neutron star is assumed, as outlined in Section II.3.2, and, when constraining one particular coupling constant, the rest are assumed to be zero.
The strongest constraints from pulsar timing arrays are on the couplings to the electron and muons. Experiments that directly use atomic clocks to search for ultralight dark matter are insensitive to the electron coupling, since changing the electron mass does not affect the spacing between energy levels Kaplan et al. (2022). In terms of (the lack of) other constraints on muons, laboratory experiments do not have enough muons to harness on earth to study the coupling of DM to them, while, in contrast, neutron stars have an abundance of muons.
In Fig. 10, there is a gray-shaded region that indicates the situation in which $A<A_{\rm grav}$, where $A$ is given by the amplitude of Eq. 41 or Eq. 45. In other words, these constraints on ultralight dark matter coupling to standard-model particles are valid when the coupling of DM to gravity can be neglected, i.e. when Kaplan et al. (2022):
$d_{i}\gtrsim\frac{4.5\times 10^{-9}}{y_{i}}\left(\frac{10^{-23}\,\text{eV}}{m_{\rm DM}}\right)\,,$ (87)
In the gray-shaded region, the interaction between gravity and DM is stronger, and thus constraints on ultralight dark matter coupling to standard-model particles alone would not be valid in that region. Furthermore, these constraints assume that scalar particles constitute all of DM, which may not be true. If they only comprise a fraction $f$ of DM, then the constraints will be weakened by $\sqrt{f}$, because of the $\sqrt{\rho_{\text{DM}}}$ term in the amplitude of these interaction signals.
Figure 10: Taken from Afzal et al. (2023). Upper limits from an analysis of NANOGrav data on the coupling of ultralight dark matter to various standard-model particles: quarks, electrons, muons, photons and gluons, shown in red for the correlated case, and red dashed for the uncorrelated case. In this model, DM would alter the moment of inertia of neutron stars and thus affect the times of arrivals of pulses in pulsar timing arrays. The constraints on one coupling constant assume that all other coupling constants are set to 0. The black dashed line, and the gray shaded region, indicate where the amplitude of the signal would be less than the amplitude of the signal arising from DM coupling only to gravity. Current constraints “Rb/Cs atomic clocks” (purple) are from Hees et al. (2016), “Al/Hg atomic clocks” (turquoise) are from Beloy et al. (2021), “MICROSCOPE” (teal) are from Bergé et al. (2018), “H/Si clock shift” (orange) are from Beloy et al. (2021), and “NS binary system” are from Kumar Poddar et al. (2019) and Dror et al. (2020).
The scalings of these constraints with the DM mass should be discussed. At high masses, the signal is essentially deterministic, and the signal-to-noise ratio scales as the square of the signal amplitude, which means that $d_{i}\propto m^{2}$. At low masses ($m<1/T_{\text{obs}})$, the signal is observed in different coherence patches, since the coherence length is shorter than the pulsar separation. The signal is no longer a sinusoid but can be expressed as a polynomial in $mt$. The first two terms of this polynomial ($mt$ and $(mt)^{2}$) are degenerated with pulsar timing terms, and thus the first observable term is $(mt)^{3}$, meaning that the signal-to-noise ratio is proportional to $m^{3}$. Thus, $d_{i}\propto 1/m$, which explains the scaling at low masses. Still, though, red noise causes the constraints to flatten as the mass decreases, which, at first, hides this scaling.
III.4.3 Constraints on vector dark matter
As discussed in Section II.4, DM could also be composed of spin-1 bosons that interact with standard model particles. In the case of pulsar timing arrays, the ultralight dark matter would cause oscillatory forces on the earth and on the pulsars themselves with strengths proportional to the charge-to-mass ratio of these objects.
Fig. 11 shows constraints on the coupling constants $g_{B}=\epsilon_{B}$ and $g_{B-L}=\epsilon_{B-L}$ as a function of the DM mass and oscillation frequency. Again, the gray-shaded region indicates the regime in which the coupling of DM to gravity becomes stronger than the coupling of DM to standard-model particles. We can see that the NANOGrav constraints greatly surpass those that come from MICROSCOPE at very low masses. Furthermore, these constraints assume that vector bosons constitute all of DM, which may not be true. If they only comprise a fraction $f$ of DM, then the constraints will be weakened by $\sqrt{f}$, because of the $\sqrt{\rho_{\text{DM}}}$ term in the amplitude of these interaction signals. Furthermore, the stochastic GW background w
Figure 11: Taken from Afzal et al. (2023). Upper limits from analyses of NANOGrav pulsar timing array data on the coupling of DM to baryon and baryon-lepton number for the correlated (solid lines) an uncorrelated (dashed line) cases. In this case, dark photon DM would couple to standard model particles in the earth and the pulsars themselves. The black dashed line, and the gray shaded region, indicates where the amplitude of the signal would be less than the amplitude of the signal arising from DM coupling only to gravity. The constraints from tests of the equivalence principle (“MICROSCOPE”) are shown in teal (Bergé et al., 2018), and the constraints previously set by the PPTA Collaboration are reported in yellow (Xue et al., 2022).
III.4.4 Constraints on conformal DM
DM that conformally (universally) couples to gravity could alter the moment of inertia of rotating pulsars via a new gravity-mediated force between neutron stars and ultralight dark matter Kuntz and Barausse (2024), which is called the Nordtvedt effect Nordtvedt (1968). If such a coupling exists, the times of arrivals of pulsars would change in a deterministic way, and would also result in a violation of the equivalence principle. This model of conformal DM has been extensively constrained by looking for impacts on the timing solutions of measured binary pulsar Damour and Taylor (1992); Kramer et al. (2006).
Ref. Smarra et al. (2024) uses European pulsar timing array data to constrain linear (Fierz-Jordan-Brans-Dicke (FJBD) theory) Fierz (1956); Jordan (1959); Brans and Dicke (1961); Dicke (1962) and quadratic (Damour-Esposito-Farese (DEF) theory) Damour and Esposito-Farese (1992, 1993) conformal DM couplings to gravity by deriving how the pulsar times of arrivals would change in the presence of the ultralight dark matter fields. In the case of FJBD conformal coupling, the residuals would appear as:
$\Delta t(t)$ $=-\int\mathrm{d}t\frac{\delta\Omega_{\text{obs}}}{\bar{\Omega}}$
$=2\alpha s_{I}\left.\frac{\sqrt{\rho_{\text{DM}}}}{M_{\text{Pl}}m^{2}}\hat{\phi}(\bm{x})\sin(m_{\rm DM}t+\theta(\bm{x}))\right|_{t_{\text{start}}-\frac{d}{c}}^{t_{\text{end }}-\frac{d}{c}},$ (88)
where $s_{I}$ is the angular moment sensitivity parameter computed in Kuntz and Barausse (2024) and is of $\mathcal{O}(1)$ in FJBD theory, $\alpha$ is the strength of the coupling, and $\hat{\phi}$ is a stochastic parameter whose value depends on whether the correlated, uncorrelated or pulsar correlated scenarios are considered. $\theta(\mathbf{x})$ is a random phase.
In Fig. 12, we show the upper limits on the linear coupling parameter $\alpha$ using European pulsar timing array data. However, Ref. Smarra et al. (2024) notes that there is some evidence for additional signal power at two masses: $m\sim 10^{-22.7}$ eV and $m\sim 10^{-21.4}$ eV, but note that other physical processes could responsible for these excesses.
Figure 12: Taken from Smarra et al. (2023). Constraints on the coupling $\log_{10}\alpha$ of conformal DM to gravity at the 95% confidence-level (solid line) for the correlated case compared to the analytical estimate obtained by using the upper limits from the European pulsar timing array search in Smarra et al. (2023) to estimate $\alpha$ (brown dashed lines). $\rho_{\text{DM}}=0.4\,\text{GeV/cm}^{3}$. Here, this model of DM would change the moment of inertia of neutron stars in the used in pulsar timing arrays, resulting in changes of the times of arrival of the pulses. If the true density of DM that conformal DM composes is less than $\rho_{\text{DM}}$, the limits will suffer, which is shown by the dotted line, for values of $\rho$ given in Smarra et al. (2023). AP4 is a particular equation-of-state of neutron stars Akmal et al. (1998).
III.4.5 Constraints on axions through polarization measurements
Changes in polarization of light by axions could also be probed in pulsar timing array experiments Xue et al. (2024); Porayko et al. (2024). Just as in ground-based interferometers, axions could induce cosmic birefrigence, an oscillation of the polarization plane of a linearly polarized source. Since pulsars constantly emit light that is beamed at earth, its polarization can be measured over time, and any changes in it could be attributed to axion-like DM/photon interactions. In Fig. 13, we show upper limits on this coupling from the Parkes Pulsar Polarization Array (PPPA), the first search of its kind. The constraints here outperform existing experiments at low axion masses.
Figure 13: Taken from Xue et al. (2024). Upper limits from analyses of European pulsar timing array data on the coupling of axion-like DM to photons.$\rho_{\text{DM}}=0.4\,\textrm{GeV}/\textrm{cm}^{3}$. The blue and red curves correspond to the limits derived using the auto-correlation-only and full-correlation signal models, respectively. Excluded portions of the axion mass /coupling parameter space lie above these curves, and constraints from different experiments are also shown. Key timescales are shown with vertical dotted lines, included $T_{\text{obs}}=18$ years. The lower panel shows the Bayes factors for each of the scenarios.
III.5 Prospects for detecting the coupling of ultralight dark matter to the interferometers
Though no searches so far have yielded conclusive evidence of any DM/standard model coupling with the interferometers, impressive constraints, relative to existing DM experiments, have been set on these interaction models. The current searches and their constraints have been framed only in terms of upper limits on dilatons, dark photons and DM that couples to gravity; however, they are actually sensitive to any kind of DM interaction that would cause a differential strain in the interferometers. In fact, to derive these constraints, we first compute the minimum detectable signal amplitude for a generic quasi-monochromatic signal as a function of frequency, and then map these values to constraints on the DM/standard model coupling constants. It is thus simply a matter of a few relations that these results are interpreted in terms of two relatively well-motivated DM models.
Of course, ultralight dark matter could take on any mass a priori, and ground-based GW detectors only allow us to probe a small region of this parameter space, from $\sim 10^{-14}-10^{-12}$ eV. Space-based GW detectors will permit being sensitive to masses a few orders of magnitude smaller, while high-frequency gravitational-wave (HFGW) detectors could see DM particles a few orders of magnitude larger Aggarwal et al. (2021). However, it should be noted that when placing constraints on ultralight dark matter, the stochasisity of the signal could affect the limits: if we happen to “get unlucky” and observe the signal when the phase offset reduces the signal amplitude to zero, as can be seen at the zero points in Fig. 2, we would have no constraint. Such considerations were made in Nakatsuka et al. (2023); Boddy et al. (2025) for ground-based interferometers and pulsar timing arrays.
The prospects for future ground-based GW interferometers, such as ET and Cosmic Explorer, to observe ultralight dark matter interactions are bright. Within a decade or so of observation, such detectors, currently a couple of orders of magnitude behind fifth-force constraints, could actually be competitive in this portion of the parameter space, thus permitting an extremely strong probe of new physics Pierce et al. (2018); Morisaki et al. (2021).
Though GW interferometers can probe extremely small DM/standard model couplings, we do not have an estimate of a minimum strength of these interactions. We thus have no guarantee, in any search that we do, that we are getting closer to probing the true coupling value between ordinary matter and the dark sector. Such a caveat is balanced by the relatively model agnostic nature of our searches, which make minimal assumptions about the signal model, and essentially just look for correlated noise. While we probably cannot believe just one model for DM/standard model interactions, we can be sure that our techniques are sensitive to a wide range of theories, even ones that have yet to be thought of.
IV Macroscopic dark matter transiting through the interferometers
IV.1 Background
As mentioned in Section I, the mass of DM can span many orders of magnitude, including being “macroscopic”, that is, from being $M_{\rm DM}\sim[1,10^{9}]$ kg Hall et al. (2018). If DM exists in this mass range, it could transit through GW interferometers, and thus be searched for in LIGO, Virgo and KAGRA data. The event rate, however, depends on the mass of DM, and there are orders of magnitude fewer heavy DM particles in any given region of space than the number of ultralight dark-matter particles Hall et al. (2018):
$\rho_{\text{DM}}=M_{\text{DM}}n_{\text{DM}}~{}~{}\Longrightarrow~{}~{}\frac{L_{\rm DM}}{10^{4}~{}\text{km}}\simeq 1.2\times\left(\frac{M_{\text{DM}}}{1~{}\text{kg}}\right)^{1/3},$ (89)
where $n_{\rm DM}$ is the number density of DM, and $L_{\rm DM}$ is a characteristic length scale for the transiting DM. For 1-kg DM, the flux of DM through a given region of space is given by $\Phi\sim n_{\rm DM}v_{0}\sim 3\times 10^{-10}$ km-2 s-1, meaning that approximately one transient event could be observed per year if the impact parameter of such an event is 10 km, which is similar in size to the three or four kilometer-long arms of LIGO, Virgo and KAGRA.
IV.2 Yukawa-like interactions
If DM only couples to gravity, it seems unlikely that GW interferometers could detect this interaction Hall et al. (2018); Du et al. (2023); however, if there is a Yukawa-like coupling between the standard model and DM, allowing for self-interactions as well, DM transiting through the interferometers could very well be detectable. The following potential $V_{i-j}$ is assumed to govern the (self-) interactions between the standard model and DM
$V_{i-j}=-M_{i}M_{j}\frac{G}{r}\Big{(}1+(-1)^{s}~{}\delta_{i}\delta_{j}\exp[-r/\lambda]\Big{)}$ (90)
[…]
where $\delta_{i,j}$ parametrizes the standard model and DM couplings, $r$ is a length scale, $\lambda$ is the screening length, $s$ is the spin, and $M_{i,j}$ are the two masses. By assuming that the size of DM is much larger than the size of GW detectors and the range of the DM force, but smaller than the average distance between the detectors, the scattering cross section from this process can be derived:
$\sigma_{\text{DM-DM}}=16\,\pi\times\frac{G^{2}\,M_{\text{DM}}^{2}\,\delta_{\text{DM}}^{4}}{v_{0}^{4}}\times\log\left[\frac{\lambda}{r_{\text{DM}}}\right],$ (91)
where $r_{\rm DM}$ is the size of the DM objects. While $|\delta_{\rm SM}|<5\times 10^{-4}$ has been strongly constrained by equivalence principle experiments Adelberger et al. (2009), the DM coupling constant could be much larger, as the only constraints come from how DM self-interaction would have influenced structure formation Spergel and Steinhardt (2000) and collisions between galaxies Harvey et al. (2015).
The lack of observed deceleration of DM in the bullet cluster collision constrained the cross-section of DM to be less than $\sim 1$ cm2/g Randall et al. (2008), which provides limits on the coupling constant of self-interacting DM to be:
$|\delta_{\text{DM}}|\lesssim 5\times 10^{9}\times\left(\frac{1~{}\text{kg}}{M_{\text{DM}}}\right)^{1/4},$ (92)
in which $\log\left[\frac{\lambda}{r_{\text{DM}}}\right]=5$. Based on the current upper limits on $\delta_{\rm SM}$ and $\delta_{\rm DM}$, the coupling of standard-model particles to DM ones could be quite large, and would thus permit an observation of a transit event in GW interferometers. Furthermore, $\delta_{\rm DM}\gg$ 1 would actually alleviate some problems with the cold DM scenario, including overly dense areas of dwarf galaxies found in simulations, and a prediction that the Milky Way produces more stars than are observed Boylan-Kolchin et al. (2011).
IV.3 Projected constraints on Yukawa-like interactions
By using the potential given in Eq. 90, the differential acceleration visible on interferometers can be computed, and from that, the matched-filter signal-to-noise ratio:
[…]
where $a(f)$ is the Fourier transform of the integral of the differential acceleration $a(t)=a_{y}^{\rm EM}-a_{y}^{\rm IM}-(a_{x}^{\rm EM}-a_{x}^{\rm IM})$ between the input mirror (IM) and the end mirror (EM) Saulson (2017), and $S_{n}(f)$ is the acceleration power spectral density of the noise. By using sensitivity curves of LIGO, Virgo and KAGRA and a threshold on $\rho>8$, the number of detections per year of transiting DM with different masses and different screening lengths can be computed, and is shown in Fig. 14 using a power spectral density for advanced LIGO at design sensitivity.
Figure 14: Taken from Hall et al. (2018). Expected number of macroscopic DM transit events per year with $\rho>8$ as measurable with advanced LIGO operating at design sensitivity, as a function of the DM mass, the screening length $\lambda$ and the coupling $g=\delta_{\text{SM}}\delta_{\text{DM}}$ between DM and the standard model.
Further work was done regarding the detection of macroscopic DM in much more detail Du et al. (2023), in which the different kinds of effects of DM on the test masses were derived. Ref. Du et al. (2023) derives the strain on the detector in terms of three observables: (1) the Doppler effect (previously considered in Hall et al. (2018)), (2) the Shapiro delay (light is curved near objects with mass and thus will take a longer time to travel down the interferometer arms, and (3) the Einstein Delay (time runs slower near more massive objects than less massive ones). Unfortunately, all effects are extremely weak because they are solely gravitational, and thus DM would have to couple to the standard model in some way to be detectable. As in Hall et al. (2018), Ref. Du et al. (2023) computes the expected constraints on the effective coupling of DM to the standard model with an effective parameter $\tilde{\alpha}\sim\delta_{i}\delta_{j}$, which are shown in Fig. 15 for a variety of GW detectors and for two choices of screening lengths: $\lambda=1$ m and $\lambda=10^{6}$ m. Existing constraints from the bullet cluster Spergel and Steinhardt (2000), MICROSCOPE Bergé et al. (2018) and neutron-star kinetic heating with and without interactions between DM and baryons that could enhance energy transfer Gresham et al. (2023). While the constraints that could be obtained from GW interferometers do not appear to beat existing ones, it should be noted that if only a part of DM can be explained by a new fifth force, the Bullet cluster constraints on self-interacting DM disappear, while GW bounds fall off linearly with this fraction of DM charged under this fifth force Du et al. (2023).
Figure 15: Taken from Du et al. (2023). 90%-projected sensitivity on the standard model/DM interaction parameter $\tilde{\alpha}\sim\delta_{\rm SM}\delta_{\rm DM}$ that can be probed from macroscopic DM transiting through several GW experiments, compared to existing constraints from others. One year of observation is assumed, and two choices for the screening length are made. Dotted lines denote existing constraints. Existing constraints are shown in dotted lines. In particular, Bullet cluster and MICROSCOPE constraints are shown in gray Spergel and Steinhardt (2000); Bergé et al. (2018), MICROSCOPE ones. The purple and red lines denote constraints from neutron star kinetic heating Gresham et al. (2023) with and without additional short-range interactions between baryons and DM that would enhance energy transfer.
Given that such deterministic signals may not be detectable in LIGO, Virgo and KAGRA, it was worth considering whether a stochastic background arising from many transiting DM events could be seen Hall et al. (2018); Du et al. (2023). However, again, it is necessary to include the Yukawa interaction to have a chance of detecting a stochastic background.
Finally, another study Lee et al. (2023) considered specifically how heavy transiting DM would appear in KAGRA, as well as the directional dependence of the signal-to-noise ratio, which turns out to be strong due to the geometry of the detector. In KAGRA, it is shown that the suspension system can be particularly sensitive to DM-standard model interactions, between [1,130] Hz.
V GWs from ultralight bosonic clouds around rotating BHs
If a new ultralight boson [^5] In the context of boson clouds around rotating black holes, this new particle could, but need not, be DM exist, it will have a particular mass, frequency and Compton wavelength. Quantum fluctuations could then allow bosons to pop into existence near a rotating BH, whose characteristic length scale is given by half of its Schwarzschild radius. Some bosons will fall in; others will scatter off, and the extent to which either of these scenarios happens depends on the properties of the black hole and the boson. If the following superradiance condition, relating the BH angular frequency $\Omega$ to the boson angular frequency $\omega$ via the magnetic quantum number $m$, is met Brito et al. (2015):
$\omega<m\Omega,$ (94)
the boson will extract energy from the BH, and thus the “bosonic wave” will be amplified. This amplification will be maximized if the Compton wavelength of this new particle comparable to the characteristic length of a BH. The massive bosons then become bound to the BH, allowing for successive scatterings, thus permitting a huge number of bosons to appear around the BH as a cloud in a given energy state (no limit on the occupation number of these particles exists a priori).. This process is called “superradiance”.
More quantitatively, solving the Klein-Gordon equation for a Kerr BH hole in the presence of a massive scalar field results in a Schrondinger-like equation that exhibits a $1/r$ potential, resulting from the gravitational interaction between the bosons and the BH. Such an equation can be solved to show that the energy states of the boson cloud are quantized analogously to those of the hydrogen atom, allowing us to describe the boson cloud as a “gravitational atom in the sky” Baumann et al. (2019a).
Qualitatively, superradiance will occur regardless of the spin of the boson; however, the timescales to build the cloud, as well as to deplete it, will differ. In all cases, however, as we will see, the GW emission timescale will be shorter than the time to build up the cloud, making these systems excellent sources of GWs.
For the black hole to be spun down such that the condition in Eq. 94 is no longer satisfied, scalar and vector clouds require Isi et al. (2019):
$\tau_{\rm inst}^{\rm(s)}\approx 27\,{\rm days}\left(\frac{M_{\rm BH}}{10\,M_{\odot}}\right)\left(\frac{0.1}{\alpha}\right)^{9}\frac{1}{\chi_{i}}\,,$ (95)
and
$\tau_{\rm inst}^{\rm(v)}\approx 2\,{\rm minutes}\left(\frac{M_{\rm BH}}{10\,M_{\odot}}\right)\left(\frac{0.1}{\alpha}\right)^{7}\frac{1}{\chi_{i}}\,,$ (96)
where $\chi_{i}$ is the BH spin at birth, and the GW fine-structure constant is
[…]
where $m_{b}$ is the mass of the boson. This new boson could have a spin of 0, 1 or 2, where spin-0 could correspond to the pseudo-scalar QCD axion (or axion-like) particle Brito et al. (2015, 2020), spin-1 could indicate dark photons arising through kinetic mixing with the ordinary photon Siemonsen et al. (2023a), or spin-2 tensor bosons that could arise from modified theories of gravity Brito et al. (2020).
After the cloud builds up, GWs can be emitted in three different ways: (1) through boson-boson annihilation, (2) through boson energy-level transitions, and (3) after a “bosenova”.
It has been shown, under a wide range of assumptions of BH populations, spins, redshift distributions, etc., that such systems would be detectable by CW searches in advanced LIGO, Virgo and KAGRA data and/or in future detectors, such as ET and Cosmic Explorer Arvanitaki et al. (2015b); Zhu et al. (2020). Furthermore, both galactic BHs Palomba et al. (2019) and remnants of mergers of BHs outside the galaxy could also be detected, the latter of which, in the scalar case, only being visible in ET and Cosmic Explorer Arvanitaki et al. (2010, 2015b); Isi et al. (2019). Annihilating vector boson cloud systems, however, emit stronger, though shorter, GW signals Brito et al. (2015); Baryakhtar et al. (2017); Siemonsen et al. (2023b); Jones et al. (2023, 2024), and potentially may contain electromagnetic counterparts if the vector boson kinetically mixes with the ordinary photon Siemonsen et al. (2023a).
The boson mass is fixed by the GW frequency, and could in principle take on any ultralight mass. That is why GW experiments across the frequency spectrum have been or are being designed to search for these boson cloud systems. While most of the literature focuses on ground-based GW interferometers, we note that HFGW detectors will also be able to probe BH superradiance of heavier bosons than those detectable by LIGO, Virgo and KAGRA Aggarwal et al. (2021), e.g. Brzeminski et al. (2024), and space-based GW interferometers will probe lighter masses Arvanitaki et al. (2015b); Brito et al. (2017), and space-based detector analyses can be used to aid in discovering boson clouds in the band of ground-based GW interferometers Ng et al. (2020).
As will be discussed, boson cloud self-interactions could have a significant impact on (1) whether the superradiance condition is saturated Yoshino and Kodama (2012, 2015a, 2015b), (2) how quickly a binary black hole system will approach merger Aurrekoetxea et al. (2024a); Takahashi et al. (2024), and (3) the rate of change of the GW frequency of an isolated boson cloud Baryakhtar et al. (2021). Self-interactions would also allow higher energy levels to grow faster. The strength of these interactions will affect the detection prospects in ET and Cosmic Explorer Collaviti et al. (2025), as well as recent search results which have not yet accounted for moderate or strong interactions.
Note that while spin-0 and spin-1 boson clouds are relatively well understood theoretically, spin-2 clouds are much more complicated. The superradiant instability of spin-2 fields occurs much faster than the others East and Siemonsen (2023), and since the backreaction of the instability has not been computed due to difficulties arising from a nonlinear coupling between the spin-2 particle and gravity, it is not yet known whether the observational signatures are similar to the vector or scalar cases, and whether there is a dependency on the specific nonlinear theory chosen.
V.1 Annihilation signal
When Eq. 94 is no longer satisfied, the cloud stops growing. Since ultralight bosons are their own anti-particles, they will start to annihilate, producing GW radiation at a frequency fixed by the boson mass and slightly redshifted by the BH mass (in the non-relativistic limit $\alpha\ll 1)$, which predominately applies to the equations shown in this work). In the scalar boson case Abbott et al. (2022d):
$f_{\text{gw}}\simeq 483\,{\rm Hz}\left(\frac{m_{\text{b}}}{10^{-12}~{}\text{eV}}\right)\times\left[1-7\times 10^{-4}\left(\frac{M_{\text{BH}}}{10M_{\odot}}\frac{m_{\text{b}}}{10^{-12}~{}\text{eV}}\right)^{2}\right].$ (98)
Qualitatively, the GWs emitted are monochromatic because the mass of the boson is fixed, and all the bosons are in the same energy state. As alluded to before, the timescales for GW emission are much longer than the time it takes to build the cloud Isi et al. (2019):
$\tau_{\rm GW}^{\rm(s)}\approx 2\times 6.5\times 10^{4}\,{\rm yr}\left(\frac{M_{\rm BH}}{10\,M_{\odot}}\right)\left(\frac{0.1}{\alpha}\right)^{15}\frac{1}{\chi_{i}}.$ (99)
$\tau_{\rm GW}^{\rm(v)}\approx 2\times 1\,{\rm day}\left(\frac{M_{\rm BH}}{10\,M_{\odot}}\right)\left(\frac{0.1}{\alpha}\right)^{11}\frac{1}{\chi_{i}}.$ (100)
where we have explicitly noted a factor of 2 that is missing from the original reference Isi et al. (2019), as pointed out in Siemonsen et al. (2023b). There will also be a classical depletion of the cloud over time, as the number of bosons starts to decrease, resulting in the following drifts in frequency over time Isi et al. (2019):
$\dot{f}_{\rm gw}^{\rm(s)}\approx 3\times 10^{-14}\,{\rm Hz/s}\left(\frac{10M_{\odot}}{M_{\rm BH}}\right)^{2}\left(\frac{\alpha}{0.1}\right)^{19}\chi_{i}^{2}\,,$ (101)
$\dot{f}_{\rm gw}^{\rm(v)}\approx 1\times 10^{-6}\,{\rm Hz/s}\left(\frac{10M_{\odot}}{M_{\rm BH}}\right)^{2}\left(\frac{\alpha}{0.1}\right)^{15}\chi_{i}^{2}\,.$ (102)
We note that relativistic computations of the frequency drift have also been performed Siemonsen et al. (2023b); May et al. (2024). In the non-relativistic limit $\alpha\ll 1$, the signals will have the corresponding amplitudes of Isi et al. (2019):
$h_{0}^{\rm(s)}\approx 8\times 10^{-28}\left(\frac{M_{\rm BH}}{10M_{\odot}}\right)\left(\frac{\alpha}{0.1}\right)^{7}\left(\frac{\rm Mpc}{r}\right)\left(\frac{\chi-\chi_{f}}{0.1}\right)$ (103)
$h_{0}^{\rm(v)}\approx 4\times 10^{-24}\left(\frac{M_{\rm BH}}{10M_{\odot}}\right)\left(\frac{\alpha}{0.1}\right)^{5}\left(\frac{\rm Mpc}{r}\right)\left(\frac{\chi-\chi_{f}}{0.1}\right)$ (104)
where the final spin $\chi_{f}$ is
$\chi_{f}=\frac{4\alpha_{f}m}{4\alpha^{2}_{f}+m^{2}}\,$ (105)
These equations indicate interesting properties for scalar and vector boson cloud annihilation signals. For bosons whose Compton wavelengths are optimally matched to the size of the BH, scalar bosons will annihilate over timescales that greatly exceed the observation time of ground- or space-based GW interferometers, and their frequencies will hardly drift over time. Thus, this signal can be thought of as a quasi-sinusoidal and persistent. On the other hand (and with the same caveat of optimal matching), the GW signal from annihilating vector boson clouds lasts significantly shorter than the scalar one, and thus can be thought of more as a “transient” CW, lasting for durations for $\mathcal{O}($hours-days), i.e. much longer than mergers of black holes routinely detected now by LIGO, Virgo and KAGRA, but much shorter than scalar boson cloud annihilation signals. Furthermore, vector signals emit much more GW power per unit time than scalar ones, but for scalar signals, analysis methods are able to integrate over the entire data collection period of LIGO, Virgo and KAGRA to improve the signal-to-noise ratio. If, in the vector case, the size of the BH does not match the Compton wavelength of the boson, the signals could resemble CWs.
V.2 Transition signal
[…]
$f_{\rm gw}\sim 15\text{ Hz}\left(\frac{m_{b}}{10^{-11}\text{ eV}/c^{2}}\right)$ (106)
and a frequency drift over time that relates to the duration of the signal $T$ and the axion decay constant $f_{a}$ Arvanitaki et al. (2015b):
$\!\!\!\!\frac{df}{dt}\!\!\simeq 10^{-11}\!\frac{\,\mathrm{Hz}\,}{\mathrm{s}}\!\left(\!\frac{f_{\text{gw}}}{90\,\mathrm{Hz}\,}\!\right)\!\!\left(\!\frac{M_{\rm BH}}{10M_{\odot}}\!\right)\!\!\left(\!\frac{10^{17}\!\,\mathrm{GeV}}{f_{a}}\!\right)^{\!2}\!\!\left(\!\frac{5\mathrm{yr}}{T_{\text{obs}}}\!\right)^{\!2}\!\!$ (107)
Unfortunately, the expected event rate of transitions in ground-based detectors is not expected to exceed one even in $\sim 3.5$ years of observing time in advanced LIGO (Fig. 5 in Arvanitaki et al. (2015b)). But, if self-interactions occur, transitions could happen and be detectable Collaviti et al. (2025).
In the vector boson case, transitions could also occur, and the overtone modes could actually grow faster than the fundamental modes at small angular numbers and for fine-structure constants and BH spin values Siemonsen and East (2020). The higher overtones could saturate much quicker than lower ones, implying that transitions to lower states could occur, thus allowing BHs to become unstable to increasing energies at lower modes. Different energy states of the boson cloud could thus be populated at the same time, enriching the kind of physics searchable via transitions.
V.3 Bosenova signal
If the self-interactions between bosons become stronger than the gravitational binding energy of the cloud, the cloud starts to deform and can no longer be described analogously to the Hydrogen atom Fukuda and Nakayama (2020); Collaviti et al. (2025); Omiya et al. (2024); Jones et al. (2024). Instead, it can collapse and explode in what is known as a “bosenova”. This phenomenon would occur if the potential energy of the cloud is approximately equal to the self-interaction energy. Bosenovas have been observed on earth in condensed matter laboratories, in which repulsive forces were switched to attractive ones in a material that formed a Bose-Einstein condensate via a tuneable external magnetic field Donley et al. (2001). There are a lot of uncertainties around whether the bosenova would occur, and it seems that including self-interactions in the cloud limits bosenovas to a corner of the allowed parameter BH-boson mass parameter space Baryakhtar et al. (2021).
In particular, in the axion case, originally proposed in Arvanitaki and Dubovsky (2011), bosenovas would likely not occur. While early papers assumed that the boson cloud reached large enough amplitudes to produce GW emission Yoshino and Kodama (2012, 2015a, 2015b), subsequent studies in the non-relativistic Baryakhtar et al. (2021) and relativistic Omiya et al. (2023) regimes confirmed that dissipative nonlinear effects would prevent large-enough cloud amplitudes from occurring to cause a bosenova. In the vector case, bosenovas could actually occur if strings are formed the Higgs-Abelian mode East (2022); East and Huang (2022); Brzeminski et al. (2024).
Though GWs from boson clouds described above may seem independent, they could in principle be happening at the same time. In particular, if ultralight scalar or vector bosons have some self-interactions, the GW signal will change. If these clouds are composed of string axions (as opposed to the QCD axion), a non-linear self-interaction may cause a bosenova if, during superradiance, the value of the scalar field approaches the axion decay constant $f_{a}$, which, for string axions, could be below the grand unification theory (GUT) scale Arvanitaki et al. (2010); Yoshino and Kodama (2012, 2015a, 2015b). This bosenova would halt the superradiance process, as about 5% of the boson cloud mass would fall back into the BH, but superradiance would immediately start again until the next bosenova is triggered. Such periodic bosenovas could sustain superradiance for much longer durations than those in the conventional scenario, permitting annihilating clouds even around extremely old BHs. Furthermore, we note that boson annilation could alter the GW of an inspiraling binary, potentially visible in space-based GW detectors Xie and Huang (2024), in a similar way that dynamical friction does.
In more general models of axions than in the axiverse, self-interactions could suppress annihilation signals as they get stronger, and transitions could occur for moderate interaction strengths. In particular, GW power sourced by annihilation would be reduced due to the smaller cloud size that can be sustained because of self-interactions, while scalar radiation, a new energy loss mechanism, consumes a portion of the energy that would source GW emission. But, multiple levels could be populated with self-interactions, allowing transitions to occur more than in the non-self-interaction scenarios Baryakhtar et al. (2021).
V.4 Clouds in binary black hole systems
We know that BHs form binaries, and if ultralight dark matter exists in the form of ultralight bosons, it should imprint some signature in the GWs arising from inspiraling binary black holes Baumann et al. (2019a). The GW signal will be modified due to tidal disruptions induced by the companion, and the multiple moments of the cloud Baumann et al. (2019a, b, 2020). These kinds of signals could be visible in future space-based detectors, since the deviations primarily occur in the early inspiral stages of the system, i.e. at frequencies much lower than currently accessible by ground-based GW interferometers. In both cases, observations of the inspiral of binary systems could indicate the presence of boson clouds around one or both of the objects. Additionally, once the binary reaches a certain stage of its evolution, the impact of both the spin-induced quadrupole moment and tidal forces will lessen, since the clouds will become disrupted by their companions. If we can observe variations in both the spin-induced quadrupole moment, parameterized as $\kappa$, and the tidal deformability, parameterized as $\Lambda$, over the course of the lifetime of the binary, it would indicate the presence of a (disrupted) boson cloud Baumann et al. (2019a). More recently, it has been show that the boson cloud could be “ionized” as the orbital separation approaches the size of the cloud. Self-interactions could also play a role in the evolution of a binary, in which two types of clouds can actually coexist around the rotating BH through mode coupling Boudon et al. (2024); Aurrekoetxea et al. (2024a); Takahashi et al. (2024). We will now describe different ways in which boson clouds could affect GWs from binary black hole systems.
V.4.1 Spin-induced quadrupole moment
The spin-induced quadrupole moment is often parameterized in terms of a dimensionless quantity $\kappa=\frac{Q}{J^{2}/M_{\rm BH}}$, which has been normalized by the object’s angular moment $J$ and mass BH. This parameter enters into the waveform at the second post-Newtonian order. Thus, the spin-induced quadrupole moment, regardless of its source, has a significant impact on the GW signal from compact objects inspiraling towards one another.
Ref. Krishnendu et al. (2017) showed that $\kappa$ can be used as a distinguishing parameter for different kinds of objects, and also a test of the no-hair theorem: as an example, $\kappa=1$ for BHs Hansen (1974), while $\kappa\sim[1.4,8]$ for neutron stars depending on the equation-of-state Laarakkers and Poisson (1999); Pappas and Apostolatos (2012). In the case of boson cloud systems, $\kappa$ will take on vastly different values, of $\mathcal{O}(1000)$, depending on the boson cloud and BH masses, and the extent to which the Kerr metric is altered in the presence of a cloud with enough mass compared to the BH.
V.4.2 Tidal disruptions
Because the boson cloud is much less compact than the BH, we expect that tidal disruptions from the companion could occur when the two objects in the binary black hole are close enough to each other Ng et al. (2020). To characterize the effects of tides in a binary system, the tidal deformability parameter $\Lambda$ is used, which depends on the mass and radius of the compact objects. For BHs, this quantity is 0, while for neutron stars, it can take on values of $\mathcal{O}(100-1000)$ Chatziioannou (2020). For boson clouds, $\Lambda\sim 10^{7}$, which is a remarkable departure from that expected from conventional astrophysical binaries Baumann et al. (2019b). In Payne et al. (2022), the authors considered the impact that superradiance would have on hierarchical BH mergers, and showed that the reduction of spin in BHs implies smaller recoil velocities of binary black hole mergers. Smaller recoil velocities imply that remnants could remain inside clusters, leading to higher chances of hierarchical formation and merger of binary black holes. Such disruptive effects of bosons on not just individual BHs but also on the BH population, imply that boson clouds, if they exist, would significantly impact the formation and evolution histories of BHs.
V.4.3 Binary-induced transitions
Over the course of the inspiral, the presence of the boson cloud will induce transitions between different energy levels of the boson clouds that would be visible in space-based GW detectors if (1) the orbital frequency matches the hyperfine splitting between two energy levels (hyperfine resonance), and (2) the growing mode mixes with decaying modes (Bohr resonance). Both kinds of transitions would also be smoking-gun signatures of the presence of boson clouds around rotating BHs in a binary Baumann et al. (2019a). More recently, by considering the slow evolution of the orbital frequency through resonances of the boson cloud system, that is, when the orbital frequency matches the energy spacing between adjacent levels, it can be shown that Landau-Zener transitions occur and significantly affect the inspiral GW signal Baumann et al. (2020).
V.4.4 Ionization
Another transition arises when the binary system separation becomes comparable to the size of the cloud: an “ionization” of the boson cloud occurs as the cloud becomes unbounded from its host BH Takahashi et al. (2022); Baumann et al. (2022a, b). The energy from the binary sources this effect, and is significantly larger than that emitted via GWs; thus, the inspiral of the binary becomes driven by the boson cloud, instead of simply perturbed by it. Ionization will also tend to circularize the binary if it formed through dynamical capture, while leaving any orbital inclinations unaffected Tomaselli et al. (2023). It has been recently found, however, that as the resonances in the boson cloud evolve during the orbit, the boson cloud is unperturbed and ordinarily visible in LISA if the cloud and binary are counter rotating; otherwise, the cloud is destroyed because of the resonances, but a distinct mark is left on the binary. That is, the binary is forced to co-rotate with the boson cloud, and its eccentricity is driven to a particular value, which allows the possibilty of doing statistics with detected binary black hole systems in LISA to probe the boson cloud hypothesis Tomaselli et al. (2024).
VI GW constraints on boson clouds
The extensive theoretical background on boson clouds around rotating BHs, coupled with the potential for them to be DM, have motivated GW probes of boson cloud/BH systems. Different GW interferometers probe vastly different mass regimes of ultralight dark matter boson clouds. While ground-based detectors are sensitive to annihilation signals from bosons of masses between $\sim 10^{-14}-10^{-12}$ eV Arvanitaki et al. (2015b), space-based detectors can probe a few orders of magnitude lower than that, $\sim 10^{-20}-10^{-16}$ eV Baumann et al. (2019a), and HFGW detectors could see systems with masses around $\sim 10^{-9}$ eV Aggarwal et al. (2021). At the moment, only ground-based GW interferometers exist, and so we focus on methods and search results from the most recent LIGO, Virgo and KAGRA observing runs.
VI.1 Constraints from the spins of detected binary black holes
We note that constraints on boson clouds can be inferred from the spin measurements of binary black hole systems Arvanitaki et al. (2015b, 2017); Brito et al. (2017); Baryakhtar et al. (2017); Cardoso et al. (2018); Ghosh and Sachdeva (2021). Recently, a Bayesian method has been developed by combining the information from each system Ng et al. (2021a), assuming a distribution for the initial BH spins and requiring that the BHs have had enough time to undergo superradiance before merging. Such a study has been applied to the second GW transient catalog (GWTC-2) Abbott et al. (2021b), and has resulted in essentially a joint posterior on the scalar boson mass and the spin distribution of BHs at their formation Ng et al. (2021b). The latter is required because BHs born with high spins that have been spun down by superradiance is partially degenerate with BHs simply born with low spins in a universe without superradiance. The study presented in Ng et al. (2021b) tentatively rules out scalar bosons in the mass range $[1.3,2.7]\times 10^{-13}$ eV.
VI.2 Search methods
While the aforementioned constraints from measured BH spins can disfavor particular masses or even confirm the presence of an ultralight boson, they rely on assumptions about the spin distribution at formation, on whether the boson clouds have time to form (which could be impacted by eccentricity at formation Peters (1964); Wen and Gair (2005)), and are a function of the (strongest) BH mergers that we see with GWs. To that end, different methods to actually search for long-lived GWs from isolated BHs (or those present in a known x-ray binary system) have been developed and used in real searches in LIGO, Virgo and KAGRA data. Detecting GW emission from an isolated BH would provide smoking-gun evidence of the existence of an ultralight bosonic field around BHs.
VI.2.1 Excess power
One method to search for boson clouds around rotating BHs lightly models the boson cloud signal as monochromatic within the entire observation time that could come from anywhere in the sky, hence inducing a Doppler modulation. This method creates time-frequency maps and corrects each of them for different sky positions, leaving behind a monochromatic signal D’Antonio et al. (2018). It employs different coherence times $T_{\text{FFT}}$ for the analysis in order to be robust against possible theoretical uncertainties in the GW signal model. Shorter $T_{\text{FFT}}$, while less sensitive than longer ones, provide larger frequency bins that allow the signal frequency to wander more, and thus are more sensitive to deviations from the monochromatic signal model.
In LABEL:fig:tfmaps, we show the core stage of this method, which involves creating the time-frequency peakmap from raw time-domain strain data, correcting for the Doppler modulation from a particular sky location, and then integrating over time to collect all peaks at a given frequency to highlight the presence of a signal. Note that while we show the color to indicate the ratio of signal to noise power, the method does not actually use it: it only requires that a given point in the time-frequency plane is above a given threshold. When moving from LABEL:fig:afterdopp to LABEL:fig:pmproj, all time-frequency points above this threshold are given a weight of “1”, and the ones are summed over time in order to create this histogram.
VI.2.2 Viterbi
Another method that is more robust against stochastic frequency variations relies on Hidden Markov Models (HMMs) to find arbitrary tracks in the time-frequency plane. The main idea is to model a time-frequency GW signal probabilistically as a Markov chain of transitions between “hidden” (unobservable) frequency states and use a detection statistic to relate “observed” frequency states with the hidden ones Quinn and Hannan (2001). In the framework for this method, the probability to jump to a new state at a given time only depends upon the previous hidden state. In total, there would be $N_{Q}^{N_{T}+1}$ possible paths through the hidden states, where $N_{Q}$ is the total number of hidden states and $N_{T}$ is the total number of times at which we have observed frequency states.The Viterbi algorithm provides a recursive, computationally efficient way of maximizing the probability that the hidden set of states is responsible for the observed sequence of states.
In LABEL:fig:sample-path, taken from Isi et al. (2019), we show the optimal path found by the Viterbi method for injected boson cloud signals in white noise that has an associated random walk of the frequency, which is meant to simulate unknown theoretical features of the signal. We can see that in both cases, with less and more variability of the signal, the Viterbi algorithm can find the appropriate track of the signal.
Recently, the Viterbi algorithm has also been adapted to search for GWs from annihilating vector boson clouds, which have timescales significantly shorter than in the scalar case Jones et al. (2023, 2024), and for which numerical waveforms must be used to gauge sensitivity Siemonsen and East (2020); Siemonsen et al. (2023b).
VI.2.3 Cross-correlation and other methods
We have already discussed in Section III.1.1 cross correlating detector data to search for ultralight dark matter that could interact directly with GW interferometers; however, it is important to note that such a method could also be used to search for boson clouds around rotating BHs, both deterministically and for a stochastic GW background composed of the superposition. Furthermore, virtually any method in the CW community could be tuned to look for such boson cloud systems because such methods already analyze almost monochromatic signals focusing on spin-downs but could also extend their parameter spaces to include spin-ups.
VI.3 Recent search results for scalar boson clouds
We can choose different BHs to point at depending on whether we are looking for scalar or vector clouds, or annihilation or transition signals. Current GW interferometer and algorithm sensitivities indicate that annihilation signals from scalar boson clouds could be observed for galactic and relatively young black holes Abbott et al. (2022d). Or, particular, known black hole systems, such as Cyg-X1, could be targeted for such clouds as well Sun et al. (2020). Finally, the stochastic GW background, composed of the superposition of signals from all of the annihilating boson cloud systems in the universe, can be probed as well through cross-correlation analyses. We go through each of these scenarios in the following subsections.
VI.3.1 All-sky search constraints
The excess power method described in Section VI.2.1 has been used to analyze LIGO O3 data to search for the presence of annihilating scalar boson clouds around isolated BHs within our galaxy Abbott et al. (2022d). This search was very computationally expensive, since each sky position had to be analyzed individually. A separate search towards the galactic center was also performed, targeting both spinning down neutron stars and spinning up boson cloud systems Abbott et al. (2022e) No evidence for scalar boson cloud systems were found in either case, and thus upper limits were placed on the presence of boson clouds with different ages, spins, and distances away from earth in two different ways Palomba et al. (2019); Dergachev and Papa (2019); Abbott et al. (2022d):
  1. 1.
    Exclude the existence of BH/boson cloud pairs certain distances away from us with particular spins and ages.
  2. 2.
    Assume mass and spin distributions for BHs in the universe, draw different BHs from these distributions, and determine the maximum distance reach that could be attained for different boson cloud masses.
The first way is agnostic towards whether such BHs exist in the galaxy, while the second one contains some (uncertain) astrophysics regarding the properties of astrophysical BHs, but allows for some physical intuition into the reach of such searches.
Constraints from the first way are shown in Fig. 18 for two different distances (1 kpc and 15 kpc), for BHs with spin of 0.9 and for a few ages of the BH/boson cloud system. The search on LIGO O3 data can thus exclude the existence of BH/boson cloud systems that have these particular combinations of parameters at certain distances away from us.
Figure 18: Taken from Abbott et al. (2022d). Using upper limits from the O3 LIGO search for GWs from annihilating boson cloud systems, exclusion region of particular boson mass /BH mass pairs are derived assuming different ages and distances (left: 1kpc; right: 15 kpc) for BHs with initial spins of $\chi_{i}=0.9$. This is the most model-agnostic way to probe boson clouds around rotating BHs using upper limits obtained from an all-sky search for CWs from isolated BHs.
Constraints from the second way are shown in Fig. 19, in which a Kroupa BH mass distribution $f(m)\propto M_{\rm BH}^{-2.3}$ Elbert et al. (2018) between $[5,~{}50]M_{\odot}$ and $[5,~{}100]M_{\odot}$ is assumed, as well as a uniform initial BH spin distribution between [0.2,0.9]. For fixed BH age, the maximum distance reach as a function of boson mass indicates that at least 5% of GWs from BH/boson cloud systems would have been detected by the search, i.e. their strain amplitudes are larger than the minimum detectable amplitudes at a given confidence level (the upper limits) of the search. As expected, on average when $50M_{\odot}$ is the maximum BH mass considered, the distance that can be reached is also smaller, since the GW amplitudes scales strongly with the BH mass.
Figure 19: Taken from Abbott et al. (2022d). Using upper limits from the O3 LIGO search for GWs from annihilating boson cloud systems, the maximum distance that the search in Abbott et al. (2022d) could reach at which at least 5% of a simulated BH population would have been detectable, that is, with a strain greater than the value of the upper limit, is shown here. To create this plot, Ref. Abbott et al. (2022d) assumed a uniform spin distribution between [0.2,0.9], different ages, and a Kroupa mass distribution with maximum mass $50M_{\odot}$ (left) and $100M_{\odot}$ (right).
We note that the search constraints present in, for example, Steltner et al. (2023); Ming et al. (2024); Covas et al. (2024), could also be cast in terms of all-sky or directed search constraints on scalar or vector boson clouds, but with particular caveats that would have be worked out depending on the actual search parameter space.
VI.3.2 Cygnus X-1 constraints
Cygnus x-1 is an x-ray binary that is close by, has a large spin, $\sim 0.9$, and has relatively well-measured orbital parameters, which make it a reasonable BH to target. Since the sky position is known, searches based on the Viterbi algorithm discussed in Section VI.2.2 can use much longer coherence times than those used in Abbott et al. (2022d), thus permitting exquisite sensitivity towards almost monochromatic signal, while also being robust to theoretical uncertainties in boson cloud source.
In Fig. 20, we show constraints on the strain amplitude arising from annihilating scalar boson cloud particles as a function of the boson mass. The search results $h_{0}^{95\%}$ are given as the black line, while the two curves represent the numerically obtained strain from this system, accounting for the spin, age, mass and distance of Cyg x-1 ($t_{\rm age}=5\times 10^{6}$ yr and $1\times 10^{5}$ yr). The region in which the search results are below the theoretical curves denote boson masses that are disfavored as having formed around Cyg x-1.
All constraints presented so far do not assume any boson self-interactions. If self interactions do occur, the available parameter space to probe would change. But, if the self-coupling is weak enough, standard searches can probe this scenario without modification. An example for the string axiverse scenario is shown in Fig. 21. The white region indicates the portion of the parameter space in which bosenovas do not occur, and thus cannot be constrained with the current search, while, again, portions of the colored parameter space above the black line indicate that those boson masses are disfavored. However, these constraints contain a number of caveats, namely: (1) the periodic bosenovas do not alter the time-frequency evolution of the GW signal, though they actually might induce oscillations every few minutes Yoshino and Kodama (2015a), and (2) the signal must last for longer than the age of the BH. Future studies are needed to determine to what extent CW searches can actually probe this particular model for string axion boson clouds around BHs, as well as to incorporate self-interactions more systematically into how CW searches are performed.
Figure 20: Taken from Sun et al. (2020), a search using O2 LIGO data for annihilating boson cloud systems around Cyg X-1. (a) Upper limits at 95% confidence on the strain amplitude of annihilating scalar boson clouds coming from the BH in Cyg X-1. The black curve indicates the upper limits from the search on O2 data, with the red stars denoting upper limits derived in frequency bands with injections. The two different curves correspond to different choices for the (unknown) age of the system (lower curve: $10^{6}$ years; upper curve: $5\times 10^{6}$ years. The GW fine structure constant is colored, and the gray region denotes boson masses disfavored under the two different age assumptions. $M=14.8M_{\odot}$, $\chi_{i}=0.99$, and $d=1.86$ kpc for the two curves.
Figure 21: Taken from Collaviti et al. (2025). Constraints on the string axiverse scenario (in blue from O2), in particular the axion decay constant. The regions are divided into three regimes: (1) strong-interactions (above the orange dashed-line), (2) intermediate interactions (between the dashed-blue and dashed orange lines) and no self-interactions (below the dashed-blue line). The three plots correspond to annihilations in different energy levels (left and middle) and to transitions between these energy levels (right). Sensitivities of future detectors are also shown on this plot for comparison.
VI.3.3 Constraints on luminous dark photons clouds
As alluded to in Section V, if the dark photon DM kinetically mixes with the ordinary photon and forms clouds around rotating BHs, it could produce both electromagnetic and GW signals, which, if detected, could enable a major multi-messenger discovery Siemonsen et al. (2023a). This additional coupling implies that soon after exponential growth of the boson cloud begins, electromagnetic fields act on charged particles in the vicinity of the BH, producing electron/positron pairs essentially out of vacuum, i.e. through the photon-assisted Schwinger mechanism Sauter (1931); Heisenberg and Euler (1936); Schwinger (1951). At this point, a pair-production cascade ensues Dunne et al. (2009); Monin and Voloshin (2010), sourced by the electromagnetic fields, creating a plasma around the BH in a state of turbulent quasi-equilibrium. This plasma emits copious amounts of electromagnetic radiation, sourced by energy injections through dissipative processes in the cloud (e.g. at magnetic reconnection sites, Landau damping, turbulence), potentially even periodically. The luminosity of these sources depends on the strength of the kinetic mixing parameter, but could be at the level of supernovae or known pulsars, and it could be periodic Siemonsen et al. (2023a). If the dark photon has a mass around $\sim 10^{-12}$ eV, it will induce the formation of clouds around stellar-mass BHs, and give off radio emission that would look remarkably similar to, and potentially indistinguishable from, known pulsars in the ATNF catalog Manchester et al. (2005).
The lack of observation of CWs from pulsars, and the potential of dark-photon clouds to act as “pulsar mimickers”, motivates a search for such clouds around targets in the ATNF catalog that could be mislabeled as pulsars. In Ref. Mirasola et al. (2025), the authors consider this possibilty by analyzing a subset of, spinning up “pulsars” from the ATNF catalog that might actually be but stellar-mass BHs with a dark-photon luminous cloud around them. Using a variety of methods – one that assumes perfect phase coherence between the observed electromagnetic radiation and GW emission, one that allows small deviations in frequency around a narrow band of the observed frequency of each compact object, and one that is most robust against theoretical uncertainties in the signal but least sensitive to GW emission –, upper limits on the dark photon/ordinary photon kinetic mixing parameter are set as a function of boson mass for the 34 sources considered in this analysis, as shown in Fig. 22. To obtain this plot, particular BH populations and electromagnetic emissions are assumed to determine the number of events on the color bar, so the bounds are not necessarily robust against changes in these choices. For the BH population, the relevant distributions are over position in the galaxy (black holes follow the stellar distribution) and mass (Salpeter), spin (uniform) and age (uniform) of the BH systems.
Though particular population models are assumed, the constraints only require ten “events” to have been observable, which is conservative with respect to the potentially $\mathcal{O}(10^{3})$ that would be expected at such coupling/mass combinations, as shown in the right-hand panel of Fig. 3 of Mirasola et al. (2025). Here, “events” just means the number of rotating BHs that could have given off a detectable signal, that is, with a GW amplitude larger than the upper limit strain amplitude, in an assumed population of $10^{8}$ BHs in the galaxy. In the end, a dark photon mass range between, roughly $[10^{-13},10^{-12}]$ eV could not have a kinetic mixing more than roughly $[10^{-9},10^{-7}]$.
Figure 22: Taken from Mirasola et al. (2025). For each of the targets of the search in Mirasola et al. (2025), the upper limit at 95% confidence on the kinetic mixing parameter $\epsilon$ is plotted as a function of the boson mass for which at least 10 systems would have been detected with signal amplitudes above the strain upper limits. The color corresponds to the number of events that should have been detected, assuming a “pessimistic” population model for BHs. The bounds reproduced from other experiments can be found in O’Hare (2020), specifically from the conversion of CMB photons into dark photons, which would cause distortions in the black body spectrum measured by COBE/FIRAS Mirizzi et al. (2009); Caputo et al. (2020) (dark gray), and from measurements by Planck and unWISE galaxies of cross correlations between temperature anisotropies in the CMB McCarthy et al. (2024) (light gray).
VI.3.4 Stochastic gravitational-wave background constraints
The superposition of GWs arising from annihilating scalar boson clouds could form a stochastic GW background detectable by current-generation GW interferometers. In Tsukada et al. (2019), the authors designed a model of the stochastic GW background arising from isolated BHs formed in two ways: (1) from remnants of core collapse supernovae, and (2) from remnants of mergers of BHs or neutron stars. The authors also performed search on LIGO O1 data using cross-correlation to look for evidence of a stochastic GW background Tsukada et al. (2019). No signal was found, and no range of boson masses could be robustly excluded unfortunately, except under optimistic assumptions about BH formation rates and spin distributions, which would disfavor the regime $2\times 10^{-13}\leq m_{b}\leq 3.8\times 10^{-13}$ eV at 95% confidence. The major reason for the inability to robustly exclude boson masses is that there exist degeneracies between the boson mass and BH astrophysical quantities, such as spins. The search was repeated in the second observing run as well, finding no evidence for any stochastic GW background of vector boson clouds around rotating BHs Tsukada et al. (2021), in particular disfavoring vector boson cloud masses between $[0.8,6]\times 10^{-13}$ eV at 95% confidence. Additionally, Ref. Yuan et al. (2022) has also disfavored bosons with masses between $[1.5,17]\times 10^{-13}$ eV depending upon what is assumed for the uniform spin distribution, with wider constraints present when higher-spinning BHs dominate the priors. Furthermore, the stochastic GW background from tensor bosons around rotating BHs has been computed and searched for in Guo et al. (2024a), and has disfavored boson masses between $[10^{-13.4},10{-11.7}]$ eV depending on the range of uniform spin distributions assumed Guo et al. (2024a).
VI.4 Prospects for GW probes of boson clouds
At the moment, despite the vast theoretical playground of ultralight dark matter boson clouds, only a few searches have actually been performed for GWs due to boson-boson annihilations. This is primarily due to the fact that the rich physics present in boson cloud binary black hole systems, and from transitions in isolated ones, can only be probed in space-based GW detectors, which are expected to be operational in the 2030s. Moreover, scalar boson clouds are by far easier to work with theoretically than vector or tensors, which has restricted the search parameter space primarily to what can be probed by CW searches. However, as the theoretical calculations of vector boson cloud waveforms improve Siemonsen and East (2020); Siemonsen et al. (2023b), new methods are being developed to search for annihilating vector boson clouds as well Jones et al. (2023, 2024). Additionally, there are also significant astrophysical uncertainties, namely: what are the distributions of BH spins, masses, distances and ages in the galaxy and beyond. Such known unknowns limit potential constraints to be quite model agnostic (not so astrophysically informative) or too tied to uncertain BH population models. These conundrums are not yet resolved, and probably will not be until we start detecting significantly more BHs than we currently do.
The chosen targets of GW searches should also be evaluated. While having accurate measurements of the mass, spin and age of remnants of binary black hole mergers would allow us to ignore astrophysical population models Arvanitaki et al. (2015a), these systems are currently too far to be detected if the clouds are composed of scalar bosons Isi et al. (2019), though could be seen up to Gpc now if they are composed of vector bosons Jones et al. (2023). On the flipside, GWs from vector boson clouds suffer from even more theoretical uncertainties than scalars Siemonsen and East (2020), and so the robustness of search results by targeting remnants would have to be evaluated.
All-sky searches for such boson cloud systems could also be enhanced. At the moment, especially in light of boson self-interactions that could alter the GW signal significantly. Potentially, new methods to track quickly-evolving waveforms, such as those in Jones et al. (2023, 2024), could be useful in a variety of contexts. Furthermore, the all-sky search specifically for scalar boson cloud systems did not consider any grid in spin-up, which limited it to be sensitive to boson clouds whose frequency drift did not exceed one frequency bin during the entire observation time. Such modifications to searches, coupled with additional theoretical understanding of the GW signal itself, could increase chances of detecting a signal.
It is also difficult to say with certainty that a particular boson mass is “ruled out”, since the constraints are either population agonistic or depend on some assumptions about BH formation, spin or mass distributions. Ref. Hoof et al. (2024) attempts to unify the way in which constraints are set in a Bayesian way by handling the subtleties associated with population assumptions. The authors encourage that posteriors on mass and spin be made publicly available in order to incorporate into their future analyses. Furthermore, Ref. LABEL:Guo:2023lbv shows that bosons could be transferred from a growing level to a depleting one if both BHs in a binary system have clouds around them, which may affect the GW waveform and hence would need to be considered in future GW searches.
VII GW probes of soliton DM
If DM has a mass of $\sim 10^{-24}-10^{-22}$ eV Hui et al. (2017), its Compton wavelength will be comparable to the size of the galactic halo, which actually limits the minimum mass of DM Sin (1994). In this mass regime DM could form so-called “solitons”, self-gravitating structures prevented from collapsing from degeneracy pressure, at the centers of galaxies Schive et al. (2014). To probe this kind of DM, these masses imply that the oscillation frequencies will be of $\mathcal{O}(\text{nHz})$, making pulsar timing arrays the ideal probe Aghaie et al. (2024). In essence, DM would affect the waveform arising from the superposition of inspiraling supermassive BHs, causing the BHs to experience dynamical friction, thus weaking the signal strain in the nHz regime. But this is not the only way to detect this mass regime of ultralight dark matter: as we will show, LISA and next-generation ground-based GW interferometers could also be used. To probe this kind of DM with conventional GW sources, low-frequency information needs to be contained in relatively higher-frequency carrier GW. The GW frequency becomes modified during its propagation between its source and the earth as it passes through the DM soliton, i.e. it is “heterodyned”, or gravitationally redshifted, analogously to the Sachs-Wolfe effect that induces fluctuations in the cosmic microwave background Sachs and Wolfe (1967).
VII.1 Impact on phase evolution of binary systems
Space-based GW detectors will observe the quasi-infinite inspiral of galactic white dwarf binaries and extreme-mass ratio inspirals of ordinary compact objects with much heavier BHs Bartolo et al. (2016); Babak et al. (2017). The long-durations of such signals permit the possibility of detecting a low-frequency modulation when passing through the ultralight dark matter soliton Wang and Zhong (2023); Brax et al. (2024); Blas et al. (2024). Ref Brax et al. (2024) has found that the impact of dynamical friction on the phase evolution of GW signals from white dwarf binaries exceeds that of the low-frequency DM-induced modulation for masses above $10^{-21}$ eV, essentially providing two orders of magnitude of mass that could be probed in the future. Unfortunately, even at these masses, the density of DM would have to be at least $10^{4}$ times higher than in the solar system in order to observe any changes to the phase evolution GW. This is broadly consistent with previous work by Wang and Zhong (2023), which assumed an enhancement of the DM density by eight orders of magnitude relative to that on earth, and showed that DM with a mass of $\sim 10^{-23}$ eV could be detected with LISA. Furthermore, Ref. Blas et al. (2024) assumes a representative population of white dwarf binaries, a fraction of which are detectable in LISA, and concludes that DM masses above $\sim 10^{-23}$ eV will remain unconstrained by LISA. However, the mass regime $\sim[2\times 10^{-22},3\times 10^{-21}]$ eV could be constrained if DM has stronger, quadratic couplings to ordinary matter.
VII.2 Imprint of DM on CW signal from neutron star
In addition to modulating the millihertz frequency of GWs from binaries in space-based GW detectors, soliton ultralight dark matter could induce changes in the CWs arising from deformed neutron stars, thus allowing the possibility of detecting this effect in future ground-based detectors Blas et al. (2024). These GWs are sourced from the rotational power of the neutron stars, thus spinning down the neutron stars as GWs are emitted Riles (2017). In ET and Cosmic Explorer, we expect to detect $\mathcal{O}(100)$ or more neutron stars spinning down due to the emission of CWs Pagliaro et al. (2023), all of which could be modulated by ultralight dark matter solitons.
Using CWs from deformed neutron stars, the soliton mass regime $\sim[2\times 10^{-22},3\times 10^{-21}]$ eV becomes accessible in ET for both linear and quadratic couplings, and improves with the number of neutron stars that are detected. The constraints will even outperform existing pulsar timing array ones, since the signal-to-noise ratio scales linearly with the carrier GW frequency. This is shown in LABEL:fig:lambda12res for LISA, ET and Cosmic Explorer, and demonstrates the powerful probe that CWs from neutron stars will be of this particular DM model.
VII.3 Pulsar-timing constraints on soliton ultralight dark matter
The evidence for a GW background from NANOGrav results can be used to put constraints on soliton DM Aghaie et al. (2024). Essentially, ultralight dark matter would induce dynamical friction in the inspiral of supermassive BHs, speeding up the orbit, which suppresses the GW strain at nanohertz frequencies. The supermassive BHs are too massive to be prevented from merging, as would be the case for lighter BHs Hui et al. (2017); Bar-Or et al. (2019). Using the most accurate limit at 3.92 nHz on the GW strain from supermassive BHs in the right-hand panel of Fig. 1 of Afzal et al. (2023), Ref. Aghaie et al. (2024) derives constraints on the soltion DM mass-to-BH mass ratio parameter space, as shown in Fig. 24. These bounds are derived by using Eq. 17 of Aghaie et al. (2024), which depends on how much energy is lost via GW emission normally and in the presence of the ultralight dark matter soliton, and on the differential density populations of supermassive BHs. The latter can be written analytically in terms of the the galaxy-stellar mass functions, galaxy merger rates and galaxy pair fractions Chen et al. (2019). It is then a matter of integrating Eq. 17 in Aghaie et al. (2024) over these different distributions (which may depend on decently uncertain astrophysics) for different choices of ultralight dark matter mass to derive the constraint on mass ratio of soliton DM to supermassive BH mass, and determine if the integrated strain exceeds the error bar on the GW strain obtained by NANOGrav.
Figure 24: Taken from Aghaie et al. (2024). Upper limits from NANOGrav data on the DM mass and solition-BH mass ratio parameter space, obtained by noting that inspiraling supermassive BH background for which NANOGrav found evidence would be modulated by the dynamical friction induced by soliton DM. Magenta lines show the scaling relations derived through DM-only numerical simulations (solid), and ones in which supermassive BHs dominate scaling relations suggested by DM-only numerical simulations and SMBH dominance (dashed).
VIII GW probes of WIMP dark matter
VIII.1 Transmuted black holes in binaries
Non-annihilating WIMPs that interact with nucleons could lose energy through single or multiple scatterings off of celestial objects Bramante et al. (2017); Bell et al. (2021), eventually becoming gravitationally bound to them over time. After repeated interactions between the WIMPs and ordinary nucleons, the WIMPs become thermalized, forming a compact core with a higher density than that of the celestial object. If the density of the particles is sufficient, gravitational collapse (in the case of bosons) or Chandrasekhar collapse (in the case of fermions) of the WIMP core will occur, thus resulting in a newborn BH. After formation, the celestial object could accrete surrounding DM on a given timescale, and if enough matter becomes bound to the object quickly enough compared to the Hawking evaporation timescale, the entire object could collapse into what is known as a transmuted black hole (TBH) Dasgupta et al. (2021); Ray (2023); Bhattacharya et al. (2023, 2024); Bhattacharya (2024). If the sum of the time to accrete new particles and the time to collapse is smaller than the age of the universe, TBHs should form.
TBHs forming from inspiraling neutron stars were first considered in Dasgupta et al. (2021). Since DM would induce neutron stars to transmute into BHs, from the observational perspective, there is no difference between electromagnetically silent neutron stars and low-mass BHs. Therefore, the non-detection of low-mass binary black holes by LIGO, Virgo and KAGRA can be used to constrain weak interaction cross-sections, since the progenitor neutron stars provide an extremely strong gravitational field to accrete and retain DM particles Bhattacharya et al. (2023); Bhattacharya (2024). In LABEL:fig:wimp-cbc-cons, we show the excluded bosonic DM mass/cross-section parameter space based on null search results from the LIGO, Virgo and KAGRA collaborations for sub-solar mass compact objects Abbott et al. (2023). Additionally, a projection for this constraint is shown for 50x more exposure time, which may be possible for the future GW network [^6] https://observing.docs.ligo.org/plan/.While analyses of LIGO, Virgo and KAGRA data cannot yet constrain new parameter space for this model of DM, future searches should be able to surpass the pulsar constraint, which comes from the fact that the nearby Gyr-old pulsar PSR 0437--4715 has not actually collapsed into a BH in its lifetime.
In contrast, stars, being less compact than neutron stars, could sustain DM cores with stronger interaction cross-sections, because of the weaker gravitational field around the star Bhattacharya et al. (2024). In LABEL:fig:lisa-wimp-cons, we show the potential excluded parameter space for null detection of TBHs arising from slowly inspiraling non-compact binaries that could be visible in space-based GW detectors. These binaries would be inspiraling for durations of order of the age of the universe, and thus exhibit CW signals in the millihertz regime. The results are parameterized in terms of $\alpha$, which denotes the fraction of stellar-mass binaries that form close enough to exist within the millihertz band. Essentially, $\alpha$ parametrizes our ignorance about how closely the population of sun-like stars form to each other.
Comparing LABEL:fig:wimp-cbc-cons and LABEL:fig:lisa-wimp-cons, we note the vastly different interaction cross-section ranges from these two complementary GW probes of WIMP DM. For both low-mass BHs and sun-like stars, fermionic DM has also been constrained in similar ways as bosonic DM has, see Bhattacharya et al. (2023, 2024) for more details.
VIII.2 GWs to probe the GeV excess
Annihilating WIMP DM and a collection of unresolved millisecond pulsars are two leading explanations for the unexpected excess of GeV gamma rays observed by the Fermi-LAT experiment coming from the Galactic Center Goodenough and Hooper (2009); Hooper and Goodenough (2011). At first, the GeV excess appeared spherically symmetric and well-fit by annihilating DM models Hooper and Linden (2011); Gordon and Macias (2013); Daylan et al. (2016); Calore et al. (2015a); Abazajian et al. (2014). But still, an astrophysical explanation of millisecond pulsars cannot yet be ruled out Abazajian (2011); Calore et al. (2015b); Yuan and Zhang (2014); Petrović et al. (2015); Ye and Fragione (2022).
GWs can be used to weigh in on this debate. If the millisecond pulsars are deformed, they will emit CWs as they rotate, and this radiation could be detectable by the LIGO, Virgo and KAGRA if the degree of deformation, called the ellipticity (defined in terms of the principle moments of inertia as $\epsilon=\frac{I_{\rm xx}-I_{\rm yy}}{I_{\rm zz}}$), is large enough Haskell et al. (2015).
First, a stochastic GW background of GWs from millisecond pulsars in the Galactic Center was considered, and projected constraints on the average ellipticity of this population were obtained as a function of the total number of millisecond pulsars in the Galactic Center Calore et al. (2019). The authors showed that $\mathcal{O}(10^{4})$ millisecond pulsars with ellipticities of $\sim 10^{-7}$ could be detectable by current detectors and be consistent with the diffuse electromagnetic radiation coming from the Galactic Center, as shown in LABEL:fig:Ntoteps. However, they did not go so far as to link GW observations with Fermi-LAT ones.
On the other hand, Ref. Miller and Zhao (2023) considered different models for how the deformations on the millisecond pulsars were sustained, either through the internal magnetic field Lander et al. (2012); Mastrano and Melatos (2012); Lander (2014) or simply as a fraction of the total electromagnetic radiation from the star. It is then assumed that the unresolved population of millisecond pulsars follows the same spin and ellipticity distribution of the known pulsars Manchester et al. (2005), and, from that, the probability of observing a CW in the search be computed, given upper limits on ellipticities from searches for isolated neutron stars or those in binary systems Abbott et al. (2022b); Miller and Zhao (2023). Once this quantity is computed, we can determine the number of millisecond pulsars in the Galactic Center by fitting parameters of luminosity functions to match the observed Galactic Center GeV excess Hooper and Linden (2016); Ploeg et al. (2020); Dinsmore and Slatyer (2022). By multiplying the predicted number of millisecond pulsars with the probability of detecting them, we can rule out luminosity function parameters that would have led to at least one millisecond pulsar being detected via GWs.
LABEL:fig:gev-cons shows the parameter space of a simplistic luminosity function, and the constraints that can be obtained through null results from CW searches, assuming particular ellipticity and frequency distributions for the millisecond pulsars. We can see that GW searches probe a complementary portion of the parameter space to what Fermi-LAT does: if there are more millisecond pulsars that are dimmer electromagnetically, this means more rotational power may go into GWs, and more of them are necessary to explain the GeV excess. Thus, GW searches would have more millisecond pulsars to find, and not seeing at least one of them rules out particular luminosity function parameters. Likewise, Fermi performs better when less millisecond pulsars are more luminous.
Most recently, Ref. Bartel and Profumo (2024) simulated a population of $\sim 40000$ millisecond pulsars in the Galactic Center that could explain the Galactic Center GeV excess, and found that the GW signal could lie between [200,1400] Hz, and would have strain amplitudes of $\mathcal{O}(10^{-30}-10^{-28})$ based on the population parameters chosen (e.g. moment of inertia around $\sim 10^{38}$ $\cdot$ m2, ellipticities mostly smaller than $10^{-9}$, and Boxy vs. spherical spatial distributions of the millisecond pulsars). It is argued that such signals are too weak to be detected by current and future GW observatories; however, this conclusion depends heavily on the assumed ellipticity distribution and the number of millisecond pulsars chosen to be in the population. If millisecond pulsars appear in larger numbers than assumed here, they would each be dimmer, since the luminosity of the GeV excess is fixed. In fact, the constraints from Miller and Zhao (2023) are strongest for larger populations than assumed in Bartel and Profumo (2024). Additionally, the ellipticity distribution employed in Bartel and Profumo (2024) only reaches a maximum of $\sim 2\times 10^{-9}$, which is consistent with the minimum ellipticity of neutron stars as predicted by Woan et al. (2018), but may be too conservative. Upper limits on ellipticities at the Galactic Center from CW searches have not yet constrained ellipticities to be less than $\sim 10^{-6}$ Abbott et al. (2022e, b); Steltner et al. (2023), which is consistent with the maximum ellipticity employed in Miller and Zhao (2023) up to $10^{-6}$.
If GWs can rule out more and more of the luminosity function parameter space to explain the GeV excess, then this will give significant weight to the annihilating WIMP DM hypothesis.
IX GW probes of DM spikes
IX.1 Background
It was first recognized in Gondolo and Silk (1999) that BHs in the DM halo accreting DM adiabatically (slowly) could allow a high-density region of DM to form, called a “spike”, which could enhance the DM annihilation rate inside the spike. Subsequent studies showed that if such spikes formed around supermassive BHs at the center of galaxies, they would would likely have been disrupted during host galaxy mergers Merritt et al. (2002), if the BH formed off-center Ullio et al. (2001), or if DM particles scattered off of nearby stars Merritt (2004); Bertone and Merritt (2005). Though, recently, upper bounds on the DM densities for supermassive BHs have been forecasted with LISA Daniel et al. (2025). However, intermediate-mass BHs (IMBHs) could sustain such spikes Zhao and Silk (2005), since their spikes may not have been torn apart through mergers over cosmic time. Furthermore, recently it has been shown that astrophysical BHs and PBHs could also host DM spikes Kohri et al. (2014); Eroshenko (2016); Boucenna et al. (2018); Hertzberg et al. (2020); Bertone et al. (2024); Dosopoulou and Silk (2025).
The density of the DM spike is characterized by the following equation Eda et al. (2013):
$\rho(r)$ $=\rho_{\rm sp}\left(\frac{r_{\rm sp}}{r}\right)^{\alpha}\,\,\,(r_{\rm min}\leq r\leq r_{\rm sp}),$ (108)
where $\rho_{\rm sp}$ is a normalization constant for the DM density of the spike, $r_{\rm min}$ is the minimum separation between the two objects, often taken to the radius at the innermost stable circular orbit, and $r_{\rm sp}$ is the spike radius. Here, $\rho_{\rm sp}=379\ M_{\odot}/\text{pc}^{3}$ and $r_{\rm sp}=0.33\text{pc}$ if the mass of the IMBH is $M_{\rm BH}=10^{3}M_{\odot}$ and the DM spike has a mass of $M_{\rm halo}=10^{6}M_{\odot}$. The power-law index $\alpha$ denotes how the concentration of DM particles falls off as one moves away from the host BH. Astrophysically speaking, measurements of $\alpha$ would allow us to glean important information regarding whether the host BH evolved from the DM halo and/or experienced mergers in the past Eda et al. (2013).
IX.2 Dephasing due to dynamical friction of IMRI systems
Ref. Eda et al. (2013) first showed that the phase evolution of an intermediate-mass ratio inspiral (a binary composed to a $\mathcal{O}(1000)M_{\odot}$ and an $\mathcal{O}(1)M_{\odot}$ object) would be significantly modified at milli-hertz frequencies due to the presence of a DM spike around the intermediate-mass BH. This occurs because the less massive object sweeps through the DM spike as it orbits around the more massive one, losing energy via dynamical friction (in addition to via GWs) and speeding up the orbit. Such effects will primarily be relevant in space-based GW detectors Duque et al. (2024); Ghodla (2024), since IMBHs inspiral and merge at much lower frequencies than stellar-mass BHs Miller (2009).
In the future, if matched filtering analyses are performed that do not account for the DM spike, dephasing with respect to the vacuum templates may occur. In Fig. 26, we show how the signal-to-noise ratio degrades as a function of the observation time, the initial frequency $f_{\rm ini}$, and the power-law index $\alpha$ of the DM spike). As described in Eda et al. (2013), dynamical friction affects the early stages of the inspiral, and as the system progresses toward, the DM spike becomes disrupted, reducing its impact on the GW signal. That is why in Fig. 26, lower initial frequencies start to lose signal-to-noise ratio at smaller values of $\alpha$. Furthermore, it is clear that above $\alpha\simeq 2$, templates that include dynamical friction will be required to successfully recover IMRI signals in space-based GW interferometers. This conclusion was significantly expanded in Eda et al. (2015), in which the authors demonstrated how accurately the power-law index $\alpha$ could be measured from a detection of an IMRI signal in LISA. This accuracy is quantified in terms of the power-law index itself and the masses of the two objects that are inspiraling. If the IMRI system consists of smaller primary and secondary masses, and a larger power-law index, the power-law index can be measured more accurately. This is because higher values of $\alpha$ induce a larger phase difference, and smaller masses in the IMBH system ensures that it spends more cycles in the LISA band before merging.
Figure 26: Taken from Eda et al. (2015). The number of cycles $N_{\text{cycle}}$ spent in the GW band $f\in[f_{\text{min}},f_{\text{max}}]$ for a primary BH with mass $M_{\rm BH}=1000M_{\odot}$ and a companion of $1M_{\odot}$. At large $\alpha,$ the DM solition causes large dephasing between the injected waveform and the one that would be used to recover the signal. Thus, the number of cycles that are captured in-band drops significantly.
PBHs could also host DM spikes, and must do so, since, at most masses, the fraction of DM that they could compose is less than one Green and Kavanagh (2021); Carr et al. (2024). Thus, Ref. Cole et al. (2023a) showed that ET and Cosmic Explorer could probe the existence DM spikes around binary PBH systems, where the primary has a mass of $[1,10]M_{\odot}$ and the system has a mass ratio of $10^{-3}$. However, just as in astrophysical binary black hole systems, these exotic mergers would be missed if the DM spike is not taken into account in matched-filtering searches in ET and Cosmic Explorer. Multi-band observation of inspirals/mergers constraining DM spikes are also possible with future deci-hertz and ground-based GW interferometers, which would enhance the constraints on the power-law index and spike-density factors by orders of magnitude relative to individual constraints from each of these detectors Tahelyani et al. (2024); Chen et al. (2024).
Recently, it has also been shown that these dephasing effects are unique enough to allow distinguishing between three BH environments: accretion disks, DM spikes, and ultralight boson clouds Cole et al. (2023b), as well as distinguishing between environmental effects and modified theories of gravity Yuan et al. (2024). However, Ref. Zwick et al. (2024) argues that practically speaking, dephasing could result from a variety of (also unmodelled) sources, and would be difficult (computationally and algorithmically speaking) to track with long-duration templates in matched filtering. Though, Ref. Wilcox et al. (2024) has indicated this could be possible within a parameterized post-Einsteinian framework. They thus propose more model-agnostic ways (i.e. periodic forces) that may influence the evolution of the binary and leave characteristic signatures on the GW signal. Such problems with matched filter, could, however, be handled by the use of alternative, non-matched filtering CW methods meant to handle intermediate-duration signals, e.g. Isi et al. (2019); Sun and Melatos (2019); Miller et al. (2018); Oliver et al. (2019); Banagiri et al. (2019); Miller et al. (2019); Guo and Miller (2022); Andrés-Carcasona et al. (2023); Alestas et al. (2024); Miller et al. (2024a); Andrés-Carcasona et al. (2024), and those discussed already in this review.
Perhaps more surprising, it has been proposed that equal-mass systems could sustain DM spikes even in the last ten orbits of binary black hole merger Aurrekoetxea et al. (2024b). Initially, it had been shown through N-body simulations that heavy, particle-like DM environments would be disrupted well before equal-mass binary black hole systems merge Bertone and Merritt (2005); Kavanagh et al. (2018). But, for wave-like, ultralight dark matter, though there have been many works that have modeled the motion of binary black holes through DM environments,Koo et al. (2024); Boudon et al. (2023); Zhong et al. (2023); Cardoso et al. (2022a, b); Vicente and Cardoso (2022); Annulli et al. (2020a, b); Brax et al. (2020) or with boson clouds Cao and Tang (2023); Traykova et al. (2023); Chia et al. (2023); Baumann et al. (2019a, 2020); Ikeda et al. (2021); Cardoso et al. (2020); Hannuksela et al. (2019); Baumann et al. (2022b); Zhang and Yang (2020); Kumar Poddar et al. (2020); Wong (2019, 2020); Kavic et al. (2020); East (2018); Siemonsen and East (2020); Baumann et al. (2022a); Siemonsen et al. (2023b); Tomaselli et al. (2023); Poddar et al. (2024); Cao et al. (2024); Guo et al. (2024b), the point during the inspiral at which the environment becomes disrupted is not clear, motivating further studies into this Bamber et al. (2023). Ref. Aurrekoetxea et al. (2024b) found that amount of dephasing that exists within the last ten orbits of a binary black hole merger depends strongly the scalar DM mass and the masses of the BHs:
$\mu\approx 5\times 10^{-17}\,\left(M/10^{6}M_{\odot}\right)^{-1}\,\mathrm{eV},$ (109)
which gives the optimal DM mass to cause the most dephasing during the last ten orbits of the binary, and determines the DM density as
$\rho_{0}\approx 10^{20}\,\left(\frac{M}{10^{6}M_{\odot}}\right)^{-2}\,\frac{M_{\odot}}{\mathrm{pc}^{3}}\,.$ (110)
which gives the DM density necessary to induce dephasing. Though the DM density presented here is larger than the average, it results from a measured dephasing over a short period of time. If the effect of the DM spike on the binary black hole system turns on at larger separations, as may be expected from previous works Cole et al. (2023a, b), then smaller densities could cause significant dephasing over the lifetime of the binary.
IX.3 Upper limits on BH environments
Because vacuum phase-evolution waveforms confirmed the detection of many binary black hole mergers by LIGO, Virgo and KAGRA, any deviations due to environmental effects, such as a DM spike, must be small enough to be inconsequential to the detected GW signals. Knowing this, Ref. Caneva Santoro et al. (2024) modeled dynamical friction and other possible environmental effects as additional (negative) PN corrections to the waveform, and placed upper limits on the density of environment around BHs by requiring the induced dephasing to be small with respect to the phase evolution of the binary black hole systems in vacuum. While the results were used to eliminate dynamical fragmentation Loeb (2016) as a possible binary formation channel, they are generic enough to apply to any BH environment. Unfortunately, the upper limits on environments affected by dynamical friction, which would be relevant for the DM spike scenario, are, at best, about seven orders of magnitude larger than the maximum density of the spike ($10^{-6}$ g/cm3 Gondolo and Silk (1999); Sadeghian et al. (2013). The upper limits are shown to drastically improve in DECIGO because the dynamical friction (and other environments) affect the early inspiral of the binary black hole system much more than at the time of merger. These findings are consistent with those found in N-Body simulations about how the dynamics of the DM spike affect the inspiral, and the detectability of these effects by LISA Kavanagh et al. (2020).
X GW probes of atomic DM
X.1 Background
It is possible that the dark sector is actually much more complicated than a single new particle with a single coupling to gravity or the standard model. In fact, a atomic DM models exist Feng et al. (2009); Kaplan et al. (2010) composed of dark protons and dark electrons that interact via a massless dark photon with a particular coupling. With such a model, DM can form bounds states analogous to the Hydrogen atom, and dissipate energy via Bremsstrahlung, recombination, and collisional excitation processes Buckley and DiFranzo (2018). If DM collects in small-enough regions of space, it will collapse and form “dark BHs Shandera et al. (2018).
Such dark BHs could form in binary systems, as astrophysical BHs do; thus, LIGO, Virgo and KAGRA observations of binary black hole mergers can constrain the existence of these objects, just as they do for PBHs Singh et al. (2021); Abbott et al. (2022f, 2023).
X.2 Constraints from observations of binary black holes
Ref. Shandera et al. (2018) first considered the prospects of detecting dark BH mergers with advanced LIGO and ET, finding that these interferometers could set stringent constraints on the fraction of DM that dark BHs could compose. If the BH merger of GW190425 Abbott et al. (2020a) is interpreted as the merger of two dark BHs, the Chandrasekhar mass Chandrasekhar (1931) can be constrained to be below $1.4M_{\odot}$ at $>99.9\%$, bounding the lower-most value mass of the dark proton as 0.95 GeV and the spacing of energy levels in a dark atom to be near $10^{-3}$ eV Singh et al. (2021).
These constraints on the mass of the dark proton and the Chandrasekhar limit of the mass of dark BHs are derived using Chandrasekhar (1931); Shandera et al. (2018); Singh et al. (2021):
$M_{\rm DC}\simeq 1.4M_{\odot}\left(\frac{m_{p}}{m_{\chi}}\right)^{2}.$ (111)
The range of the dark proton mass $m_{\chi}$ can be obtained by setting $M_{\rm DC}=M_{\rm min}$ for the values of $M_{\rm min}$ for which the fraction of DM that dark BHs could compose is less than 1:
$\sqrt{\frac{1.4M_{\odot}}{M_{\rm min}^{\rm max}}}m_{p}<m_{\chi}<\sqrt{\frac{1.4M_{\odot}}{M_{\rm max}^{\rm max}}}m_{p}$ (112)
$M_{\rm min}$ is the minimum mass in a distribution of dark BHs masses that is assumed, as described below.
Further constraints on the fraction of DM composed of dark BHs are presented in three cases, assuming that (1) none of the detected BHs by LIGO, Virgo and KAGRA are dark, (2) half are dark, and (3) GW190425 and/or GW190814 Abbott et al. (2020b) are dark. Here, we limit ourselves to case (1) and case (3). These constraints rely on an assumed mass distribution of dark BHs based on Population III star formation works Bromm and Larson (2004), which is chosen to be $P(m)\propto m^{-b}$ along with a mass range of $[M_{\rm min},rM_{\rm min}]$. The parameters $b=[-1,2],r=[2,100]$ determine the mass distribution of dark BHs. With these assumptions, we show in Fig. 27 the constraint on the fraction of DM that dark BHs could compose as a function of the average mass of the BHs considered, in relation to other limits, in case (1) and in case (3). Assuming all BHs are not dark, the constraint beats existing ones by several orders of magnitude; assuming that GW190425 and GW190814 are dark BHs, a region in this parameter space is carved out of possible average mass/ DM fractions. More observations of dark BHs would be needed to further narrow this region. However, it must be emphasized that while GW190425 and GW190814 being dark BHs are not yet excluded by other experiments, this work does not provide a way of distinguishing astrophysical BHs from dark BHs.
Figure 27: Taken from Singh et al. (2021). Constraints on the fraction of DM that dark BHs could compose relative to other experimental upper limits using all binary black holes detected in the first three observing runs of LIGO, Virgo and KAGRA. These plots assume an atomic model of DM that forms dark BHs. Even in the absence of any dark BHs, the constraints obtained in Singh et al. (2021) significantly outperform existing microlensing experiments Allsman et al. (2001); Tisserand et al. (2007); Niikura et al. (2019) and those from dynamics of dwarf galaxies Brandt (2016); Koushiappas and Loeb (2017). Additional constraints arise from PBH searches Abbott et al. (2019); Gaggero et al. (2017); Raidal et al. (2017, 2019); Ali-Haimoud et al. (2017). SB stands for stochastic background and refers to analyses of the first and second observing runs of LIGO, Virgo and KAGRA data for stochastic GW backgrounds from generic systems.
X.3 Constraints from sub-solar mass binary black hole searches
The lack of smoking-gun evidence for dark BHs in current LIGO, Virgo and KAGRA observations motivates the need to consider sub-solar mass BHs, whose origins, if detected, cannot be explained by astrophysical means. Searches for GWs from sub-solar mass compact objects in binary systems have grown in popularity in recent years, covering mass ranges of roughly $[10^{-11},1]M_{\odot}$ Nitz and Wang (2021); Phukon et al. (2021); Abbott et al. (2022f); Miller et al. (2022b); Nitz and Wang (2022); Abbott et al. (2023, 2022b); Miller et al. (2024b), and, traditionally, have put constraints on the fraction of DM that PBHs could compose. However, as discussed in Singh et al. (2021), these searches do not assume a particular formation mechanism for these compact objects, and the binaries can thus be interpreted as arising from not only PBHs, but also dark BHs.
In the most recent analysis of LIGO, Virgo and KAGRA data from the second half of the third observing run, constraints on dark BHs with masses below $1M_{\odot}$ were placed, as shown in Fig. 28, using different matched filtering methods, again assuming a power-law dependence for the mass distribution of dark BHs and as a function of the minimum mass of this distribution. Nuisance parameters in the mass distribution are marginalized over. These limits do not go below $\sim 10^{-2}M_{\odot}$ because the maximum mass probed is between $[2,1000]M_{\odot}$, and so the distribution tends to disfavor smaller masses. Moreover, the limits do not go above a solar mass because the search is only for sub-solar mass objects.
From this work Abbott et al. (2023), the mass of the dark proton has been constrained to be [0.66,8.8] GeV, which is actually broader than the constraint in Singh et al. (2021) of [0.91,4.76] GeV since there are no assumed detections of dark BHs. Additionally, the energy spacing between dark atomic energy levels can also be constrained, assuming values for the mass of the dark fermions, the dark fine structure constant, and the Jeans mass, and using Eq. 1 of Singh et al. (2021) and Eq. 4 of Shandera et al. (2018).
Figure 28: Taken from Abbott et al. (2023). Constraints on the fraction of DM that dark BHs could compose in the atomic DM scenario. A mass distribution for dark BHs is assumed with the shown minimum mass on the $x$-axis, and the constraints are shown for three different matched filtering algorithms: GstLAL (dotted), MBTA (dashed) and PyCBC (solid). Constraints from a previous search for sub-solar mass compact objects in the first half of the third observing run of LIGO, Virgo and KAGRA are also shown  Abbott et al. (2022f).
XI Conclusions
There is no shortage of observable signatures that DM could leave on GW sources when the particle, mass and couplings of DM the standard model are varied. We do not assume to have covered every single DM model in this work, but we hope to have provided an overview for the interested reader that wishes to pursue more details about different ways that GWs can be used to probe the existence of particle DM.
In this work, we have explored (1) how ultralight dark matter could couple to GW interferometers, (2) how macroscopic DM could transit through the interferometers, (3) how boson cloud systems formed through a superradiance process around rotating BHs could emit GWs by annihilation, (4) how soliton DM could induce changes in GWs arising from binary systems and isolated neutron stars, (5) how WIMP DM could collect around celestial objects and induce them to collapse into BHs detectable via GW emission, (6) how environmental effects of DM spikes around BHs could affect GW signals in future detectors, and (7) how GW observations and non-observations can be used to probe atomic models of DM. Some of these types for DM have been extensively constrained through analysis of terrestrial GW detector data; others will require the advent of space-based and next-generation ground-based GW interferometers to obtain insight into these types of DM.
Despite the “WIMP miracle” announced decades ago, WIMPs have not yet been detected, and are only one of numerous possible ways that DM could exist in the Universe. The lack of detection of WIMPs over the years has opened up both the experimental and theoretical playgrounds for new probes of DM, one of which is through its impact on GW signals, or as a signal in interferometer data itself. Moreover, it has been assumed in this work, and almost all those cited, that DM is the way to solve the puzzling cosmological observations on a variety of scales. However, modified theories of gravity are real contenders to solve this problem as well, though their community is a lot smaller and not quite as popular as the DM one. These two communities – DM and modified gravity – act entirely independently, choosing to break, respectively, one of the two assumptions of “classical” physics: (1) the matter in the universe is mostly baryonic and luminous, and (2) general relativity and, in the limit of slow-moving objects, Newtonian theory, are correct Martens and Lehmkuhl (2020). It may be that these assumptions need to be broken simultaneously, or that these two distinct fields have more in common than it appears, as argued in Martens and Lehmkuhl (2020).
We only briefly mentioned PBHs in this review, which have been reviewed extensively elsewhere Green and Kavanagh (2021); Miller (2024). PBHs could also comprise a fraction of or the totality of DM, and recent observations of binary black holes by the LIGO, Virgo and KAGRA with low spins whose merging rates are consistent with PBHs have reignited interested in this hypothesis Bird et al. (2016); Clesse and García-Bellido (2017); Sasaki et al. (2016). While analyses of PBHs typically receive criticism for using monochromatic mass functions to place constraints, all of the results presented here implicitly assume that DM is only composed of particles that are being constrained [^7] The exception to this statement is superradiance probes of BHs– these bosons need not be DM.. It could be that multiple forms of DM co-exist, e.g. in the atomic DM models, or that there is a whole dark sector.
Nevertheless, GWs provide a new window to probe the existence of DM in a variety of ways. Most of the actual analyses performed on GW data do not assume particular models for DM, and only constrain them later when one interprets null (or actual) results in terms of a DM model. GW searches are therefore powerful, relatively model-agnostic probes of DM that permit the detection of a variety of DM signatures.
Acknowledgments
We thank William East, Alexandre Gottel, Dana Jones, Jun’ya Kume, Sachiko Kuroyanagi, Ling Sun, and Yue Zhao for reviewing this manuscript.
This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation
This research has made use of data, software and/or web tools obtained from the Gravitational Wave Open Science Center (https://www.gw-openscience.org/ ), a service of LIGO Laboratory, the LIGO Scientific Collaboration and the Virgo Collaboration. LIGO Laboratory and Advanced LIGO are funded by the United States National Science Foundation (NSF) as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale della Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain.
We would like to thank all of the essential workers who put their health at risk during the COVID-19 pandemic, without whom we would not have been able to complete this work.

### v3

@@
 Or, particle DM may (or may not) exist alongside PBHs, which may comprise a portion or the totality of DM Hawking (1971); Carr et al.
+(2010); Dai et al.
 (2009); Clesse and García-Bellido (2015); Carr et al.
@@
 Despite these efforts, however, WIMPS remain undetected [^1] DAMA claims to see a periodic oscillation of a DM signal, though this result has not yet been reproduced Petriello and Zurek (2008), and so it is worthwhile to ask whether other models could better explain DM.
-Thus, the experimental development from the particle physics community has widened to include probing ultralight DM through torsion balance experiments, such as Eöt-Wash Su et al.
+Thus, the experimental development from the particle physics community has widened to include probing ultralight DM through (1) torsion balance experiments, such as Eöt-Wash Su et al.
 (1994); Schlamminger et al.
@@
 (2012); Bergé et al.
-(2018), whose results can be interpreted as constraints on DM, and resonant cavity experiments, such as ADMX Du et al.
-(2018), among others.
+(2018), whose results can be interpreted as constraints on DM, (2) resonant cavity experiments, such as ADMX Du et al.
+(2018), (3) the Event Horizon telescope Doeleman et al.
+(2009); Akiyama et al.
+(2019), which could measure photons lensed by boson clouds and thus affect BH images Chen et al.
+(2023), and which could measure changes in the polarization of the electric field when the light passes through the axion background Chen et al.
+(2022, 2022), and (4) astrometry experiments, such as Gaia Prusti et al.
+(2016) and the Roman Space Telescope Akeson et al.
+(2019), that can meausre stochastic metric perturbations sourced by ultralight dark matter Kim (2024).
 It is therefore also worth asking whether high-precision GW interferometers, such as LIGO, Virgo and KAGRA Aasi et al.
@@
 (2015); Aso et al.
-(2013), can be used to search for DM, and in particular, whether continuous gravitational-wave (CWs) can be used as the probe.
-While not yet as extensive as particle physics experiments, GW probes of DM have already, in many cases, surpassed limits from existing experiments, or will beat them as the detectors, both in ground and in space, develop.
+(2013), can be used to search for DM, which has also been briefly reviewed in Bertone et al.
+(2020).
+Axions and other forms of DM could be probed via their interaction with binary neutron star inspirals that could be detected in the future with LIGO, Virgo and KAGRA Hook and Huang (2018); Croon et al.
+(2018).
+The observational signature of axions on binary neutron star inspirals would be a dephasing effect induced by each object having a scalar charge, and the mass of the scalar field Hook and Huang (2018); Zhang et al.
+(2021).
+Furthermore, a variety of follow-up studies Sagunski et al.
+(2018); Huang et al.
+(2019) have been performed regarding the detectability of the effects of axions on binary neutron star inspirals, with Ref.
+Zhang et al.
+(2021) placing the first-ever constraints on the nuclear coupling of axion-like particles with masses below $10^{-11}$ eV using the detection of GWs from the binary neutron star GW170817, effectively ruling out axion decay constants of [$1.6\times 10^{16}$,$10^{18}$] GeV.
+Furthermore, even the very slow orbit of the Hulse-Taylor binary and other known binaries constrains the axion decay constant to be below certain values depending on the system Kumar Poddar et al.
+(2020) to be $f_{a}<\mathcal{O}(10^{11})$ GeV, and resonances could appear during the orbit when the orbital frequency of the binary matches that of the axion Kadota et al.
+(2024); Kim and Yang (2024).
+Finally, axions and the DM energy density can also be constrained via the evidence for stochastic GW background from pulsar timing arrays Guo et al.
+(2024a); Chen et al.
+(2024a) and also explain the mass gap in observed BHs Croon and Sakstein (2023).
+In particular, in this review, we would like to explore how continuous gravitational-wave (CWs) can be used as a probe of DM.
 CWs are quasi-monochromatic, quasi-infinite-duration signals that canonically arise from asymmetrically rotating, lumpy neutron stars.
@@
 (2019); Miller and Zhao (2023); Bartel and Profumo (2024).
-Finally, axions can be constrained using the very slow orbit of the Hulse-Taylor binary and other known binaries to be below to be $f_{a}<\mathcal{O}(10^{11})$ GeV, depending on the system Kim and Yang (2024).
+In LABEL:fig:summary, we show a landscape of the kinds of particle DM that could be probed with current and future interferometers, both by using GWs and and through direct detection of particles interacting with the instruments.
+There are thus so many different kinds of DM that can be searched for by using instruments that were not strictly designed to probe the existence of DM.
 This review article is broken into different parts depending both on the type of DM considered and constraints on DM interactions with gravity or standard model particles.
@@
 Ultralight scalar DM will also cause deviations in the times of arrivals of pulses as given in Khmelnitsky and Rubakov (2014), though the mass of such particles are 10 orders of magnitude lower than those probed with ground-based interferometers.
+It can also induce stochastic perturbations in the metric, which would also induce time-varying changes in the times of arrivals of pulses Kim and Mitridate (2024), but generally could be seen in interferometers with long baselines Kim (2023).
 In Section II, we discussed various models of scalar DM that can be constrained with pulsar timing arrays.
@@
 III.4.1 Gravitationally-interacting ultralight dark matter
-As discussed before, DM could interact only via gravity, and alter the distances between the pulsars and earth, analgous to GWs.
+As discussed before, DM could interact only via gravity, and alter the distances between the pulsars and earth, analogously to GWs.
 Using European pulsar timing array data, a search was performed in the residuals for this kind of DM interaction.
@@
 If they only comprise a fraction $f$ of DM, then the constraints will be weakened by $\sqrt{f}$, because of the $\sqrt{\rho_{\text{DM}}}$ term in the amplitude of these interaction signals.
-Furthermore, the stochastic GW background w
 Figure 11: Taken from Afzal et al.
@@
 The constraints from tests of the equivalence principle (“MICROSCOPE”) are shown in teal (Bergé et al., 2018), and the constraints previously set by the PPTA Collaboration are reported in yellow (Xue et al., 2022).
+The recent NANOGRAV evidence for a stochastic GW background is also consistent with a vector ultralight dark matter hypothesis.
+Ref.
+Chowdhury et al.
+(2024) has found that ultralight dark matter that couples to muons ($L_{\mu}-L_{\tau})$ is actually more consistent with NANOGRAV’s observations than a stochastic GW background.
+This type of ultralight dark matter would cause the pulsars to oscillate (because they are composed of muons) and induce a spin-dependent frequency shift in the pulses that would change the angular correlations with respect to the stochastic GW background case (because of the gravitational potential of the ultralight dark matter).
+The mass and coupling constants of this ultralight dark matter are not ruled out by any other experiment, and lie in the region
 III.4.4 Constraints on conformal DM
@@
 (2023b); Jones et al.
-(2023, 2024), and potentially may contain electromagnetic counterparts if the vector boson kinetically mixes with the ordinary photon Siemonsen et al.
+(2023, 2025), and potentially may contain electromagnetic counterparts if the vector boson kinetically mixes with the ordinary photon Siemonsen et al.
 (2023a).
@@
 The superradiant instability of spin-2 fields occurs much faster than the others East and Siemonsen (2023), and since the backreaction of the instability has not been computed due to difficulties arising from a nonlinear coupling between the spin-2 particle and gravity, it is not yet known whether the observational signatures are similar to the vector or scalar cases, and whether there is a dependency on the specific nonlinear theory chosen.
+Such ultralight bosons could also be used to probe the quantum nature of BHs Mitra et al.
+(2024).
 V.1 Annihilation signal
@@
 (2019a).
-The GW signal will be modified due to tidal disruptions induced by the companion, and the multiple moments of the cloud Baumann et al.
-(2019a, b, 2020).
+The GW signal will be modified due to tidal disruptions induced by the companion, the multiple moments of the cloud Baumann et al.
+(2019a, b, 2020) or through a dynamical friction drag forced induced by the ultralight dark matter.
 These kinds of signals could be visible in future space-based detectors, since the deviations primarily occur in the early inspiral stages of the system, i.e.
 at frequencies much lower than currently accessible by ground-based GW interferometers.
-In both cases, observations of the inspiral of binary systems could indicate the presence of boson clouds around one or both of the objects.
+In both cases, observations of the inspiral of binary systems, and in particular extreme mass ratio inspiral (EMRI) systems Zhang and Yang (2019); Brito and Shah (2023), could indicate the presence of boson clouds around one or both of the objects if there is some dephasing with respect to the vacuum .
 Additionally, once the binary reaches a certain stage of its evolution, the impact of both the spin-induced quadrupole moment and tidal forces will lessen, since the clouds will become disrupted by their companions.
@@
 Krishnendu et al.
-(2017) showed that $\kappa$ can be used as a distinguishing parameter for different kinds of objects, and also a test of the no-hair theorem: as an example, $\kappa=1$ for BHs Hansen (1974), while $\kappa\sim[1.4,8]$ for neutron stars depending on the equation-of-state Laarakkers and Poisson (1999); Pappas and Apostolatos (2012).
+(2017) and Ref.
+Rahman and Bhattacharyya (2023) showed that $\kappa$ can be used as a distinguishing parameter for different kinds of objects, and also a test of the no-hair theorem: as an example, $\kappa=1$ for BHs Hansen (1974), while $\kappa\sim[1.4,8]$ for neutron stars depending on the equation-of-state Laarakkers and Poisson (1999); Pappas and Apostolatos (2012).
 In the case of boson cloud systems, $\kappa$ will take on vastly different values, of $\mathcal{O}(1000)$, depending on the boson cloud and BH masses, and the extent to which the Kerr metric is altered in the presence of a cloud with enough mass compared to the BH.
@@
 Recently, the Viterbi algorithm has also been adapted to search for GWs from annihilating vector boson clouds, which have timescales significantly shorter than in the scalar case Jones et al.
-(2023, 2024), and for which numerical waveforms must be used to gauge sensitivity Siemonsen and East (2020); Siemonsen et al.
+(2023, 2025), and for which numerical waveforms must be used to gauge sensitivity Siemonsen and East (2020); Siemonsen et al.
 (2023b).
 VI.2.3 Cross-correlation and other methods
-We have already discussed in Section III.1.1 cross correlating detector data to search for ultralight dark matter that could interact directly with GW interferometers; however, it is important to note that such a method could also be used to search for boson clouds around rotating BHs, both deterministically and for a stochastic GW background composed of the superposition.
+We have already discussed in Section III.1.1 cross correlating detector data to search for ultralight dark matter that could interact directly with GW interferometers; however, it is important to note that such a method could also be used to search for boson clouds around rotating BHs, both deterministically and for a stochastic GW background composed of the superposition of cosmological or astrophysical GW signals.
 Furthermore, virtually any method in the CW community could be tuned to look for such boson cloud systems because such methods already analyze almost monochromatic signals focusing on spin-downs but could also extend their parameter spaces to include spin-ups.
@@
 Furthermore, the stochastic GW background from tensor bosons around rotating BHs has been computed and searched for in Guo et al.
-(2024a), and has disfavored boson masses between $[10^{-13.4},10{-11.7}]$ eV depending on the range of uniform spin distributions assumed Guo et al.
-(2024a).
+(2024b), and has disfavored boson masses between $[10^{-13.4},10{-11.7}]$ eV depending on the range of uniform spin distributions assumed Guo et al.
+(2024b).
 VI.4 Prospects for GW probes of boson clouds
@@
 (2023b), new methods are being developed to search for annihilating vector boson clouds as well Jones et al.
-(2023, 2024).
+(2023, 2025).
 Additionally, there are also significant astrophysical uncertainties, namely: what are the distributions of BH spins, masses, distances and ages in the galaxy and beyond.
@@
 Potentially, new methods to track quickly-evolving waveforms, such as those in Jones et al.
-(2023, 2024), could be useful in a variety of contexts.
+(2023, 2025), could be useful in a variety of contexts.
 Furthermore, the all-sky search specifically for scalar boson cloud systems did not consider any grid in spin-up, which limited it to be sensitive to boson clouds whose frequency drift did not exceed one frequency bin during the entire observation time.
@@
 Furthermore, Ref.
-LABEL:Guo:2023lbv shows that bosons could be transferred from a growing level to a depleting one if both BHs in a binary system have clouds around them, which may affect the GW waveform and hence would need to be considered in future GW searches.
+Guo et al.
+(2024c) shows that bosons could be transferred from a growing level to a depleting one if both BHs in a binary system have clouds around them, which may affect the GW waveform and hence would need to be considered in future GW searches.
 VII GW probes of soliton DM
@@
 (2023, 2024) for more details.
+In contrast to TBHs, we note that if DM collects inside neutron stars, a different signature would be present.
+Depending on what equation-of-state assumed for baryonic or fermionic DM, neutron stars could actually be DM admixed neutron stars, which would have a particular equation-of-state that differs from the (currently unknown) canonical one.
+Evidence for such an effect could be found in changes to the tidal deformability parameter that could be measured by GW detectors in the future Collier et al.
+(2022).
 VIII.2 GWs to probe the GeV excess
@@
 (2001), or if DM particles scattered off of nearby stars Merritt (2004); Bertone and Merritt (2005).
-Though, recently, upper bounds on the DM densities for supermassive BHs have been forecasted with LISA Daniel et al.
+Though, recently, upper bounds on the DM densities for supermassive BHs have been forecast with LISA Daniel et al.
 (2025).
-However, intermediate-mass BHs (IMBHs) could sustain such spikes Zhao and Silk (2005), since their spikes may not have been torn apart through mergers over cosmic time.
+However, intermediate-mass BHs (IMBHs) could sustain such spikes Zhao and Silk (2005); Kim et al.
+(2023), since their spikes may not have been torn apart through mergers over cosmic time.
 Furthermore, recently it has been shown that astrophysical BHs and PBHs could also host DM spikes Kohri et al.
@@
 (2020); Bertone et al.
-(2024); Dosopoulou and Silk (2025).
+(2024); Bertone (2024); Dosopoulou and Silk (2025).
 The density of the DM spike is characterized by the following equation Eda et al. (2013):
@@
 (2013) first showed that the phase evolution of an intermediate-mass ratio inspiral (a binary composed to a $\mathcal{O}(1000)M_{\odot}$ and an $\mathcal{O}(1)M_{\odot}$ object) would be significantly modified at milli-hertz frequencies due to the presence of a DM spike around the intermediate-mass BH.
-This occurs because the less massive object sweeps through the DM spike as it orbits around the more massive one, losing energy via dynamical friction (in addition to via GWs) and speeding up the orbit.
+This occurs because the less massive object sweeps through the DM spike as it orbits around the more massive one, losing energy via dynamical friction (in addition to via GWs) and speeding up the orbit Speeney et al.
+(2022).
 Such effects will primarily be relevant in space-based GW detectors Duque et al.
+(2024); Rahman et al.
+(2024); Chowdhuri et al.
 (2024); Ghodla (2025), since IMBHs inspiral and merge at much lower frequencies than stellar-mass BHs Miller (2009).
@@
 (2024); Chen et al.
-(2024).
+(2024b).
 Recently, it has also been shown that these dephasing effects are unique enough to allow distinguishing between three BH environments: accretion disks, DM spikes, and ultralight boson clouds Cole et al.
@@
 (2018).
-But, for wave-like, ultralight dark matter, though there have been many works that have modeled the motion of binary black holes through DM environments,Koo et al.
+But, for wave-like, ultralight dark matter, though there have been many works that have modeled the motion of binary black holes through DM environments Koo et al.
 (2024); Boudon et al.
@@
 (2020a, b); Brax et al.
+(2020); Bromley et al.
 (2024) or with boson clouds Cao and Tang (2023); Traykova et al.
@@
 (2024); Guo et al.
-(2024b), the point during the inspiral at which the environment becomes disrupted is not clear, motivating further studies into this Bamber et al.
+(2024d), the point during the inspiral at which the environment becomes disrupted is not clear, motivating further studies into this Bamber et al.
 (2023).
@@
 The upper limits are shown to drastically improve in DECIGO because the dynamical friction (and other environments) affect the early inspiral of the binary black hole system much more than at the time of merger.
-These findings are consistent with those found in N-Body simulations about how the dynamics of the DM spike affect the inspiral, and the detectability of these effects by LISA Kavanagh et al.
+These findings are consistent with those found in N-Body simulations about how the dynamics of the DM spike affect the inspiral, and the detectability of these effects by LISA Cardoso and Maselli (2020); Kavanagh et al.
 (2020).
@@
 (2016); Clesse and García-Bellido (2017); Sasaki et al.
+(2016); Croon et al.
 (2023).

### v4

@@
 Our knowledge of physics has been incomplete for decades, if not for over a century Bertone and Hooper (2018).
-Dark energy and dark matter (DM) could comprise 69% and 26% of the universe, and yet we have not been able to directly detect either.
-The experimental evidence for the existence of some “invisible” or DM is overwhelming: stars orbiting the center of the Milky Way would have flown out of their orbits if DM did not hold them in place Freeman (1970); Rogstad and Shostak (1972); Whitehurst and Roberts (1972); Roberts and Rots (1973); anisotropies in the cosmic microwave background could result from DM potential wells in the early universe Peebles (1982); Burles and Tytler (1998); DM could seed the formation of large-scale structures visible today, which would have been impossible with ordinary matter alone Pagels and Primack (1982); Davis et al.
+Dark energy and dark matter (DM) comprise 69% and 26% of the universe, respectively, and yet we have not been able to directly detect either.
+The experimental evidence for the existence of DM is overwhelming: stars orbiting the center of the spiral galaxies would have flown out of their orbits if DM did not hold them in place Freeman (1970); Rogstad and Shostak (1972); Whitehurst and Roberts (1972); Roberts and Rots (1973); anisotropies in the cosmic microwave background can result from DM potential wells in the early universe Peebles (1982); Burles and Tytler (1998); DM can seed the formation of large-scale structures visible today, which would have been impossible with ordinary matter alone Pagels and Primack (1982); Davis et al.
 (1982); Blumenthal et al.
@@
 (1985); Navarro et al.
-(1996); and gravitational lensing of light by galaxies in between us and the light source could be explained by invisible mass inside the galaxy Paczynski (1986); Alcock et al.
+(1996); and gravitational lensing of light by galaxies in between us and the light source can be explained by DM inside the galaxy Paczynski (1986); Alcock et al.
 (1993); Aubourg et al.
@@
 (2007).
-While the observational evidence indicates that DM may exist, it does not explain what it is made of.
-Could it be a single new particle, such as the axion Peccei and Quinn (1977a, b); Weinberg (1978) or dark photon Galison and Manohar (1984); Holdom (1986); Pierce et al.
-(2018); Filippi and De Napoli (2020), a collection of new particles, or macroscopic, in the form of PBHs formed within $\sim 1$ second of the Big Bang Green and Kavanagh (2021)?
-Or, particle DM may (or may not) exist alongside PBHs, which may comprise a portion or the totality of DM Hawking (1971); Carr et al.
+While the observational evidence shows that DM exists, it does not explain what it is made of.
+Can it be a single new particle, such as the axion Peccei and Quinn (1977a, b); Weinberg (1978) or dark photon Galison and Manohar (1984); Holdom (1986); Pierce et al.
+(2018); Filippi and De Napoli (2020), or macroscopic, in the form of PBHs formed within $\sim 1$ second of the Big Bang Green and Kavanagh (2021)?
+Or, can particle DM exist alongside PBHs, which may comprise a portion or the totality of DM Hawking (1971); Carr et al.
 (2010); Dai et al.
 (2009); Clesse and García-Bellido (2015); Carr et al.
-(2016); Clesse and García-Bellido (2017); Clesse and Garcia-Bellido (2022); Miller (2024).
-The numerous hypotheses regarding the existence of DM suggest that the mass of its constituents can vary across hundreds of orders of magnitude Belenchia et al.
+(2016); Clesse and García-Bellido (2017); Clesse and Garcia-Bellido (2022); Miller (2024a)?
+The numerous hypotheses regarding the origins of DM suggest that the mass of its constituents can vary across hundreds of orders of magnitude Belenchia et al.
 (2022).
@@
 (2008, 2011); Aprile et al.
-(2012, 2016, 2024), and LUX Akerib et al.
-(2013, 2016).
+(2012, 2016, 2024), and LZ (LUX-Zeppelin) Malling et al.
+(2011); Akerib et al.
+(2013, 2016); Mount et al.
+(2017); Akerib et al.
+(2020).
 Other experiments have tried to detect annihilating DM indirectly through its gamma-ray emission in the Galactic Center, galaxy clusters, or other areas, e.g.
@@
 (2012).
-Despite these efforts, however, WIMPS remain undetected [^1] DAMA claims to see a periodic oscillation of a DM signal, though this result has not yet been reproduced Petriello and Zurek (2008), and so it is worthwhile to ask whether other models could better explain DM.
-Thus, the experimental development from the particle physics community has widened to include probing ultralight DM through (1) torsion balance experiments, such as Eöt-Wash Su et al.
+Despite these efforts, however, WIMPS remain undetected [^1] DAMA claims to see a periodic oscillation of a DM signal, though this result has not yet been reproduced Petriello and Zurek (2008), and so it is worthwhile to ask whether other models can better explain DM.
+Thus, the experimental development from the particle physics community has widened to include probing ultralight ($\ll 1$ eV) DM through (1) torsion balance experiments, such as Eöt-Wash Su et al.
 (1994); Schlamminger et al.
@@
 (2009); Akiyama et al.
-(2019), which could measure photons lensed by boson clouds and thus affect BH images Chen et al.
-(2023), and which could measure changes in the polarization of the electric field when the light passes through the axion background Chen et al.
+(2019), which can measure photons lensed by boson clouds and thus affect BH images Chen et al.
+(2023), and which can measure changes in the polarization of the electric field when the light passes through the axion background Chen et al.
 (2022, 2022), and (4) astrometry experiments, such as Gaia Prusti et al.
 (2016) and the Roman Space Telescope Akeson et al.
-(2019), that can meausre stochastic metric perturbations sourced by ultralight dark matter Kim (2024).
-It is therefore also worth asking whether high-precision GW interferometers, such as LIGO, Virgo and KAGRA Aasi et al.
+(2019), that can measure stochastic metric perturbations sourced by ultralight dark matter Kim (2024).
+In light of the diversity of experiments mentioned, it is also worth asking whether high-precision GW interferometers, such as LIGO, Virgo, KAGRA and GEO600 Affeldt et al.
+(2014); Dooley et al.
+(2016); Aasi et al.
 (2015); Acernese et al.
 (2015); Aso et al.
-(2013), can be used to search for DM, which has also been briefly reviewed in Bertone et al.
-(2020).
-Axions and other forms of DM could be probed via their interaction with binary neutron star inspirals that could be detected in the future with LIGO, Virgo and KAGRA Hook and Huang (2018); Croon et al.
-(2018).
-The observational signature of axions on binary neutron star inspirals would be a dephasing effect induced by each object having a scalar charge, and the mass of the scalar field Hook and Huang (2018); Zhang et al.
-(2021).
-Furthermore, a variety of follow-up studies Sagunski et al.
-(2018); Huang et al.
-(2019) have been performed regarding the detectability of the effects of axions on binary neutron star inspirals, with Ref.
-Zhang et al.
-(2021) placing the first-ever constraints on the nuclear coupling of axion-like particles with masses below $10^{-11}$ eV using the detection of GWs from the binary neutron star GW170817, effectively ruling out axion decay constants of [$1.6\times 10^{16}$,$10^{18}$] GeV.
-Furthermore, even the very slow orbit of the Hulse-Taylor binary and other known binaries constrains the axion decay constant to be below certain values depending on the system Kumar Poddar et al.
-(2020) to be $f_{a}<\mathcal{O}(10^{11})$ GeV, and resonances could appear during the orbit when the orbital frequency of the binary matches that of the axion Kadota et al.
-(2024); Kim and Yang (2024).
-Finally, axions and the DM energy density can also be constrained via the evidence for stochastic GW background from pulsar timing arrays Guo et al.
-(2024a); Chen et al.
-(2024a) and also explain the mass gap in observed BHs Croon and Sakstein (2023).
-In particular, in this review, we would like to explore how continuous gravitational-wave (CWs) can be used as a probe of DM.
+(2013), can be used to search for DM, as previously reviewed in Bertone et al.
+(2020); Bertone (2024).
+While it has been demonstrated that constraints on axions can be set using current observations of binary neutron star inspirals (GW170817) Hook and Huang (2018); Croon et al.
+(2018); Zhang et al.
+(2021) and the Hulse-Taylor binary Kumar Poddar et al.
+(2020), in this review, we would like to explore how continuous gravitational-wave (CWs) can be used as a probe of DM.
 CWs are quasi-monochromatic, quasi-infinite-duration signals that canonically arise from asymmetrically rotating, lumpy neutron stars.
@@
 (1998); Cutler (2002), changes in the magnetosphere Lyne et al.
-(2010), etc.) complicates these searches, and thus methods that are robust not only to non-Gaussian noise disturbances but also to theoretical uncertainties had to be developed to look for such systems Krishnan et al.
+(2010), etc.) complicate these searches, and thus methods that are robust not only to non-Gaussian noise disturbances but also to theoretical uncertainties had to be developed to look for such systems Krishnan et al.
 (2004); Dhurandhar et al.
@@
 (2019).
-Until now, such searches have only yielded constraints on the maximum possible deformation of the size of known Abbott et al.
+Until recently, such searches have only yielded constraints on the maximum possible deformation of the size of known Abbott et al.
 (2021a, 2022a); Ashok et al.
 (2021, 2024); Abac et al.
-(2025) and unknown neutron stars Abbott et al.
+(2025a) and unknown neutron stars Abbott et al.
 (2022b); Steltner et al.
 (2023).
-It turns out that with CWs, we can not only probe neutron stars, but also a wide range of DM models.
-If DM particles couple to standard model ones, we could see them in GW data via their interactions with detector components, which would give rise to a narrow-band, stochastic “correlated-noise” signal Pierce et al.
+In this review, we show how CW methods have been generalized to search for the following diverse set of DM models:
+  1. 1.
+    Ultralight bosons (scalar, vector, or tensor): Oscillating fields can couple to standard-model particles, leading to periodic modulations in the interferometer arm lengths, refractive indices, neutron-star moment of inertias, or timings of atomic clocks Khmelnitsky and Rubakov (2014); Pierce et al.
 (2018); Guo et al.
@@
 (2023).
-Separately, if a new boson that may or may not be DM exists and couples to gravity, and is of just the right mass, it will form clouds around rotating BHs, extracting mass and spin from them as they grow, and emitting CWs as the DM particles annihilate Brito et al.
+CW methods are sensitive to these narrowband, quasi-monochromatic signals and can be applied to searches for this kind of DM in ground-based interferometers, pulsar timing arrays, or future space-based antennas.
+Depending on the particular interferometer, ultralight dark matter masses between $10^{-22},10^{-11}]$ eV can be probed, showing how instruments not even designed to search for DM are able to cover a wide range of masses and thus a variety of DM models.
+  2. 2.
+    Macroscopic dark matter objects: Compact dark objects of masses of $\mathcal{O}(1-10^{9})$ kg passing through or near GW interferometers could induce transient but quasi-periodic disturbances in the test masses Hall et al.
+(2018); Du et al.
 (2023).
-Heavy, strongly-interacting DM could induce collapses of stars that form in binaries, and as these objects orbit each other, they will emit GWs at particular frequencies distinct from if they were ordinary stars of the same mass Singh et al.
+Searches using matched filtering can place constraints on Yukawa-type interactions or on the rate of such transient events.
+Even though individually, each transient would manifest as short-lived perturbation of the detector, if multiple encounters were to occur over the observation time, the resulting signature would appear as a stochastic background of transients.
+Potentially, CW methods could be applied to look for the periodicity or stochastic background of such transients.
+Moreover, networks of interferometers could help discriminate these events from terrestrial noise, providing both spatial and temporal correlations that enhance detection prospects.
+  3. 3.
+    Boson clouds around rotating black holes: Ultralight bosons can grow around black holes via superradiance, accelerated when the Compton wavelength of the boson matches the size of the BH, forming clouds that extract energy and momentum from the BH and emit nearly monochromatic GWs through annihilation or level transitions Arvanitaki and Dubovsky (2011); Arvanitaki et al.
+(2015a); Brito et al.
+(2015, 2017).
+CW searches can target these long-lived signals and test the existence of bosons in the $10^{-13},10^{-11}]$ eV mass range in ground-based GW interferometers, and between $10^{-17},10^{-15}]$ eV in future space-based GW antennas.
+  4. 4.
+    Soliton dark matter: Gravitational interactions can lead to the formation of solitonic cores of ultralight dark matter at the centers of galaxies Schive et al.
+(2014); Hui et al.
+(2017).
+Compact binaries or isolated neutron stars moving through solitons experience additional gravitational potentials that can modify the phase evolution of the GWs as binary systems inspiral, and as neutron stars spin down, respectively.
+For binaries, this manifests as a long-lived, cumulative dephasing effect during the inspiral Aghaie et al.
+(2024), while for isolated nonaxisymmetric neutron stars, frequency modulations in the emitted CWs Blas et al.
+(2025) are induced.
+In both cases, the effect is coherent over long timescales, making CW search methods particularly well suited to test these scenarios.
+Depending on the soliton mass and size, signatures would be detectable with pulsar timing array data (probing supermassive binary inspirals crossing galactic cores) or with next-generation terrestrial detectors such as ET and Cosmic Explorer (probing stellar-mass binaries and isolated pulsars embedded in soliton potentials).
+  5. 5.
+    WIMP dark matter: Even though WIMPs are not inherently wave-like, their capture in stars can induce their collapse into BHs Singh et al.
 (2021); Bhattacharya et al.
 (2024).
-Such ultralight particles could also interact with the material inside neutron stars, changing their moments of inertia and thus their CW emission Khmelnitsky and Rubakov (2014).
-CWs can be used to probe the millisecond pulsar hypothesis to the GeV excess, which, as it is more constrained, lends more strength to the WIMP DM annihilation hypothesis for the GeV excess Calore et al.
-(2019); Miller and Zhao (2023); Bartel and Profumo (2024).
-In LABEL:fig:summary, we show a landscape of the kinds of particle DM that could be probed with current and future interferometers, both by using GWs and and through direct detection of particles interacting with the instruments.
-There are thus so many different kinds of DM that can be searched for by using instruments that were not strictly designed to probe the existence of DM.
+CW searches can probe this scenario indirectly, by looking for the long-lived inspirals of BHs space-based GW interferometers.
+Additionally, null results from CW searches from millisecond pulsars in the Galactic Center can indirectly preference WIMP DM as responsible for the gamma-ray GeV excess emanating from the center of the Milky Way Calore et al.
+(2019); Miller and Zhao (2023).
+  6. 6.
+    Dark matter spikes around compact objects: Dense DM profiles near BHs inspiraling can generate dynamical friction as the secondary object flows through the DM spike of the first, leading to small but measurable dephasing in these long-lived inspirals Gondolo and Silk (1999); Eda et al.
+(2013, 2015); Kavanagh et al.
+(2020); Coogan et al.
+(2022); Cole et al.
+(2023a, b); Bertone (2024).
+While originally such effects were considered around extreme-mass ratio inspirals (EMRIs), intermediate-mass ratio inspirals (IMRIs) and ordinary binaries have been shown to allow the presence of DM spikes, and such dephasing could be measurable in ET, Cosmic Explorer or space-based GW interferometers, when inspiraling binaries will spend hours, days or years in-band.
+CW methods can thus test the presence and properties of such spikes.
+  7. 7.
+    Atomic dark matter: If DM possesses a composite structure with a dark sector analogue of electromagnetism, it could form bound states or “dark atoms” Feng et al.
+(2009); Kaplan et al.
+(2010); Buckley and DiFranzo (2018); Shandera et al.
+(2018).
+Such atomic DM can exhibit dissipative dynamics, allowing it to cool and collapse differently from standard cold DM.
+This can lead to the formation of compact dark objects, modify binary black hole merger populations, or even produce sub-solar mass binaries that would be invisible electromagnetically.
+While GW searches have constrained only the presence of such atomic DM using matched filtering searches, planetary- or asteroid-mass systems (masses $0.1M_{\odot}$) could also have been formed from atomic DM Fan et al.
+(2013a, b); Buckley and DiFranzo (2018), and would emit longer-duration GWs.
+Thus, CW searches can be applied to look for such systems Miller et al.
+(2022a); Abbott et al.
+(2022b); Miller et al.
+(2024).
+In each of aforementioned scenarios, considering DM within the framework of CWs allows us to adapt methods originally designed to probe neutron stars to search for DM.
+This is especially important given the variety and complexity of the underlying physics in different DM models, as well as the non-stationary, non-Gaussian noise characteristic of GW detectors.
+Traditionally, matched filtering has been the primary technique for detecting GW signals from binary black hole and binary neutron star mergers, in which known templates are correlated with observed data to extract weak signals embedded in noise.
+However, this approach becomes less effective when dealing with exotic scenarios, such as those discussed in this review, where the signals may not conform to standard templates, or where the creation of non-traditional templates is computationally infeasible.
+Importantly, we consider two general scenarios for the detection of DM using GW interferometers: (1) DM interacts either gravitationally or via couplings to standard model particles, leaving a direct imprint on interferometers, similar to traditional direct detection experiments; and (2) DM sources or modifies GW emission itself, which is then observed by interferometers.
+In Fig.
+1, we present a landscape of the types of particle DM that can be probed with current and future interferometers in the context of the first scenario.
+For each DM candidate, the figure maps out the interferometer component or astrophysical system it affects, the resulting observable, and the classes of interferometers capable of detecting the effect.
+Note that scenario (1) does not lead to GW signals; rather, the “signal” arises from the coupling of DM to the standard model.
+In contrast, Fig.
+2 summarizes the second scenario, showing how GW signals can be sourced or modified by the presence of DM.
+Together, these plots provide a high-level overview of the key models and observables discussed throughout this review.
+[…]
+Figure 2: Similar to Fig.
+1.
+This plot shows the kinds of DM that can be probed via their generation of GWs, both now and in the future.
+The logic is as follows, moving from left to right: type of DM $\rightarrow$ astrophysical source affected $\rightarrow$ observable $\rightarrow$ which detector(s) the signal could be seen in.
+We note that the mass range of particle DM probed here ranges many orders of magnitude, from $\mathcal{O}(10^{-23}-10^{-21})$ eV with pulsar timing arrays to $\mathcal{O}(10^{-16}-10^{-13})$ eV with space-based detectors to $\mathcal{O}(10^{-13}-10^{-11})$ eV with ground-based interferometers, and includes $10^{3},10^{9}]$ GeV WIMP DM, probes of large-scale DM structure via solitons and DM spikes, and atomic DM.
+Matching colors imply that the source would induce those physical signatures.
 This review article is broken into different parts depending both on the type of DM considered and constraints on DM interactions with gravity or standard model particles.
@@
 Section IV focuses on how macroscopic DM could transit through the interferometers, giving rise to a measurable differential acceleration of the mirrors.
-In Section V, we describe how scalar and vector bosons (that could be DM) form around rotating BHs, and the types of GW signals that could be emitted if these BHs are isolated or are in binary systems.
+In Section V, we describe how scalar and vector bosons form around rotating BHs, and the types of GW signals that could be emitted if these BHs are isolated or are in binary systems.
 Search results for annihilating scalar and vector boson cloud systems are explained in Section VI.
 Next, we discuss how soliton DM would impact GW signals from binary systems and neutron stars in Section VII.
-The collection of infamous WIMPs around BHs and in the Galactic Center are discussed in Section VIII in the context of how they can be probed with GWs.
+The collection of WIMPs around BHs and in the Galactic Center are discussed in Section VIII in the context of how they can be probed with GWs.
 We explain how DM spikes around rotating BHs could alter GW signals in Section IX.
@@
 We conclude in Section XI about the future for the burgeoning field of GW probes of particle DM.
-II Ultralight dark matter interacting with the interferometers
-II.1 Generic features of ultralight dark matter signal
-If it exists, cold, ultralight DM could interact with standard-model particles in model-dependent ways.
+Note that we use natural units throughout this review: $c=\hbar=\epsilon_{0}=1$.
+II Ultralight DM interacting with GW interferometers
+Numerous models for DM exist throughout the literature, and one review cannot possibly cover all of them.
+Instead, in this section, we focus specifically on ultralight dark matter models that have been shown to cause observable signals in GW interferometers.
+Over the last decade, the impact of more and more DM models on ground- and space-based instruments has been considering, allowing the field of “direct DM searches with GW interferometers” to grow.
+The models we present here should be taken as the minimum number of ways in which DM could affect GW interferometers.
+In Section II.1, we outline some general properties of all ultralight dark matter considered in this review, and then delve into four types of DM that couple to the standard model: axions (Section II.2), dilatons (Section II.3), vector, dark photon DM (Section II.4), tensor DM (Section II.5).
+We end with a scalar DM field that couples only to gravity (Section II.6).
+II.1 Generic features of ultralight dark matter signals
+Cold, ultralight DM could interact with standard-model particles in many model-dependent ways.
 Such interactions would cause macroscopic differences in the materials, and, depending on the type of DM, would lead to an observable signal in different components of GW interferometers.
@@
 First, the number of ultralight DM particles in a given region in space is gigantic, and can be calculated by attributing the DM energy density $\rho_{\text{DM}}$ to result from ultralight DM:
-$N_{0}$ $=$ $\lambda_{\rm C}^{3}\frac{\rho_{\text{DM}}}{m_{\rm DM}c^{2}}=\left(\frac{2\pi\hbar}{m_{\rm DM}v_{0}}\right)^{3}\frac{\rho_{\text{DM}}}{{m_{\rm DM}c^{2}}},$ (1)
+$N_{0}$ $=$ $\lambda_{\rm C}^{3}\frac{\rho_{\text{DM}}}{m_{\rm DM}}=\left(\frac{2\pi}{m_{\rm DM}v_{0}}\right)^{3}\frac{\rho_{\text{DM}}}{{m_{\rm DM}}},$ (1)
 $\approx$ $1.69\times 10^{54}\left(\frac{10^{-12}\text{ eV}}{m_{\rm DM}}\right)^{4},$
 where $v_{0}\simeq 220$ km/s is the virial velocity Smith et al.
-(2007), $\hbar$ is Planck’s reduced constant, $c$ is the speed of light, $m_{\rm DM}$ is the mass of the ultralight dark matter particle, $\lambda_{\rm C}$ is the Compton wavelength of DM, and $\rho_{\text{DM}}=0.3\,\mathrm{GeV}/$m3 is the DM energy density de Salas and Widmark (2021) (unless otherwise noted).
-The large occupation number $N_{0}$ implies that ultralight DM can be treated as a coherent region, where within that region, DM can be approximated as a plane wave that oscillates at a fixed frequency $f_{0}$, which is given by the DM mass:
-$f_{0}=\frac{m_{\rm DM}c^{2}}{2\pi\hbar}\simeq 241\text{ Hz}\left(\frac{m_{\rm DM}}{10^{-12}\text{ eV}/c^{2}}\right),$ (2)
-In other words, the observable effect is that the interferometer components will forever oscillate at a fixed frequency since they always sit in the DM field Hall et al.
+(2007), $m_{\rm DM}$ is the mass of the ultralight dark matter particle, $\lambda_{\rm C}$ is the Compton wavelength of DM, and $\rho_{\text{DM}}=0.3\,\mathrm{GeV}/$cm3 is the DM energy density de Salas and Widmark (2021).
+The large occupation number $N_{0}$ implies that ultralight DM can be approximated as a plane wave that oscillates at a fixed frequency $f_{0}$, which is given by the DM mass:
+$f_{0}=\frac{m_{\rm DM}}{2\pi}\simeq 241\text{ Hz}\left(\frac{m_{\rm DM}}{10^{-12}\text{ eV}}\right),$ (2)
+In other words, the observable effect is that the interferometer components will forever oscillate at a fixed frequency since they always sit in the DM field Pierce et al. (2018).
+Despite the fixed frequency implied by Eq.
+2, we note that in reality, the DM field is not a perfectly monochromatic wave because the constituent particles have a finite Maxwell-Boltzmann velocity dispersion set by the virial velocity $v_{0}$ of the galactic halo.
+This spread in velocities causes a small but finite spread in the oscillation frequency, which in turn limits the coherence of the field.
+As a result, the DM wave can be treated as coherent only over a characteristic time scale $T_{\text{coh}}$, after which the accumulated phase dispersion between different velocity components causes the signal to lose coherence Carney et al.
 (2021).
-The DM wave also has a finite coherence time $T_{\text{coh}}$ that relates to the virial velocity of DM:
-$T_{\text{coh}}=\frac{4\pi\hbar}{m_{\rm DM}v_{0}^{2}}=1.4\times 10^{4}\text{ s}\left(\frac{10^{-12}\text{ eV}}{m_{\rm DM}}\right),$ (3)
-If we observe for a duration longer than the coherence time, the signal will no longer be monochromatic but have its power stochastically distributed at frequencies slightly higher than the characteristic frequency of the DM particle:
-$\Delta f_{v}=\frac{1}{2}\left(\frac{v_{0}}{c}\right)^{2}f_{0}\approx 2.94\times 10^{-7}f_{0}.$ (4)
-This frequency shift arises because in a given coherent region, the number of DM particles is huge, and all are traveling at different velocities that are Maxwell-Boltzmann distributed around the virial velocity.
-Furthermore, the DM signal has an associated coherence length $L_{\text{coh}}$:
-$L_{\text{coh}}=\frac{2\pi\hbar}{m_{\rm DM}v_{0}}=1.6\times 10^{9}\text{ m}\left(\frac{10^{-12}\text{ eV}/c^{2}}{m_{\rm DM}}\right).$ (5)
-If observing the interaction of DM with standard-model model particles in detectors separated by distances less than $L_{\text{coh}}$, which is the case for ground-based detectors, we expect to observe a correlated signal, again at an almost fixed frequency.
-In Fig.
-2, we show an example of what a simulated vector DM signal interacting with one of the LIGO detectors looks like.
+The wave coherent time $T_{\text{coh}}$ is given by Carney et al.
+(2021):
+$T_{\text{coh}}=\frac{4\pi}{m_{\rm DM}v_{0}^{2}}=1.4\times 10^{4}\text{ s}\left(\frac{10^{-12}\text{ eV}}{m_{\rm DM}}\right),$ (3)
+Practically speaking, we can observe the interaction of DM with ground-based GW interferometers for as long as the detectors are on.
+Since observing runs of LIGO, Virgo and KAGRA span around one year, we will always be in a regime in which the signal will not be monochromatic and thus have its power stochastically distributed at frequencies slightly higher than the characteristic frequency of the DM particle.
+The following expression sets the range of frequencies for which this happens:
+$\Delta f_{v}=\frac{1}{2}v_{0}^{2}f_{0}\approx 2.94\times 10^{-7}f_{0}.$ (4)
+Additionally, a finite coherence time also implies a finite coherence length $L_{\text{coh}}$:
+$L_{\text{coh}}=\frac{2\pi}{m_{\rm DM}v_{0}}=1.6\times 10^{9}\text{ m}\left(\frac{10^{-12}\text{ eV}}{m_{\rm DM}}\right).$ (5)
+This length scale characterizes the maximum separation over which two detectors can still observe the DM-induced oscillations with the same phase.
+Since the arm lengths and baselines of current ground-based GW interferometers are much smaller than $L_{\text{coh}}$, we expect them to record correlated signals at the same frequencies.
+However, $L_{\text{coh}}$ increases with decreasing $m_{\rm DM}$, which will not be a problem for spaced-based GW interferometers ($L_{\text{coh}}$ will still exceed arm separation and $T_{\text{coh}}$ will greatly exceed the observation time), but does require careful treatment when using pulsar timing arrays, in which $L_{\text{coh}}$ is of the same order as the separation between pulsars.
+This will be discussed further in Section III.3.
+The finite $L_{\text{coh}}$ and $T_{\text{coh}}$ naturally imply that cross-correlation and other CW methods are natural choices to look for this kind of DM, which will be detailed in Section III.1.
+Additionally, we show a concrete example of a potential ultralight dark-matter signal in Fig.
+3.
+Here, simulated vector DM particles are interacting with one of the LIGO detectors, which produces a stochastic, narrow-band signal in the frequency domain.
 Figure 3: Taken from Miller et al.
-(2021).
+(2021a).
 We show the strain time series $h(t)$ (left) and the Fourier transform of it (right) of a simulated dark photon DM signal.
-The structure in the amplitude spectrum density in the right-hand plot arises because we have taken the length of the Fourier transform to exceed that of the coherence time of the signal, and the signal is a superposition of $1000$ dark photons traveling with distinct Maxwell-Boltzmann-distributed velocities, which cause small frequency deviations away from the minimal frequency $f_{0}=740.436$ Hz ($m_{A}=3.06\times 10^{-12}$ eV$/c^{2}$).
+The structure in the amplitude spectrum density in the right-hand plot arises because we have taken the length of the Fourier transform to exceed that of the coherence time of the signal.
+This dark photon DM signal is a superposition of $1000$ dark photons traveling with distinct Maxwell-Boltzmann-distributed velocities, which cause small frequency deviations away from the minimal frequency $f_{0}=740.436$ Hz ($m_{A}=3.06\times 10^{-12}$ eV).
 This plot shows why it is important to choose $T_{\text{FFT}}\sim T_{\text{coh}}$ for the analysis of DM signals.
@@
 We will assume that we are working within $T_{\text{coh}}$, in order to avoid technical difficulties from working with particles of slightly different frequencies.
-However, we note that in practice, this consideration is necessary when modelling the signal and assessing the sensitivity of GW interferometers to DM/standard-model interactions.
+However, we note that in practice, this consideration is necessary when assessing the sensitivity of GW interferometers to DM/standard-model interactions Pierce et al.
+(2018); Miller et al.
+(2021a); Vermeulen et al.
+(2021); Göttel et al.
+(2024).
 II.2 Axions
-Axions could couple to photons, given by the following Chern-Simons interaction Lagrangian:
+Axions were originally proposed as a solution to the strong CP problem in QCD Peccei and Quinn (1977a, b), and more generally, axion-like particles arise in many extensions of the standard model.
+If sufficiently light and weakly interacting, they can also constitute a viable component of DM.
+A key feature of axions is their ability to couple to standard-model fields, in particular photons, through a Chern–Simons interaction.
+This coupling is described by the Lagrangian Chern and Simons (1974):
 $\mathcal{L}_{I}$ $=\frac{g_{\text{a}\gamma}}{4}a(t)F_{\mu\nu}\tilde{F}^{\mu\nu}$
-[…]
-where $a(t)$ is the axion field, $g_{\text{a}\gamma}$ is the axion-photon coupling constant, $F_{\mu\nu}$ is field strength of electromagnetic field, $\tilde{F}^{\mu\nu}\equiv\epsilon^{\mu\nu\rho\sigma}F_{\rho\sigma}/2$ is its Hodge dual with the Levi-Civita anti-symmetric tensor $\epsilon^{\mu\nu\rho\sigma}$, the dot denotes the time derivative, and $A_{\mu}$ is the electromagnetic vector potential.
+$=g_{\text{a}\gamma}\dot{a}(t)\epsilon_{ijk}A_{i}\partial_{j}A_{k},$ (6)
+where $a(t)$ is the axion field, $g_{\text{a}\gamma}$ is the axion-photon coupling constant, $F_{\mu\nu}$ is field strength of electromagnetic field, $\tilde{F}^{\mu\nu}\equiv\epsilon^{\mu\nu\rho\sigma}F_{\rho\sigma}/2$ is its Hodge dual with the Levi-Civita anti-symmetric tensor $\epsilon^{\mu\nu\rho\sigma}$, the dot denotes the time derivative, $\epsilon_{ijk}$ is Levi-Civita tensor, and $A_{\mu}$ is the electromagnetic vector potential.
 $F_{\mu\nu}\equiv\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$.
-All spatial derivatives here are being neglected.
+Because the coherence length of the axion is much larger than typical experimental setups, the axion field can be treated as homogeneous; thus, all spatial derivatives of the axion field here are neglected.
 In Fourier space, we can write $A_{i}$ in terms of two circular polarization modes:
 $A_{i}(t,\bm{x})=\sum_{\lambda=\text{L},\text{R}}\int\dfrac{\mathrm{d}^{3}k}{(2\pi)^{3}}A_{\lambda}(t,\bm{k})\,e^{\lambda}_{i}(\hat{\bm{k}})\,e^{i\bm{k}\cdot\bm{x}}\ ,$ (7)
-where $\bm{k}$ is the wave number vector, and “L” and “R” denote the left- and right-hand polarizations, and $\bm{x}$ is a spatial point at which the DM field is measured.
-By completely fixing gauge degrees of freedom ($\partial_{i}A_{i}=0$ and $A_{0}=0$, respectively), and noting how the polarization vectors $e^{\lambda}_{i}(\hat{\bm{k}})$ in different directions relate to each other Nagano et al.
-(2019), the equations of motion for $A_{\text{L/R}}(t,\bm{k})$ can be derived:
+where $\bm{k}$ is the wave number vector, $k\equiv|\bm{k}|$, and “L” and “R” denote the left- and right-hand polarizations, and $\bm{x}$ is a spatial point at which the DM field is measured.
+By completely fixing gauge degrees of freedom in which the divergence of the spatial part of the field is zero and its temporal component vanishes ($\partial_{i}A_{i}=0$ and $A_{0}=0$), and noting that the following conditions are satisfied:
+  1. 1.
+[…]
+  2. 2.
+    The two circularly polarized states are orthonormal: $e^{\lambda}_{i}(\hat{\bm{k}})e^{\lambda^{\prime}*}_{i}(\hat{\bm{k}})=\delta^{\lambda\lambda^{\prime}}$
+  3. 3.
+[…]
+the equations of motion for $A_{\text{L/R}}(t,\bm{k})$ can be derived Nagano et al. (2019):
 $\ddot{A}_{\text{L/R}}+\omega^{2}_{\text{L/R}}A_{\text{L/R}}$ $=0$ (8)
@@
 Noting that the axion field can be written as (assuming to work within one $T_{\text{coh}}$:
-$a(t)=a_{0}\cos(m_{\rm DM}t)=\left(\frac{\hbar\sqrt{2\rho_{\text{DM}}}}{m_{\rm DM}c}\right)\cos(m_{\rm DM}t)$ (10)
+$a(t)=a_{0}\cos(m_{\rm DM}t)=\left(\frac{\sqrt{2\rho_{\text{DM}}}}{m_{\rm DM}}\right)\cos(m_{\rm DM}t)$ (10)
 where $a_{0}$ is the axion field amplitude normalized by the DM energy density, and $t$ is time.
@@
 in which the laser light is assumed to have a wavelength that relates to $k$ as: $\lambda=2\pi/k$.
-For our purposes, in the presence of an axion field, the phase velocities of circularly polarized photons in the interferometers are modulated Nagano et al.
+Eq.
+11 implies that in the presence of an axion field, the phase velocities of circularly polarized photons in interferometers are modulated.
+In order to detect this effect in LIGO, Virgo and KAGRA instruments, additional but simplistic optical components are required in order to measure the optical path difference between two orthogonal polarizations of light ($p$- and $s$-polarizations) Nagano et al.
 (2021).
-This signal can only be seen if additional but simplistic optical components are added to measure the optical path difference between two orthogonal polarizations of light ($p$- and $s$-) Nagano et al.
-(2021).
-In practice, GW detectors work with linearly polarized light ($p$-) that will be modulated by axions, making $s$- polarized light.
-Furthermore, such deviations in polarizations can also be measured in pulsar timing arrays Xue et al.
+In practice, GW interferometers use linearly polarized light ($p$-polarized), which becomes partially $s$-polarized due to axion-induced modulation.
+Similar deviations in polarization can also be probed with pulsar timing arrays Xue et al.
 (2024); Porayko et al.
-(2024).
+(2025), and experimental constraints on axions from them will be discussed in Section III.3.5.
+Notably, ground-based interferometers and pulsar timing arrays are sensitive to very different axion mass ranges, roughly $10^{-13},10^{-11}]$ eV and $10^{-23},10^{-21}]$ eV, respectively, but can achieve comparable absolute constraints on the axion-photon coupling $g_{\text{a}\gamma}$, which vary strongly with the axion mass.
+However, no constraints exist yet for this effect from ground-based interferometers.
 II.3 Dilaton scalar dark matter
-Dilaton-like DM Preskill et al.
+Another class of ultralight dark matter candidates are dilatons, scalar fields that can couple to standard-model particles and induce time-dependent variations in fundamental constants Preskill et al.
 (1983); Abbott and Sikivie (1983); Dine and Fischler (1983); Cho and Keum (1998); Cho and Kim (2009); Arvanitaki et al.
-(2015a); Stadnik and Flambaum (2015a, b, 2016) could cause the rest mass of the electron and other physical constants to oscillate with time.
-In other words, atoms’ Bohr radii of different interferometer components would be modulated at the dilaton frequency Grote and Stadnik (2019), and would induce changes in the size and index of refraction of the beam splitter Vermeulen et al.
-(2021) and the arm mirrors Göttel et al.
-(2024).
-Since beam-splitter size oscillations would asymmetrically affect both rays, DM effects would lead to a differential phase shift (i.e.
-a measurable signal).
-In fact, at first the beam splitter effect was assumed to be the only one that mattered, but recently, the finite differences between the thicknesses of the mirrors in each arm of the interferometer lead to non-negligible differences in the path lengths of the light in the interferometer Göttel et al.
-(2024).
-Similarly to the axion, the scalar ultralight dark-matter field $\phi$ can be written as Arvanitaki et al. (2015a); Derevianko (2018); Vermeulen et al. (2021):
-$\phi(t,\vec{r})=\left(\frac{\hbar\sqrt{2\rho_{\rm{DM}}}}{m_{\rm DM}c}\right)\cos\left(\omega t-\vec{k}\cdot\vec{r}+\Psi\right),$ (13)
-where $\vec{r}$ is a position vector, $\omega=(m_{\rm DM}c^{2})/\hbar$ is the angular Compton frequency, $\vec{k}=(m_{\rm DM}\vec{v}_{\text{obs}})/\hbar$ is the wave vector,$\Psi$ is a phase factor, and $\vec{v}_{\text{obs}}$ is the velocity of the DM relative to the observer.
-The Lagrangian $\mathcal{L}_{\mathrm{int}}$ for this scalar field $\phi$ is Vermeulen et al. (2021):
+(2015b); Stadnik and Flambaum (2015a, b, 2016).
+In particular, a dilaton field can cause the electron mass and other physical constants to oscillate at the dilaton frequency.
+Similarly to the axion, the scalar ultralight dark-matter field $\phi$ can be written as Arvanitaki et al.
+(2015b); Derevianko (2018); Vermeulen et al.
+(2021):
+$\phi(t,\vec{x})=\left(\frac{\sqrt{2\rho_{\rm{DM}}}}{m_{\rm DM}}\right)\cos\left(m_{\rm DM}t-\vec{k}\cdot\vec{x}+\Psi\right),$ (13)
+where $\vec{x}$ is a position vector, $\vec{k}=m_{\rm DM}\vec{v}_{0}$ is the wave vector, and $\Psi$ is a phase factor, and has a corresponding Lagrangian $\mathcal{L}_{\mathrm{int}}$ Vermeulen et al.
+(2021):
 $\mathcal{L}_{\mathrm{int}}\supset\frac{\phi}{\Lambda_{\gamma}}\frac{F_{\mu\nu}F^{\mu\nu}}{4}-\frac{\phi}{\Lambda_{e}}m_{e}\bar{\psi}_{e}\psi_{e},$ (14)
 where $\psi_{e}$ and $\bar{\psi}_{e}$ are the standard-model electron field and its Dirac conjugate, $m_{e}$ is the mass of the electron, and $\Lambda_{\gamma}$ and $\Lambda_{e}$ denote the scalar DM coupling parameters to the photon and electron, respectively.
-For comparison, we write a part of the standard model Lagrangian as:
-$\mathcal{L}_{\textrm{SM}}\supset-\frac{F_{\mu\nu}F^{\mu\nu}}{4}-m_{e}\bar{\psi}_{e}\psi_{e}-eJ_{\mu}A^{\mu}$ (15)
-and we can see that the rest mass of the electron effectively changes in the presence of DM:
+We will now show how the fundamental constants will change in response to the dilaton field. To begin, we write a part of the standard model Lagrangian that will be compared to Eq. 14:
+$\mathcal{L}_{\textrm{SM}}\supset-\frac{F_{\mu\nu}F^{\mu\nu}}{4}-m_{e}\bar{\psi}_{e}\psi_{e}-eJ_{\mu}A^{\mu},$ (15)
+where $J^{\mu}$ is a current term. From comparing Eq. 14 and Eq. 15 term-wise, we can see that the rest mass of the electron effectively changes in the presence of DM:
 $m_{e}^{\prime}=m_{e}\left(1+\frac{\phi}{\Lambda_{e}}\right)$ (16)
@@
 $\alpha^{\prime}=\frac{e^{\prime 2}}{4\pi}=\alpha\left(1-\frac{\phi}{\Lambda_{\gamma}}\right)^{-1}\simeq\alpha\left(1+\frac{\phi}{\Lambda_{\gamma}}\right)$ (19)
+Thus, the fundamental constants oscillate in response to the dilaton field.
+In the following subsections, we will discuss how these variations induce measurable strain in ground-based interferometers and pulsar timing arrays through their effects on solids, neutron stars, and atomic clocks.
 II.3.1 Changes to size and indices of refraction of solids
-The changes in fundamental constants will lead to a change in length of, or strain on, solid bodies.
-Noting that the length $l$ of an object is roughly $l\sim N_{A}a_{B}$, where $N_{A}$ is the number of atoms in the body and $a_{B}=(m_{e}\alpha)^{-1}$ is the Bohr radius, we can obtain the strain on a solid body in general:
-$\delta l=-N_{A}(m_{e}\alpha)^{-2}(m_{e}\delta\alpha+\alpha\delta m_{e})$ (20)
-$\frac{\delta l}{l}=-\left(\frac{\delta\alpha(t)}{\alpha}+\frac{\delta m_{e}(t)}{m_{e}}\right)$ (21)
+The oscillations in fundamental constants induced by a dilaton field lead to modulations of the Bohr radii of atoms in interferometer components Grote and Stadnik (2019), leading to measurable changes in the size and refractive index of the beam splitter Vermeulen et al.
+(2021), and the arm mirrors Göttel et al.
+(2024).
+Oscillations in the beam splitter size cause photons traveling down each arm to traverse slightly different paths along its surface, producing a differential phase shift.
+Initially, it was assumed that the beam-splitter effect was the dominant contribution Vermeulen et al.
+(2021), but more recent work has shown that small differences in the thicknesses of the mirrors in each arm also lead to non-negligible path-length variations in the interferometer Göttel et al.
+(2024).
+We will now calculate the impact of dilatons on GW interferometers.
+A typical solid length can be estimated as roughly $l\sim N_{A}a_{B}$, where $N_{A}$ is the number of atoms in the body and $a_{B}=(m_{e}\alpha)^{-1}$ is the Bohr radius.
+Small changes in the electron mass or fine-structure constant then produce a fractional change in length, or strain, given by
+$\delta l$ $=-N_{A}(m_{e}\alpha)^{-2}(m_{e}\delta\alpha+\alpha\delta m_{e})$
+$\frac{\delta l}{l}$ $=-\left(\frac{\delta\alpha(t)}{\alpha}+\frac{\delta m_{e}(t)}{m_{e}}\right)$
 $\delta l$ $=l\phi\left(\frac{1}{\Lambda_{\gamma}}+\frac{1}{\Lambda_{e}}\right)$ (20)
@@
 $\delta\alpha$ $=\alpha^{\prime}-\alpha=\alpha\frac{\phi}{\Lambda_{\gamma}}$ (22)
-Then, the differential displacement $\delta(L_{x}-L_{y})$ on the beam splitter in GW interferometers can be derived Grote and Stadnik (2019).
-Along the $x$-arm, assuming that light is incident on the beam splitter at 45 degrees, a reflectivity of the mirror of 50%, the change in the $x$-arm length, $\delta L_{x}$, is calculated as:
+In the context of GW interferometers, this effect translates into a differential displacement of the beam splitter along the two arms.
+For the $x$-arm, assuming light incidence at 45 degrees and 50% mirror reflectivity, the induced change in arm length, $\delta L_{x}$, can be computed as in Grote and Stadnik (2019):
 $\delta L_{x}\approx\delta[\sqrt{2}nl-l/(2\sqrt{2})-w/2],$ (23)
-where $w$ is the thickness of the mirrors. Along the $y$-arm, the change in arm length is:
+where $w$ is the thickness of the mirrors, $n$ is the index of the beamsplitter, and $l$ is the length of the beamsplitter. Along the $y$-arm, the change in arm length is Vermeulen et al. (2021):
 $\delta L_{y}=-\delta l/(2\sqrt{2})-\delta w/2,$ (24)
@@
 $\delta(L_{x}-L_{y})$ $\approx\sqrt{2}\left[\left(n-\frac{1}{2}\right)\delta l\right]$ (25)
-$\approx\left(\frac{1}{\Lambda_{\gamma}}+\frac{1}{\Lambda_{e}}\right)\left(\frac{n\,l\,\hbar\,\sqrt{2\,\rho_{\mathrm{DM}}}}{m_{\rm DM}\,c}\right)\cos\left(\omega t\right),$ (28)
-where $n$ and $l$ are the index of refraction and length of the beam splitter, respectively.
-We only consider the $\delta l$ contribution, not one from $\delta n$, because it is much larger than than the $\delta n$ one.
-Additionally, dilatons could be visible in a different part of the interferometers, that is, in the reference cavity.
+$\approx\left(\frac{1}{\Lambda_{\gamma}}+\frac{1}{\Lambda_{e}}\right)\left(\frac{n\,l\,\sqrt{2\,\rho_{\mathrm{DM}}}}{m_{\rm DM}\,}\right)\cos\left(m_{\rm DM}t\right),$ (26)
+We have only considered the contribution arising from the change in length of the beamsplitter, not from the change in index of refraction, because the former is much larger than the latter Grote and Stadnik (2019).
+In addition to dilatons interacting with atoms in the beamsplitter, the reference cavity of the detector can provide a region in which dilatons interact appreciably with light, allowing for potential detection.
 Laser light locked to an optical cavity made out of a solid material will have its frequency modulated due to the presence of a dilaton DM field, compared to light locked to a free-space suspended cavity, whose frequency would remain stable Derevianko (2018); Geraci et al.
 (2019); Hall and Aggarwal (2022).
-The strain amplitude of such a signal can be written as:
-${h}_{\text{DM}}(t)=\left[\frac{\hbar}{m_{\rm DM}c^{3}}\sqrt{8\pi\rho_{\text{DM}}G}(d_{e}+d_{m_{e}})\right]\cos\omega t$ (29)
+The strain amplitude of such a signal, which would be distinct from that in Eq.
+26, can be written as:
+${h}_{\text{DM}}(t)=\left[\frac{1}{m_{\rm DM}}\sqrt{8\pi\rho_{\text{DM}}G}(d_{e}+d_{m_{e}})\right]\cos m_{\rm DM}t,$ (27)
 where $d_{e}$ and $d_{m_{e}}$ relate to the other couplings as:
@@
 where $M_{\mathrm{Pl}}$ is the Planck mass, and $m_{e}$ is the electron mass.
+The effects described in Eq.
+26 and Eq.
+27 can both produce measurable signals in interferometers.
+Currently, however, only the spatial strain affecting mirrors and the beamsplitter has motivated dedicated search methods for dilaton dark matter (see Section III.1) and has been constrained using data from GEO600 and LIGO (see Section III.2).
+A search strategy targeting the temporal modulation of the laser frequency is still under development Hall and Aggarwal (2022).
 II.3.2 Changes to neutron-star moment of inertia
-Fluctuations in fundamental constants could also cause other effects, such as changing the moment of inertia of neutron stars Kaplan et al.
-(2022), or changing the times measured cesium clocks Kaplan et al.
-(2022).
-These effects will be relevant when we describe how ultralight dark matter could be detected in pulsar timing arrays.
-To study the former, we must also consider interactions in QCD Damour and Donoghue (2010); Kaplan et al.
+Beyond laboratory-based detectors, dilaton-induced fluctuations in fundamental constants could also have observable consequences in astrophysical systems.
+For example, they could alter moment of inertia of neutron stars Kaplan et al.
+(2022), which would produce deviations in the arrival time of pulses, to be discussed in Section III.3.
+In order to quantify the extent to which the neutron star moment of inertia will change, we must know how protons and neutrons will change in response to dilatons.
+This motivates us to consider the following interactions in QCD of quarks in the presence of dilatons Damour and Donoghue (2010); Kaplan et al.
 (2022):
 $\mathcal{L}_{\phi,\rm QCD}\supset\frac{\phi}{\Lambda}\left(\frac{d_{g}\beta_{3}}{2g_{3}}G_{\mu\nu}^{A}G_{A}^{\mu\nu}-\sum_{q=u,d}(d_{q}+\gamma_{q}d_{g})m_{q}\bar{q}q\right)$ (29)
-where $\beta_{3}$ is the QCD beta function, $g_{3}$ is the QCD gauge coupling, $G_{\mu\nu}^{A}$ is the gluon field strength tensor, $\gamma_{q}$ are the light quark anomalous dimensions, $\Lambda=M_{\text{Pl}}/\sqrt{4\pi}$, and $d_{g}$, and $d_{q}$ are the dilaton coupling constants to the gluon and the quarks, respectively.
-Reading off the Lagrangian, and noting the difference from the standard model, the change in masses of quarks is:
-$\frac{\delta m_{q}}{m_{q}}$ $=\frac{d_{q}}{\Lambda}\phi\,,$ (32)
-while in nucleon masses:
+where $\beta_{3}$ is the QCD beta function, $g_{3}$ is the QCD gauge coupling, $G_{\mu\nu}^{A}$ is the gluon field strength tensor, $\gamma_{q}$ are the light quark anomalous dimensions, $\Lambda=M_{\text{Pl}}/\sqrt{4\pi}$, $m_{q}$ is the quark mass, and $d_{g}$ and $d_{q}$ are the dilaton coupling constants to the gluon and the quarks, respectively.
+Similarly to Section II.3, we can write the standard-model Lagrangian for quarks:
+$\mathcal{L}_{\text{quark}}^{\text{SM}}=-\sum_{q}m_{q}\bar{q}q,$ (30)
+which, upon comparing with Eq. 29, allows us to read off the change in quark mass as
+$\frac{\delta m_{q}}{m_{q}}$ $=\frac{d_{q}}{\Lambda}\phi.$ (31)
+The shift in nucleon masses can then be obtained by noting that nucleons are composite objects whose masses arise mostly from QCD binding energy, with a smaller contribution from the quark masses.
+The dilaton couples to both the quark masses and the gluon field strength, so the fractional change in a nucleon mass is a weighted sum of these effects Damour and Donoghue (2010); Kaplan et al.
+(2022):
 $\frac{\delta m_{p,n}}{m_{p,n}}\simeq\frac{1}{\Lambda}\Big{(}d_{g}+C_{n}\,d_{\hat{m}}\Big{)}\phi\,$, (32)
-where $C_{n}=0.048$ Damour and Donoghue (2010), $m_{p}$ is proton mass, $m_{n}$ is the neutron mass, and the symmetric combination of the quark mass couplings has been defined to be:
-$d_{\hat{m}}\equiv\frac{d_{u}m_{u}+d_{d}m_{d}}{m_{u}+m_{d}}\,.$ (34)
-Moving back to the neutron star, conservation of angular momentum requires that the fluctuations in the spin frequency $\delta\omega$ relate to changes in the pulsar’s moment of inertia $\delta I$ as:
+where $C_{n}$ accounts for the fraction of the nucleon mass due to the light quarks, and
+$d_{\hat{m}}=\frac{m_{u}d_{u}+m_{d}d_{d}}{m_{u}+m_{d}}$ (33)
+is the effective quark coupling.
+In other words, the nucleon mass responds both to the direct change in quark masses induced by the dilaton and to the change in the gluon’s binding energy, allowing us to propagate the fundamental dilaton couplings to macroscopic observables such as the neutron-star moment of inertia.
+Eqs.
+31 and 32 show how dilaton DM will alter the masses of quarks and nucleons, which are the primary constitutions of neutron stars.
+Thus, these changed masses will impact the rotational frequency and moment of inertia of neutron stars, which will affect the times of arrivals (ToAs) of pulses that come from them.
+To calculate these effects, let us recall that conservation of angular momentum requires that the fluctuations in the spin frequency $\delta\omega$ relate to changes in the pulsar’s moment of inertia $\delta I$ as:
 $\frac{\delta\omega}{\omega_{0}}=-\frac{\delta I}{I_{0}}\,.$ (34)
 where $I_{0}$ is the unperturbed moment of inertia and $\omega_{0}$ is the unperturbed angular frequency of the neutron star.
-Based on the above discussion, DM interactions can cause changes in the masses of fundamental particles, which, in turn, cause time-dependent oscillations in the moment of inertia.
-In the case of neutron stars, the mass of the neutron will oscillate, and thus, indirectly, the mass of the neutron star itself will change as well.
 Assuming a simplistic model of a non-rotating, spherically symmetric neutron star composed only of neutrons with a polytropic equation-of-state, the radius and moment of inertia can be written as Shapiro and Teukolsky (1983):
@@
 $I_{0}$ $\propto M_{0}R^{2}=M^{1/3}m_{n}^{-16/3},$ (36)
-where $R$ and $M_{0}$ are the radius and unperturbed mass of the neutron star.
-Solving explicitly the Tolman–Oppenheimer–Volkoff (TOV) equations for a realistic equation-of-state would cause $\mathcal{O}(1)$ deviations from the simple scalings.
-These “unknown” scalings can be parameterized in terms of $\eta$ and $\delta\eta$
+where $R$ and $M_{0}$ are the radius and unperturbed mass of the neutron star, respectively.
+Plugging in Eqs.
+35 and 36 into Eq.
+34, we can derive how the moment of inertia changes as a function of the neutron-star mass and neutron mass:
 $\frac{\delta I}{I_{0}}=\eta\frac{\delta M}{M_{0}}+\delta\eta\frac{\delta m_{n}}{m_{n}}\,.$ (37)
-which are set to $\eta=1/3$ and $\delta\eta=-16/3$ in line with the simplistic model, and to illustrate the concept.
+where $\delta M$ is the change in the pulsar’s mass, and $\eta=1/3$ and $\delta\eta=-16/3$ in this simplistic model [^2] Solving explicitly the Tolman–Oppenheimer–Volkoff (TOV) equations for a realistic equation-of-state would cause $\mathcal{O}(1)$ deviations from the simple scalings.
+These “unknown” scalings can be parameterized in terms of $\eta$ and $\delta\eta$.
 Because of the changes in the pulsar’s moment of inertia, the pulsar’s spin frequency will change.
-For pulsar timing array experiments, the measured quantities are the times of arrivals of pulses from each pulsar, which will change in response to the oscillating moment of inertia.
-In practice, the observable here is the residual $h$ of the times of arrivals of pulses from the neutron star, which is given by:
-[…]
+For pulsar timing array experiments, the measured quantities are the ToAs of pulses from each pulsar, which will change in response to the oscillating moment of inertia.
+In practice, the observable here is the residual $h$ of the ToAs of pulses from the neutron star, which is given by:
+$h=-\int\frac{\delta\omega}{\omega_{0}}dt=\int\frac{\delta I}{I_{0}}dt\,.$ (38)
 Now, the changes in the particle mass must be related to the macroscopic change in the pulsar mass
-$\frac{\delta M}{M_{0}}$ $=\sum_{f\in\{e,\mu,p,n\}}Y_{f}\frac{m_{f}}{m_{n}}\frac{\delta m_{f}}{m_{f}}\,,$ (40)
-where $\delta M$ is the change in the pulsar’s mass.
-The sum is over the different types of particles present in the pulsar: electrons, muons, protons and neutrons, $Y_{f}\equiv N_{f}/(N_{n}+N_{p})$ is a weighting function that gives the number of $f$ particles relative to the sum of protons and neutrons (which dominate the pulsar’s mass).
-For neutron star observed by pulsar timing arrays, the following is assumed for $Y_{f}$: $Y_{n}=0.9$, $Y_{p}=Y_{e}=0.1$ and $Y_{\mu}=0.05$ Bell et al.
+$\frac{\delta M}{M_{0}}$ $=\sum_{f\in\{e,\mu,p,n\}}Y_{f}\frac{m_{f}}{m_{n}}\frac{\delta m_{f}}{m_{f}}\,.$ (39)
+The sum is over the different types of particles present in the pulsar: electrons, muons, protons and neutrons, $Y_{f}\equiv N_{f}/(N_{n}+N_{p})$ is a weighting function that gives the number of $f$ particles relative to the sum of protons $N_{p}$ and neutrons $N_{n}$ (which dominate the pulsar’s mass).
+For neutron stars observed by pulsar timing arrays, the following is assumed for $Y_{f}$: $Y_{n}=0.9$, $Y_{p}=Y_{e}=0.1$ and $Y_{\mu}=0.05$ Bell et al.
 (2019); Kaplan et al.
@@
 16 to Eq.
-40, and subsequently integrating over time in Eq.
-39, the changes in the timing residuals due to the ultralight dark-matter particles can be obtained:
-$h(t)=$ $\frac{\sqrt{2\rho_{\phi}}}{m_{\rm DM}^{2}\Lambda}\left(\vec{y}\cdot\vec{d}\,\right)\hat{\phi}_{P}\sin{\big{(}m_{\rm DM}t+\gamma(\vec{x}_{P})\big{)}}\,,$ (41)
+39, and subsequently integrating over time in Eq.
+38, the changes in the timing residuals due to the ultralight dark-matter particles can be obtained:
+$h(t)=$ $\frac{\sqrt{2\rho_{\text{DM}}}}{m_{\rm DM}^{2}\Lambda}\left(\vec{y}\cdot\vec{d}\,\right)\hat{\phi}_{P}\sin{\big{(}m_{\rm DM}t+\gamma(\vec{x}_{P})\big{)}}\,$, (40)
 where $\vec{x}_{P}$ is the location of the pulsar, $\hat{\phi}_{P}$ is the pulsar normalized signal amplitude, $\gamma$ is a phase, and $\vec{y}\cdot\vec{d}\equiv\sum_{i}y_{i}d_{i}$ where $i\in\{m_{e},m_{\mu},g,\hat{m}\}$ and:
-[…]
-This type of DM interaction would leave imprints on pulsar timing array experiments Porayko et al. (2018); Smarra et al. (2023); Afzal et al. (2023), which will be discussed later.
+$\begin{split}\{y_{g},\,y_{\hat{m}},\,y_{\mu},\,y_{e}\}=&\,\eta\left\{1,\,C_{n},\,6\times 10^{-3},\,5\times 10^{-5}\right\}\\ +&\delta\eta\left\{1,\,C_{n},\,0,\,0\right\}\,.\end{split}$ (41)
+Using the above formalism, dilatons have been constrained with recent pulsar timing array experiments Porayko et al.
+(2018); Smarra et al.
+(2023); Afzal et al.
+(2023), which will be extensively discussed in Section III.3.
 II.3.3 Changes to atomic clock times
 Changes in fundamental constants would also affect the way in which atomic clocks measure time.
-Since pulsar timing array experiments reference their measurements of the times of arrivals of pulses to terrestrial time (TT) using atomic clocks composed primarily of cesium, any oscillation of the fundamental constants will affect the atomic clocks, and thus the times measured Flambaum et al.
+Since pulsar timing array experiments reference their measurements of the ToAs of pulses to terrestrial time (TT) using atomic clocks composed primarily of cesium, any oscillation of the fundamental constants will affect the atomic clocks, and thus the times measured Flambaum et al.
 (2004).
 Likewise, a shift in the frequency of the atomic clock will give rise to an apparent shift in the measured pulsar spin frequency, given by Eq.
-39.
+38.
 From Flambaum et al. (2004); Kaplan et al. (2022), the frequency of an atomic clock is:
@@
 where $Z$ is the nuclear charge, $F_{\rm rel}(Z\alpha)$ is a relativistic correction to an atom’s energy levels, $\mu$ is the nuclear magnetic moment, and $\zeta=1$ if the clocks that have hyperfine transitions (the case of cesium), and $\zeta=0$ for optical transitions [^3] This equation arises from solving the non-relativistic Schrodinger equation for the valence electron wavefunction and evaluating it at the nucleus of the atom..
-The Casmir effect $F_{\rm rel}(Z\alpha)$ is a correction that comes from solving the relativistic wave equation for an electron near the nucleus
-The fluctuations in fundamental constants will cause fluctuations in the frequencies of atomic clocks composed of atoms $A$:
-$\begin{split}\frac{\delta f_{A}}{f_{A}}\simeq&\Bigg{[}\frac{\delta m_{e}}{m_{e}}+(4+K_{A})\frac{\delta\alpha}{\alpha}\\ &+\zeta\Bigg{(}\frac{\delta m_{e}}{m_{e}}+C_{A}\sum_{q=u,d}\frac{\delta m_{q}}{m_{q}}-\frac{\delta m_{p}}{m_{p}}\Bigg{)}\Bigg{]}\,,\end{split}$ (44)
-where the derivative of the Casmir effect term is $\delta F_{\rm rel}/F_{\rm rel}=K_{A}$, and the derivative of the nuclear magnetic moment is $\delta\mu/\mu=C_{A}\,\delta m_{q}/m_{q}$.
-For Cesium, $K_{A}=0.83$ depends nontrivially on $Z$ and $\alpha$, and $C_{A}=0.110$ Flambaum et al.
+The fluctuations in fundamental constants will cause modulations in the frequencies of atomic clocks composed of atoms $A$.
+Using previously derived expressions for variations in the electron mass, fine-structure constant, and quark and nucleon masses, the fractional frequency shift can be obtained Flambaum et al.
+(2004); Kaplan et al.
+(2022):
+[…]
+where $\delta F_{\rm rel}/F_{\rm rel}=K_{A}$, and the derivative of the nuclear magnetic moment is $\delta\mu/\mu=C_{A}\,\delta m_{q}/m_{q}$.
+For Cesium, $K_{A}=0.83$ depends on $Z$ and $\alpha$, and $C_{A}=0.110$ Flambaum et al.
 (2004).
-$m_{q}$ is a quark mass, and the index $q$ runs over up and down quarks.
-Finally, by substituting the aforementioned equations for fluctuations of the fundamental constants into Eq.
-44, and then further plugging this result into Eq.
-39, the timing residuals induced by the ultralight dark-matter are :
-$h(t)=\frac{\sqrt{2\rho_{\text{DM}}}}{m_{\rm DM}^{2}\Lambda}\left(\vec{y}\cdot\vec{d}\right)\hat{\phi}_{E}\sin{\big{(}m_{\rm DM}t+\gamma\big{)}}\,.$ (45)
+The index $q$ runs over up and down quarks.
+Finally, by substituting the aforementioned equations for fluctuations of the fundamental constants (Eqs.
+21, 22, 31 and 32) into Eq.
+43, and then further plugging this result into Eq.
+38, the timing residuals induced by the ultralight dark-matter can be obtained:
+$h(t)=\frac{\sqrt{2\rho_{\text{DM}}}}{m_{\rm DM}^{2}\Lambda}\left(\vec{y}\cdot\vec{d}\right)\hat{\phi}_{E}\sin{\big{(}m_{\rm DM}t+\gamma(\vec{x}_{E})\big{)}}\,.$ (44)
 where $\vec{x}_{E}$ is the position of the Earth, and the sensitivity parameters of this search are given by:
@@
 (2023); Afzal et al.
-(2023), which will be discussed later.
+(2023), which will be discussed in Section III.3.
+We would like to draw a parallel between Eqs.
+40 and 45.
+Both equations have a similar form, though the underlying physics is different, as indicated by the phase factors: the former depends on the phase at the neutron star, while the latter depends on the phase on earth (at the atomic clocks).
+Of course, the sensitivity parameters of the search are different, which encode which couplings that each effect depends on.
+These two effects highlight an important point that will be made throughout this review: while the physical signatures of different kinds of DM are different, often the impact on the signal (the residuals in pulsar timing arrays, the phase shifts in ground- and spaced-based interferometers) are similar, and only the interpretation of the lack of a detection is different.
 II.4 Vector dark matter
-DM could be spin-1, i.e.
+So far, we have considered explain how spin-0 particles would affect GW interferometers; however, DM can also have nonzero spins.
+Here, we consider a model for DM that has spin-1, i.e.
 vectors, which we refer to as “dark photons”.
-Dark photons could explain the entirety of the relic abundance of DM, which could arise from the misalignment mechanism Nelson and Scholtz (2011); Arias et al.
+Like dilatons, dark photons could explain the entirety of the relic abundance of DM, which could arise from the misalignment mechanism Nelson and Scholtz (2011); Arias et al.
 (2012); Graham et al.
@@
 (2019), or from cosmic string network decays Long and Wang (2019).
-They could couple to standard model particles, either to baryon or baryon-lepton number.
-In the interferometers, these interactions would occur everywhere, but would be detectable in the four interferometer mirrors.
+They could couple to standard-model particles, either to baryon or baryon-lepton number.
+In the interferometers, these interactions would occur everywhere, but would be most pronounced in the four interferometer test masses in the main resonant cavities of the instrument.
 The dark photons [^4] These are different dark photons than those that kinetically mix with the ordinary photon, which will be discussed later in the context of luminous superradiance Siemonsen et al.
@@
 (2018); Miller et al.
-(2021).
-Dark photons can be described analogously to ordinary photons, i.e. with a four-vector potential, which can be written as:
-[…]
-where $(A_{0})_{\mu}$ is the four-amplitude of $A_{\mu}$ $\Upsilon$ is a random phase, and $\vec{x}$ is the position at which $A_{\mu}$ is measured.
+(2021a), as will be shown below.
+To derive the impact of dark photons on GW interferometers, we will begin by describing dark photons analogously to ordinary photons, i.e. with a four-vector potential:
+$A_{\mu}(t,\vec{x})=(A_{0})_{\mu}\sin(m_{\rm DM}t-\vec{k}\cdot\vec{x}+\Upsilon),$ (46)
+where $(A_{0})_{\mu}$ is the four-amplitude of $A_{\mu}$ $\Upsilon$ is a random phase, $\vec{k}$ is the wavevector, and $\vec{x}$ is the position at which $A_{\mu}$ is measured.
 The index $\mu={0,1,2,3}$ refers to both time and spatial components.
@@
 After making this choice, we can compare the time component to magnitude of the spatial components of the four-potential:
-$\frac{(A_{0})_{0}}{|\vec{A}_{0}|}=\frac{v_{0}}{c}\simeq 7.667\times 10^{-4},$ (48)
-where $|\vec{A}_{0}|=\left(\frac{\hbar\sqrt{2\rho_{\text{DM}}}}{m_{\rm DM}c^{2}}\frac{1}{\sqrt{\epsilon_{0}}}\right)$ is the magnitude of the spatial components of $A_{\mu}$, normalized by the present DM energy density of the universe, and $\epsilon_{0}$ is the permittivity of vacuum.
+$\frac{(A_{0})_{0}}{|\vec{A}_{0}|}=v_{0}\simeq 7.667\times 10^{-4},$ (47)
+where $|\vec{A}_{0}|=\left(\frac{\sqrt{2\rho_{\text{DM}}}}{m_{\rm DM}}\right)$ is the magnitude of the spatial components of $A_{\mu}$, normalized by the present DM energy density of the universe.
 In Eq.
-48, we see that the dark scalar potential is suppressed by about three orders of magnitude compared to the the dark three-vector potential.
+47, we see that the dark scalar potential is suppressed by about three orders of magnitude compared to the the dark three-vector potential.
 Therefore, the time-component of the four-vector potential can be safely neglected, leaving only the three-vector potential:
-$\vec{A}=\vec{A}_{0}\sin\left(\omega t-\vec{k}\cdot\vec{x}+\Upsilon\right),$ (49)
-The Lagrangian $\mathcal{L}$ that characterizes the dark photon coupling to a number current density $J^{\mu}$ of baryons or baryons minus leptons is:
-$\mathcal{L}=-\frac{1}{4\mu_{0}}F^{\mu\nu}F_{\mu\nu}+\frac{1}{2\mu_{0}}\left(\frac{m_{\rm DM}c}{\hbar}\right)^{2}A^{\mu}A_{\mu}-\epsilon eJ^{\mu}A_{\mu},\$ (50)
-where $\mu_{0}$ is the magnetic permeability in vacuum and $\epsilon$ is the strength of the particle/dark photon coupling normalized by the electromagnetic coupling constant.
-Using Eq. 47, the “dark” electric and magnetic fields can be derived:
-$\vec{E}$ $=\partial_{0}\vec{A}-\vec{\nabla}A_{0}\simeq\omega\vec{A}_{0}\cos(\omega t-\vec{k}\cdot\vec{x}+\phi),$ (51)
-$\vec{B}$ =×$=\vec{\nabla}\times\vec{A}=-\vec{k}\times\vec{A}_{0}\cos(\omega t-\vec{k}\cdot\vec{x}+\phi),$ (52)
-and we note that the electric field is much stronger than the magnetic one:
-$\frac{|\vec{E}|}{|c\vec{B}|}\sim\frac{\omega}{|\vec{k}|}=\frac{c^{2}}{|\vec{v}|}\sim 10^{3}c,$ (53)
+$\vec{A}=\vec{A}_{0}\sin\left(m_{\rm DM}t-\vec{k}\cdot\vec{x}+\Upsilon\right),$ (48)
+With the four-potential, we can write the Lagrangian $\mathcal{L}$ that characterizes the dark photon coupling to a number current density $J^{\mu}$ of baryons or baryons minus leptons:
+$\mathcal{L}=-\frac{1}{4}F^{\mu\nu}F_{\mu\nu}+\frac{1}{2}m_{\rm DM}^{2}A^{\mu}A_{\mu}-\epsilon eJ^{\mu}A_{\mu},\\$ (49)
+where $\epsilon$ is the strength of the particle/dark photon coupling normalized by the electromagnetic coupling constant.
+Using Eqs. 46 and 48, we can compute the “dark” electric and magnetic fields, analogously to electromagnetism, which will subsequently allow us to compute the “dark” force:
+$\vec{E}$ $=\partial_{0}\vec{A}-\vec{\nabla}A_{0}\simeq m_{\rm DM}\vec{A}_{0}\cos(m_{\rm DM}t-\vec{k}\cdot\vec{x}+\Upsilon),$ (50)
+$\vec{B}$ $=\vec{\nabla}\times\vec{A}=-\vec{k}\times\vec{A}_{0}\cos(m_{\rm DM}t-\vec{k}\cdot\vec{x}+\Upsilon),$ (51)
+It is important to understand the relative magnitudes of the electric and magnetic fields. Comparing them, we see that the electric field is much stronger than the magnetic one:
+$\frac{|\vec{E}|}{|\vec{B}|}\sim\frac{m_{\rm DM}}{|\vec{k}|}=\frac{1}{|\vec{v}|}\sim 10^{3},$ (52)
 which implies that we need to only consider the electric field.
 This dark electric field causes the test masses to oscillate at a quasi-fixed frequency given by the dark photon mass.
-The acceleration of a given mirror can thus be derived Guo et al.
+The acceleration of a given mirror can thus be derived using Eqs.
+49 and 48 Guo et al.
 (2019); Pierce et al.
 (2018):
-$\vec{a}_{j}(t,\vec{x}_{j})$ $=\frac{\vec{F}_{j}(t,\vec{x}_{j})}{M_{j}}\simeq\epsilon e\frac{Q_{D,j}}{M_{j}}\omega|\vec{A}_{0}|\hat{A}\cos(\omega t-\vec{k}\cdot\vec{x}_{j}+\phi),$
-$\epsilon^{2}$ $=\frac{\alpha_{\text{DP}}}{\alpha}$ (54)
-where $\alpha_{\text{DP}}$ is the dark photon fine structure constant, $\epsilon^{2}$ is the ratio between the dark photon fine structure constant and the electromagnetic fine structure constant $\alpha$, and $Q_{D,j}$ is the total charge in the $j$th mirror of mass $M_{j}$.
-$\hat{A}$ is a unit vector, and $\phi$ is a function of the spatial coordinates.
+$\vec{a}_{j}(t,\vec{x}_{j})$ $=\simeq\epsilon e\frac{Q_{D,j}}{M_{j}}m_{\rm DM}|\vec{A}_{0}|\hat{A}\cos(m_{\rm DM}t-\vec{k}\cdot\vec{x}_{j}+\Upsilon),$
+where $Q_{D,j}$ is the total charge in the $j$th mirror of mass $M_{j}$, and $\hat{A}$ is a unit vector of the vector potential.
 If dark photons couple to the baryon number, $q_{j}$ is the number of protons and neutrons in each mirror; if they couple to the difference between the baryon and lepton numbers, $q_{j}$ is the number of neutrons in each mirror.
@@
 Thus, the interferometer experiences a measurable differential acceleration, and thus a differential strain.
-Assuming that all mirrors are the same, integrating Eq.
-54 twice over time, and averaging over random polarization and propagation directions, the strain on the interferometers caused by a dark photon DM signal is Pierce et al.
+Assuming that all mirrors have the same charge-to-mass ratio, integrating LABEL:eqn:accel twice over time, and averaging over random polarization and propagation directions, the strain on the interferometers caused by a dark photon DM signal can be computed Pierce et al.
 (2018):
-$\sqrt{\langle h^{2}_{D}\rangle}$ $=C\frac{Q_{D}}{M}\frac{\hbar e}{c^{4}\sqrt{\epsilon_{0}}}\sqrt{2\rho_{\text{DM}}}v_{0}\frac{\epsilon}{f_{0}},$
+$\sqrt{\langle h^{2}_{D}\rangle}$ $=C\frac{Q}{M}e\sqrt{2\rho_{\text{DM}}}v_{0}\frac{\epsilon}{f_{0}},$
 $\simeq 6.56\times 10^{-26}\left(\frac{\epsilon}{10^{-22}}\right)\left(\frac{100\text{ Hz}}{f_{0}}\right),$ (54)
-where $C=\sqrt{2}/3$ (in LIGO) is a geometrical factor obtained by averaging over all possible dark photon propagation and polarization directions, and different detector geometricians (the calculation for $C$ is shown in the appendix of Pierce et al.
+where $C=\sqrt{2}/3$ (for an interferometer with two perpendicular arms) is a geometrical factor obtained by averaging over all possible dark photon propagation and polarization directions, and different detector geometries (the calculation for $C$ is shown in the appendix of Pierce et al.
 (2018) and in Morisaki et al.
@@
 This strain can actually be larger than that given in Eq.
-55:
+54 and can be derived Morisaki et al.
+(2021):
 $\sqrt{\langle h_{C}^{2}\rangle}$ $=\frac{\sqrt{3}}{2}\sqrt{\langle h_{D}^{2}\rangle}\frac{2\pi f_{0}L}{v_{0}},$
 $\simeq 6.58\times 10^{-26}\left(\frac{\epsilon}{10^{-23}}\right).$ (55)
-The total strain is: $\langle h_{{\rm total}}^{2}\rangle=\langle h_{D}^{2}\rangle+\langle h_{C}^{2}\rangle$.
+The total power is: $\langle h_{{\rm total}}^{2}\rangle=\langle h_{D}^{2}\rangle+\langle h_{C}^{2}\rangle$.
+Both of these effects can be constrained by analyzing data from LIGO, Virgo and KAGRA and pulsar timing arrays, which will be discussed in Sections III.2 and III.3, though in different mass regimes.
 II.5 Tensor dark matter
@@
 Specifically, bimetric gravity Hassan and Rosen (2012); Schmidt-May and von Strauss (2016), a theory in which a massless and massive spin-2 field interact, provides a plausible candidate for ultralight dark matter Marzola et al.
-(2018), produced by the misalignment mechanism Preskill et al.
 (2018).
-Here, we consider a massive spin-2 field $M_{\mu\nu}$ described by the Fierz-Pauli Lagrangian density Armaleo et al. (2020):
+Here, we consider a massive spin-2 field $M_{\mu\nu}$ described by the Fierz-Pauli Lagrangian density Fierz and Pauli (1939); Armaleo et al. (2020):
 ${\mathcal{L}}$ $\,:=\frac{1}{2}M_{\mu\nu}\mathcal{E}^{\mu\nu\rho\sigma}M_{\rho\sigma}-\frac{1}{4}m_{\rm DM}^{2}\left(M_{\mu\nu}M^{\mu\nu}-M^{2}\right)\,$, (56)
@@
 (2018).
-The field spatial component can be written in a similar way as for dark photons and scalar bosons at given position $\mathbf{x}$:
-$M_{ij}$ $=\frac{\sqrt{2\rho_{\text{DM}}(\vec{x})}}{m}\cos{\left(\frac{mc^{2}}{\hbar}t+\vec{k}\cdot\vec{x}+\Upsilon(\vec{x})\right)}\varepsilon_{ij}(\vec{x})\,,$ (58)
-where $\Upsilon$ is a random phase, and $\varepsilon_{ij}$ encodes information about the five polarizations of the massive spin-2 field.
-The strain on the GW detectors arises analogously to that from GWs: a stretching of space-time in the presence of the field, since the massive spin-2 metric and its coupling constant can be absorbed into the definition of the massless spin-2 metric in the linear regime of the coupling constant $\alpha_{\rm TB}$ [^4] In the literature, $\alpha_{\rm TB}=\alpha$ simply, but to disentangle it from the fine structure constants, we rename it.
+The field spatial component can be written in a similar way as for dark photons and scalar bosons at given position:
+$M_{ij}$ $=\frac{\sqrt{2\rho_{\text{DM}}(\vec{x})}}{m_{\rm DM}}\cos{\left(m_{\rm DM}t+\vec{k}\cdot\vec{x}+\Upsilon(\vec{x})\right)}\varepsilon_{ij}(\vec{x})\,$, (57)
+where and $\varepsilon_{ij}$ encodes the polarizations of the massive spin-2 field.
+The strain on GW detectors arises analogously to that from GWs: a stretching of space-time in the presence of the field, since the massive spin-2 metric and its coupling constant can be absorbed into the definition of the massless spin-2 metric in the linear regime of the coupling constant $\alpha_{\rm TB}$ [^5] In the literature, $\alpha_{\rm TB}=\alpha$ simply, but to disentangle it from the fine structure constants, we rename it.
 TB is tensor boson.
@@
 (2020); thus, the perturbation $h_{ij}$ is:
-$h_{ij}(t)$ $=\frac{\alpha_{\rm TB}}{M_{\text{Pl}}}M_{ij}(t)=\frac{\alpha_{\rm TB}\sqrt{2\rho_{\text{DM}}}}{mM_{\text{Pl}}}\cos{\left(\frac{DMc^{2}}{\hbar}t+\Upsilon\right)}\varepsilon_{ij}(\mathbf{x})\,,$ (59)
-where $\alpha$ is a dimensionless constant that quantifies the difference between the strengths of each of the spin-2 fields, and $M_{\text{Pl}}$ is the Planck mass.
-Contracting $h_{ij}$ with the GW detector response function $D^{ij}$ in the tensor case, the resulting strain is:
-$h(t)=D^{ij}h_{ij}=\frac{\alpha_{\rm TB}\sqrt{\rho_{\text{DM}}}}{\sqrt{2}mM_{\text{Pl}}}\cos{\left(\frac{m_{\rm DM}c^{2}}{\hbar}t+\Upsilon\right)}\Delta\varepsilon,$ (60)
-where $\Delta\varepsilon:=\,\varepsilon_{ij}(n^{i}n^{j}-m^{i}m^{j})$.
-II.6 Gravitational interactions of dark matter
+$h_{ij}(t)$ $=\frac{\alpha_{\rm TB}}{M_{\text{Pl}}}M_{ij}(t)=\frac{\alpha_{\rm TB}\sqrt{2\rho_{\text{DM}}}}{m_{\rm DM}M_{\text{Pl}}}\cos{\left(m_{\rm DM}t+\Upsilon\right)}\varepsilon_{ij}(\vec{x})\,$, (58)
+where $\alpha_{\rm TB}$ is a dimensionless constant that quantifies the difference between the strengths of each of the spin-2 fields, and $M_{\text{Pl}}$ is the Planck mass.
+To obtain the strain on GW interferometers, we note that the detector response function can be parameterized in terms of a tensor, often written as $D^{ij}$ Romano and Cornish (2017).
+By contracting $h_{ij}$ with $D^{ij}$, the resulting strain can be derived Armaleo et al.
+(2021).
+$h(t)=D^{ij}h_{ij}=\frac{\alpha_{\rm TB}\sqrt{\rho_{\text{DM}}}}{\sqrt{2}mM_{\text{Pl}}}\cos{\left(m_{\rm DM}t+\Upsilon\right)}\Delta\varepsilon,$ (59)
+where $\Delta\varepsilon:=\,\varepsilon_{ij}(n^{i}n^{j}-m^{i}m^{j})$ and $n,m$ are unit vectors pointing along directions $i,j$.
+We note that an interesting discrepancy arises when comparing the results from Refs.
+Aoki and Mukohyama (2016); Manita et al.
+(2024) with those from Refs.
+Marzola et al.
+(2018); Armaleo et al.
+(2021) regarding the strain induced by tensor dark matter.
+The two sets of works differ cause a factor of 2 difference in the strain, which can be traced back to a factor of eight difference in their respective Lagrangians.
+This difference stems from variations in the normalization factors and scaling of the tensor field amplitude, as well as differences in how the tensor field’s energy density is treated.
+Such discrepancies are not uncommon in the study of dark matter and may have implications for the interpretation of experimental constraints and the modeling of astrophysical signals.
+However, we note that no constraints on tensor DM have been placed as of yet using GW interferometers.
+II.6 Gravitationally interacting dark matter
 The previous sections assumed a coupling of DM to standard-model particles.
@@
 This is, in fact, the simplest assumption that can be made, since we only have evidence for the existence of DM through gravitational interactions.
-One such hypothetical scalar DM particle has a Lagrangian of:
+Below, we consider two kinds of scalar DM that couple only to gravity.
+II.6.1 Minimally coupled dark matter
+In minimal coupling models, DM interacts with the standard model exclusively through gravity by contributing to the stress-energy tensor.
+The gravitational field responds to this energy density and pressure according to Einstein’s field equations, where the presence of dark matter modifies the curvature of spacetime in the same manner as any other form of energy or matter.
+Here, ultralight dark matter can source time-varying gravitational potentials, leading to observable effects such as oscillations in pulsar timing residuals, which will be discussed in Section III.3.
+A commonly studied example of a minimally coupled scalar DM field $\phi$ is described by the Lagrangian Khmelnitsky and Rubakov (2014):
 $\mathcal{L}=\sqrt{-g}\left[\frac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{2}m_{\rm DM}^{2}\phi^{2}\right]\,.$ (60)
-and Khmelnitsky and Rubakov (2014):
+where $g^{\mu\nu}$ is the metric tensor, and $g$ is the determinant of the metric tensor. In this model, the scalar field evolves:
 $\phi(\vec{x},t)=\frac{\sqrt{2\rho_{\text{DM}}}}{m_{\rm DM}}\hat{\phi}(\vec{x})\cos\left(m_{\rm DM}t+\gamma(\vec{x})\right),$ (61)
-where $\gamma(\vec{x})$ is a phase that depends on position $\vec{x}$, and $\hat{\phi}(\vec{x})$ describes the interference pattern caused by the ultralight dark matter.
+where $\gamma(\vec{x})$ is a phase that depends on position $\vec{x}$, and $\hat{\phi}(\vec{x})$ describes the variation in space of the ultralight dark-matter field.
 As before, the scalar field density is normalized by $\rho_{\text{DM}}$.
-Scalar DM could cause the arrival times of pulses from millisecond pulsars on earth to change, in an analogous way that GWs do Khmelnitsky and Rubakov (2014).
+Scalar DM can cause the arrival times of pulses from millisecond pulsars on earth to change, in an analogous way that GWs do Khmelnitsky and Rubakov (2014), i.e.
+introduce stochastic fluctuations of the metric Kim and Mitridate (2024); Kim (2023).
 Such a difference in the arrival times can be written as Khmelnitsky and Rubakov (2014); Porayko et al.
@@
 $\Psi_{\text{c}}(\vec{x})\approx 6.52\times 10^{-18}\left(\frac{10^{-22~}\text{eV}}{m_{\rm DM}}\right)^{2}\left(\frac{\rho_{\phi}}{0.4~\text{GeV}/\text{cm}^{3}}\right),$ (63)
-where $\gamma_{\text{P}}\equiv 2\gamma(\vec{x}_{\text{p}})-2m_{\rm DM}d_{p}/c$ ($\gamma_{\text{E}}\equiv 2\gamma(\vec{x}_{\text{e}})$) parameterize random phases evaluated at the pulsar (P) or Earth (E), and $d_{p}$ is the distance between the pulsar and the Earth.
-The DM energy density background is assumed to be constant when calculating $\Psi_{\text{c}}(\vec{x})$, though the possibility of deviations, due to the coherently oscilalting ultralight dark-matter field, are parametrized in terms of two phase factors: one for the pulsar $\hat{\phi}^{2}(\vec{x}_{\text{p}})\equiv\hat{\phi}^{2}_{\text{P}}$ and one for the Earth ($\hat{\phi}^{2}(\vec{x}_{\text{e}})\equiv\hat{\phi}^{2}_{\text{E}}$).
-Since all the pulsars are separated by at least a kpc, assuming a constant DM energy density is valid Porayko et al.
-(2018).
-III Observational constraints on dark matter directly interacting with gravitational-wave interferometers
-Section II focused on the currently-known types of DM that could cause a signal in GW interferometer data.
-While we have indicated the strain amplitudes of such signals for some of the DM models considered, we have not yet discussed whether the interferometers are actually sensitive to such weak signals.
-In this section, we will describe how searches for DM interactions with GW detectors are performed, show current constraints on various models, and explain what future constraints will be on others.
-Each method discussed below can achieve a certain sensitivity, that is, a minimum detectable strain amplitude on the detector at a given confidence level.
-This strain amplitude, the left-hand side of many equations in Section II, is then mapped to constraints on the physical coupling constant of DM to the standard model.
-The methods themselves are agnostic towards the particular model of DM that is being searched for.
+[…]
+II.6.2 Non-minimally coupled (conformal) dark matter
+Another form of DM that couples only to gravity would be conformal DM.
+This kind of DM typically involve scalar fields that couple to the trace of the energy-momentum tensor.
+These models are closely related to the dilaton-type scalar DM discussed in Section II.3, but with specific couplings motivated by scalar–tensor theories of gravity, such as Fierz–Jordan–Brans–Dicke (FJBD) and Damour–Esposito–Farese (DEF) theories Fierz and Pauli (1939); Fierz (1956); Dicke (1962); Jordan (1959); Damour and Esposito-Farese (1992, 1993).
+In such models, DM couples universally to the gravitational sector through a conformal rescaling of the metric.
+This coupling can induce a gravity-mediated force between neutron stars and the ultralight dark matter field, altering the moment of inertia of rotating pulsars – a manifestation of the Nordtvedt effect Nordtvedt (1968).
+These variations lead to deterministic changes in pulsar ToA and constitute a violation of the equivalence principle.
+This kind of conformal scalar DM has been extensively studied and constrained using binary pulsar timing measurements Damour and Taylor (1992); Kramer et al.
+(2006), which will be discussed inSection III.3.
+In the case of a linear FJBD coupling, the pulsar timing residuals take the form:
+$\Delta t(t)$ $=2\alpha s_{I}\left.\frac{\sqrt{\rho_{\text{DM}}}}{M_{\text{Pl}}m_{\rm DM}^{2}}\hat{\phi}(\vec{x})\sin(m_{\rm DM}t+\theta(\vec{x}))\right|_{t_{\text{start }}-d}^{t_{\text{end }}-d},$ (64)
+where $s_{I}$ is the angular moment sensitivity parameter computed in Kuntz and Barausse (2024) and is of $\mathcal{O}(1)$ in FJBD theory, $\alpha$ is the strength of the coupling, $\theta(\vec{x})$ is a random phase, and $\hat{\phi}$ is a stochastic parameter whose value depends on whether the correlated, uncorrelated or pulsar correlated scenarios are considered (see Section III.3 for a description of these terms).
+In summary, both dilaton-type and conformal scalar DM predict couplings that can modify either the properties of matter or the effective gravitational interaction, leading to observable signatures.
+In ground-based interferometers, these effects manifest primarily as strain signals induced by oscillations in material properties or cavity frequencies, while in pulsar timing arrays they appear as deterministic modulations in pulse arrival times due to variations in neutron-star moment of inertias or equivalence-principle–violating forces.
+These complementary channels allow terrestrial and astrophysical detectors to probe overlapping but distinct regions of parameter space.
+In the next section, we review the current observational constraints on such models, focusing on limits derived from GW interferometer data.
+III Observational constraints on DM directly interacting with GW interferometers
+Section II outlined several theoretically motivated models of DM that could leave detectable imprints in GW interferometer data, and we showed how these models map on to induced strains on the detectors.
+The natural next question is whether current and future interferometers are actually sensitive to such weak signals.
+In this section, we turn to the search strategies themselves: the methods used to look for DM interactions with gravitational-wave detectors, the observational constraints that have already been obtained, and the prospects for upcoming instruments.
+Each method introduced in Section III.1 can be characterized by a sensitivity, i.e.
+the minimum detectable strain amplitude at a given confidence level.
+This quantity is what ultimately gets translated into limits on the physical couplings between DM and the standard model, and does not typically depend on the DM model considered.
+In other words, these search techniques are model-agnostic: they look for correlated, excess power at each frequency the dataset, independent of the underlying nature of DM.
+The interpretation comes later, when the observed sensitivities are framed as constraints on particular coupling constants and interaction scenarios.
+After describing the methods, we present current observational constraints set by ground- and space-based instruments in Section III.2, and those set by pulsar timing arrays in Section III.3.
+We conclude in Section III.4 by discussing how future GW interferometers could set tighter constraints on, or potentially detect, a wider variety of DM models, as outlined in Section II.
 III.1 Methods
@@
 However, the stochastic frequency variations prevent us from doing that.
-If we observe for a duration $T_{\text{obs}}$ longer than $T_{\text{coh}}$, the signal will not be sinusoidal and thus its power will be spread among different frequency bins, which would inhibit a possible detection.
-Thus, the following methods have been designed to combine signal power across chunks of data $T_{\text{FFT}}\ll T_{\text{obs}}$ in a way that optimizes sensitivity towards particle DM interactions with the standard model and gravity.
+If we observe for a duration $T_{\text{obs}}$ longer than $T_{\text{coh}}$, the signal will not be sinusoidal and thus its power will be spread among different frequency bins, as shown in Fig.
+3, which would inhibit a possible detection.
+Thus, the following methods have been designed to combine signal power across chunks of data of length $T_{\text{FFT}}\ll T_{\text{obs}}$ in such a way that avoid power loss and thus optimizes sensitivity towards particle DM interactions with the standard model and gravity.
+Essentially, these methods vary the length of the chunk of data $T_{\text{FFT}}$ to match the coherence time of DM $T_{\text{coh}}$.
+Each of the methods described below formulates its sensitivity, detection statistic, and search strategy differently.
+As such, applying multiple methods to search for the same type of DM is worthwhile, since each independently defines how to assess the significance of candidates, compute upper limits, set detection thresholds, and veto outliers that historically have arisen from instrumental or environmental noise.
+Using diverse approaches provides a valuable cross-check and increases the robustness of any potential detection or constraint.
 III.1.1 Cross-correlation
-Cross-correlation Pierce et al.
+Conceptually, cross-correlation Pierce et al.
 (2018); Guo et al.
 (2019) requires at least two separate time-domain datasets that are Fourier transformed and multiplied together to compute the cross-power in each frequency bin.
-The cross power is then divided by the auto-power of each detector, and then summed over all the Fourier transforms to arrive at a measure of power at each frequency.
-It is then divided by the standard deviation of the noise to compute the signal-to-noise ratio.
-The cross-correlated signal strength for detector pair $IJ$ is:
+The cross power is then divided by the auto-power (the power spectral density) of each detector, and then summed over all the Fourier transforms to arrive at a measure of power at each frequency.
+It is then divided by the standard deviation of the noise to compute the signal-to-noise ratio (SNR).
+Mathematically, the cross-correlated signal strength for detector pair $IJ$ is Pierce et al. (2018); Guo et al. (2019):
 $S_{IJ,j}=\frac{1}{N_{\rm FFT}}\sum_{i=1}^{N_{\rm FFT}}\frac{z_{I,ij}z_{J,ij}^{*}}{P_{I,ij}P_{J,ij}},$ (65)
 in the $j^{\rm th}$ frequency bin at the $i^{\rm th}$ time.
-“*” denotes the complex conjugate, and $z_{1,ij}$ and $z_{2,ij}$ is the Fourier transform of the time-domain data from detectors 1 and 2, respectively, and $N_{\text{FFT}}$ is the total number of FFTs taken over the observing run.
-If the observing run lasts $T_{\text{obs}}$ and the FFT length is $T_{\text{FFT}}$, then $N_{\text{FFT}}=[T_{\text{obs}}/T_{\text{FFT}}]$.
-$P_{1,ij}$ and $P_{2,ij}$ are the individual detector power spectral densities, which are typically estimated using a running median of each FFT.
-Typically $T_{\text{FFT}}=1800$ s, independently of the DM mass, which ensures a relatively easy analysis but implies the sensitivity towards DM interactions is not optimal at each mass analyzed.
-To obtain a detection statistic, i.e. the signal-to-noise ratio, we have to divide the signal strength by the variance of the data, which is:
+“*” denotes the complex conjugate, and $z_{I,ij}$ and $z_{J,ij}$ is the Fourier transform of the time-domain data from detectors II and $J$, respectively, and $N_{\text{FFT}}$ is the total number of FFTs taken over the observing run.
+If the observing run lasts $T_{\text{obs}}$ and the FFT length is $T_{\text{FFT}}$, then $N_{\text{FFT}}=T_{\text{obs}}/T_{\text{FFT}}$.
+$P_{I,ij}$ and $P_{J,ij}$ are the individual detector power spectral densities, which are typically estimated using a running median over each FFT.
+Typically, $T_{\text{FFT}}=1800$ s, independently of the DM mass, which ensures a relatively straightforward analysis but implies the sensitivity towards DM interactions is not optimal at each mass analyzed.
+In addition to the signal strength, we have to obtain an estimate of the variance of the noise data. This can be computed from the individual power spectral densities in each FFT as Guo et al. (2019):
 $\sigma_{IJ,j}^{2}=\frac{1}{N_{\rm FFT}}\left\langle\frac{1}{2P_{I,ij}P_{J,ij}}\right\rangle_{N_{\rm FFT}},$ (66)
@@
 We note that the average is over inverse noise-weighted power spectral densities, (analogous to adding resistors in parallel), which helps to suppress spurious power due to large noise disturbances.
-The signal-to-noise ratio (SNR), the dection statistic, is then simply:
+The SNR, the detection statistic in each frequency bin, is then simply:
 ${\rm SNR}_{IJ,j}=\frac{S_{IJ,j}}{\sigma_{IJ,j}}.$ (67)
 In the presence of pure Gaussian noise, the SNR will follow a normal distribution, with a mean of 0 and a standard deviation of 1.
-If the SNR exceeds a certain threshold, which is set both theoretically and by the trials factor, then a particular frequency is classified as being “significant”.
-Depending on the type of DM searched for, the signal-to-noise ratio for spin-0 (labeled SD), spin-1 (DP) and spin-2 (DG) cross-correlation searches can be written as Manita et al. (2024):
-${\rm SNR}^{\rm SD}_{IJ,j}$ $\approx\left|\frac{2\sqrt{2}\alpha_{A}^{2}f_{\rm SD}T_{\rm eff}({\cal B}^{\prime}_{IJ}\Gamma_{IJ}+{\cal C}^{\prime}\gamma_{IJ})}{\sqrt{S_{n,I}(\frac{m_{\rm DM}}{2\pi})S_{n,J}(\frac{m_{\rm DM}}{2\pi})}}\right|\,.$ (68)
-${\rm SNR}^{\rm DP}_{IJ,j}$ $\approx\left|\frac{2\sqrt{2}\epsilon_{D}^{2}f_{\rm DP}T_{\rm eff}({\cal B}_{IJ}\Gamma_{IJ}+{\cal C}(1+\frac{1}{3}\Omega^{z}_{\rm DP})\gamma_{IJ})}{\sqrt{S_{n,I}(\frac{m_{\rm DM}}{2\pi})S_{n,J}(\frac{m_{\rm DM}}{2\pi})}}\right|\,,$ (69)
-${\rm SNR}^{\rm DG}_{IJ,j}$ $\approx\left|\frac{2\sqrt{2}\alpha^{2}f_{\rm DG}T_{\rm eff}{\cal A}\gamma_{IJ}}{\sqrt{S_{n,I}(\frac{m_{\rm DM}}{2\pi})S_{n,J}(\frac{m_{\rm DM}}{2\pi})}}\right|\,,$ (70)
-For spin-0: $f_{\rm SD}\equiv\rho_{\rm SD}/\rho_{\text{DM}}$ is the ratio of the energy density of the spin-0 DM to the total energy density of DM.
-$\alpha_{A}$ is the dimensionless coupling constant that depends on the coupling constants of the scalar DM particles to standard model ones, e.g.
-the electron and photon, and are given by Eq.
-44-47 in Manita et al.
+If the SNR exceeds a certain threshold, which is set both theoretically assuming Gaussian noise and by the trials factor (accounting for the size of the parameter space), then a particular frequency is classified as being “significant”.
+Depending on the type of ultralight dark matter being searched for, the detectability of a signal in cross-correlation searches between interferometers depends on the spin of the DM: spin-0 (scalar), spin-1 (vector), or spin-2 (tensor) Manita et al.
 (2024).
-$\gamma_{IJ}$ is the overlap-reduction function that quantities the sensitivity loss by cross correlating detectors that are not co-located and co-aligned, and is given in Tab.
-I in Manita et al.
-(2024) for different pairs of detectors.
-It specifically arises from the spatial displacement of mirrors by DM.
-$\Gamma_{IJ}$ is the overlap-reduction function that arises from the finite-light time travel effects of the signal on the mirrors Morisaki et al.
+The SNR for a given detector pair is determined by the DM energy density fraction in that spin state, the strength of its couplings to standard-model particles, the observation time, and the relative orientation, location and geometry of the detectors.
+The latter is characterized by an “overlap reduction function” (ORF), which determines the sensitivity loss when applying cross-correlation to non-collocated, non-aligned detector pairs.
+For these models of DM, two kinds of ORFs arise: one from the spatial displacement of mirrors, and another from the finite-light time travel effects of the signal on the mirrors, a previously explained in Section II.4 Morisaki et al.
 (2021).
-For spin-1: $f_{\rm DP}\equiv\rho_{\rm DP}/\rho_{\text{DM}}$ is the ratio of the energy density of the spin-1 DM to the total energy density of DM.
-$\Omega^{z}_{\rm DP}$ is a term that depends on the polarization of DM, and $\Omega^{z}_{\rm DP}=1/3$ is when the polarizations are randomized.
-For spin-2: The SNR is independent of how the DM energy density is distributed in each polarization mode.
-$f_{\rm DG}\equiv\rho_{\rm DG}/\rho_{\text{DM}}$ is the ratio of the energy density of the spin-2 DM to the total energy density of DM.
-In all the expressions $T_{\rm eff}=T_{\text{obs}}$ if $T_{\text{obs}}<T_{\text{coh}}$, and $T_{\rm eff}=\sqrt{T_{\text{obs}}T_{\text{FFT}}}$ if $T_{\text{obs}}>T_{\text{coh}}$.
-Furthermore, the definitions of the dimensionless coefficients ${\cal C}$, ${\cal B}_{IJ}$ and ${\cal A}$ are:
-${\cal C}$ $\equiv\frac{\rho_{\rm DM}e^{2}v_{0}^{2}}{5\omega^{2}}\left(\frac{Q_{D}}{M}\right)^{2}\,,$ (71)
-${\cal B}_{IJ}$ $\equiv\frac{8\rho_{\rm DM}e^{2}}{3\omega^{4}L_{I}L_{J}}\left(\frac{Q_{D}}{M}\right)^{2}\sin^{2}\left(\frac{\omega L_{I}}{2}\right)\sin^{2}\left(\frac{\omega L_{J}}{2}\right)\,,$ (72)
-${\cal A}$ $\equiv\frac{\rho_{\rm DM}}{5\omega^{2}M_{\text{Pl}}^{2}}$ (73)
-where $\omega=2\pi f$ is the angular frequency of the ultralight dark matter, $L_{I}$ and $L_{J}$ are the arm-lengths of each interferometer,
-III.1.2 Excess power time-frequency method
+In particular, cross-correlation is not nearly as sensitive to the finite-light travel time effect as to the spatial effect when it was applied to search for spin-1 DM, which will be shown in Section III.2.
+Larger couplings, longer effective observation times, and favorable overlap between detectors increase the SNR and thus the probability of detecting a signal.
+Conversely, weaker couplings, shorter observations, or misaligned detectors reduce the SNR, making a signal more difficult to distinguish from instrumental noise.
+After an SNR is computed via Eq.
+68, it can be translated directly from SNR to coupling strengths using Eq.
+120-122 in Manita et al.
+(2024)).
+It is also useful to calculate a minimum detectable strain amplitude, which can also be used to place constraints on the DM coupling constants for scalar, vector, or tensor ultralight dark matter.
+Given the SNR, the corresponding strain amplitude can be estimated as:
+$h_{0,j}=\sqrt{\frac{2\,\mathrm{SNR}\,P_{I,j}\,P_{J,j}}{\gamma^{2}\,T_{\text{obs}}\,T_{\text{FFT}}}},$ (68)
+where $\gamma$ is the ORF that depends on the way that different kinds of DM couple to the interferometers Manita et al.
+(2024).
+By inserting the SNR returned from a search (Eq.
+67), we can obtain an estimate of the minimum detectable strain amplitude for a potential DM signal, and then map those amplitude values to constraints on DM couplings in spin-0, spin-1 or spin-2 using the strain amplitude equations derived in Section II, analogously to translating the SNR directly to coupling strengths in Manita et al.
+(2024).
+III.1.2 Excess power method
 Cross-correlation is able to capture the phase information in each FFT; however, at the moment the software has not yet been implemented to change $T_{\text{FFT}}$ to match $T_{\text{coh}}$, which would result in optimal sensitivity to each DM mass.
 Thus, another method Miller et al.
-(2021) was developed that can vary $T_{\text{FFT}}$ as a function of frequency, and that is more robust against noise disturbances.
-It relies on creating time-frequency spectrograms in which frequencies in each FFT are only kept if their powers are (1) above a given threshold and (2) local maxima.
-After applying these cuts, a time-frequency “peakmap” is created, which is a collection of ones that indicates particular time/frequency points at which the aforementioned two conditions are met.
+(2021a) was developed that can vary $T_{\text{FFT}}$ as a function of frequency, and that is more robust against noise disturbances.
+It relies on creating time-frequency spectrograms in which frequencies in each FFT are only kept if their equalized powers are (1) above a given threshold $\theta_{\rm thr}=2.5$ and (2) local maxima.
+The equalized power is computed at each time and frequency bin by calculating the ratio $R_{ij}$ of the square modulus of the FFT with a running median estimation of the power spectral density, as in Section III.1.1:
+$R_{ij}=\frac{|FFT|_{I,ij}^{2}}{P_{I,ij}}$ (69)
+which, on average, takes on $\mathcal{O}(1)$ values.
+After applying these cuts, a time-frequency “peakmap” is created, which is a collection of ones that indicates particular time/frequency points at which the aforementioned two conditions are met (the value $R_{ij}$ does not enter into any subsequent part of the analysis).
 A peakmap can be created every Hz, and $T_{\text{FFT}}$ can be varied to match $T_{\text{coh}}$ in each 1 Hz window.
@@
 Thus, we can sum, in each frequency bin, the ones that are present, in essence creating a histogram.
-On this histogram, at each frequency, we compute a detection statistic called the “critical ratio”:
+On this histogram, at each frequency, we compute a detection statistic called the “critical ratio” $CR$:
 $CR=\frac{n-\mu}{\sigma}$ (70)
@@
 The $CR$, like the SNR, follows a normal distribution.
-The minimum detectable strain amplitude of this method is:
-$h_{\text{0,min}}$ $\approx\frac{\mathcal{G}}{N_{\text{FFT}}^{1/4}\theta^{1/2}_{\rm thr}}\sqrt{\frac{S_{n}(f)}{T_{\text{FFT,max}}}}$
-×$\times\left(\frac{p_{0}(1-p_{0})}{p_{1}^{2}}\right)^{1/4}\sqrt{CR_{\text{thr}}-\sqrt{2}\mathrm{erfc}^{-1}(2\Gamma)},$
-$N_{\text{FFT}}$ $=\frac{T_{\text{obs}}}{T_{\text{FFT,max}}},$
-$p_{0}$ $=e^{-\theta_{\text{thr}}}-e^{-2\theta_{\text{thr}}}+\frac{1}{3}e^{-3\theta_{\text{thr}}}=0.0755\text{ for $\theta_{\text{thr}}=2.5$,}$
-$p_{1}$ $=e^{-\theta_{\text{thr}}}-2e^{-2\theta_{\text{thr}}}+e^{-3\theta_{\text{thr}}}\text{ }=0.0692\text{ for $\theta_{\text{thr}}=2.5$,}$ (75)
-where $\mathcal{G}=[1.3,2.8,2.08]$ for scalar, vector or tensor DM interactions, respectively.
-The $\mathcal{O}(1)$ differences arise from averages over the spin-1 or spin-2 polarizations.
-$\theta_{\text{thr}}$ is a threshold on the equalized power in the time-frequency spectrogram, and $\Gamma$ is the confidence level.
+Using the $CR$ and assuming Gaussian noise, the minimum strain amplitude of a sinusoidal signal can be derived that would, in a frequentest interpretation, produce a detectable signal in a fraction $\geq\Gamma$ of a repeated number of experiments Astone et al.
+(2014); Miller et al.
+(2021a):
+$h_{\text{0,min}}$ $\approx\frac{\mathcal{G}}{T_{\text{obs}}^{1/4}T_{\text{FFT}}^{1/4}}\sqrt{\frac{P_{I}(f)}{\theta_{\text{thr}}}}$
+$\times\left(\frac{p_{0}(1-p_{0})}{p_{1}^{2}}\right)^{1/4}\sqrt{CR_{\text{thr}}-\sqrt{2}\mathrm{erfc}^{-1}(2\Gamma)},$
+$p_{0}$ $=e^{-\theta_{\text{thr}}}-e^{-2\theta_{\text{thr}}}+\frac{1}{3}e^{-3\theta_{\text{thr}}},$
+$p_{1}$ $=e^{-\theta_{\text{thr}}}-2e^{-2\theta_{\text{thr}}}+e^{-3\theta_{\text{thr}}}$ (71)
+where $\mathcal{G}$ depends on whether DM is scalar, vector or tensor:
+$\mathcal{G}$ $=\sqrt{\frac{2\pi}{2.4308}}\frac{\sqrt{2}}{3}\times\begin{cases}1,\text{\quad\, dilatons}\\ \sqrt{\frac{9}{2}},\text{ dark photons}\\ \sqrt{\frac{5}{2}},\text{ tensor bosons}\\ \frac{3}{\sqrt{2}}\frac{5}{2},\text{ CWs}\end{cases}$ (72)
+$=\begin{cases}1.31,\text{ \, dilatons}\\ 2.78,\text{ dark photons}\\ 2.08,\text{ tensor bosons}\\ 4.02,\text{ CWs}\end{cases}$ (73)
+The $\mathcal{O}(1)$ differences arise from averages over the spin-1 or spin-2 polarizations, and are valid for ground-based L-shaped GW interferometers (for triangular shapes, the averages will be slightly different).
+The $\frac{\sqrt{2}}{3}$ factor comes from an average over the DM Maxwell-Boltzmann distributed velocities as discussed in Section II.4, while the $\frac{2\pi}{2.4308}$ factor comes from convolving a sinusoid with a rectangular window function when computing the sensitivity of GW interferometers to a sinusoidal signal Astone et al.
+(2014).
 III.1.3 Logarithmic power spectral density (LPSD) method
-To truly maximize the sensitivity towards DM signals, one must, with infinite precision, analyze chunks of data with $T_{\text{coh}}$ as the DM mass changes.
-LPSD Tröbs and Heinzel (2006); Vermeulen et al.
+To maximize sensitivity to ultralight dark-matter signals, it is important to let $T_{\text{FFT}}$ match $T_{\text{coh}}$ as the DM mass varies.
+The logarithmic power spectral density (LPSD) method Tröbs and Heinzel (2006); Vermeulen et al.
 (2021); Göttel et al.
-(2024) is an attempt to do this, which creates spectra that are logarithmically spaced in frequency.
-It is extremely computationally intensive, since it requires an implementation separate from the discrete FFT.
-The benefit of using such an algorithm is its ability to estimate power spectral densities logarithmically, thereby capturing important features of the signals that could be missed when estimating the power spectral density across a linear frequency range.
-Given that the ultralight dark matter causes frequency variations of $\mathcal{O}(10^{-6})$, it is worthwhile to use LPSD to capture these tiny fluctuations in frequency.
-Accounting for the strain on the test masses, the loss of sensitivity due to the number of round-trips of the laser light, as well, the induced strain can also be written in terms of transfer functions between scalar DM and the interferometers:
-$h(\omega)\approx\left(\frac{1}{\Lambda_{\gamma}}+\frac{1}{\Lambda_{e}}\right)\cdot\left(\frac{\hbar\,\sqrt{2\,\rho_{\mathrm{DM}}}}{m_{\rm DM}\,c}\right)A_{\rm cal}^{-1}(\omega)$ (76)
-where $A_{\rm cal}$ is a function given in Göttel et al.
-(2024) that encapsulates the transfer functions between the DM and the beam splitter and test masses, as well as the thicknesses and indices of refraction of the mirrors, the mirrors’ transmissivities, and the GW/ detector transfer function.
-Then, $h(\omega)$ can be estimated from the data using Eq.
-11 and 12 in Göttel et al.
-(2024), and Eq.
-16 in Cowan et al.
-(2011), which can be mapped to constraints on the coupling constants using Eq.
-76.
+(2024) implements this idea by constructing spectra that are logarithmically spaced in frequency, in contrast to the excess power method in Section III.1.2 that use linearly spaced spectra every 1 Hz.
+While this approach is computationally intensive due to the need for a separate implementation from the standard discrete FFT, it captures subtle features in the data that could otherwise be missed.
+In particular, since ultralight dark matter induces fractional frequency variations on the order of $\mathcal{O}(10^{-6})$, a logarithmic spectral estimate is well-suited to resolving such tiny modulations.
+Once the LPSD is computed, it can be combined with the interferometer response to convert measured strains into constraints on DM couplings.
+Accounting for the strain induced on the test masses and beam splitter (see Section II.3.1), the expected signal can be written as
+$h(\omega)\approx\left(\frac{1}{\Lambda_{\gamma}}+\frac{1}{\Lambda_{e}}\right)\cdot\left(\frac{\sqrt{2\,\rho_{\mathrm{DM}}}}{m_{\rm DM}}\right)A_{\rm cal}^{-1}(\omega)\,$, (74)
+where $A_{\rm cal}(\omega)$ encodes the full interferometer transfer function.
+This function accounts for the response of the beam splitter and test masses to the DM signal, which is amplified by the difference in mirror thicknesses as discussed in Section II.3.1, mirror transmissivities, and the overall canonical GW detector transfer function.
+By applying Eq.
+74 to the logarithmically estimated spectra, one can extract an estimate of the strain $h(\omega)$ from the data.
+This estimate can then be mapped to constraints on the coupling constants $\Lambda_{\gamma}$ and $\Lambda_{e}$, providing a direct connection between the interferometer measurement and the underlying ultralight dark matter model Göttel et al.
+(2024); Cowan et al.
+(2011).
+In this way, the LPSD method offers a way to detect logarithmically-spaced frequency modulations that might be missed by linear spectral analyses.
 III.1.4 Stochastic summing method
-This method leverages the fact that the signal power will be stochastically spread at frequencies above the characteristic one if $T_{\text{FFT}}>T_{\text{coh}}$, and essentially sums the power at those frequencies.
-Afterwards, a detection statistic, like the signal-to-noise ratio, is calculated by dividing by the noise power spectral density:
-$\rho(f_{c})\equiv\sum_{i}^{N_{\text{FFT}}}\sum_{f_{0}\leq f_{n}\leq f_{0}(1+\kappa^{2}v^{2}_{0}/c^{2})}\frac{4|\tilde{d}(f_{n};t_{i})|^{2}}{T_{\text{FFT}}S(f_{n};t_{i})},$ (77)
-where $\tilde{d}(f_{n};t_{i})$ is the FFT of $i^{\rm th}$ data chunk, $S(f_{n};t_{i})$ is the one-sided noise power spectral density at $t=t_{i}$, and $\kappa$ = 3.17 to ensure that no more than 1% of signal power is lost by restricting the inner sum.
+The stochastic summing method Nakatsuka et al.
+(2023) exploits the fact that when $T_{\text{FFT}}>T_{\text{coh}}$, the signal power from ultralight dark matter is spread stochastically over a range of frequencies rather than concentrated at a single frequency.
+Instead of searching for a narrow peak, this method sums the power across all frequency bins where the signal is expected, producing a detection statistic that captures the total available signal power.
+$\rho(f_{c})\equiv\sum_{i}^{N_{\text{FFT}}}\sum_{f_{0}\leq f_{n}\leq f_{0}(1+\kappa^{2}v^{2}_{0})}\frac{4|\tilde{d}(f_{n};t_{i})|^{2}}{T_{\text{FFT}}S(f_{n};t_{i})},$ (75)
+where $\tilde{d}(f_{n};t_{i})$ is the FFT of $i^{\rm th}$ data chunk, $S(f_{n};t_{i})$ is the one-sided noise power spectral density at $t=t_{i}$, and $\kappa$ = 3.17 is chosen to ensure that no more than 1% of signal power is lost by limiting the frequencies summed over.
 The outer sum is over all chunks of length $T_{\text{FFT}}$; the inner sum is over the frequency spread of the signal $\delta f_{v}$.
-The sensitivity of this method has been estimated in Nakatsuka et al.
-(2023) for both axions and dark photons, analytically and numerically in the case of $T_{\text{FFT}}=T_{\text{obs}}$.
-In the absence of a signal, assuming Gaussian noise, a detection threshold can be established as a function of the false alarm probability $\alpha$ (Eq.
-47 in Nakatsuka et al.
-(2023)), and this threshold can be fed back into an integral over the likelihood function (Eq.
-46 in Nakatsuka et al.
-(2023)), to compute, at a given confidence level, the minimum detectable signal amplitude $\bar{\lambda}_{\mathrm{up}}$:
-$\bar{\lambda}_{\mathrm{up}}(T_{\text{obs}})=\begin{cases}(\Delta_{X,\mathrm{tot}})^{-1/2}\sqrt{\frac{\ln(\alpha)}{\ln(\beta)}-1},&\mathrm{for}~{}T_{\text{obs}}<\tau/\kappa,\\ (\Delta_{X,\mathrm{tot}})^{-1/2}\sqrt{M_{\alpha}+M_{1-\beta}}\left(\kappa T_{\text{obs}}/T_{\text{coh}}\right)^{1/4}&\mathrm{for}~{}T_{\text{obs}}\gg T_{\text{coh}}.\end{cases}$ (78)
-$\Delta_{X,\text{tot}}$ encodes the deterministic part of the spectral shape of the signal, and depends on the relative motion of the DM velocity with respect to the interferometer arms.
-$\beta$ is the detection probability.
-In principle, Eq.
-11 and 17 of Nakatsuka et al.
-(2023) give these relative motion terms; in practice, however, conservative ($\vec{v}=[0,0,v_{0}]$) and optimistic ($\vec{v}=[v_{0},0,0]]$) choices of the direction of the DM velocity with respect to the interferometers result in values of [0.38,0.81], respectively (see Eq.
-52 in Nakatsuka et al.
-(2023)).
-$M_{\chi}$ represents the relation between the peak width and the area of the Gaussian distribution defined by
-[…]
-Depending on the type of DM signal assumed to be interacting with the interferometers, $\bar{\lambda}_{\mathrm{up}}$ can be equated to the following physical normalized (by the noise) DM signal amplitudes to obtain the upper limit on the corresponding coupling constant:
-$\bar{\lambda}_{\mathrm{axion}}$ $=\frac{T_{\text{obs}}}{\sqrt{T_{\text{obs}}S_{n}}}g_{\alpha\gamma}\sqrt{\frac{\rho_{\rm DM}}{m^{2}}}\frac{m}{2k},\quad$
-$\bar{\lambda}_{\mathrm{time}}$ $=\epsilon e\frac{2T_{\text{obs}}}{\sqrt{T_{\text{obs}}S_{n}}}\sqrt{\frac{2\rho_{\rm DM}}{3m^{2}}}\frac{(Q/M)_{\mathrm{in}}}{mL}\sin^{2}\left(\frac{mL}{2}\right),\quad$
-$\bar{\lambda}_{\mathrm{space}}$ $=\epsilon e\frac{2T_{\text{obs}}}{\sqrt{T_{\text{obs}}S_{n}}}\sqrt{\frac{2\rho_{\rm DM}}{3m^{2}}}\frac{(Q/M)_{\mathrm{in}}\bar{v}}{2{\sqrt{2}}},\quad$
-$\bar{\lambda}_{\mathrm{charge}}$ $=\epsilon e\frac{2T_{\text{obs}}}{\sqrt{T_{\text{obs}}S_{n}}}\sqrt{\frac{2\rho_{\rm DM}}{3m^{2}}}\frac{|(Q/M)_{e}-(Q/M)_{\mathrm{in}}|}{2Lm}.$ (80)
-where $L$ is the length of the interferometers and $(Q/M)_{e},(Q/M)_{\mathrm{in}}$ are the charge-to-mass ratios of the end mirrors and input mirrors, respectively.
-The above approach works if $T_{\text{coh}}\neq T_{\text{obs}}$, however in practice, this does happen, since the data are typically broken in to smaller chunks with $T_{\text{FFT}}\sim T_{\text{coh}}$.
-The likelihood equation derived in Nakatsuka et al.
-(2023) is actually unstable in this region, and so a numerical approach is used to estimate sensitivity and derive upper limits as in Abac et al.
+The sensitivity of the stochastic summing method has been studied for both axions and dark photons Nakatsuka et al.
+(2023), using a combination of analytical estimates and numerical simulations.
+In the absence of a signal and assuming Gaussian noise, a detection threshold can be established based on a chosen false-alarm probability.
+This threshold can then be used to determine, at a given confidence level, the minimum detectable signal amplitude, similar to Eq.
+71 in Section III.1.2.
+The exact sensitivity depends on the spectral shape of the signal, which in turn is influenced by the relative motion of the DM with respect to the interferometer.
+Conservative and optimistic assumptions about the direction of the DM velocity lead to $\mathcal{O}(1)$ variations in the sensitivity, which are accounted for in this method Nakatsuka et al.
 (2023).
+Once the minimum detectable strain amplitude is determined, it can be translated into upper limits on the relevant coupling constants for each type of dark matter interaction, similarly to what has been discussed in Sections III.1.1 and III.1.2.
+These include axion-photon couplings as well as scalar or vector couplings that induce displacements or charge effects on the interferometer mirrors.
+In practice, because data are typically divided into segments with $T_{\text{FFT}}\sim T_{\text{coh}}$, the analytic likelihood approach can become unstable.
+Consequently, numerical simulations are used to estimate the sensitivity and derive robust upper limits, as done in a search on O3 KAGRA data Abac et al.
+(2024), which will be discussed in Section III.2.
 III.1.5 Distinguishing amongst dark-matter models
-Only a couple of works Miller et al.
-(2022a); Manita et al.
-(2024) have attempted to find ways of determining whether a particular type of DM interacted with GW detectors.
+Only a few works Miller et al.
+(2022b); Manita et al.
+(2024) have explored ways to determine whether a particular type of DM interacted with GW detectors.
 In Miller et al.
-(2022a), the authors employ the Wiener filter Wiener et al.
-(1964) to compute the cross power across different detector pairs in the follow-up stages of the analysis, using different $T_{\text{FFT}}$, to exploit the ways that types of DM will couple to different interferometer baselines.
-Not only was the Wiener filter shown to produce low residuals when the signal in the data matched the model, and high values when the data were filtered with the wrong signal but also it could, with high confidence, veto candidates returned in a real search of O3 data.
-Ref.
+(2022b), the authors employ the Wiener filter Wiener et al.
+(1964) to compute the cross power across different detector pairs during follow-up stages, using different $T_{\text{FFT}}$ to exploit the fact that different types of DM couple differently to different detector baselines.
+In Wiener filtering, the goal is to find a waveform that, when subtracted from the data, leaves only noise.
+The residual quantifies the mismatch between the data, $h(t)+n(t)$, and a candidate signal, $h(t)$.
+Low residuals indicate that the candidate waveform closely matches the data, whereas high residuals indicate poor agreement.
+Miller et al.
+(2022b) showed that the Wiener filter produces small residuals when the data contain a signal consistent with the model and large residuals when the wrong signal is applied.
+This property allowed the method to confidently veto spurious candidates in real O3 data.
+Conversely, this means that the method has the potential to confirm the detection of a certain type of DM interaction signal.
+This is in contrast to other methods and previous searches Guo et al.
+(2019); Miller et al.
+(2021a); Abbott et al.
+(2022c), which focused mostly on rejecting outliers that were shown to be compatabile with noise.
+After the work of Miller et al.
+(2022b), Ref.
 Manita et al.
 (2024) uses standard cross correlation to argue that the overlap reduction function can be used as way to distinguish amongst different spins of DM particles, since the cross power will be different.
-In both cases, cross power is important, since the individual power spectra are more or less indistinguishable for all types of ultralight dark matter, which arises because the phase evolutions for the considered types of ultralight dark matter here are the same.
-III.2 Ground-based interferometers
+In particular, Ref.
+Manita et al.
+(2024) how the cross-correlation SNR would change as a function of the type of DM and which baseline was considered.
+In both cases, cross power is important, since the individual power spectra from one detector are indistinguishable for all types of ultralight dark matter, which arises because the phase evolutions for the considered types of ultralight dark matter here are the same.
+III.2 Ground- and space-based interferometers
 Searches for DM particles interacting with the LIGO, Virgo, KAGRA and GEO600 interferometers have been performed using data from the most recent observing run, O3, resulting in constraints on dilatons and dark photons Abbott et al.
@@
 (2024).
-We summarize the results that have arisen from each of the existing ground-based interferometers.
+Furthermore, LISA Pathfinder data were also used to set constraints on dark photon DM Miller and Mendes (2023); Frerick et al.
+(2024).
+The theoretical framework for each particle is described in Section II.3 and Section II.4, respectively.
+In this subsection, we summarize key constraints from existing ground-based interferometers on these two particles.
 III.2.1 LIGO, Virgo and GEO600
-In Abbott et al.
-(2022c), two methods were used to search for ultralight dark matter: cross-correlation and excess power.
+In Ref.
+Abbott et al.
+(2022c), two complementary methods were employed to search for spin-1 ultralight dark matter that can couple to baryons in the test masses of LIGO in O3 data – cross-correlation and excess power, both of which are described in Section III.1.1 and Section III.1.2, respectively.
 Cross-correlation benefits from utilizing the phase information of the signal, something that is lost when using the time-frequency excess power method.
-On the other hand, the excess power method matches its coherence time $T_{\text{FFT}}$ to the signal coherence time $T_{\text{coh}}$ in every Hz band analyzed, allowing an improved sensitivity across the frequency range compared to cross-corelation, which fixed $T_{\text{FFT}}=1800$ s.
-In Fig.
-3, we show the upper limits on the coupling constant $\epsilon^{2}$ for both methods, in comparison to existing experiments.
+On the other hand, the excess power method matches its coherence time $T_{\text{FFT}}$ to the signal coherence time $T_{\text{coh}}$ in every one-Hz band analyzed, allowing an improved sensitivity across the frequency range compared to cross-corelation, which fixed $T_{\text{FFT}}=1800$ s.
+Both searches returned no significant DM candidate signals; thus, 95% confidence-levle upper limits were set on the coupling strength $\epsilon^{2}$ of dark photons to baryons for both methods, as shown in Fig.
+4..
 We can see, across a broad frequency range, that results from GW interferometers surpass by orders of magnitude constraints from Eöt-Wash and the MICROSCOPE experiments.
-These upper limits were derived for cross-correlation and excess power by following the procedure outlined in Section III.1.1 and Section III.1.2, respectively.
+These upper limits were derived for cross-correlation and excess power following the procedure outlined in Section III.1.1 and Section III.1.2, respectively.
 Both methods employed the Feldman-Cousins approach Feldman and Cousins (1998) to set upper limits, which is robust against non-Gaussian noise and has been shown to produce conservative upper limits with respect to those that would have been obtained through simulations Miller et al.
-(2021).
+(2021a).
 Figure 4: Adapted from Abbott et al.
 (2022c).
-Upper limits obtained from analyzing LIGO O3 data on the coupling of dark photons to baryons $U(1)_{\rm B}$ in the LIGO mirrors.
+Upper limits obtained from analyzing LIGO O3 data on the square of the coupling of dark photons to baryons $U(1)_{\rm B}$ in the LIGO mirrors.
+The physics behind this form of DM has been discussed in Section II.4.
+The limits derived from each of the methods discussed in Sections III.1.1 and III.1.2 are show in red (cross-correlation) and cyan (excess power), respectively.
 MICROSCOPE Bergé et al.
-(2018) and Eot-wash torsion balance upper limits are plotted as a comparison to the results here Schlamminger et al.
+(2018) and Eöt-Wash torsion balance upper limits are plotted as a comparison to the results here Schlamminger et al.
 (2008).
-To produce limits on dark photon/baryon-lepton coupling, $U(1)_{\rm B-L}$ in the LIGO mirrors, these limits should be multiplied by four.
-Additionally, a recent search for scalar bosons interacting with LIGO was performed Göttel et al.
-(2024) using LPSD, resulting in tight constraints on the coupling constants given in Eq.
-27.
+To produce limits on the square of the dark photon/baryon-lepton coupling, $U(1)_{\rm B-L}$ in the LIGO mirrors, these limits should be multiplied by four.
+While cross-correlation and excess power analyses were used for dark photon searches, the LPSD technique (described in Section III.1.3) was employed in Ref.
+Göttel et al.
+(2024) to search for scalar dilaton dark matter in the LIGO O3 dataset.
+This analysis placed strong limits on the coupling of scalar dark matter to electrons and photons.
 In Fig.
-4, we can see, in blue, the constraints that have been derived, which supersede, at least a low masses, those that were derived in a GEO600 search for the same kind of DM Vermeulen et al.
+5, we can see, in blue, the constraints that have been derived, which supersede, at least at low masses, those that were derived in a GEO600 search for the same kind of DM Vermeulen et al.
 (2021).
 The enhancement comes from the improved low-frequency sensitivity of LIGO relative to GEO600, as well as incorporating the interaction of scalar DM with the mirrors, as well as the beam splitter, into the search.
-These upper limits were derived for by following the procedure outlined in Section III.1.3 and Section III.1.2.
+However, at high frequencies, we can see that GEO600 still outperforms LIGO, primarily because of sophisticated quantum technologies that were employed in its design.
+These upper limits were derived by following the procedure outlined in Section III.1.3 and Section III.1.2.
 Figure 5: Taken from Göttel et al.
@@
 95% confidence-level upper limits on $\Lambda_{i}^{-1}$, the coupling of dilaton DM to electrons or photons, as a function of mass and frequency from LIGO O3 data, which would have cause time-dependent oscillations of the sizes and indices of refraction of the beamsplitter and the LIGO mirrors.
-a) and b) show these constraints compared to existing ones on $\Lambda_{e}$ and $\Lambda_{\gamma}$, respectively.
+The physics behind this form of DM has been discussed in Section II.3.1.a) and b) show these constraints compared to existing ones on $\Lambda_{e}$ and $\Lambda_{\gamma}$, respectively.
 The results from the LIGO O3 search are shown by the thick blue line, constraints from direct experimental searches for DM Vermeulen et al.
@@
 (2021); Zhang et al.
-(2023); Fukusumi et al.
+(2023a); Fukusumi et al.
 (2023); Sherrill et al.
@@
 III.2.2 KAGRA
-KAGRA is another interferometer in Japan that, while not yet as sensitive to GWs as LIGO and Virgo, can be used to search for DM.
-In particular, the different materials used in the mirrors in KAGRA can be exploited to be sensitive to vector DM that couples to baryon-lepton number Michimura et al.
-(2020), since the signal-to-noise ratio is enhanced for objects with different charge-to-mass ratios, as indicated by Eq.
-80.
-Furthermore, this kind of DM can actually be seen in non-GW channels, such as the auxiliary length channels PRCL and MICH.
-The raw outputs of PRCL and MICH are time-varying displacement sof the forms (we also write the GW channel for comparison, $\delta L_{\rm DARM}$) Abac et al.
-(2024):
-$\delta L_{\rm DARM}$ $=$ $\delta(L_{\rm x}-L_{\rm y}),$ (81)
+KAGRA offers unique opportunities to search for DM.
+In particular, the different materials used in KAGRA’s mirrors result in significantly different charge-to-mass ratios for baryon-lepton number, enhancing the sensitivity to vector DM coupling to this quantum number Michimura et al.
+(2020).
+While LIGO and Virgo currently set the strongest limits on dark photon coupling to baryon-lepton number – obtainable from the results in Fig.
+4 by multiplying the limits on baryon number coupling by four Abbott et al.
+(2022c) – this difference in mirror materials means that once KAGRA reaches comparable sensitivity levels, it will outperform LIGO in constraining such couplings Michimura et al.
+(2020).
+This enhancement is particularly apparent in KAGRA’s auxiliary length control channels, PRCL and MICH, which record displacements of mirrors made from both fused silica (beam splitter and power recycling mirrors) and cryogenic sapphire (test masses).
+The baryon-lepton charge-to-mass ratios differ by approximately $9\times 10^{-3}$ between these materials, compared to only about $4\times 10^{-5}$ for baryon number alone, significantly amplifying the DM interaction signal in these channels – see Eq.
+54.
+The raw outputs of PRCL and MICH correspond to time-varying displacements of the form Abac et al. (2024):
 $\delta L_{\rm MICH}$ $=\delta(l_{\rm x}-l_{\rm y}),$ (76)
-$\delta L_{\rm PRCL}$ $=$ $\delta[(l_{\rm x}+l_{\rm y})/2+l_{\rm p}],$ (83)
-In KAGRA, $L_{x}=L_{y}=3000$m is the arm length, $l_{x}=26.7$m is the distance between the beam splitter and the input $x$ mirror, $l_{y}=23.3$m is the distance between the beam splitter and the input $y$ mirror, and $l_{p}=41.6$m is the the length of the power recycling cavity.
-All of these lengths are much shorter than the DM coherence length $L_{\text{coh}}$.
+$\delta L_{\rm PRCL}$ $=\delta\left[\frac{l_{\rm x}+l_{\rm y}}{2}+l_{\rm p}\right],$ (77)
+where $l_{x}=26.7\,\text{m}$ and $l_{y}=23.3\,\text{m}$ are the distances between the beam splitter and the input mirrors, and $l_{p}=41.6\,\text{m}$ is the length of the power recycling cavity.
+All these lengths are much shorter than the DM coherence length $L_{\text{coh}}$.
 See Fig.
@@
 (2024) for a schematic of the interferometer.
-These two channels contain both fused silica mirrors used for the beam splitter and power recycling mirrors, and cryogenic sapphire test masses.
-The charge-to-mass ratios for baryon-lepton number differ by $\sim 9\times 10^{-3}$ for fused silica and sapphire, while for baryon number alone, these quantities differ by only $\sim 4\times 10^{-5}$, showing a large enhancement of the baryon-lepton DM interaction signal.
 No significant signals were found in this analysis of KAGRA O3 data Abac et al.
-(2024), so upper limits, as shown in Fig.
-5, were calculated following the procedure outlined in Section III.1.4.
-The upper limits are shown for each channel analyzed, and many narrow-band noise artifacts are visible, which tends to limit the sensitivity of this search at low freqeuncies.
-The constraints are not as strong as those that come from LIGO; however, future searches with KAGRA are expected to yield powerful constraints on this type of DM due to the difference in charge-to-mass ratios of the mirrors Michimura et al.
+(2024), and upper limits on the coupling strength were derived following the procedure outlined in Section III.1.4.
+These limits, shown in Fig.
+6 for each channel, exhibit many narrow-band noise artifacts that limit sensitivity at low frequencies.
+Although not yet competitive with LIGO constraints, future KAGRA searches are expected to provide powerful limits on this type of DM due to the pronounced difference in baryon-lepton charge-to-mass ratios among its mirrors Michimura et al.
 (2020).
@@
 (2024).
-95% confidence-level upper limit on the strength of the coupling between dark photons and baryon-lepton number using KAGRA O3 data, derived from two particular interferometer channels: MICH (blue line) and PRCL (orange line).
-Because the KAGRA input mirrors, end mirrors, and power recycling mirrors are made of different material, the coupling strength of dark photons to baryon-lepton number is enhanced, and is more visible in these particular channels than in the ordinary strain one.
-In the low-mass (low-frequency) range, many narrow-band noise disturbances are seen, which arise due to unknown sources.
-III.3 Space-based interferometers
-Spaced-based GW interferometers, such as LISA Danzmann and Team (2003); Babak et al.
-(2021), Taiji Hu and Wu (2017), Tianqin Luo et al.
-(2016) and DECIGO Hu and Wu (2017), will hopefully fly within the next 10-15 years.
-The exquisite sensitivity of GWs in the $\mu$Hz to mHz band also allows the prospect to probe ultralight dark matter with masses of [$10^{-19},10^{-14}$] eV, a few orders of magnitude lower than those currently searched for in ground-based interferometers Kim (2023).
-Until now, however, only a protoype for LISA, called LISA Pathfinder, has been flown as a “proof-of-concept” mission Armano et al.
-(2016); Vetrugno (2017); Armano et al.
+95% confidence-level upper limit on the strength of the coupling between dark photons and baryon-lepton number using KAGRA O3 data and the method described in Section III.1.4, derived from two particular interferometer channels: MICH (blue line) and PRCL (orange line).
+The physics behind this form of DM has been discussed in Section II.4.
+Because the KAGRA input mirrors, end mirrors, and power recycling mirrors are made of different material, the coupling strength of dark photons to baryon-lepton number is enhanced (Eq.
+54), and appears more strongly in these particular channels than in the ordinary strain one.
+In the low-mass (low-frequency) range, many narrow-band noise disturbances of unknown origin are seen.
+III.2.3 LISA Pathfinder
+[…]
+Such a search was preformed using LISA Pathfinder data Miller and Mendes (2023) using the excess power method described in Section III.1.2 for ultralight dark photons coupling to baryons, however, no physical constraints on the coupling strength were obtained, since the noise level was too high, the arm length was too short (only $40$ cm compared to the expected LISA arm lengths of $\mathcal{O}(10^{6})$ km), and only one arm existed.
+However, the search procedure designed will serve as a roadmap for future analyses of spaced-based detector data, since many problems, including limited sampling at low frequency, gaps, noise non-stationarities and glitches had to be handled in this analysis of data from a space-based detector.
+After this search was performed, however, it was realized that data from a different channel would produce tighter constraints, particularly in the case of baryon-lepton coupling.
+Here, the relative acceleration of the spacecraft and one of the test masses would produce a much stronger signal than that which was searched for previously because the charge-to-mass ratios of each are different Frerick et al.
 (2024).
-Nevertheless, pilot searches can be performed on LISA Pathfinder data as a way to prepare for when space-based detectors fly, which allows for the development of robust data analysis pipelines and for handling of the peculiarities of future data, e.g.
-glitches, downtimes, etc.
-Such a search was preformed using LISA Pathfinder data Miller and Mendes (2023) using the excess power method described in Section III.1.2 for ultralight dark photons coupling to baryons, however, no physical constraints on the coupling strength were obtained, since the noise level was too high, the arm length was too short (only $40$ cm compared to $\mathcal{O}(10^{6})$ km), and only one arm existed.
-However, the search procedure designed will serve as a roadmap for future analyses of spaced-based detector data, since many problems, including limited sampling at low frequency, gaps, noise non-stationarities and glitches had to be handled in this analysis of data from a space-based detector.
-After this search was performed, however, it was realized that the relative acceleration of the spacecraft and one of the test masses would produce a much stronger signal than that which was searched for previously, which was the relative acceleration of the two test masses Frerick et al.
-(2024).
+This echoes the logic presented in Section III.2.2 to use KAGRA to search for dark photons coupling to baryon-lepton number.
 A conservative upper limit on the baryon-lepton coupling was set without performing a search, assuming no signal would have been detected in LISA Pathfinder data, as shown in Fig.
-6, by using a reference acceleration power spectral density of the spacecraft with respect to one of the test masses.
-We can see that the upper limits from the relative acceleration of space-craft/ test mass are much better than that from Miller and Mendes (2023), labeled .(decoherence).
+7, by using a reference acceleration power spectral density of the spacecraft with respect to one of the test masses.
+We can see that the upper limits from the relative acceleration of space-craft/ test mass are much better than that from Miller and Mendes (2023), labeled (decoherence).
 The constraint on B-L coupling is stronger with respect to that of baryon coupling because the test mass and spacecraft are primarily made of gold, and gold/carbon, respectively.
@@
 (2024).
-Upper limits on the coupling of dark photons to baryon-lepton number considering the relative acceleration of the spacecraft and one of the test masses in LISA Pathfinder.
-The constraints derived from a search for the induced acceleration of the two test masses by dark photons in LISA Pathfinder data is shown as “decoherence” .
+Upper limits on the coupling of dark photons to baryon-lepton number considering the relative acceleration of the spacecraft and one of the test masses in LISA Pathfinder (dark blue region).
+The physics behind this form of DM has been discussed in Section II.4, and constrains the same model as in Figs.
+6 and 4, though at much lower boson masses.
+The constraints derived from a search for the induced acceleration of the two test masses by dark photons in LISA Pathfinder data is shown as “decoherence” (dark red region).
 Using the relative acceleration caused by dark photon interaction with baryon-lepton number allows for stringent constraints to be placed on the coupling, since the spacecraft and test masses are made of different materials.
-Constraints from other experiments and future projections for LISA are shown as well.
+Constraints from other experiments (Eöt-Wash and MICROSCOPE) and future projections for LISA, MAGIS-100 and torision balance are shown as well.
+Details on the curves coming from STE-QUEST and Asteroids can be found in O’HARE (2020).
 Only B-L constraints are shown from Frerick et al.
-(2024) because those arising from coupling to baryons are not yet strong enough to probe a tighter constraint than provided by MICROSCOPE.
+(2024) because those arising from coupling to baryons are not yet strong enough to probe a tighter constraint than that provided by MICROSCOPE.
 However, many conservative assumptions were made in Frerick et al.
 (2024), including the composition of the spacecraft and the use of a reference acceleration power spectral density, both of which could, if studied in more detail, lead to more stringent constraints for both the baryon and B-L coupling scenarios.
-III.4 Pulsar timing arrays
-Because pulsars are extremely accurate clocks, they can also be used to search for GWs, which will distort the spacetime between earth and pulsars.
-Practically speaking, GWs will cause a change in the times of arrivals of pulses with respect to that in the absence of GWs.
-Of course, other effects between the pulsars and earth must be taken into account, e.g.
-interstellar medium dispersion and scattering, instabilities in the rotation of the pulsar itself, which makes these measurements particularly challenging.
-Nevertheless, pulsar timing arrays can be formed by cross correlating the times of arrivals from multiple pulsars in the galaxy, mitigating many of these effects.
-Ultralight scalar DM will also cause deviations in the times of arrivals of pulses as given in Khmelnitsky and Rubakov (2014), though the mass of such particles are 10 orders of magnitude lower than those probed with ground-based interferometers.
-It can also induce stochastic perturbations in the metric, which would also induce time-varying changes in the times of arrivals of pulses Kim and Mitridate (2024), but generally could be seen in interferometers with long baselines Kim (2023).
-In Section II, we discussed various models of scalar DM that can be constrained with pulsar timing arrays.
-We again divide this section in the different types of DM that can be probed with pulsar timing arrays.
-III.4.1 Gravitationally-interacting ultralight dark matter
-As discussed before, DM could interact only via gravity, and alter the distances between the pulsars and earth, analogously to GWs.
+III.3 Pulsar timing arrays
+Because pulsars are extremely stable astrophysical clocks, they can be used to search for GWs, which perturb the spacetime between Earth and pulsars.
+In practical terms, GWs induce fluctuations in the ToAs of radio pulses, relative to what would be expected in their absence.
+However, interpreting these ToA shifts is complicated by other astrophysical and instrumental effects – such as dispersion and scattering in the interstellar medium, and intrinsic rotational instabilities of the pulsars themselves – which makes precision measurements challenging.
+Nevertheless, these effects can be mitigated by cross-correlating data from many pulsars in the sky to form a pulsar timing array.
+As described in Section II.3, ultralight scalar DM can also induce variations in pulse ToAs Khmelnitsky and Rubakov (2014), though the relevant mass range lies many orders of magnitude below that probed by ground-based interferometers.
+Such scalar fields can cause coherent oscillations in the gravitational potential or effective constants of nature, and may also generate stochastic perturbations in the metric Kim and Mitridate (2024), which could lead to further timing variations over long baselines Kim (2023).
+Throughout this section, we will see that, independently of the physical mechanism of coupling of DM to gravity or the standard model, the shapes of the constraints on the coupling constant appear similar.
+This is because in these cases, we are looking for an almost sinusoidal signal embedded in pulsar timing array data with particular noise characteristics that are independent of the DM model.
+Though the underlying physics behind each DM model is different, the observable – shifts in ToAs– is identical, though with different amplitudes that relate to the coupling constant of DM.
+This connects to our previous comments in the introductory paragraphs of Section III: the signal arising from DM coupling to gravity or the standard model is the same – a quasi-sinusoidal oscillation of spacetime, of materials, of test masses, etc.
+– but the cause and therefore the interpretation of time-varying ToA in pulsar timing arrays or time-varying displacements of test masses in ground- or space-based GW interferometers is fundamentally different, which leads to different constraint plots whose curves look similar.
+In Section II.3, we discussed the theoretical mechanisms by which scalar DM could leave signatures on pulsar timing array data, and in Section III.2 we described how related effects could be searched for with ground-based interferometers.
+In contrast to that section, here we organize existing constraints from pulsar timing arrays according to the specific physical mechanism by which scalar DM couples to the standard model or to gravity.
+Each mechanism leads to a distinct signature in the pulsar timing residuals.
+We treat these mechanisms independently: if we put constraints on DM coupling to dilatons in the context of one mechanism, we assume that all other mechanisms do not contribute
+III.3.1 Gravitationally-interacting ultralight dark matter
+As discussed in Section II.6, DM could interact only via gravity, and alter the distances between the pulsars and earth, analogously to GWs.
 Using European pulsar timing array data, a search was performed in the residuals for this kind of DM interaction.
@@
 This means that each pulsar experiences the same phase of the ultralight dark-matter signal.
-In this case, the same coherence patch and local DM energy density are measured.
+In this case, the same coherence patch and local DM energy density are measured: the amplitude of the ultralight dark-matter interaction signal can be attributed completely to measurement of $\rho_{\text{DM}}$ coming from galaxy rotation curves [^6] This case is directly analogous to measurements of ultralight dark-matter couplings in ground- and space-based GW interferometers, discussed in Section III.2, where the arm lengths are negligible compared to $L_{\text{coh}}$, so all test masses experience the same ultralight dark-matter phase..
   3. 3.
-    “Pulsar-correlated” means that $L_{\text{coh}}$ is smaller than the radius of the galactic center sampled by experiments that derive galaxy rotation curves.
-However, estimates of the DM energy density are from different coherence patches and thus measurements averaged over different patches.
+    “Pulsar-correlated” means that $L_{\text{coh}}$ is larger than the earth-pulsar and inter-pulsar separations, so $\hat{\phi}_{P}^{2}=\hat{\phi}_{E}^{2}$, but smaller than the galactocentric radius sampled by galaxy rotation curves.
+In this case, pulsars and the Earth probe the same coherence patch, but the ultralight dark-matter amplitude measured with pulsars cannot be identified one-to-one with the DM energy density inferred from galaxy rotation curves, since the latter averages over many independent coherence patches.
 Each type of analysis optimally probes a different mass regime depending on the DM coherence length.
@@
 (2023).
-95% confidence-level upper limits from analyses of European pulsar timing array data on ultralight dark matter interacting only gravitationally and changing the times of arrivals of pulses in an analogous way as GWs do.
+95% confidence-level upper limits from analyses of European pulsar timing array data on ultralight dark matter interacting only gravitationally and changing the ToAs of pulses in an analogous way that GWs do.
+The physics behind this form of DM has been discussed in Section II.6.
 The panels show (a) the dimensionless strain amplitude $\Psi_{\text{c}}$ and (b) the fraction of the DM energy density that scalar DM could compose as a function of DM mass and frequency.
 $\rho_{\text{DM}}=0.4\text{GeV}/\text{cm}^{3}$.
-For comparison, upper limits on these quantities are shown from previous searches Porayko and Postnov (2014); Porayko et al.
+For comparison, upper limits on these quantities are shown from previous searches (PPTA Bayes 2018, NANOGrav 2018) Porayko and Postnov (2014); Porayko et al.
 (2018).
+Furthermore, the dashed line indicates the signal amplitude assuming that this form of DM comprises all of DM; thus, curves that lie above this dashed line do not represent physical constraints.
+Upper limits on the amplitude are shown for the three cases outlined in Section III.3.1.
 In the right panel, we zoom in on the excluded ultralight dark matter masses.
 The horizontal dotted line represents the scenario in which ultralight dark matter is responsible for all of the local DM energy density.
-Scalar ultralight dark-matter particles with mass $-24.0<\text{log}_{10}\text{~{}}(m_{\rm DM}/\text{eV})<-23.7$ can comprise at most $30-40$ of $\rho_{\text{DM}}$, while particles with masses $-23.7<\text{log}_{10}\text{~{}}(m_{\rm DM}/\text{eV})<-23.3$ could comprise up to $\sim 70$ of $\rho_{\text{DM}}$.
-Additionally, NANOGrav data can be used to constrain this kind of ultralight dark matter.
+The right panel shows that scalar ultralight dark-matter particles with mass $24.0<\text{log}_{10}\text{~}(m_{\rm DM}/\text{eV})<-23.7$ can comprise at most $30-40\%$ of $\rho_{\text{DM}}$, while particles with masses $23.7<\text{log}_{10}\text{~}(m_{\rm DM}/\text{eV})<-23.3$ could comprise up to $\sim 70$ % of $\rho_{\text{DM}}$.
+Additionally, NANOGrav data has been be used to constrain this kind of ultralight dark matter Afzal et al.
+(2023), though the constraints are weaker than from the EPTA.
+In particular, NANOGrav cannot yet probe a physical constraint for the fraction of DM energy density that this model of DM could compose, likely because their dataset only spans 15 years, while the European pulsar timing array collaboration analyzed almost 25 years of data.
+III.3.2 Constraints on scalar conformal dark matter
+While the previous subsection considered generic strain-like gravitational interactions of scalar ultralight dark matter, scalar conformal couplings represent a distinct class of models in which ULDM interacts through the trace of the energy–momentum tensor, As discussed in Section II.6.
+This model of DM can also be constrained with pulsar timing array data, and in particular Ref.
+Smarra et al.
+(2024) uses European pulsar timing array data to constrain both linear (FJBD-type) and quadratic (DEF-type) conformal couplings.
 In Fig.
-9, a similar constraint is shown on a quantity that relates to the fraction of the DM energy density that gravitationally-interacting DM could compose.
-$S_{\phi}$ is the spin of DM, which can be 0 or 1 in this model.
-NANOGrav writes the timing residuals for the $I^{\rm th}$ pulsar as:
-$h_{I}(t)$ $=\sum_{i}A_{E,I}^{i}(\bm{x}_{E})\sin(\omega t+\gamma_{E}^{i})$ (84)
-$+A_{P,I}^{i}(\bm{x}_{P,I})\sin(\omega t+\gamma_{P,I}^{i})\,,$ (85)
-where $i$ indicates the polarization ($=0$ for spin-0 and $=1,2,3$ for spin-1).
-This expression is divided into “earth” and “pulsar” terms with amplitudes and arbitrary phases of $A_{E,I}^{i}(\bm{x}_{E})$ and phase $\gamma_{E}^{i}$, and $A_{P,I}^{i}(\bm{x}_{P,I})$ and time-independent phase $\gamma_{P,I}^{i}$, respectively.
-These phases depend on the positions of the earth and the pulsar with respect to the ultralight dark-matter field (see Eq.
-60 and 61 of Afzal et al.
-(2023)).
-Like the EPTA, NANOGrav assumes a model of ultralight dark matter that induces metric perturbations and therefore changes in the times of arrivals of pulses Khmelnitsky and Rubakov (2014).
-The amplitude of the signal that is constrained is similar to Eq.
-64:
-$A_{E,I}^{i}(\bm{x}_{E})=A_{P,I}^{i}(\bm{x}_{E})=A_{\rm grav}=\left(2S_{\phi}+1\right)\displaystyle\frac{\pi G\rho_{\text{DM}}}{2m_{\rm DM}^{2}}\hat{\phi}^{2}(\bm{x})$ (86)
-where $S_{\phi}=0$ or 1 for scalar and vector DM, respectively.
-Fig.
-9 shows the NANOGrav constraints on this amplitude, which can be compared to the right-hand panel of Fig.
-8.
-NANOGrav cannot yet probe a physical constraint for the fraction of DM energy density that this model of DM could compose, likely because their dataset only spans 15 years, while the European pulsar timing array collaboration analyzed almost 25 years of data.
-Figure 9: Taken from Afzal et al.
+9, we show the upper limits on the linear coupling parameter $\alpha$ using European pulsar timing array data.
+However, despite showing upper limits (that assume no signal), Ref.
+Smarra et al.
+(2024) notes that there is some evidence for additional signal power at two masses: $m\sim 10^{-22.7}$ eV and $m\sim 10^{-21.4}$ eV, though other physical processes could be responsible for these excesses.
+It is worth noting that the effective change in the neutron-star moment of inertia in this scenario arises from the universal rescaling of the metric due to the conformal coupling, and is therefore distinct from the composition-dependent moment of inertia variations discussed in Section II.3, where scalar fields couple directly to quarks and nucleons.
+The constraints arising from the interaction of dilatons with standard-model particles in neutron stars will be discussed in the next subsection.
+Figure 9: Taken from Smarra et al.
 (2024).
-Upper limits on the local ULDM density for the correlated (solid line) and uncorrelated (dashed line) signals at the $95\%$ credible level using data from the NANOGrav pulsar timing array.
-The model of DM considered here is the same as that in Fig.
-8.
-The gray dashed line indicates the predicted DM abundance.
-III.4.2 Variation of fundamental constants constraints
-From the most recent NANOGrav dataset, constraints on ultralight dark matter that causes fundamental constants to oscillate Afzal et al.
-(2023) can be set as well for masses below $10^{-22}$ eV.
-In the case that ultralight dark matter causes changes in the moment of inertia of the pulsar, Fig.
-10 depicts constraints on the couplings of the scalar DM particle to up- and down- quarks weighted by their masses ($\hat{m})$, electrons $(d_{e})$, muons $(d_{\mu})$, photons $(d_{\gamma})$ and gluons $(d_{g})$.
-To derive these constraints, the simplest model of the equation-of-state of the neutron star is assumed, as outlined in Section II.3.2, and, when constraining one particular coupling constant, the rest are assumed to be zero.
+Constraints on the coupling $\log_{10}\alpha$ of conformal DM to gravity at the 95% confidence-level (solid line) for the correlated case compared to the analytical estimate obtained by using the upper limits from the European pulsar timing array search in Smarra et al.
+(2024) to estimate $\alpha$ (brown dashed lines).
+The physics behind this form of DM has been discussed in Section II.6.
+$\rho_{\text{DM}}=0.4\,\text{GeV/cm}^{3}$.
+Here, this model of DM would change the moment of inertia of neutron stars used in pulsar timing arrays, resulting in changes of the ToA of the pulses.
+If conformal DM does not comprise all of DM, the limits will weaken, which is shown by the dotted line, for values of $\rho$ given in Smarra et al.
+(2023).
+AP4 indicates that these limits are derived assuming the AP4 equation-of-state of neutron stars Akmal et al.
+(1998).
+III.3.3 Variation of fundamental constants constraints
+[…]
 The strongest constraints from pulsar timing arrays are on the couplings to the electron and muons.
-Experiments that directly use atomic clocks to search for ultralight dark matter are insensitive to the electron coupling, since changing the electron mass does not affect the spacing between energy levels Kaplan et al.
+This is in contrast to constraints that come from atomic clock experiments, which also aim to find ultralight dark matter but are insensitive to the electron coupling, since changing the electron mass does not affect the spacing between energy levels Kaplan et al.
 (2022).
@@
 40 or Eq.
-45.
+44.
 In other words, these constraints on ultralight dark matter coupling to standard-model particles are valid when the coupling of DM to gravity can be neglected, i.e.
@@
 In the gray-shaded region, the interaction between gravity and DM is stronger, and thus constraints on ultralight dark matter coupling to standard-model particles alone would not be valid in that region.
-Furthermore, these constraints assume that scalar particles constitute all of DM, which may not be true.
-If they only comprise a fraction $f$ of DM, then the constraints will be weakened by $\sqrt{f}$, because of the $\sqrt{\rho_{\text{DM}}}$ term in the amplitude of these interaction signals.
+Furthermore, these constraints assume that scalar particles constitute all of DM, but if they only comprise a fraction $f$ of DM, then the constraints will be weakened by $\sqrt{f}$, because of the $\sqrt{\rho_{\text{DM}}}$ term in the amplitude of these interaction signals.
 Figure 10: Taken from Afzal et al.
@@
 Upper limits from an analysis of NANOGrav data on the coupling of ultralight dark matter to various standard-model particles: quarks, electrons, muons, photons and gluons, shown in red for the correlated case, and red dashed for the uncorrelated case.
-In this model, DM would alter the moment of inertia of neutron stars and thus affect the times of arrivals of pulses in pulsar timing arrays.
+In this model, DM would alter the moment of inertia of neutron stars and thus affect the ToAs of pulses in pulsar timing arrays.
+The physics behind this form of DM has been discussed in Section II.3.2.
 The constraints on one coupling constant assume that all other coupling constants are set to 0.
@@
 (2020).
+The shape of these constraints is similar to those in Fig.
+11, since both searches look for nearly sinusoidal signals embedded in pulsar timing array data with similar noise characteristics, even though the underlying DM coupling mechanisms differ.
 The scalings of these constraints with the DM mass should be discussed.
-At high masses, the signal is essentially deterministic, and the signal-to-noise ratio scales as the square of the signal amplitude, which means that $d_{i}\propto m^{2}$.
-At low masses ($m<1/T_{\text{obs}})$, the signal is observed in different coherence patches, since the coherence length is shorter than the pulsar separation.
-The signal is no longer a sinusoid but can be expressed as a polynomial in $mt$.
-The first two terms of this polynomial ($mt$ and $(mt)^{2}$) are degenerated with pulsar timing terms, and thus the first observable term is $(mt)^{3}$, meaning that the signal-to-noise ratio is proportional to $m^{3}$.
-Thus, $d_{i}\propto 1/m$, which explains the scaling at low masses.
+At high masses, the signal is essentially deterministic, and the SNR scales as the square of the signal amplitude, which means that $d_{i}\propto m_{\rm DM}^{2}$.
+At low masses ($m_{\rm DM}<1/T_{\text{obs}})$, the signal is observed in different coherence patches, since the coherence length is shorter than the pulsar separation.
+The signal is no longer a sinusoid but can be expressed as a polynomial in $mt$, where $t$ is time.
+The first two terms of this polynomial ($m_{\rm DM}t$ and $(m_{\rm DM}t)^{2}$) are degenerated with pulsar timing terms, and thus the first observable term is $(m_{\rm DM}t)^{3}$, meaning that the SNR is proportional to $m_{\rm DM}^{3}$.
+Thus, $d_{i}\propto 1/m_{\rm DM}$, which explains the scaling at low masses.
 Still, though, red noise causes the constraints to flatten as the mass decreases, which, at first, hides this scaling.
-III.4.3 Constraints on vector dark matter
+III.3.4 Constraints on vector dark matter
 As discussed in Section II.4, DM could also be composed of spin-1 bosons that interact with standard model particles.
@@
 We can see that the NANOGrav constraints greatly surpass those that come from MICROSCOPE at very low masses.
-Furthermore, these constraints assume that vector bosons constitute all of DM, which may not be true.
-If they only comprise a fraction $f$ of DM, then the constraints will be weakened by $\sqrt{f}$, because of the $\sqrt{\rho_{\text{DM}}}$ term in the amplitude of these interaction signals.
+Furthermore, these constraints assume that vector bosons constitute all of DM, but if they only comprise a fraction $f$ of DM, then the constraints will be weakened by $\sqrt{f}$.
 Figure 11: Taken from Afzal et al.
 (2023).
-Upper limits from analyses of NANOGrav pulsar timing array data on the coupling of DM to baryon and baryon-lepton number for the correlated (solid lines) an uncorrelated (dashed line) cases.
-In this case, dark photon DM would couple to standard model particles in the earth and the pulsars themselves.
-The black dashed line, and the gray shaded region, indicates where the amplitude of the signal would be less than the amplitude of the signal arising from DM coupling only to gravity.
-The constraints from tests of the equivalence principle (“MICROSCOPE”) are shown in teal (Bergé et al., 2018), and the constraints previously set by the PPTA Collaboration are reported in yellow (Xue et al., 2022).
-The recent NANOGRAV evidence for a stochastic GW background is also consistent with a vector ultralight dark matter hypothesis.
-Ref.
+Upper limits from analyses of NANOGrav pulsar timing array data on the coupling of DM to baryon and baryon-lepton number, shown for the correlated (solid lines) and uncorrelated (dashed lines) cases.
+In this scenario, dark photon DM couples to standard model particles in both the Earth and the pulsars.
+The physics behind this form of DM has been discussed in Section II.4.
+The black dashed line and gray shaded region indicate where the amplitude of the signal would be less than that from gravitational coupling alone.
+Constraints from equivalence principle tests (“MICROSCOPE”) are shown in teal (Bergé et al., 2018), and previous limits from the PPTA Collaboration are shown in yellow (Xue et al., 2022).
+The similarity in shape between the curves in this figure and those in Fig.
+10 arises because both searches target nearly sinusoidal signals embedded in pulsar timing array data with similar noise characteristics.
+Although the underlying DM models differ—scalar versus vector, moment of inertia changes versus direct forces—the observable effect in each case is the same: modulations in pulse ToAs.
+The coupling strength to standard-model particles determines the amplitude of these signals.
+Furthermore, recent evidence for a stochastic GW background reported by NANOGrav has also been interpreted in terms of vector ultralight dark matter.
+In particular, Ref.
 Chowdhury et al.
-(2024) has found that ultralight dark matter that couples to muons ($L_{\mu}-L_{\tau})$ is actually more consistent with NANOGRAV’s observations than a stochastic GW background.
-This type of ultralight dark matter would cause the pulsars to oscillate (because they are composed of muons) and induce a spin-dependent frequency shift in the pulses that would change the angular correlations with respect to the stochastic GW background case (because of the gravitational potential of the ultralight dark matter).
-The mass and coupling constants of this ultralight dark matter are not ruled out by any other experiment, and lie in the region
-III.4.4 Constraints on conformal DM
-DM that conformally (universally) couples to gravity could alter the moment of inertia of rotating pulsars via a new gravity-mediated force between neutron stars and ultralight dark matter Kuntz and Barausse (2024), which is called the Nordtvedt effect Nordtvedt (1968).
-If such a coupling exists, the times of arrivals of pulsars would change in a deterministic way, and would also result in a violation of the equivalence principle.
-This model of conformal DM has been extensively constrained by looking for impacts on the timing solutions of measured binary pulsar Damour and Taylor (1992); Kramer et al.
-(2006).
-Ref.
-Smarra et al.
-(2024) uses European pulsar timing array data to constrain linear (Fierz-Jordan-Brans-Dicke (FJBD) theory) Fierz (1956); Jordan (1959); Brans and Dicke (1961); Dicke (1962) and quadratic (Damour-Esposito-Farese (DEF) theory) Damour and Esposito-Farese (1992, 1993) conformal DM couplings to gravity by deriving how the pulsar times of arrivals would change in the presence of the ultralight dark matter fields.
-In the case of FJBD conformal coupling, the residuals would appear as:
-$\Delta t(t)$ $=-\int\mathrm{d}t\frac{\delta\Omega_{\text{obs}}}{\bar{\Omega}}$
-$=2\alpha s_{I}\left.\frac{\sqrt{\rho_{\text{DM}}}}{M_{\text{Pl}}m^{2}}\hat{\phi}(\bm{x})\sin(m_{\rm DM}t+\theta(\bm{x}))\right|_{t_{\text{start}}-\frac{d}{c}}^{t_{\text{end }}-\frac{d}{c}},$ (88)
-where $s_{I}$ is the angular moment sensitivity parameter computed in Kuntz and Barausse (2024) and is of $\mathcal{O}(1)$ in FJBD theory, $\alpha$ is the strength of the coupling, and $\hat{\phi}$ is a stochastic parameter whose value depends on whether the correlated, uncorrelated or pulsar correlated scenarios are considered.
-$\theta(\mathbf{x})$ is a random phase.
-In Fig.
-12, we show the upper limits on the linear coupling parameter $\alpha$ using European pulsar timing array data.
-However, Ref.
-Smarra et al.
-(2024) notes that there is some evidence for additional signal power at two masses: $m\sim 10^{-22.7}$ eV and $m\sim 10^{-21.4}$ eV, but note that other physical processes could responsible for these excesses.
-Figure 12: Taken from Smarra et al.
-(2023).
-Constraints on the coupling $\log_{10}\alpha$ of conformal DM to gravity at the 95% confidence-level (solid line) for the correlated case compared to the analytical estimate obtained by using the upper limits from the European pulsar timing array search in Smarra et al.
-(2023) to estimate $\alpha$ (brown dashed lines).
-$\rho_{\text{DM}}=0.4\,\text{GeV/cm}^{3}$.
-Here, this model of DM would change the moment of inertia of neutron stars in the used in pulsar timing arrays, resulting in changes of the times of arrival of the pulses.
-If the true density of DM that conformal DM composes is less than $\rho_{\text{DM}}$, the limits will suffer, which is shown by the dotted line, for values of $\rho$ given in Smarra et al.
-(2023).
-AP4 is a particular equation-of-state of neutron stars Akmal et al.
-(1998).
-III.4.5 Constraints on axions through polarization measurements
-Changes in polarization of light by axions could also be probed in pulsar timing array experiments Xue et al.
+(2024) showed that a vector ultralight dark matter model coupled to muons ($L_{\mu}-L_{\tau}$) can provide an even better fit to NANOGrav’s observations than a conventional stochastic GW background.
+In this scenario, pulsars undergo oscillations due to their muon content, leading to a spin-dependent frequency shift in the pulses.
+This produces angular correlations that differ from the standard Hellings–Downs curve expected for a stochastic GW background, thereby offering a way to test the vector ultralight dark-matter hypothesis against future pulsar timing array data.
+III.3.5 Constraints on axions through polarization measurements
+As described in Section II.2, axion-like dark matter induces an oscillatory modulation of the polarization plane of linearly polarized light—cosmic birefringence—via its coupling to photons.
+Recent analyses of European pulsar timing array data, particularly from the Parkes Pulsar Polarization Array (PPPA), have performed the first dedicated searches for these axion-induced polarization oscillations Xue et al.
 (2024); Porayko et al.
 (2025).
-Just as in ground-based interferometers, axions could induce cosmic birefrigence, an oscillation of the polarization plane of a linearly polarized source.
-Since pulsars constantly emit light that is beamed at earth, its polarization can be measured over time, and any changes in it could be attributed to axion-like DM/photon interactions.
-In Fig.
-13, we show upper limits on this coupling from the Parkes Pulsar Polarization Array (PPPA), the first search of its kind.
-The constraints here outperform existing experiments at low axion masses.
+By tracking polarization angles over long time baselines, these searches place upper limits on the axion-photon coupling constant, $g_{a\gamma}$, for axion masses in the ultra-light regime.
+Fig.
+12 shows the resulting constraints from this analysis.
+The red and blue curves correspond to different statistical treatments of the data, including models that account for spatial correlations between pulsars.
+Notably, these PTA constraints improve upon previous laboratory and astrophysical bounds at very low axion masses (below approximately $10^{-22}$ eV), where the long coherence time of the axion field aligns well with the decade-scale observation windows of PTAs.
+These results highlight the complementary sensitivity of PTAs compared to ground-based interferometers, which probe higher axion masses but with comparable constraints on $g_{a\gamma}$.
+Together, they cover a broad mass range, making multi-messenger searches a powerful strategy for testing axion dark matter scenarios.
 Figure 12: Taken from Xue et al.
 (2024).
-Upper limits from analyses of European pulsar timing array data on the coupling of axion-like DM to photons.$\rho_{\text{DM}}=0.4\,\textrm{GeV}/\textrm{cm}^{3}$.
+Upper limits from analyses of European pulsar timing array data on the coupling of axion-like DM to photons.
+$\rho_{\text{DM}}=0.4\,\textrm{GeV}/\textrm{cm}^{3}$.
+The physics behind this form of DM has been discussed in Section II.2.
 The blue and red curves correspond to the limits derived using the auto-correlation-only and full-correlation signal models, respectively.
@@
 Key timescales are shown with vertical dotted lines, included $T_{\text{obs}}=18$ years.
-The lower panel shows the Bayes factors for each of the scenarios.
-III.5 Prospects for detecting the coupling of ultralight dark matter to the interferometers
+Smoothed constraints are shown as dashed lines in cyan, yellow, green, and purple, representing limits from polarization angle (PA) variations observed in the Crab Nebula by POLARBEAR Adachi et al.
+(2024), analyses of PPTA DR1 and Crab Nebula data from QUIJOTE Castillo et al.
+(2022), measurements of CMB polarization by SPT-3G Ferguson et al.
+(2022), and studies of axion-like DM-induced washout effects during recombination from the Planck mission Fedderke et al.
+(2019), respectively.
+Additional constraints from the CAST experiment Anastassopoulos et al.
+(2017), as well as observations of SN1987A Payez et al.
+(2015) and X-ray spectral distortions in quasar H1821+643 Reynés et al.
+(2021), are shown as black dashed lines.
+The lower panel shows the Bayes factors (BFs, essentially, the preference for DM to be in the data versus noise) for each of the scenarios.
+The BF stays around zero, although some significant spikes are apparent, which has been found to be due to imprecise noise modelling, not astrophysical signals.
+III.4 Prospects for detecting the coupling of ultralight DM to GW interferometers
 Though no searches so far have yielded conclusive evidence of any DM/standard model coupling with the interferometers, impressive constraints, relative to existing DM experiments, have been set on these interaction models.
 The current searches and their constraints have been framed only in terms of upper limits on dilatons, dark photons and DM that couples to gravity; however, they are actually sensitive to any kind of DM interaction that would cause a differential strain in the interferometers.
-In fact, to derive these constraints, we first compute the minimum detectable signal amplitude for a generic quasi-monochromatic signal as a function of frequency, and then map these values to constraints on the DM/standard model coupling constants.
-It is thus simply a matter of a few relations that these results are interpreted in terms of two relatively well-motivated DM models.
+In fact, to derive these constraints, we first compute the minimum detectable signal amplitude for a generic quasi-monochromatic signal as a function of frequency, as outlined in Section III.1, and then map these values to constraints on the DM/standard model coupling constants.
+It is thus simply a choice that these results are interpreted in terms of two relatively well-motivated DM models.
 Of course, ultralight dark matter could take on any mass a priori, and ground-based GW detectors only allow us to probe a small region of this parameter space, from $\sim 10^{-14}-10^{-12}$ eV.
@@
 (2021).
-However, it should be noted that when placing constraints on ultralight dark matter, the stochasisity of the signal could affect the limits: if we happen to “get unlucky” and observe the signal when the phase offset reduces the signal amplitude to zero, as can be seen at the zero points in Fig.
-2, we would have no constraint.
-Such considerations were made in Nakatsuka et al.
+However, it should be noted that when placing constraints on ultralight dark matter, the stochasticity of the signal could affect the limits: if we happen to “get unlucky” and observe the signal when the phase offset reduces the signal amplitude to zero, as can be seen at the zero points in Fig.
+3, we would have no constraint.
+Refs.
+Nakatsuka et al.
 (2023); Boddy et al.
-(2025) for ground-based interferometers and pulsar timing arrays.
+(2025) have shown that for ground- and space-based interferometers, that this effect has an $\mathcal{O}(1)$ impact on upper limits, but needs to be mitigated for pulsar timing arrays by averaging over the signals coming from many pulsars.
 The prospects for future ground-based GW interferometers, such as ET and Cosmic Explorer, to observe ultralight dark matter interactions are bright.
-Within a decade or so of observation, such detectors, currently a couple of orders of magnitude behind fifth-force constraints, could actually be competitive in this portion of the parameter space, thus permitting an extremely strong probe of new physics Pierce et al.
+Within a decade or so of observation, such detectors could actually be competitive at masses between $10^{-15},10^{-14}]$ eV, thus permitting an extremely strong probe of new physics Pierce et al.
 (2018); Morisaki et al.
@@
 Such a caveat is balanced by the relatively model agnostic nature of our searches, which make minimal assumptions about the signal model, and essentially just look for correlated noise.
-While we probably cannot believe just one model for DM/standard model interactions, we can be sure that our techniques are sensitive to a wide range of theories, even ones that have yet to be thought of.
-IV Macroscopic dark matter transiting through the interferometers
-IV.1 Background
-As mentioned in Section I, the mass of DM can span many orders of magnitude, including being “macroscopic”, that is, from being $M_{\rm DM}\sim[1,10^{9}]$ kg Hall et al.
+While we probably cannot believe any one model for DM/standard model interactions, we can be sure that our techniques are sensitive to a wide range of theories, even ones that have yet to be thought of.
+IV Macroscopic DM transiting through GW interferometers
+As mentioned in Section I, the mass of DM could span many orders of magnitude.
+In particular, macroscopic dark matter, with masses in the range $M_{\rm DM}\sim[1,10^{9}]$ kg, has been proposed as a viable candidate Hall et al.
 (2018).
-If DM exists in this mass range, it could transit through GW interferometers, and thus be searched for in LIGO, Virgo and KAGRA data.
-The event rate, however, depends on the mass of DM, and there are orders of magnitude fewer heavy DM particles in any given region of space than the number of ultralight dark-matter particles Hall et al.
-(2018):
-$\rho_{\text{DM}}=M_{\text{DM}}n_{\text{DM}}~{}~{}\Longrightarrow~{}~{}\frac{L_{\rm DM}}{10^{4}~{}\text{km}}\simeq 1.2\times\left(\frac{M_{\text{DM}}}{1~{}\text{kg}}\right)^{1/3},$ (89)
-where $n_{\rm DM}$ is the number density of DM, and $L_{\rm DM}$ is a characteristic length scale for the transiting DM.
-For 1-kg DM, the flux of DM through a given region of space is given by $\Phi\sim n_{\rm DM}v_{0}\sim 3\times 10^{-10}$ km-2 s-1, meaning that approximately one transient event could be observed per year if the impact parameter of such an event is 10 km, which is similar in size to the three or four kilometer-long arms of LIGO, Virgo and KAGRA.
-IV.2 Yukawa-like interactions
-If DM only couples to gravity, it seems unlikely that GW interferometers could detect this interaction Hall et al.
+These candidates could include composite objects such as nuclear-density nuggets Itoh (1970); Witten (1984) or Q-balls Coleman (1985),
+If such objects were to transit through or near a GW interferometer, they could exert measurable gravitational forces on the test masses.
+In most models, the dominant interaction would be gravitational: the passage of a massive, compact object would cause a time-varying, localized gravitational acceleration of the test masses, leading to a transient displacement signal in the detector.
+The resulting signature would be a short-duration, impulsive signal, potentially localized in time and space depending on the trajectory of the DM object.
+Unlike continuous-wave or stochastic DM signals discussed in Section II, this effect would appear as a transient, with a characteristic waveform determined by the mass, size, velocity, and trajectory of the object.
+The event rate of such transits is inversely proportional to the mass of the dark matter particles, and the number density $n_{\rm DM}$ depends on $\rho_{\text{DM}}$:
+$\rho_{\text{DM}}=M_{\rm DM}\cdot n_{\text{DM}},$ (79)
+This equation implies that the characteristic length scale $L_{\rm DM}$ associated with such objects (assuming constant density), i.e. the distance between DM objects, is given by:
+$\frac{L_{\rm DM}}{10^{4}~\text{km}}\simeq 1.2\times\left(\frac{M_{\text{DM}}}{1~\text{kg}}\right)^{1/3}.$ (80)
+For $M_{\rm DM}=1$ kg, this corresponds to a flux of $\Phi\sim n_{\rm DM}v_{0}\sim 3\times 10^{-10}$ km-2 s-1, assuming a typical DM velocity $v_{0}\sim 220$ km/s.
+Moreover, the size of DM, $r_{\rm DM}$, would be tiny with respect to $L_{\rm DM}$: $r_{\rm DM}\sim(M_{\rm DM}/(\frac{4}{3}\pi\rho_{\rm nugget}))^{1/3}\sim 3\mu\text{m}$ for $\rho_{\rm nugget}\sim 10^{17}\text{ kg/m}^{3}$.
+This implies that one transient event per year would be expected in a detector with an effective cross-sectional area777the area in the plane perpendicular to the DM flux within which a passing object would produce a transient of order a few hundred km2, i.e.
+DM would have an impact parameter of $\sim 10$ km, which is of the same order as the length of the interferometer arms (4 km).
+However, gravitationally interacting DM would not be able to produce such an event rate Hall et al.
 (2018); Du et al.
-(2023); however, if there is a Yukawa-like coupling between the standard model and DM, allowing for self-interactions as well, DM transiting through the interferometers could very well be detectable.
-The following potential $V_{i-j}$ is assumed to govern the (self-) interactions between the standard model and DM
+(2023), so we will need to consider interactions of DM with the standard model to enhance the event rate.
+In Section IV.1, we will discuss the physics of Yukawa-like interactions that macroscopic DM can have with the standard model, which will enhance the strength of the transiting DM object.
+Then, in Section IV.2, we show the projected constraints on Yukawa-like interactions using future data from ground-based GW interferometers.
+IV.1 Yukawa-like interactions
+If there is a Yukawa-like coupling between the standard model and DM, allowing for self-interactions as well, DM transiting through the interferometers could be detectable.
+The interaction potential between DM and standard-model particles can be described by the following form:
 $V_{i-j}=-M_{i}M_{j}\frac{G}{r}\Big{(}1+(-1)^{s}~\delta_{i}\delta_{j}\exp[-r/\lambda]\Big{)}$ (81)
-[…]
-where $\delta_{i,j}$ parametrizes the standard model and DM couplings, $r$ is a length scale, $\lambda$ is the screening length, $s$ is the spin, and $M_{i,j}$ are the two masses.
-By assuming that the size of DM is much larger than the size of GW detectors and the range of the DM force, but smaller than the average distance between the detectors, the scattering cross section from this process can be derived:
+$\text{where }~i,j=\text{SM,DM.}$~~~~~~~~~~~~
+where $\delta_{i,j}$ represents the coupling strengths between particles $i,j$, and $i,j$ could be either standard-model or DM particles.
+$r$ is a length scale, $\lambda$ is the screening length, $s$ is the spin of the interaction mediator, and $M_{i,j}$ are the two masses.
+By noting that the physical size of the detector and range of the interaction force are much larger than $r_{\rm DM}$, but much smaller than the separation between DM particles $L_{\rm DM}$, the interaction can be modeled in terms of an effective geometric cross section and derived using the Born approximation Born (1926):
+$\frac{d\sigma}{d\Omega}=\frac{M_{\rm DM}^{2}}{4\pi^{2}}\left|\int V(\vec{r})e^{i(\vec{k}_{i}-\vec{k}_{f})\cdot\vec{r}}d^{3}\vec{r}\right|^{2}=\frac{M_{\rm DM}^{2}}{4\pi^{2}}\left|\tilde{V}(\vec{k}_{i}-\vec{k}_{f})\right|^{2}$ (82)
+where $\vec{k}_{i}-\vec{k}_{f}$ is the momentum transfer between the initial and final states.
+The integral is simply the Fourier transform of the potential in Eq.
+81 (considering only the interaction term):
+$\tilde{V}=\frac{4\pi GM_{\rm DM}^{2}\delta_{\rm DM}^{2}}{(M_{\rm DM}v_{0}\sin({\theta/2}))^{2}+\lambda^{-2}},$ (83)
+where $\theta$ is the scattering angle, i.e.
+the angle between the incoming and outgoing momenta $\vec{k}_{i}$ and $\vec{k}_{f}$, and $\vec{k}_{i}$ and $\vec{k}_{f}$ have been written in terms of $v_{0}$ and $M_{\rm DM}$.
+Plugging in Eq.
+81 into Eq.
+82 and integrating over solid angle, the cross-section can be derived:
 $\sigma_{\text{DM-DM}}=16\,\pi\times\frac{G^{2}\,M_{\text{DM}}^{2}\,\delta_{\text{DM}}^{4}}{v_{0}^{4}}\times\log\left[\frac{\lambda}{r_{\text{DM}}}\right],$ (84)
-where $r_{\rm DM}$ is the size of the DM objects.
 While $|\delta_{\rm SM}|<5\times 10^{-4}$ has been strongly constrained by equivalence principle experiments Adelberger et al.
@@
 (2015).
-The lack of observed deceleration of DM in the bullet cluster collision constrained the cross-section of DM to be less than $\sim 1$ cm2/g Randall et al.
+In particular, the lack of observed deceleration of DM in the bullet cluster collision constrained the cross-section of DM to be less than $\sim 1$ cm2/g Randall et al.
 (2008), which provides limits on the coupling constant of self-interacting DM to be:
 $|\delta_{\text{DM}}|\lesssim 5\times 10^{9}\times\left(\frac{1~\text{kg}}{M_{\text{DM}}}\right)^{1/4},$ (85)
-in which $\log\left[\frac{\lambda}{r_{\text{DM}}}\right]=5$.
-Based on the current upper limits on $\delta_{\rm SM}$ and $\delta_{\rm DM}$, the coupling of standard-model particles to DM ones could be quite large, and would thus permit an observation of a transit event in GW interferometers.
+in which $\log\left[\frac{\lambda}{r_{\text{DM}}}\right]=5$, chosen only as a benchmark, corresponding to a force range much larger $r_{\rm DM}$.
+Based on the current upper limits on $\delta_{\rm SM}$ and $\delta_{\rm DM}$, could allow an observation of a transit event in GW interferometers.
 Furthermore, $\delta_{\rm DM}\gg$ 1 would actually alleviate some problems with the cold DM scenario, including overly dense areas of dwarf galaxies found in simulations, and a prediction that the Milky Way produces more stars than are observed Boylan-Kolchin et al.
 (2011).
-IV.3 Projected constraints on Yukawa-like interactions
-By using the potential given in Eq. 90, the differential acceleration visible on interferometers can be computed, and from that, the matched-filter signal-to-noise ratio:
-[…]
-where $a(f)$ is the Fourier transform of the integral of the differential acceleration $a(t)=a_{y}^{\rm EM}-a_{y}^{\rm IM}-(a_{x}^{\rm EM}-a_{x}^{\rm IM})$ between the input mirror (IM) and the end mirror (EM) Saulson (2017), and $S_{n}(f)$ is the acceleration power spectral density of the noise.
-By using sensitivity curves of LIGO, Virgo and KAGRA and a threshold on $\rho>8$, the number of detections per year of transiting DM with different masses and different screening lengths can be computed, and is shown in Fig.
-14 using a power spectral density for advanced LIGO at design sensitivity.
+IV.2 Projected constraints on Yukawa-like interactions
+In order to search for transiting macroscopic DM, it has been proposed in Hall et al. (2018) to use matched filtering. To do so, firstly we must calculate the matched filtering SNR $\rho$:
+$\rho^{2}=4\int\frac{|a(f)|^{2}}{S_{n}(f)}\,df;$ (86)
+Here, $a(f)$ is the Fourier transform of the differential acceleration $a(t)$ between the test masses, and $S_{n}(f)$ is the power spectral density of the detector.
+The differential acceleration $a(t)$ arises from the DM transiting through the interferometer test masses, and is calculated with the difference in accelerations between the input mirror and end mirror Saulson (2017) in each arm.
+The form of $a(t)$ can be derived by taking the gradient of Eq.
+81:
+$\vec{a}(r)=-\frac{GM_{\rm DM}}{r^{2}}\left[1+(-1)^{s}\,\delta_{\rm DM}\,\delta_{\rm SM}\,e^{-r/\lambda}\left(1+\frac{r}{\lambda}\right)\right]\hat{r}$, (87)
+and then the individual accelerations along the $x$ and $y$ arms can be computed as:
+$\Delta a_{x}$ $=a(x=L_{x},0,0)-a(0,0,0)$ (88)
+$\Delta a_{y}$ $=a(0,0=L_{y},0)-a(0,0,0),$ (89)
+$\Delta a_{x}$ $=-GM_{\rm DM}\Bigg{[}\frac{L_{x}-x_{\rm DM}}{r_{1}^{3/2}}-\frac{-x_{\rm DM}}{r_{\rm DM}^{3/2}}\Bigg{]}$
+$\times\Bigg{[}1+(-1)^{s}\,\delta_{\rm DM}\,\delta_{\rm SM}\,\exp\Big{(}-\frac{r_{1}}{\lambda}\Big{)}\Big{(}1+\frac{r_{1}}{\lambda}\Big{)}\Bigg{]},$ (90)
+$\Delta a_{y}$ $=-GM_{\rm DM}\Bigg{[}\frac{L_{y}-y_{\rm DM}}{r_{2}^{3/2}}-\frac{-y_{\rm DM}}{r_{\rm DM}^{3/2}}\Bigg{]}$
+$\times\ \Bigg{[}1+(-1)^{s}\,\delta_{\rm DM}\,\delta_{\rm SM}\,\exp\Big{(}-\frac{r_{2}}{\lambda}\Big{)}\Big{(}1+\frac{r_{2}}{\lambda}\Big{)}\Bigg{]},$ (91)
+where
+$r_{1}$ $=\sqrt{(L_{x}-x_{\rm DM})^{2}+y_{\rm DM}^{2}+z_{\rm DM}^{2}},$ (92)
+$r_{2}$ $=\sqrt{x_{\rm DM}^{2}+(L_{y}-y_{\rm DM})^{2}+z_{\rm DM}^{2}},$ (93)
+$r_{\rm DM}$ $=\sqrt{x_{\rm DM}^{2}+y_{\rm DM}^{2}+z_{\rm DM}^{2}},$ (94)
+and $x_{\rm DM}$, $y_{\rm DM}$ and $z_{\rm DM}$ are the spatial coordinates of the DM object.
+The acceleration depends directly on the coupling of DM to standard model particles; thus, the SNR will be proportional to this coupling as well.
+In searches with GW data, a threshold must be set that delineates candidates that are most likely noise with those that could be of astrophysical origin.
+To estimate the number of detectable transiting DM events, an SNR threshold of $\rho>8$ is applied, which is standard in GW searches Maggiore (2008), and the detection rate as a function of DM mass, and coupling $g\equiv\delta_{\rm SM}\delta_{\rm DM}$ is computed:
+$\mathcal{R}\sim\Phi\sigma_{\rm DM-SM}A$ (95)
+where $A$ is the effective area of the detector through which DM transits.
+The effective area for macroscopic DM transits is determined by the physical area of the detector, the angle of incidence of the DM on the detector, and the interaction cross-section between DM and the interferometer.
+For larger cross sections (larger couplings), the SNR of the signal will increase.
+Thus, the SNR threshold determines a minimum detectable cross section, which in turns fixes the rate of transiting DM:
+We show potential constraints on the rates of DM transiting through advanced LIGO at design sensitivity in Fig.
+13.
+These results demonstrate that macroscopic DM can be detectable, depending on the parameters of the interaction.
 Figure 13: Taken from Hall et al.
 (2018).
-Expected number of macroscopic DM transit events per year with $\rho>8$ as measurable with advanced LIGO operating at design sensitivity, as a function of the DM mass, the screening length $\lambda$ and the coupling $g=\delta_{\text{SM}}\delta_{\text{DM}}$ between DM and the standard model.
-Further work was done regarding the detection of macroscopic DM in much more detail Du et al.
-(2023), in which the different kinds of effects of DM on the test masses were derived.
-Ref.
+Expected number of macroscopic DM transit events per year detectable by advanced LIGO at design sensitivity, as a function of DM mass, screening length $\lambda$, and coupling $g=\delta_{\rm SM}\delta_{\rm DM}$.
+The physics behind this form of DM has been discussed in Section IV.1.
+The blue, red and green colors indicate different screening lengths $\lambda$ of the Yukawa-like force.
+The three panels show the rate of transits for DM objects with three different masses.
+The rate is calculated using Eq.
+95 and the preceding equations, assuming a detection threshold on the SNR of 8.
+The threshold fixes the minimum detectable coupling accessible to a matched filtering search.
+Further work in Du et al.
+(2023) has expanded on this analysis, incorporating additional effects of DM transits, such as the Doppler effect, Shapiro delay, and Einstein delay, each of which contributes to the strain observed in the interferometer.
+These effects arise from the interaction of the DM object with the interferometer’s test masses, leading to small deviations in the expected strain signal.
+  1. 1.
+    Doppler effect: As the DM object moves relative to the detector, its velocity induces a periodic shift in the frequency of the signal received by the interferometer.
+This is similar to the classical Doppler shift seen in sound or light waves, but in this case, it is gravitational.
+The Doppler effect results in a frequency-dependent modulation of the strain signal, with the amplitude of the strain varying as the velocity of the DM object changes.
+  2. 2.
+    Shapiro delay: This effect arises due to the curving of spacetime near massive objects.
+In particular, light traveling near a massive object, such as a transiting DM object, will experience a delay in arrival time.
+This delay would be observed as a time shift in the signal received by the detectors, as the light takes longer to travel through the curved spacetime near the DM object.
+  3. 3.
+    Einstein delay: Similar to the Shapiro delay, the Einstein delay occurs due to the effects of gravitational time dilation.
+Time is slower near more massive objects.
+Thus, the passage of a massive DM object through the interferometer’s arms would result in a slight delay in the signal, since the clocks at the test masses would tick slower as the object passes by.
+While these gravitational effects are real, they are typically very weak unless DM interacts with standard-model particles.
+To make the effects measurable by GW interferometers, Ref.
 Du et al.
-(2023) derives the strain on the detector in terms of three observables: (1) the Doppler effect (previously considered in Hall et al.
-(2018)), (2) the Shapiro delay (light is curved near objects with mass and thus will take a longer time to travel down the interferometer arms, and (3) the Einstein Delay (time runs slower near more massive objects than less massive ones).
-Unfortunately, all effects are extremely weak because they are solely gravitational, and thus DM would have to couple to the standard model in some way to be detectable.
-As in Hall et al.
-(2018), Ref.
-Du et al.
-(2023) computes the expected constraints on the effective coupling of DM to the standard model with an effective parameter $\tilde{\alpha}\sim\delta_{i}\delta_{j}$, which are shown in Fig.
-15 for a variety of GW detectors and for two choices of screening lengths: $\lambda=1$ m and $\lambda=10^{6}$ m.
-Existing constraints from the bullet cluster Spergel and Steinhardt (2000), MICROSCOPE Bergé et al.
-(2018) and neutron-star kinetic heating with and without interactions between DM and baryons that could enhance energy transfer Gresham et al.
+(2023) confirmed that DM must couple with standard-model particles, agreeeing with the findings of Hall et al.
 (2018).
-While the constraints that could be obtained from GW interferometers do not appear to beat existing ones, it should be noted that if only a part of DM can be explained by a new fifth force, the Bullet cluster constraints on self-interacting DM disappear, while GW bounds fall off linearly with this fraction of DM charged under this fifth force Du et al.
+The strength of this coupling is given in Du et al.
+(2023) by the effective parameter $\tilde{\alpha}\sim\delta_{\rm SM}\delta_{\rm DM}$, where $\tilde{\alpha}=g$ in Hall et al.
 (2018).
+Thus, detecting these weak gravitational signals requires a sufficiently large value of $\tilde{\alpha}$, which can be constrained by observations of DM transits.
+Fig.
+14 shows the projected 90% sensitivity on $\tilde{\alpha}$, which can be probed by GW detectors for macroscopic DM.
+The results, shown for two different screening lengths ($\lambda=1\,\text{m}$ and $10^{6}\,\text{m}$), are compared to existing constraints from the Bullet Cluster Spergel and Steinhardt (2000), MICROSCOPE Bergé et al.
+(2018), and neutron-star kinetic heating experiments Gresham et al.
+(2023).
+While GW detectors may not surpass these existing constraints, they provide complementary information, especially if part of the DM population interacts with the standard model via a new fifth force.
 Figure 14: Taken from Du et al.
 (2023).
-90%-projected sensitivity on the standard model/DM interaction parameter $\tilde{\alpha}\sim\delta_{\rm SM}\delta_{\rm DM}$ that can be probed from macroscopic DM transiting through several GW experiments, compared to existing constraints from others.
-One year of observation is assumed, and two choices for the screening length are made.
-Dotted lines denote existing constraints.
-Existing constraints are shown in dotted lines.
-In particular, Bullet cluster and MICROSCOPE constraints are shown in gray Spergel and Steinhardt (2000); Bergé et al.
-(2018), MICROSCOPE ones.
-The purple and red lines denote constraints from neutron star kinetic heating Gresham et al.
-(2023) with and without additional short-range interactions between baryons and DM that would enhance energy transfer.
-Given that such deterministic signals may not be detectable in LIGO, Virgo and KAGRA, it was worth considering whether a stochastic background arising from many transiting DM events could be seen Hall et al.
+Projected sensitivity at 90% confidence to the standard-model/DM interaction parameter $\tilde{\alpha}\sim\delta_{\rm SM}\delta_{\rm DM}$ from macroscopic DM transits through several GW detectors.
+The physics behind this form of DM has been discussed in Section IV.1.
+The figure compares the projected sensitivity to existing constraints, including those from the Bullet Cluster, MICROSCOPE, and neutron-star kinetic heating, for two choices of screening length.
+Here, $T_{\text{obs}}=1$, and $\lambda=1$ m and $\lambda=10^{6}$ m, and an SNR threshold of 2 are assumed.
+The gray line represents combined limits from Bullet Cluster observations Spergel and Steinhardt (2000); Kahlhoefer et al.
+(2014) and the MICROSCOPE experiment Bergé et al.
+(2018); Fayet (2019).
+Purple and red lines indicate bounds from neutron star kinetic heating Gresham et al.
+(2023), with and without additional short-range DM–baryon interactions, respectively.
+While deterministic signals might be challenging to detect due to the weak gravitational interactions, there is also potential for detecting a stochastic background arising from numerous transiting DM events.
+This stochastic background could be visible if the Yukawa interaction is sufficiently strong, as explored in Hall et al.
 (2018); Du et al.
 (2023).
-However, again, it is necessary to include the Yukawa interaction to have a chance of detecting a stochastic background.
-Finally, another study Lee et al.
-(2023) considered specifically how heavy transiting DM would appear in KAGRA, as well as the directional dependence of the signal-to-noise ratio, which turns out to be strong due to the geometry of the detector.
-In KAGRA, it is shown that the suspension system can be particularly sensitive to DM-standard model interactions, between [1,130] Hz.
-V GWs from ultralight bosonic clouds around rotating BHs
-If a new ultralight boson [^5] In the context of boson clouds around rotating black holes, this new particle could, but need not, be DM exist, it will have a particular mass, frequency and Compton wavelength.
-Quantum fluctuations could then allow bosons to pop into existence near a rotating BH, whose characteristic length scale is given by half of its Schwarzschild radius.
-Some bosons will fall in; others will scatter off, and the extent to which either of these scenarios happens depends on the properties of the black hole and the boson.
-If the following superradiance condition, relating the BH angular frequency $\Omega$ to the boson angular frequency $\omega$ via the magnetic quantum number $m$, is met Brito et al.
+Future work will focus on whether a stochastic signal could be distinguished from the background noise in upcoming GW observations.
+V GWs from ultralight boson clouds around rotating black holes
+If new ultralight bosons [^8] In the context of boson clouds around rotating black holes, this new particle could, but need not, be DM exist, quantum fluctuations could allow them to pop into existence near a rotating BH.
+Some bosons will fall in; others will scatter off, and the extent to which either of these scenarios happens depends on the mass and spin of both the black hole and the new boson.
+If the following so-called “superradiance condition”, which relates the BH angular frequency $\Omega$ to the boson angular frequency $\omega$ via the magnetic quantum number $m$, is met Brito et al.
 (2015):
@@
 the boson will extract energy from the BH, and thus the “bosonic wave” will be amplified.
-This amplification will be maximized if the Compton wavelength of this new particle comparable to the characteristic length of a BH.
-The massive bosons then become bound to the BH, allowing for successive scatterings, thus permitting a huge number of bosons to appear around the BH as a cloud in a given energy state (no limit on the occupation number of these particles exists a priori)..
+This amplification will be maximized if the Compton wavelength of this new particle is comparable to the radius of a BH.
+The massive bosons then become bound to the BH because they sit in the potential well of the BH, allowing for successive scatterings, thus permitting a huge number of bosons to appear around the BH as a cloud in a given energy state (no limit on the occupation number of these particles exists).
 This process is called “superradiance”.
 More quantitatively, solving the Klein-Gordon equation for a Kerr BH hole in the presence of a massive scalar field results in a Schrondinger-like equation that exhibits a $1/r$ potential, resulting from the gravitational interaction between the bosons and the BH.
-Such an equation can be solved to show that the energy states of the boson cloud are quantized analogously to those of the hydrogen atom, allowing us to describe the boson cloud as a “gravitational atom in the sky” Baumann et al.
+The solution of this equation shows that the energy states of the boson cloud are quantized analogously to those of the hydrogen atom, allowing us to describe the boson cloud as a “gravitational atom in the sky” Baumann et al.
 (2019a).
@@
 In all cases, however, as we will see, the GW emission timescale will be shorter than the time to build up the cloud, making these systems excellent sources of GWs.
-For the black hole to be spun down such that the condition in Eq. 94 is no longer satisfied, scalar and vector clouds require Isi et al. (2019):
+The cloud will continue to build up until the BH is spun down enough such that the condition in Eq.
+96 is no longer satisfied.
+In particular, the timescales $\tau_{\rm inst}^{\rm(s)}$ and $\tau_{\rm inst}^{\rm(v)}$ for which this occurs scalar and vector clouds, respectively, are Isi et al.
+(2019) [^9] Note that we do not consider tensor boson clouds in this work:
 $\tau_{\rm inst}^{\rm(s)}\approx 27\,{\rm days}\left(\frac{M_{\rm BH}}{10\,M_{\odot}}\right)\left(\frac{0.1}{\alpha}\right)^{9}\frac{1}{\chi_{i}}\,$, (97)
@@
 $\tau_{\rm inst}^{\rm(v)}\approx 2\,{\rm minutes}\left(\frac{M_{\rm BH}}{10\,M_{\odot}}\right)\left(\frac{0.1}{\alpha}\right)^{7}\frac{1}{\chi_{i}}\,$, (98)
-where $\chi_{i}$ is the BH spin at birth, and the GW fine-structure constant is
-[…]
-where $m_{b}$ is the mass of the boson.
+where $\chi_{i}$ is the BH spin at birth, and the GW fine-structure constant $\alpha$ is
+$\alpha$ $=GM_{\rm BH}m_{b}$ (99)
+$=0.11\left(\frac{M_{\rm BH}}{10M_{\odot}}\right)\left(\frac{m_{b}}{1.47\times 10^{-12}\,\mathrm{eV}}\right)$ (100)
+where $m_{b}$ is the mass of the boson and $M_{\rm BH}$ is the mass of the BH.
 This new boson could have a spin of 0, 1 or 2, where spin-0 could correspond to the pseudo-scalar QCD axion (or axion-like) particle Brito et al.
 (2015, 2020), spin-1 could indicate dark photons arising through kinetic mixing with the ordinary photon Siemonsen et al.
-(2023a), or spin-2 tensor bosons that could arise from modified theories of gravity Brito et al.
+(2023a) [^10] Note that this dark photon is not the dark photon described in Section II.4., or spin-2 tensor bosons that could arise from modified theories of gravity Brito et al.
 (2020).
-After the cloud builds up, GWs can be emitted in three different ways: (1) through boson-boson annihilation, (2) through boson energy-level transitions, and (3) after a “bosenova”.
-It has been shown, under a wide range of assumptions of BH populations, spins, redshift distributions, etc., that such systems would be detectable by CW searches in advanced LIGO, Virgo and KAGRA data and/or in future detectors, such as ET and Cosmic Explorer Arvanitaki et al.
-(2015b); Zhu et al.
+After the cloud builds up in an isolated BH system, GWs can be emitted in three different ways: (1) through boson-boson annihilation, (2) through boson energy-level transitions, and (3) after a “bosenova”, all of which will be discussed in Sections V.1, V.2 and V.4.
+Furthermore, the GW frequency is fixed by the boson mass, which could in principle take on any ultralight mass; thus, it is important to note that GW interferometers, from pulsar timing arrays to HFGW instruments, will allow an extensive mass range to be probed, from $10^{-22},10^{-9}]\,\mathrm{eV}$ Aggarwal et al.
+(2021); Brzeminski et al.
+(2025); Arvanitaki et al.
+(2015a); Brito et al.
+(2017); Ng et al.
 (2020).
+It has been shown, under a wide range of assumptions of BH populations, spins, redshift distributions, etc., that such systems could be detectable by CW searches in advanced LIGO, Virgo and KAGRA data and/or in future detectors, such as ET and Cosmic Explorer Arvanitaki et al.
+(2015a); Zhu et al.
+(2020); and if such sources are not seen, particular boson masses can be disfavored – this will be the subject of Section VI.
 Furthermore, both galactic BHs Palomba et al.
 (2019) and remnants of mergers of BHs outside the galaxy could also be detected, the latter of which, in the scalar case, only being visible in ET and Cosmic Explorer Arvanitaki et al.
-(2010, 2015b); Isi et al.
-(2019).
+(2010, 2015a); Isi et al.
+(2019), which will be discussed more in Section VI.1.
 Annihilating vector boson cloud systems, however, emit stronger, though shorter, GW signals Brito et al.
@@
 (2023a).
-The boson mass is fixed by the GW frequency, and could in principle take on any ultralight mass.
-That is why GW experiments across the frequency spectrum have been or are being designed to search for these boson cloud systems.
-While most of the literature focuses on ground-based GW interferometers, we note that HFGW detectors will also be able to probe BH superradiance of heavier bosons than those detectable by LIGO, Virgo and KAGRA Aggarwal et al.
-(2021), e.g.
-Brzeminski et al.
-(2025), and space-based GW interferometers will probe lighter masses Arvanitaki et al.
-(2015b); Brito et al.
-(2017), and space-based detector analyses can be used to aid in discovering boson clouds in the band of ground-based GW interferometers Ng et al.
+In this section, we will discuss the three main ways in which GWs can be emitted from isolated BH/ boson cloud systems: annihilation, transitions and bosenovae (Sections V.1, V.2 and V.4).
+We will also detail how self-interactions between the bosons could impact the superradiance process and the resulting GWs in Section V.3.
+Finally, we will explain some of the rich physics that can occur when boson clouds form around binary black hole systems in Section V.5.
+V.1 Annihilation signal
+Ultralight bosons are their own anti-particles, which means that they continuously annihilate (even as the cloud is growing), producing GW radiation at a frequency fixed by the boson mass and slightly redshifted by the BH mass (in the non-relativistic limit $\alpha\ll 1$, which predominately applies to the equations shown in this review).
+In the scalar boson case Abbott et al.
+(2022d):
+$f_{\text{gw}}$ $\simeq 483\,{\rm Hz}\left(\frac{m_{\text{b}}}{10^{-12}~\text{eV}}\right)$
+$\times\left[1-7\times 10^{-4}\left(\frac{M_{\text{BH}}}{10M_{\odot}}\frac{m_{\text{b}}}{10^{-12}~\text{eV}}\right)^{2}\right].$ (101)
+Qualitatively, the GWs emitted are monochromatic because the mass of the boson is fixed, and all the bosons are in the same energy state.
+As alluded to before, the timescales for GW emission are much longer than the time it takes to build the cloud Isi et al.
+(2019):
+$\tau_{\rm GW}^{\rm(s)}\approx 2\times 6.5\times 10^{4}\,{\rm yr}\left(\frac{M_{\rm BH}}{10\,M_{\odot}}\right)\left(\frac{0.1}{\alpha}\right)^{15}\hskip-2.0pt\frac{1}{\chi_{i}}.$ (102)
+$\tau_{\rm GW}^{\rm(v)}\approx 2\times 1\,{\rm day}\left(\frac{M_{\rm BH}}{10\,M_{\odot}}\right)\left(\frac{0.1}{\alpha}\right)^{11}\hskip-2.0pt\frac{1}{\chi_{i}}.$ (103)
+where we have explicitly noted a factor of 2 that is missing from the original reference Isi et al. (2019), as pointed out in Siemonsen et al. (2023b).
+As the number of bosons starts to decrease, there will also be a classical depletion of the cloud over time.
+This reduction in bosons causes the cloud to contract, resulting in the following positive drifts in frequency over time Isi et al.
+(2019) [^11] We note that relativistic computations of the frequency drift have also been performed Siemonsen et al.
+(2023b); May et al.
+(2025), which are useful for real searches of GWs arising from boson clouds but not less important for our discussions here.:
+$\dot{f}_{\rm gw}^{\rm(s)}\approx 3\times 10^{-14}\,{\rm Hz/s}\left(\frac{10M_{\odot}}{M_{\rm BH}}\right)^{2}\left(\frac{\alpha}{0.1}\right)^{19}\chi_{i}^{2}\,$, (104)
+$\dot{f}_{\rm gw}^{\rm(v)}\approx 1\times 10^{-6}\,{\rm Hz/s}\left(\frac{10M_{\odot}}{M_{\rm BH}}\right)^{2}\left(\frac{\alpha}{0.1}\right)^{15}\chi_{i}^{2}\,.$ (105)
+In the non-relativistic limit $\alpha\ll 1$, the signals will have the corresponding amplitudes of Isi et al. (2019):
+$h_{0}^{\rm(s)}\approx 8\times 10^{-28}\left(\frac{M_{\rm BH}}{10M_{\odot}}\right)\left(\frac{\alpha}{0.1}\right)^{7}\left(\frac{\rm Mpc}{r}\right)\left(\frac{\chi-\chi_{f}}{0.1}\right)$ (106)
+$h_{0}^{\rm(v)}\approx 4\times 10^{-24}\left(\frac{M_{\rm BH}}{10M_{\odot}}\right)\left(\frac{\alpha}{0.1}\right)^{5}\left(\frac{\rm Mpc}{r}\right)\left(\frac{\chi-\chi_{f}}{0.1}\right),$ (107)
+where the final spin $\chi_{f}$ is
+$\chi_{f}=\frac{4\alpha_{f}m}{4\alpha^{2}_{f}+m^{2}}\,.$ (108)
+These equations indicate interesting properties for scalar and vector boson cloud annihilation signals.
+For bosons whose Compton wavelengths are optimally matched to the size of the BH, scalar bosons will annihilate over timescales that greatly exceed the observation time of ground- or space-based GW interferometers (Eq.
+102), and their frequencies will hardly drift over time.
+Thus, this signal can be thought of as a quasi-sinusoidal and persistent.
+On the other hand (and with the same caveat of optimal matching), the GW signal from annihilating vector boson clouds lasts significantly shorter than the scalar one (Eq.
+103), and thus can be thought of more as a “transient” CW, lasting for durations for $\mathcal{O}($hours-days), i.e.
+much longer than mergers of black holes routinely detected now by LIGO, Virgo and KAGRA, but much shorter than scalar boson cloud annihilation signals.
+Furthermore, vector signals emit much more GW power per unit time than scalar ones, but for scalar signals, analysis methods are able to integrate over the entire data collection period of LIGO, Virgo and KAGRA to improve the SNR, as will be discussed in Section VI.2.
+However, in the vector case, it is possible to have CW emission for particular combinations of boson mass and BH mass, such as large BHs and small boson masses, where the spin-up of the signal, as described by Eq.
+105, becomes comparable to that in the scalar case (Eq.
+104), thus making the signal almost monochromatic.
+V.2 Transition signal
+In the scalar boson case, just as electrons jump between discrete energy levels in atoms, bosons can also transition between discrete energy states, provided these states are populated.
+The energy states of the boson cloud are quantized, with each state corresponding to specific quantum numbers.
+Significant transition signals are expected when the cloud reaches a population that allows transitions between particular energy levels, such as from the 5g to the 6g states, where “g” refers to the orbital angular momentum quantum number of the state.
+These transitions between specific energy levels are expected to generate detectable gravitational wave signals that could last for durations much longer than the typical observation times of ground-based gravitational wave interferometers.
+The frequency of these signals is given by Arvanitaki et al.
+(2015a):
+$f_{\rm gw}\sim 15\,\text{Hz}\left(\frac{m_{b}}{10^{-11}\,\text{eV}}\right)$ (109)
+The frequency of the signal would drift over time according to the duration of the signal and the axion decay constant $f_{a}$ Arvanitaki et al. (2015a):
+$\frac{df}{dt}\simeq 10^{-11}\,\frac{\text{Hz}}{\text{s}}\left(\frac{f_{\rm gw}}{90\,\text{Hz}}\right)\left(\frac{M_{\rm BH}}{10\,M_{\odot}}\right)\left(\frac{10^{17}\,\text{GeV}}{f_{a}}\right)^{2}\left(\frac{5\,\text{yr}}{T_{\rm obs}}\right)^{2}$ (110)
+We note here that the axion decay constant, arising from self-interactions, affects the spin-up of the signal.
+In the previous discussions surrounding Eqs.
+104 and 105, self-interactions were assumed to be zero.
+While the expected event rate of these transitions in ground-based detectors is unlikely to exceed one event in around 3.5 years of observation with advanced LIGO Arvanitaki et al.
+(2015a), the presence of self-interactions could potentially enhance the transition rates, making these signals detectable Collaviti et al.
 (2025).
-As will be discussed, boson cloud self-interactions could have a significant impact on (1) whether the superradiance condition is saturated Yoshino and Kodama (2012, 2015a, 2015b), (2) how quickly a binary black hole system will approach merger Aurrekoetxea et al.
-(2024a); Takahashi et al.
-(2024), and (3) the rate of change of the GW frequency of an isolated boson cloud Baryakhtar et al.
+In the vector boson case, transitions can also occur, and interestingly, overtone modes may grow faster than the fundamental modes, particularly for small angular numbers, fine-structure constants, and black hole spin values Siemonsen and East (2020).
+These higher overtones could saturate more quickly than the fundamental modes, allowing transitions to lower states and increasing the overall energy of the cloud, potentially leading to detectable gravitational wave emission.
+V.3 Impact of boson self-interactions on superradiance
+While the formation of ultralight boson clouds around rotating BHs is already expected to result in observable GWs, the strength of the self-interactions within the boson cloud can play a crucial role in determining the nature and detectability of these signals.
+Specifically, the self-interactions between bosons in the cloud can influence key dynamical processes, such as:
+  1. 1.
+    Saturation of Superradiance: Self-interactions modify the growth timescale, potentially leading to a faster saturation of the superradiance process.
+This could result in stronger GW emission once the cloud reaches its maximum energy state.
+The timescale for superradiance saturation is sensitive to the coupling strength between bosons, and self-interactions may allow the cloud to reach a critical size faster, enhancing the CW signal emitted during this phase Yoshino and Kodama (2012, 2015a, 2015b).
+  2. 2.
+    Merger Dynamics: The presence of self-interactions can affect how quickly the boson cloud dissipates or decays, influencing the coalescence of the black hole system.
+This decay modifies the timescale of the inspiral phase and the subsequent merger, altering the frequency evolution of the emitted GWs.
+A faster decay due to self-interactions could lead to a sharper rise in the frequency of the GW signal Boudon et al.
+(2024); Aurrekoetxea et al.
+(2024); Takahashi et al.
 (2024).
-Self-interactions would also allow higher energy levels to grow faster.
-The strength of these interactions will affect the detection prospects in ET and Cosmic Explorer Collaviti et al.
-(2025), as well as recent search results which have not yet accounted for moderate or strong interactions.
+  3. 3.
+    Enhanced Annihilation of Boson Clouds in Isolated Black Holes: In a non-interacting system, the annihilation rate of bosons is limited by the density and distribution of particles in the cloud.
+Self-interactions, however, can increase the density of the cloud by accelerating the process of boson accumulation and scattering.
+As a result, the rate of boson annihilation can be significantly enhanced.
+This increased annihilation rate leads to stronger and potentially more detectable GW signals from the cloud Baryakhtar et al.
+(2021).
+  4. 4.
+    Enhanced Energy-Level Transitions: The energy levels of bosons in the cloud are quantized, similar to the energy levels of electrons in an atom.
+Self-interactions can facilitate the transition of bosons between higher energy states more rapidly, increasing the amplitude of the emitted gravitational waves.
+This could make energy-level transitions a more significant source of detectable GW signals, especially if the cloud is in a high-energy state Baryakhtar et al.
+(2021).
+These effects become particularly relevant when considering the potential detection of boson clouds in future GW observatories, such as Cosmic Explorer and ET Collaviti et al.
+(2025).
+Recent searches for these signals have not accounted for these interactions, and their inclusion may shift current detection prospects by enhancing the signal strength or changing the emission characteristics.
+In this section, we will explore how self-interactions influence the dynamics of boson clouds and their implications for gravitational wave astronomy.”
 Note that while spin-0 and spin-1 boson clouds are relatively well understood theoretically, spin-2 clouds are much more complicated.
@@
 (2024).
-V.1 Annihilation signal
-When Eq.
-94 is no longer satisfied, the cloud stops growing.
-Since ultralight bosons are their own anti-particles, they will start to annihilate, producing GW radiation at a frequency fixed by the boson mass and slightly redshifted by the BH mass (in the non-relativistic limit $\alpha\ll 1)$, which predominately applies to the equations shown in this work).
-In the scalar boson case Abbott et al.
-(2022d):
-$f_{\text{gw}}\simeq 483\,{\rm Hz}\left(\frac{m_{\text{b}}}{10^{-12}~{}\text{eV}}\right)\times\left[1-7\times 10^{-4}\left(\frac{M_{\text{BH}}}{10M_{\odot}}\frac{m_{\text{b}}}{10^{-12}~{}\text{eV}}\right)^{2}\right].$ (98)
-Qualitatively, the GWs emitted are monochromatic because the mass of the boson is fixed, and all the bosons are in the same energy state.
-As alluded to before, the timescales for GW emission are much longer than the time it takes to build the cloud Isi et al.
-(2019):
-$\tau_{\rm GW}^{\rm(s)}\approx 2\times 6.5\times 10^{4}\,{\rm yr}\left(\frac{M_{\rm BH}}{10\,M_{\odot}}\right)\left(\frac{0.1}{\alpha}\right)^{15}\frac{1}{\chi_{i}}.$ (99)
-$\tau_{\rm GW}^{\rm(v)}\approx 2\times 1\,{\rm day}\left(\frac{M_{\rm BH}}{10\,M_{\odot}}\right)\left(\frac{0.1}{\alpha}\right)^{11}\frac{1}{\chi_{i}}.$ (100)
-where we have explicitly noted a factor of 2 that is missing from the original reference Isi et al.
-(2019), as pointed out in Siemonsen et al.
-(2023b).
-There will also be a classical depletion of the cloud over time, as the number of bosons starts to decrease, resulting in the following drifts in frequency over time Isi et al.
-(2019):
-$\dot{f}_{\rm gw}^{\rm(s)}\approx 3\times 10^{-14}\,{\rm Hz/s}\left(\frac{10M_{\odot}}{M_{\rm BH}}\right)^{2}\left(\frac{\alpha}{0.1}\right)^{19}\chi_{i}^{2}\,,$ (101)
-$\dot{f}_{\rm gw}^{\rm(v)}\approx 1\times 10^{-6}\,{\rm Hz/s}\left(\frac{10M_{\odot}}{M_{\rm BH}}\right)^{2}\left(\frac{\alpha}{0.1}\right)^{15}\chi_{i}^{2}\,.$ (102)
-We note that relativistic computations of the frequency drift have also been performed Siemonsen et al.
-(2023b); May et al.
-(2025).
-In the non-relativistic limit $\alpha\ll 1$, the signals will have the corresponding amplitudes of Isi et al.
-(2019):
-$h_{0}^{\rm(s)}\approx 8\times 10^{-28}\left(\frac{M_{\rm BH}}{10M_{\odot}}\right)\left(\frac{\alpha}{0.1}\right)^{7}\left(\frac{\rm Mpc}{r}\right)\left(\frac{\chi-\chi_{f}}{0.1}\right)$ (103)
-$h_{0}^{\rm(v)}\approx 4\times 10^{-24}\left(\frac{M_{\rm BH}}{10M_{\odot}}\right)\left(\frac{\alpha}{0.1}\right)^{5}\left(\frac{\rm Mpc}{r}\right)\left(\frac{\chi-\chi_{f}}{0.1}\right)$ (104)
-where the final spin $\chi_{f}$ is
-$\chi_{f}=\frac{4\alpha_{f}m}{4\alpha^{2}_{f}+m^{2}}\,$ (105)
-These equations indicate interesting properties for scalar and vector boson cloud annihilation signals.
-For bosons whose Compton wavelengths are optimally matched to the size of the BH, scalar bosons will annihilate over timescales that greatly exceed the observation time of ground- or space-based GW interferometers, and their frequencies will hardly drift over time.
-Thus, this signal can be thought of as a quasi-sinusoidal and persistent.
-On the other hand (and with the same caveat of optimal matching), the GW signal from annihilating vector boson clouds lasts significantly shorter than the scalar one, and thus can be thought of more as a “transient” CW, lasting for durations for $\mathcal{O}($hours-days), i.e.
-much longer than mergers of black holes routinely detected now by LIGO, Virgo and KAGRA, but much shorter than scalar boson cloud annihilation signals.
-Furthermore, vector signals emit much more GW power per unit time than scalar ones, but for scalar signals, analysis methods are able to integrate over the entire data collection period of LIGO, Virgo and KAGRA to improve the signal-to-noise ratio.
-If, in the vector case, the size of the BH does not match the Compton wavelength of the boson, the signals could resemble CWs.
-V.2 Transition signal
-[…]
-$f_{\rm gw}\sim 15\text{ Hz}\left(\frac{m_{b}}{10^{-11}\text{ eV}/c^{2}}\right)$ (106)
-and a frequency drift over time that relates to the duration of the signal $T$ and the axion decay constant $f_{a}$ Arvanitaki et al. (2015b):
-$\!\!\!\!\frac{df}{dt}\!\!\simeq 10^{-11}\!\frac{\,\mathrm{Hz}\,}{\mathrm{s}}\!\left(\!\frac{f_{\text{gw}}}{90\,\mathrm{Hz}\,}\!\right)\!\!\left(\!\frac{M_{\rm BH}}{10M_{\odot}}\!\right)\!\!\left(\!\frac{10^{17}\!\,\mathrm{GeV}}{f_{a}}\!\right)^{\!2}\!\!\left(\!\frac{5\mathrm{yr}}{T_{\text{obs}}}\!\right)^{\!2}\!\!$ (107)
-Unfortunately, the expected event rate of transitions in ground-based detectors is not expected to exceed one even in $\sim 3.5$ years of observing time in advanced LIGO (Fig.
-5 in Arvanitaki et al.
-(2015b)).
-But, if self-interactions occur, transitions could happen and be detectable Collaviti et al.
-(2025).
-In the vector boson case, transitions could also occur, and the overtone modes could actually grow faster than the fundamental modes at small angular numbers and for fine-structure constants and BH spin values Siemonsen and East (2020).
-The higher overtones could saturate much quicker than lower ones, implying that transitions to lower states could occur, thus allowing BHs to become unstable to increasing energies at lower modes.
-Different energy states of the boson cloud could thus be populated at the same time, enriching the kind of physics searchable via transitions.
-V.3 Bosenova signal
-If the self-interactions between bosons become stronger than the gravitational binding energy of the cloud, the cloud starts to deform and can no longer be described analogously to the Hydrogen atom Fukuda and Nakayama (2020); Collaviti et al.
+V.4 Bosenova signal
+If the self-interactions between bosons become stronger than the gravitational binding energy of the cloud, the cloud begins to destabilize and cannot be described analogously to the hydrogen atom Fukuda and Nakayama (2020); Collaviti et al.
 (2025); Omiya et al.
@@
 (2025).
-Instead, it can collapse and explode in what is known as a “bosenova”.
-This phenomenon would occur if the potential energy of the cloud is approximately equal to the self-interaction energy.
-Bosenovas have been observed on earth in condensed matter laboratories, in which repulsive forces were switched to attractive ones in a material that formed a Bose-Einstein condensate via a tuneable external magnetic field Donley et al.
+The cloud may undergo a significant reconfiguration, leading to a dramatic release of energy, a process referred to as a “bosenova.” This occurs when the potential energy of the boson cloud becomes comparable to or exceeds the self-interaction energy, leading to the rapid rearrangement of the bosonic matter.
+The phenomenon of a bosenova has been observed in condensed matter systems, in which a Bose-Einstein condensate undergoes a similar process due to a change in interactions induced by an external magnetic field Donley et al.
 (2001).
-There are a lot of uncertainties around whether the bosenova would occur, and it seems that including self-interactions in the cloud limits bosenovas to a corner of the allowed parameter BH-boson mass parameter space Baryakhtar et al.
+However, the occurrence of a bosenova in the context of ultralight bosons around black holes remains highly uncertain, with studies suggesting that self-interactions may limit bosenovas to a restricted region of the allowed parameter space, particularly for certain combinations of black hole and boson masses Baryakhtar et al.
 (2021).
@@
 Though GWs from boson clouds described above may seem independent, they could in principle be happening at the same time.
-In particular, if ultralight scalar or vector bosons have some self-interactions, the GW signal will change.
-If these clouds are composed of string axions (as opposed to the QCD axion), a non-linear self-interaction may cause a bosenova if, during superradiance, the value of the scalar field approaches the axion decay constant $f_{a}$, which, for string axions, could be below the grand unification theory (GUT) scale Arvanitaki et al.
+In particular, if ultralight scalar or vector bosons have some self-interactions, and If these clouds are composed of string axions (as opposed to the QCD axion), a non-linear self-interaction may cause a bosenova if, during superradiance, the value of the scalar field approaches the axion decay constant $f_{a}$, which could be below the grand unification theory (GUT) scale Arvanitaki et al.
 (2010); Yoshino and Kodama (2012, 2015a, 2015b).
 This bosenova would halt the superradiance process, as about 5% of the boson cloud mass would fall back into the BH, but superradiance would immediately start again until the next bosenova is triggered.
-Such periodic bosenovas could sustain superradiance for much longer durations than those in the conventional scenario, permitting annihilating clouds even around extremely old BHs.
-Furthermore, we note that boson annilation could alter the GW of an inspiraling binary, potentially visible in space-based GW detectors Xie and Huang (2024), in a similar way that dynamical friction does.
-In more general models of axions than in the axiverse, self-interactions could suppress annihilation signals as they get stronger, and transitions could occur for moderate interaction strengths.
-In particular, GW power sourced by annihilation would be reduced due to the smaller cloud size that can be sustained because of self-interactions, while scalar radiation, a new energy loss mechanism, consumes a portion of the energy that would source GW emission.
-But, multiple levels could be populated with self-interactions, allowing transitions to occur more than in the non-self-interaction scenarios Baryakhtar et al.
-(2021).
-V.4 Clouds in binary black hole systems
-We know that BHs form binaries, and if ultralight dark matter exists in the form of ultralight bosons, it should imprint some signature in the GWs arising from inspiraling binary black holes Baumann et al.
+Such periodic bosenovas could sustain superradiance for much longer durations than those in the conventional scenario, e.g.
+in Arvanitaki and Dubovsky (2011), permitting annihilating clouds even around extremely old BHs, and even affect GWs from binary black hole systems Xie and Huang (2024) .
+V.5 Clouds in binary black hole systems
+If boson clouds form around binary black hole systems, they should imprint some signature in the GWs arising from their inspiral Baumann et al.
 (2019a).
-The GW signal will be modified due to tidal disruptions induced by the companion, the multiple moments of the cloud Baumann et al.
-(2019a, b, 2020) or through a dynamical friction drag forced induced by the ultralight dark matter.
+In particular, the GW signal will be modified due to tidal disruptions induced by the companion, the multiple moments of the cloud Baumann et al.
+(2019a, b, 2020) or through a dynamical friction drag force induced by the boson clouds.
 These kinds of signals could be visible in future space-based detectors, since the deviations primarily occur in the early inspiral stages of the system, i.e.
-at frequencies much lower than currently accessible by ground-based GW interferometers.
-In both cases, observations of the inspiral of binary systems, and in particular extreme mass ratio inspiral (EMRI) systems Zhang and Yang (2019); Brito and Shah (2023), could indicate the presence of boson clouds around one or both of the objects if there is some dephasing with respect to the vacuum .
+at frequencies much lower than currently accessible by ground-based GW interferometers Guo et al.
+(2024a).
+In both cases, observations of the inspiral of binary systems, and in particular EMRI systems Zhang and Yang (2019); Brito and Shah (2023), could indicate the presence of boson clouds around one or both of the objects if there is some dephasing with respect to the waveform in vacuum.
 Additionally, once the binary reaches a certain stage of its evolution, the impact of both the spin-induced quadrupole moment and tidal forces will lessen, since the clouds will become disrupted by their companions.
-If we can observe variations in both the spin-induced quadrupole moment, parameterized as $\kappa$, and the tidal deformability, parameterized as $\Lambda$, over the course of the lifetime of the binary, it would indicate the presence of a (disrupted) boson cloud Baumann et al.
+If we can observe variations in both the spin-induced quadrupole moment and the tidal deformability over the course of the lifetime of the binary, it would indicate the presence of a (disrupted) boson cloud Baumann et al.
 (2019a).
-More recently, it has been show that the boson cloud could be “ionized” as the orbital separation approaches the size of the cloud.
-Self-interactions could also play a role in the evolution of a binary, in which two types of clouds can actually coexist around the rotating BH through mode coupling Boudon et al.
-(2024); Aurrekoetxea et al.
-(2024a); Takahashi et al.
-(2024).
-We will now describe different ways in which boson clouds could affect GWs from binary black hole systems.
-V.4.1 Spin-induced quadrupole moment
-The spin-induced quadrupole moment is often parameterized in terms of a dimensionless quantity $\kappa=\frac{Q}{J^{2}/M_{\rm BH}}$, which has been normalized by the object’s angular moment $J$ and mass BH.
-This parameter enters into the waveform at the second post-Newtonian order.
+More recently, it has been shown that as the orbital separation in a binary black hole system decreases and becomes comparable to the size of the boson cloud, the boson cloud could be “ionized” Takahashi et al.
+(2022); Baumann et al.
+(2022a, b).
+This ionization occurs when the energy from the binary black hole system becomes large enough to unbind the cloud from its host BH.
+As a result, the inspiral of the binary becomes significantly influenced by the dynamics of the boson cloud, rather than merely being perturbed by it.
+We will now describe four different ways in which boson clouds could affect GWs from binary black hole systems.
+V.5.1 Spin-induced quadrupole moment
+The spin-induced quadrupole moment is often parameterized in terms of a dimensionless quantity $\kappa=\frac{Q}{J^{2}/M_{\rm BH}}$, where $Q$ is the quadrupole moment of the BH, that has been normalized by the angular moment $J$ and mass of the BH.
+This parameter enters into the waveform at the second post-Newtonian (2PN) order, i.e.
+as corrections of order $(v/c)^{4}$ relative to the leading quadrupole radiation.
 Thus, the spin-induced quadrupole moment, regardless of its source, has a significant impact on the GW signal from compact objects inspiraling towards one another.
@@
 In the case of boson cloud systems, $\kappa$ will take on vastly different values, of $\mathcal{O}(1000)$, depending on the boson cloud and BH masses, and the extent to which the Kerr metric is altered in the presence of a cloud with enough mass compared to the BH.
-V.4.2 Tidal disruptions
+V.5.2 Tidal disruptions
 Because the boson cloud is much less compact than the BH, we expect that tidal disruptions from the companion could occur when the two objects in the binary black hole are close enough to each other Ng et al.
@@
 To characterize the effects of tides in a binary system, the tidal deformability parameter $\Lambda$ is used, which depends on the mass and radius of the compact objects.
-For BHs, this quantity is 0, while for neutron stars, it can take on values of $\mathcal{O}(100-1000)$ Chatziioannou (2020).
+For BHs in vacuum, this quantity is 0, while for neutron stars, it can take on values of $\mathcal{O}(100-1000)$ Chatziioannou (2020).
 For boson clouds, $\Lambda\sim 10^{7}$, which is a remarkable departure from that expected from conventional astrophysical binaries Baumann et al.
@@
 Smaller recoil velocities imply that remnants could remain inside clusters, leading to higher chances of hierarchical formation and merger of binary black holes.
-Such disruptive effects of bosons on not just individual BHs but also on the BH population, imply that boson clouds, if they exist, would significantly impact the formation and evolution histories of BHs.
-V.4.3 Binary-induced transitions
-Over the course of the inspiral, the presence of the boson cloud will induce transitions between different energy levels of the boson clouds that would be visible in space-based GW detectors if (1) the orbital frequency matches the hyperfine splitting between two energy levels (hyperfine resonance), and (2) the growing mode mixes with decaying modes (Bohr resonance).
-Both kinds of transitions would also be smoking-gun signatures of the presence of boson clouds around rotating BHs in a binary Baumann et al.
+Such disruptive effects of bosons on not just individual BHs but also on the BH population, imply that boson clouds would significantly impact the formation and evolution histories of BHs.
+V.5.3 Binary-induced transitions
+During the inspiral of a binary black hole system that contains boson clouds, the interaction between the cloud and the BHs can cause transitions between different energy levels in the clouds.
+These transitions can potentially be detected by space-based GW interferometers if any of the following mechanisms is present:
+  1. 1.
+    Hyperfine Resonance: The orbital frequency of the binary system matches the energy spacing between two adjacent energy levels in the boson cloud.
+This condition is analogous to a resonance in atomic physics, where the energy difference between two states is exactly matched by an external driving frequency.
+In this case, the system stays in equilibrium, and energy is transferred between energy levels that are in resonance Baumann et al.
 (2019a).
-More recently, by considering the slow evolution of the orbital frequency through resonances of the boson cloud system, that is, when the orbital frequency matches the energy spacing between adjacent levels, it can be shown that Landau-Zener transitions occur and significantly affect the inspiral GW signal Baumann et al.
+  2. 2.
+    Bohr Resonance: In addition to hyperfine resonance, the growing mode (the mode that is being excited by the inspiral) mixes with decaying modes of the cloud.
+This interaction allows energy to be transferred between different states, influencing the GW signal Baumann et al.
+(2019a).
+  3. 3.
+    Landau-Zener Transitions: When the orbital frequency of the inspiral matches the energy spacing between adjacent energy levels, the system undergoes a non-adiabatic shift in energy states.
+This transition process, characterized by the Landau-Zener mechanism, significantly alters the frequency and amplitude evolution of the GW signal.
+These non-adiabatic transitions can provide observable changes in the signal, offering additional insights into the properties of the boson cloud Baumann et al.
 (2020).
-V.4.4 Ionization
-Another transition arises when the binary system separation becomes comparable to the size of the cloud: an “ionization” of the boson cloud occurs as the cloud becomes unbounded from its host BH Takahashi et al.
+These effects, all of which could be present simultaneously, would provide “smoking-gun” signatures of the presence of boson clouds around rotating black holes in a binary system Baumann et al.
+(2019a).
+V.5.4 Ionization
+When the binary system separation becomes comparable to the size of the cloud, an “ionization” of the boson cloud occurs as the cloud becomes unbounded from its host BH Takahashi et al.
 (2022); Baumann et al.
 (2022a, b).
-The energy from the binary sources this effect, and is significantly larger than that emitted via GWs; thus, the inspiral of the binary becomes driven by the boson cloud, instead of simply perturbed by it.
+The energy from the binary drives this effect, and is significantly larger than that emitted via GWs; thus, the inspiral of the binary becomes dominated by the boson cloud, instead of simply perturbed by it.
 Ionization will also tend to circularize the binary if it formed through dynamical capture, while leaving any orbital inclinations unaffected Tomaselli et al.
@@
 It has been recently found, however, that as the resonances in the boson cloud evolve during the orbit, the boson cloud is unperturbed and ordinarily visible in LISA if the cloud and binary are counter rotating; otherwise, the cloud is destroyed because of the resonances, but a distinct mark is left on the binary.
-That is, the binary is forced to co-rotate with the boson cloud, and its eccentricity is driven to a particular value, which allows the possibilty of doing statistics with detected binary black hole systems in LISA to probe the boson cloud hypothesis Tomaselli et al.
+That is, the binary is forced to co-rotate with the boson cloud, and its eccentricity is driven to a particular value, which allows the possibility of doing statistics with detected binary black hole systems in LISA to probe the boson cloud hypothesis Tomaselli et al.
 (2024).
+Taken together, the various phenomena discussed above—from level transitions and resonances to the dramatic ionization of the cloud—highlight the rich interplay between boson clouds and binary black holes.
+Each mechanism offers a distinct pathway by which the presence of a boson cloud can alter binary evolution and leave characteristic imprints on the emitted GWs.
+These theoretical insights provide the foundation for future searches with space-based GW data Bertone et al.
+(2020); Bertone (2024).
+In the following section, however, we turn to current ground-based GW interferometer data, which so far constrain only annihilation signals from boson clouds, as discussed in Section V.1.
 VI GW constraints on boson clouds
@@
 While ground-based detectors are sensitive to annihilation signals from bosons of masses between $\sim 10^{-14}-10^{-12}$ eV Arvanitaki et al.
-(2015b), space-based detectors can probe a few orders of magnitude lower than that, $\sim 10^{-20}-10^{-16}$ eV Baumann et al.
+(2015a), space-based detectors can probe a few orders of magnitude lower than that, $\sim 10^{-20}-10^{-16}$ eV Baumann et al.
 (2019a), and HFGW detectors could see systems with masses around $\sim 10^{-9}$ eV Aggarwal et al.
@@
 At the moment, only ground-based GW interferometers exist, and so we focus on methods and search results from the most recent LIGO, Virgo and KAGRA observing runs.
-VI.1 Constraints from the spins of detected binary black holes
-We note that constraints on boson clouds can be inferred from the spin measurements of binary black hole systems Arvanitaki et al.
-(2015b, 2017); Brito et al.
-(2017); Baryakhtar et al.
-(2017); Cardoso et al.
-(2018); Ghosh and Sachdeva (2021).
-Recently, a Bayesian method has been developed by combining the information from each system Ng et al.
-(2021a), assuming a distribution for the initial BH spins and requiring that the BHs have had enough time to undergo superradiance before merging.
-Such a study has been applied to the second GW transient catalog (GWTC-2) Abbott et al.
-(2021b), and has resulted in essentially a joint posterior on the scalar boson mass and the spin distribution of BHs at their formation Ng et al.
-(2021b).
-The latter is required because BHs born with high spins that have been spun down by superradiance is partially degenerate with BHs simply born with low spins in a universe without superradiance.
-The study presented in Ng et al.
-(2021b) tentatively rules out scalar bosons in the mass range $[1.3,2.7]\times 10^{-13}$ eV.
+In particular, only GWs from annihilating boson clouds around isolated BHs, discussed in Section V.1, can be searched for and constrained with current GW data.
+This section is broken into four parts: in Section VI.1, we will describe potential sources of GWs from annihilating boson clouds that would be visible in LIGO, Virgo and KAGRA.
+Next, in Section VI.2, we will discuss the CW methods employed to do searches for each of the targets.
+We will then show recent constraints from different searches for GWs from annihilating boson clouds in Section VI.3, and conclude in Section VI.4 with prospects for future searches for boson clouds with current and future data.
+VI.1 Search targets
+In Section V, we have explained the different mechanisms that could generate boson clouds around rotating BHs, and thus source GWs through different physical processes – annihilation, transitions, resonances in binary systems, etc.
+From a practical point of view, we must also discuss where and how we should look for boson clouds.
+Currently, only ground-based GW interferometers have been built; thus, we are essentially limited to looking for only one of the aforementioned GW emission mechanisms: annihilation, as described in Section V.1.
+Thus, the next logical question would be: what kind of systems would exhibit annihilation of boson clouds, and where in the sky can they be found?
+In principle, there are four targets to consider
+  1. 1.
+    Known galactic BHs (e.g. Cygnus X-1 Fabian et al. (1989))
+  2. 2.
+    Known remnants of binary black hole mergers (e.g. GW150914 Abbott et al. (2016))
+  3. 3.
+    Known binary black hole inspirals and mergers Abac et al. (2025b)
+  4. 4.
+    Unknown galactic BHs
+Each of these sources has its own benefits and drawbacks: known BHs have many of their intrinsic parameters measured – mass, spin, inclination angle, etc.
+–; however, they are old and thus the strength of GW emission from annihilating boson clouds would be significantly diminished with respect to that when they were first born.
+On the contrary, remnants of BH mergers are extremely young: their clouds build up quickly and they could begin emitting days or weeks after formation, as described in Section V.
+Their age is thus precisely measured, as well as the final mass and spin (within $10\%$ in the case of GW150914 Abbott et al.
+(2016)).
+These estimates can be even more precise than those that arise from some x-ray binaries.
+However, remnants of binary black hole mergers are far: GW150914 merged at $0.410$ Mpc, and other detected systems have been seen out to Gpcs.
+Thus, the scaling of GW amplitude with distance implies that we could not detect remnants of mergers with annihilating scalar boson clouds around them.
+We note, however, that rapidly annihilating vector boson clouds could be detected out to Gpc Jones et al.
+(2023, 2025) in current-generation GW interferometers, making merger remnants a viable target even now.
+On the other hand, with known binary black hole inspirals, boson clouds could alter the GW signal in a variety of ways, as detailed in Section V.5.
+Thus, using the spins and masses inferred from LIGO, Virgo and KAGRA measurements, the allowed boson mass can be constrained: if certain BH masses are detected with low spins, this could be evidence that boson clouds have spun down BHs.
+On the contrary, if BHs with high spins are detected, this could rule out the presence of boson clouds existing in a particular mass range Ng et al.
+(2021a).
+Such an analysis allows us to obtain some evidence for or rule out the presence of boson clouds around rotating binary black holes, but does not necessarily permit us to claim confidently that we have detected a boson cloud.
+Finally, unknown galactic BHs could take on a wide range of masses, ages and spins depending on when and how they formed.
+The number of expected BHs in the galaxy is $\mathcal{O}(10^{9})$, which comes from considering the age of the galaxy ($\sim 10^{11}$ years) and the supernova rate ($\sim 1$ per century) Madau and Dickinson (2014).
+Thus, the BHs we observe are such a small portion of the total number that exist in the local universe, and may not be representative of the full population.
+However, searches for such systems are entirely blind: we do not know any BH mass or spin parameters, nor do we know the location of such systems.
+While this may seem to be a daunting task, computationally efficient methods from the CW community have been leveraged to do these kinds of searches D’Antonio et al.
+(2018); Isi et al.
+(2019); Jones et al.
+(2023, 2025).
+We will now list the methods that have been developed to search for these sources, and then delve into the constraints that have been set using real LIGO, Virgo and KAGRA data.
 VI.2 Search methods
-While the aforementioned constraints from measured BH spins can disfavor particular masses or even confirm the presence of an ultralight boson, they rely on assumptions about the spin distribution at formation, on whether the boson clouds have time to form (which could be impacted by eccentricity at formation Peters (1964); Wen and Gair (2005)), and are a function of the (strongest) BH mergers that we see with GWs.
-To that end, different methods to actually search for long-lived GWs from isolated BHs (or those present in a known x-ray binary system) have been developed and used in real searches in LIGO, Virgo and KAGRA data.
+Different methods to search for long-lived GWs from isolated BHs (or those present in a known x-ray binary system) have been developed and used in real searches in LIGO, Virgo and KAGRA data.
 Detecting GW emission from an isolated BH would provide smoking-gun evidence of the existence of an ultralight bosonic field around BHs.
-VI.2.1 Excess power
-One method to search for boson clouds around rotating BHs lightly models the boson cloud signal as monochromatic within the entire observation time that could come from anywhere in the sky, hence inducing a Doppler modulation.
-This method creates time-frequency maps and corrects each of them for different sky positions, leaving behind a monochromatic signal D’Antonio et al.
+Each has been designed to look for a quasi-monochromatic (scalar boson) or rapidly evolving long-transient (vector boson) signal arising from specific targets, such as Cygnus X-1, or originating from anywhere in the sky.
+While searches for known systems are significantly less expensive computationally than all-sky ones, they are limited to the few BHs about which we have measurements of distance, spin, mass and age.
+Thus, a multi-pronged approach has been taken within the community to probe both known and unknown sources of GWs.
+VI.2.1 Excess power method
+One approach to searching for boson clouds around rotating BHs is the excess-power method, which models the expected signal as approximately monochromatic over the entire observation time.
+Because the source could lie anywhere on the sky, the signal is modulated by the Earth’s motion, leading to a Doppler-shifted frequency.
+To recover such a signal, the data are transformed into time–frequency maps and corrected for different sky positions, so that a truly monochromatic signal would appear as a narrow line D’Antonio et al.
 (2018).
-It employs different coherence times $T_{\text{FFT}}$ for the analysis in order to be robust against possible theoretical uncertainties in the GW signal model.
-Shorter $T_{\text{FFT}}$, while less sensitive than longer ones, provide larger frequency bins that allow the signal frequency to wander more, and thus are more sensitive to deviations from the monochromatic signal model.
-In LABEL:fig:tfmaps, we show the core stage of this method, which involves creating the time-frequency peakmap from raw time-domain strain data, correcting for the Doppler modulation from a particular sky location, and then integrating over time to collect all peaks at a given frequency to highlight the presence of a signal.
+This strategy is similar to that described in Section III.1.2, but here the modulation arises from the Doppler effect rather than from the random velocity distribution of individual DM particles.
+In practice, the analysis is repeated using different coherence times $T_{\text{FFT}}$, which control the width of the frequency bins.
+Shorter $T_{\text{FFT}}$ values are less sensitive overall but more robust to small frequency drifts, making the search less dependent on specific signal-model assumptions.
+In Fig.
+15, we show the core stage of this method, which involves creating the time-frequency peakmap from raw time-domain strain data (discussed previously in Section III.1.2), correcting for the Doppler modulation from a particular sky location, and then integrating over time to collect all peaks at a given frequency to highlight the presence of a signal.
 Note that while we show the color to indicate the ratio of signal to noise power, the method does not actually use it: it only requires that a given point in the time-frequency plane is above a given threshold.
-When moving from LABEL:fig:afterdopp to LABEL:fig:pmproj, all time-frequency points above this threshold are given a weight of “1”, and the ones are summed over time in order to create this histogram.
+When moving from Fig.
+15 to Fig.
+15, all time-frequency points above this threshold are given a weight of one, and the ones are summed over time in order to create this histogram.
+Figure 15: This plot shows how the method described in Section VI.2.1 works for a simulated GW signal from an annihilating scalar boson cloud system.
+The physics behind this form of DM has been discussed in Section V.1.
+(a) The Doppler-modulated time-frequency peakmap of the GW signal coming from a particular position in the sky is created by Fourier transforming the GW strain data in chunks of length $T_{\text{FFT}}$, dividing the square of the FFT by the estimation of the power spectral density, and selecting local maxima above a threshold.
+(b) For a fixed sky position, the time-frequency peakmap is the Doppler corrected, which makes the signal monochromatic.
+(c) The projected time-frequency map is shown here, which corresponds to an integral over time in each frequency bin.
+Note that each time-frequency point is labeled simply as “one” if it is above a threshold, and is zero otherwise.
+This histogram thus is a count of the number of times at which the equalized power exceeded a given threshold.
 VI.2.2 Viterbi
-Another method that is more robust against stochastic frequency variations relies on Hidden Markov Models (HMMs) to find arbitrary tracks in the time-frequency plane.
+Another method that is more robust against stochastic frequency variations relies on Hidden Markov Models (HMMs) to find tracks in the time-frequency plane.
 The main idea is to model a time-frequency GW signal probabilistically as a Markov chain of transitions between “hidden” (unobservable) frequency states and use a detection statistic to relate “observed” frequency states with the hidden ones Quinn and Hannan (2001).
 In the framework for this method, the probability to jump to a new state at a given time only depends upon the previous hidden state.
-In total, there would be $N_{Q}^{N_{T}+1}$ possible paths through the hidden states, where $N_{Q}$ is the total number of hidden states and $N_{T}$ is the total number of times at which we have observed frequency states.The Viterbi algorithm provides a recursive, computationally efficient way of maximizing the probability that the hidden set of states is responsible for the observed sequence of states.
-In LABEL:fig:sample-path, taken from Isi et al.
-(2019), we show the optimal path found by the Viterbi method for injected boson cloud signals in white noise that has an associated random walk of the frequency, which is meant to simulate unknown theoretical features of the signal.
-We can see that in both cases, with less and more variability of the signal, the Viterbi algorithm can find the appropriate track of the signal.
+In total, there would be $N_{Q}^{N_{T}+1}$ possible paths through the hidden states, where $N_{Q}$ is the total number of hidden states and $N_{T}$ is the total number of times at which we have observed frequency states.
+The Viterbi algorithm provides a recursive, computationally efficient way of maximizing the probability that the hidden set of states is responsible for the observed sequence of states.
+In Fig.
+16, taken from Isi et al.
+(2019), we show the optimal path found by the Viterbi method for injected boson cloud signals in Gaussian noise that has an associated random walk of the frequency, which is meant to simulate unknown theoretical features of the signal.
+We can see that in both cases, with less and more variability of the signal, the Viterbi algorithm can find the appropriate track.
 Recently, the Viterbi algorithm has also been adapted to search for GWs from annihilating vector boson clouds, which have timescales significantly shorter than in the scalar case Jones et al.
@@
 (2023b).
+(a) Less variability
+(b) More variability
+Figure 16: Taken from Isi et al.
+(2019).
+This figure illustrates how the Viterbi HMM algorithm described in Section VI.2.2 works to identify tracks in the time-frequency plane that would arise from GW emission from annihilating boson clouds around rotating BHs.
+The physics behind this form of DM has been discussed in Section V.1.
+In particular, these plots show two examples of injected and recovered boson cloud signals with random walk frequency variations $|\delta f|\leq 0.1\Delta f$ (weak variation) and (b) $|\delta f|\leq 0.5\Delta f$ (strong variation), where $\Delta f=1/T_{\rm drift}$ is the frequency bin size for the coherence time of the analysis $T_{\rm drift}=8$ days.
+The top panels show the injected random walk frequency evolutions, which cannot be seen by eye in the time-frequency tracks.
+The Viterbi optimal path results in a good match to the injected signal path.
+Each step in time on the $x$-axis is 8 days.
+The signal has a frequency of $201.2$ Hz, a spin-up of $10^{-12}$ Hz/s, a randomly chosen sky position, and a strain amplitude of $5\times 10^{-26}$.
 VI.2.3 Cross-correlation and other methods
 We have already discussed in Section III.1.1 cross correlating detector data to search for ultralight dark matter that could interact directly with GW interferometers; however, it is important to note that such a method could also be used to search for boson clouds around rotating BHs, both deterministically and for a stochastic GW background composed of the superposition of cosmological or astrophysical GW signals.
-Furthermore, virtually any method in the CW community could be tuned to look for such boson cloud systems because such methods already analyze almost monochromatic signals focusing on spin-downs but could also extend their parameter spaces to include spin-ups.
-VI.3 Recent search results for scalar boson clouds
-We can choose different BHs to point at depending on whether we are looking for scalar or vector clouds, or annihilation or transition signals.
-Current GW interferometer and algorithm sensitivities indicate that annihilation signals from scalar boson clouds could be observed for galactic and relatively young black holes Abbott et al.
-(2022d).
-Or, particular, known black hole systems, such as Cyg-X1, could be targeted for such clouds as well Sun et al.
-(2020).
-Finally, the stochastic GW background, composed of the superposition of signals from all of the annihilating boson cloud systems in the universe, can be probed as well through cross-correlation analyses.
-We go through each of these scenarios in the following subsections.
+Furthermore, virtually any method in the CW community Riles (2023) could be tuned to look for such boson cloud systems because such methods already analyze almost monochromatic signals focusing on spin-downs but could also extend their parameter spaces to include spin-ups.
+VI.3 Search results for scalar boson clouds
+Data from the first, second and third observing runs of LIGO, Virgo and KAGRA have been utilized to place constraints on the presence of boson clouds around rotating BHs within our galaxy.
+One search adopted a general approach, targeting systems across the entire sky.
+Another focused specifically on GW emission from a boson cloud around Cygnus X-1, and disfavored a range of boson masses.
+A third search investigated whether luminous dark photon clouds around stellar-mass BHs could mimic known pulsars, setting constraints on the dark-photon mass and its kinetic mixing with the photon.
+A fourth analysis estimated the stochastic GW backgrounds resulting from the combined effects of scalar, vector, and tensor boson clouds, examining the data for each component individually.
+Lastly, a fifth study used the masses and spins of observed binary black hole mergers to constrain the BH spin distribution and the mass of the boson.
+Each of these searches is detailed in the subsections that follow.
 VI.3.1 All-sky search constraints
 The excess power method described in Section VI.2.1 has been used to analyze LIGO O3 data to search for the presence of annihilating scalar boson clouds around isolated BHs within our galaxy Abbott et al.
-(2022d).
+(2022d, e).
 This search was very computationally expensive, since each sky position had to be analyzed individually.
-A separate search towards the galactic center was also performed, targeting both spinning down neutron stars and spinning up boson cloud systems Abbott et al.
-(2022e) No evidence for scalar boson cloud systems were found in either case, and thus upper limits were placed on the presence of boson clouds with different ages, spins, and distances away from earth in two different ways Palomba et al.
+No evidence for scalar boson cloud systems was found in either case, and thus upper limits were placed on the presence of boson clouds with different ages, spins, and distances away from earth in two different ways Palomba et al.
 (2019); Dergachev and Papa (2019); Abbott et al.
@@
     Assume mass and spin distributions for BHs in the universe, draw different BHs from these distributions, and determine the maximum distance reach that could be attained for different boson cloud masses.
-The first way is agnostic towards whether such BHs exist in the galaxy, while the second one contains some (uncertain) astrophysics regarding the properties of astrophysical BHs, but allows for some physical intuition into the reach of such searches.
+The first way is agnostic towards whether such BHs exist in the galaxy, while the second one contains some (uncertain) astrophysics regarding the mass and spin distributions of astrophysical BHs, but allows for some physical intuition into the reach of such searches.
 Constraints from the first way are shown in Fig.
@@
 Using upper limits from the O3 LIGO search for GWs from annihilating boson cloud systems, exclusion region of particular boson mass /BH mass pairs are derived assuming different ages and distances (left: 1kpc; right: 15 kpc) for BHs with initial spins of $\chi_{i}=0.9$.
-This is the most model-agnostic way to probe boson clouds around rotating BHs using upper limits obtained from an all-sky search for CWs from isolated BHs.
+The physics behind this form of DM has been discussed in Section V.1.
+This is the most model-agnostic way to probe boson clouds around rotating BHs: the limits do not assume distributions over the mass, spin, distance of BHs in the Galaxy.
 Constraints from the second way are shown in Fig.
-19, in which a Kroupa BH mass distribution $f(m)\propto M_{\rm BH}^{-2.3}$ Elbert et al.
-(2018) between $[5,~{}50]M_{\odot}$ and $[5,~{}100]M_{\odot}$ is assumed, as well as a uniform initial BH spin distribution between [0.2,0.9].
-For fixed BH age, the maximum distance reach as a function of boson mass indicates that at least 5% of GWs from BH/boson cloud systems would have been detected by the search, i.e.
-their strain amplitudes are larger than the minimum detectable amplitudes at a given confidence level (the upper limits) of the search.
+18, in which a Kroupa BH mass distribution $f(m)\propto M_{\rm BH}^{-2.3}$ Elbert et al.
+(2018) between $5,~50]M_{\odot}$ and $5,~100]M_{\odot}$ is assumed, as well as a uniform initial BH spin distribution between [0.2,0.9].
+For a fixed BH age, the maximum distance reach as a function of boson mass indicates that at least 5% of GWs from BH/boson cloud systems would have been detected by the search, i.e.
+their strain amplitudes are larger than the minimum detectable amplitudes at a given confidence level (the upper limits) of the search (an example of this quantity is in Section III.1.2, Eq.
+71.
+).
 As expected, on average when $50M_{\odot}$ is the maximum BH mass considered, the distance that can be reached is also smaller, since the GW amplitudes scales strongly with the BH mass.
@@
 (2022d) could reach at which at least 5% of a simulated BH population would have been detectable, that is, with a strain greater than the value of the upper limit, is shown here.
+The physics behind this form of DM has been discussed in Section V.1.
 To create this plot, Ref.
@@
 (2022d) assumed a uniform spin distribution between [0.2,0.9], different ages, and a Kroupa mass distribution with maximum mass $50M_{\odot}$ (left) and $100M_{\odot}$ (right).
+These limits depend on the distributions assumed and are thus less agnostic than those presented in Fig.
+17.
 We note that the search constraints present in, for example, Steltner et al.
@@
 Since the sky position is known, searches based on the Viterbi algorithm discussed in Section VI.2.2 can use much longer coherence times than those used in Abbott et al.
-(2022d), thus permitting exquisite sensitivity towards almost monochromatic signal, while also being robust to theoretical uncertainties in boson cloud source.
+(2022d), thus permitting exquisite sensitivity towards almost monochromatic signals, while also being robust against theoretical uncertainties in the GW emission from Cygnus X-1, e.g.
+due to uncertainties on the measured semi-major axis, etc., due to the presence of an accretion disk, etc.
 In Fig.
-20, we show constraints on the strain amplitude arising from annihilating scalar boson cloud particles as a function of the boson mass.
-The search results $h_{0}^{95\%}$ are given as the black line, while the two curves represent the numerically obtained strain from this system, accounting for the spin, age, mass and distance of Cyg x-1 ($t_{\rm age}=5\times 10^{6}$ yr and $1\times 10^{5}$ yr).
-The region in which the search results are below the theoretical curves denote boson masses that are disfavored as having formed around Cyg x-1.
+19, we show constraints on the strain amplitude arising from annihilating scalar boson cloud particles as a function of the boson mass using LIGO O2 data.
+The search results $h_{0}^{95\%}$ are given as the black line, while the two curves represent the numerically obtained strain from this system, accounting for the spin, age, mass and distance of Cygnus X-1 ($t_{\rm age}=5\times 10^{6}$ yr and $1\times 10^{5}$ yr).
+The region in which the search results are below the theoretical curves denote boson masses that are disfavored as having formed around Cygnus X-1.
 All constraints presented so far do not assume any boson self-interactions.
 If self interactions do occur, the available parameter space to probe would change.
-But, if the self-coupling is weak enough, standard searches can probe this scenario without modification.
-An example for the string axiverse scenario is shown in Fig.
-21.
-The white region indicates the portion of the parameter space in which bosenovas do not occur, and thus cannot be constrained with the current search, while, again, portions of the colored parameter space above the black line indicate that those boson masses are disfavored.
-However, these constraints contain a number of caveats, namely: (1) the periodic bosenovas do not alter the time-frequency evolution of the GW signal, though they actually might induce oscillations every few minutes Yoshino and Kodama (2015a), and (2) the signal must last for longer than the age of the BH.
-Future studies are needed to determine to what extent CW searches can actually probe this particular model for string axion boson clouds around BHs, as well as to incorporate self-interactions more systematically into how CW searches are performed.
+But, if the self-coupling is weak enough, standard methods discussed in Section VI.2 can probe this scenario without modification.
+We show in Fig.
+20 constraints on GWs arising from a scalar boson cloud around Cygnus X-1 that have self-interactions.
+Each panel corresponds to a particular physical mechanism: annihilation in the 211 energy level (left); annihilation in the 322 energy level (middle); and transitions between 322 and 211 energy levels (right).
+The dashed curves divide the plot into different self-interacting regimes, from top to bottom: strong self-interactions, intermediate self-interactions, and the gravitational regime (negligible self-interactions).
+The energy level labeling correspond to the quantum numbers $n,l,m$, respectively, where $n$ is the principal quantum number, $l$ is the total angular momentum quantum number, and $m$ is the azimuthal quantum number.
+The color indicates the expected signal amplitude, which, for different detectors (different colored lines), is mapped to projected constraints on the axion decay constant.
+The white line indicates the axion decay constant that is fixed by the axion mass.
+Constraints on self-interacting DM using the search results on O2 data in Sun et al.
+(2020) are shown only in the left-hand plot, since that search was designed to analyze 211 annihilation from Cygnus X-1.
+Furthermore, current and future ground-based interferometers will be able to place constraints on the axion decay constant at frequencies of $\mathcal{O}(100)$ Hz (left plot), while space-based interferometers, such as DECIGO and MAGIS-Space are required to probe transition signals at frequencies below 1 Hz (right panel).
+The figure also highlights the possibility of multi-band GW astronomy: it would be possible to detect a transition from the 322 to 211 energy levels in space-based observatories while simultaneously observing the CW arising from annihilation of bosons in the 211 energy level in ground-based interferometers Collaviti et al.
+(2025).
+Annihilation signals from the 322 level (middle plot) would be too weak to be detected by any of the considered current or future GW interferometers.
 Figure 19: Taken from Sun et al.
-(2020), a search using O2 LIGO data for annihilating boson cloud systems around Cyg X-1.
-(a) Upper limits at 95% confidence on the strain amplitude of annihilating scalar boson clouds coming from the BH in Cyg X-1.
+(2020).
+Upper limits at 95% confidence on the strain amplitude of annihilating scalar boson clouds coming from the BH in Cygnus X-1.
+The physics behind this form of DM has been discussed in Section V.1.
 The black curve indicates the upper limits from the search on O2 data, with the red stars denoting upper limits derived in frequency bands with injections.
-The two different curves correspond to different choices for the (unknown) age of the system (lower curve: $10^{6}$ years; upper curve: $5\times 10^{6}$ years.
-The GW fine structure constant is colored, and the gray region denotes boson masses disfavored under the two different age assumptions.
-$M=14.8M_{\odot}$, $\chi_{i}=0.99$, and $d=1.86$ kpc for the two curves.
+The two different curves correspond to the signal amplitude with different choices for the (unknown) age of the system (lower curve: $10^{6}$ years; upper curve: $5\times 10^{6}$ years.
+The GW fine structure constant is colored, and the gray regions denote boson masses disfavored under the two different age assumptions.
+The parameters of Cygnux X-1 used to compute the amplitude curves are: $M=14.8M_{\odot}$, $\chi_{i}=0.99$, and $d=1.86$ kpc.
 Figure 20: Taken from Collaviti et al.
 (2025).
-Constraints on the string axiverse scenario (in blue from O2), in particular the axion decay constant.
-The regions are divided into three regimes: (1) strong-interactions (above the orange dashed-line), (2) intermediate interactions (between the dashed-blue and dashed orange lines) and no self-interactions (below the dashed-blue line).
-The three plots correspond to annihilations in different energy levels (left and middle) and to transitions between these energy levels (right).
-Sensitivities of future detectors are also shown on this plot for comparison.
-VI.3.3 Constraints on luminous dark photons clouds
-As alluded to in Section V, if the dark photon DM kinetically mixes with the ordinary photon and forms clouds around rotating BHs, it could produce both electromagnetic and GW signals, which, if detected, could enable a major multi-messenger discovery Siemonsen et al.
+Parameter space reachable by current and next-generation GW interferometers for (left) 211 annihilation, (middle) 322 annihilation, and (right) 322 $\rightarrow$ 211 transition signals from a scalar boson cloud around Cygnus X-1, assuming an age of $10^{5}$ yr and the parameters for Cygnus X-1 given in Fig.
+19.
+The numbers 211, 322, and the 322 $\rightarrow$ 211 transition refer to specific energy levels of the cloud.
+The physics behind this form of DM has been discussed in Sections V.1 and V.2.
+The dashed curves divide the parameter space into three regimes: (1) top, above orange-dashed line: the strong self-interaction regime, (2) middle, between dashed-orange and dashed-blue lines: the moderate self-interaction regime, and (3) bottom, below the dashed-blue line: the gravitational (negligible self-interactions) regime.
+The white dotted line shows the expected relation between the axion and its decay constant.
+Darker colors here correspond to weaker GW signals primarily due to smaller occupation numbers in the energy states of the boson clouds.
+Constraints on 211 annihilation come from ground-based interferometers, while constraints from transitions will come from space-based interferometers.
+The 322 annihilation signal is too weak to be detected by any current our future observatory.
+The O2 constraints come from a real analysis of GW data and are thus representative of both the actual power spectral density of that dataset and an exclusion region on the plot.
+VI.3.3 Constraints on luminous dark photon clouds
+As alluded to in Section V, if dark photon DM kinetically mixes with the ordinary photon and forms clouds around rotating BHs, it would produce both electromagnetic and GW signals, which, if detected, could enable a major multi-messenger discovery Siemonsen et al.
 (2023a).
@@
 This plasma emits copious amounts of electromagnetic radiation, sourced by energy injections through dissipative processes in the cloud (e.g.
-at magnetic reconnection sites, Landau damping, turbulence), potentially even periodically.
-The luminosity of these sources depends on the strength of the kinetic mixing parameter, but could be at the level of supernovae or known pulsars, and it could be periodic Siemonsen et al.
+at magnetic reconnection sites, Landau damping, turbulence).
+The luminosity of these sources depends on the strength of the kinetic mixing parameter, but could be at the level of supernovae or known pulsars, and and also periodic Siemonsen et al.
 (2023a).
-If the dark photon has a mass around $\sim 10^{-12}$ eV, it will induce the formation of clouds around stellar-mass BHs, and give off radio emission that would look remarkably similar to, and potentially indistinguishable from, known pulsars in the ATNF catalog Manchester et al.
+If the dark photon has a mass around $\sim 10^{-12}$ eV, it will induce the formation of clouds around stellar-mass BHs, and give off radio emission that would look remarkably similar to that from known pulsars in the ATNF catalog Manchester et al.
 (2005).
-The lack of observation of CWs from pulsars, and the potential of dark-photon clouds to act as “pulsar mimickers”, motivates a search for such clouds around targets in the ATNF catalog that could be mislabeled as pulsars.
-In Ref.
+The absence of observed CWs from pulsars, combined with the possibility that dark-photon clouds could act as “pulsar mimickers,” motivates searches for such clouds around sources in the ATNF catalog that might be misidentified as pulsars.
+Ref.
 Mirasola et al.
-(2025), the authors consider this possibilty by analyzing a subset of, spinning up “pulsars” from the ATNF catalog that might actually be but stellar-mass BHs with a dark-photon luminous cloud around them.
-Using a variety of methods – one that assumes perfect phase coherence between the observed electromagnetic radiation and GW emission, one that allows small deviations in frequency around a narrow band of the observed frequency of each compact object, and one that is most robust against theoretical uncertainties in the signal but least sensitive to GW emission –, upper limits on the dark photon/ordinary photon kinetic mixing parameter are set as a function of boson mass for the 34 sources considered in this analysis, as shown in Fig.
-22.
-To obtain this plot, particular BH populations and electromagnetic emissions are assumed to determine the number of events on the color bar, so the bounds are not necessarily robust against changes in these choices.
-For the BH population, the relevant distributions are over position in the galaxy (black holes follow the stellar distribution) and mass (Salpeter), spin (uniform) and age (uniform) of the BH systems.
-Though particular population models are assumed, the constraints only require ten “events” to have been observable, which is conservative with respect to the potentially $\mathcal{O}(10^{3})$ that would be expected at such coupling/mass combinations, as shown in the right-hand panel of Fig.
+(2025) explores this scenario by analyzing a subset of spinning-up “pulsars” in the catalog that could instead be stellar-mass BHs surrounded by luminous dark-photon clouds.
+Three analysis strategies are considered: one assuming perfect phase coherence between the electromagnetic emission and any associated GW signal; one allowing for small frequency deviations around the measured pulsar frequency; and one that is most robust to theoretical uncertainties but least sensitive to GW emission.
+From these methods, the authors derive upper limits on the dark-photon–photon kinetic mixing parameter as a function of boson mass for the 34 sources studied, shown in Fig.
+21.
+To produce these bounds, assumptions are made about both the BH population and their electromagnetic signatures, so the precise results are not fully robust against variations in these choices.
+The adopted BH distributions include a stellar-tracing spatial distribution, a Salpeter mass function, uniform spin and age distributions, and a total galactic population of $10^{8}$ BHs.
+Even under these assumptions, the analysis requires only $\sim 10$ detectable “events” (i.e., BHs with GW amplitudes exceeding the strain upper limit), which is conservative compared to the $\mathcal{O}(10^{3})$ events expected at certain coupling–mass values (see Fig.
 3 of Mirasola et al.
-(2025).
-Here, “events” just means the number of rotating BHs that could have given off a detectable signal, that is, with a GW amplitude larger than the upper limit strain amplitude, in an assumed population of $10^{8}$ BHs in the galaxy.
-In the end, a dark photon mass range between, roughly $[10^{-13},10^{-12}]$ eV could not have a kinetic mixing more than roughly $[10^{-9},10^{-7}]$.
+(2025)).
+Ultimately, the study excludes dark-photon masses in the range $\sim[10^{-13},10^{-12}]$ eV with kinetic mixing larger than roughly $10^{-9},10^{-7}]$.
 Figure 21: Taken from Mirasola et al.
 (2025).
-For each of the targets of the search in Mirasola et al.
+For each of the radio pulsar mimickers targeted in Mirasola et al.
 (2025), the upper limit at 95% confidence on the kinetic mixing parameter $\epsilon$ is plotted as a function of the boson mass for which at least 10 systems would have been detected with signal amplitudes above the strain upper limits.
-The color corresponds to the number of events that should have been detected, assuming a “pessimistic” population model for BHs.
+Each vertical line corresponds to one radio pulsar, whose frequency fixes the mass of the dark photon.
+The color corresponds to the number of events that should have been detected, assuming a “optimistic” population model for BHs, i.e.
+using a larger maximum mass and spin for the assumed BH population with respect to the “pessimistic” population also constrained.
 The bounds reproduced from other experiments can be found in O’Hare (2020), specifically from the conversion of CMB photons into dark photons, which would cause distortions in the black body spectrum measured by COBE/FIRAS Mirizzi et al.
@@
 (2024) (light gray).
+The upper limits arising from this search surpass those from existing experiments, and indicate that between $10$ and $10^{4}$ events could have been detected at couplings of between $\sim[10^{-10},10^{-7}]$.
 VI.3.4 Stochastic gravitational-wave background constraints
@@
 In Tsukada et al.
-(2019), the authors designed a model of the stochastic GW background arising from isolated BHs formed in two ways: (1) from remnants of core collapse supernovae, and (2) from remnants of mergers of BHs or neutron stars.
-The authors also performed search on LIGO O1 data using cross-correlation to look for evidence of a stochastic GW background Tsukada et al.
+(2019), the authors designed two models of the source of this stochastic GW background: (1) from remnants of core collapse supernovae, and (2) from remnants of mergers of BHs or neutron stars.
+The authors also performed search on LIGO O1 data using cross-correlation to look for evidence of a stochastic GW background from scalar boson clouds Tsukada et al.
 (2019).
 No signal was found, and no range of boson masses could be robustly excluded unfortunately, except under optimistic assumptions about BH formation rates and spin distributions, which would disfavor the regime $2\times 10^{-13}\leq m_{b}\leq 3.8\times 10^{-13}$ eV at 95% confidence.
-The major reason for the inability to robustly exclude boson masses is that there exist degeneracies between the boson mass and BH astrophysical quantities, such as spins.
+Here, “optimistic’ means assuming rapidly spinning and frequently forming BHs in the galaxy.
+This limitation highlighted the degeneracy between boson mass and poorly constrained astrophysical parameters, particularly BH spins.
 The search was repeated in the second observing run as well, finding no evidence for any stochastic GW background of vector boson clouds around rotating BHs Tsukada et al.
-(2021), in particular disfavoring vector boson cloud masses between $[0.8,6]\times 10^{-13}$ eV at 95% confidence.
-Additionally, Ref.
+(2021), in particular disfavoring vector boson cloud masses between $0.8,6]\times 10^{-13}$ eV at 95% confidence.
+Ref.
 Yuan et al.
-(2022) has also disfavored bosons with masses between $[1.5,17]\times 10^{-13}$ eV depending upon what is assumed for the uniform spin distribution, with wider constraints present when higher-spinning BHs dominate the priors.
-Furthermore, the stochastic GW background from tensor bosons around rotating BHs has been computed and searched for in Guo et al.
-(2024b), and has disfavored boson masses between $[10^{-13.4},10{-11.7}]$ eV depending on the range of uniform spin distributions assumed Guo et al.
-(2024b).
+(2022) has also disfavored bosons with masses between $1.5,17]\times 10^{-13}$ eV depending upon what is assumed for the uniform spin distribution, with wider constraints present when higher-spinning BHs dominate the priors.
+Beyond scalar and vector cases, Ref.
+Guo et al.
+(2024b) investigated tensor bosons, computing the corresponding stochastic GW background and performing a dedicated search.
+Their analysis disfavored boson masses between $10^{-13.4}$ and $10^{-11.7}$ eV, again with the precise range depending on the assumed range of uniform spin distribution .
+Taken together, these studies demonstrate that the stochastic GW background provides a powerful probe of ultralight scalar, vector, and tensor clouds, but the strength of the resulting constraints depends strongly on assumptions about black-hole formation, mass, and spin distributions.
+This mirrors the caveats discussed in Section VI.3.1, where similar assumptions underlie the derived limits.
+Moving forward, it is therefore essential to develop robust and consistent methods for setting constraints that minimize reliance on specific black-hole population parameters while retaining astrophysical interpretability.
+VI.3.5 Constraints from the spins of detected binary black holes
+We note that constraints on boson clouds can be inferred from the spin measurements of binary black hole systems Arvanitaki et al.
+(2015a, 2017); Brito et al.
+(2017); Baryakhtar et al.
+(2017); Cardoso et al.
+(2018); Ghosh and Sachdeva (2021).
+Recently, a Bayesian method has been developed by combining the information from each system Ng et al.
+(2021b), assuming a distribution for the initial BH spins and requiring that the BHs have had enough time to undergo superradiance before merging.
+Such a study has been applied to the second GW transient catalog (GWTC-2) Abbott et al.
+(2021b), and has resulted in essentially a joint posterior on the scalar boson mass and the spin distribution of BHs at their formation Ng et al.
+(2021a).
+The latter is required because BHs born with high spins that have been spun down by superradiance is partially degenerate with BHs simply born with low spins in a universe without superradiance.
+The study presented in Ng et al.
+(2021a) tentatively rules out scalar bosons in the mass range $1.3,2.7]\times 10^{-13}$ eV.
+These constraints from measured BH spins can disfavor particular masses or even confirm the presence of an ultralight boson; however, they rely on assumptions about the spin distribution at formation, on whether the boson clouds have time to form (which could be impacted by eccentricity at formation Peters (1964); Wen and Gair (2005)), and are a function of the (strongest) BH mergers that we see with GWs.
+Thus, these kinds of analyses complement those previously that look for CWs around isolated, rotating BHs.
+Both have their advantages: CW analyses described above allow us to probe unknown sources and known, nearby sources whose signals we expect to be stronger than those arising from remnants of mergers.
+Likewise, the spins and masses of BHs formed in binary system are well measured, and the constraints obtained in Ng et al.
+(2021a) will only improve as we detect more and more BHs.
 VI.4 Prospects for GW probes of boson clouds
@@
 While having accurate measurements of the mass, spin and age of remnants of binary black hole mergers would allow us to ignore astrophysical population models Arvanitaki et al.
-(2015a), these systems are currently too far to be detected if the clouds are composed of scalar bosons Isi et al.
+(2015b), these systems are currently too far to be detected if the clouds are composed of scalar bosons Isi et al.
 (2019), though could be seen up to Gpc now if they are composed of vector bosons Jones et al.
@@
 On the flipside, GWs from vector boson clouds suffer from even more theoretical uncertainties than scalars Siemonsen and East (2020), and so the robustness of search results by targeting remnants would have to be evaluated.
-All-sky searches for such boson cloud systems could also be enhanced.
-At the moment, especially in light of boson self-interactions that could alter the GW signal significantly.
+All-sky searches for such boson cloud systems could also be enhanced, especially in light of boson self-interactions that could alter the GW signal significantly as discussed in Section V.3.
 Potentially, new methods to track quickly-evolving waveforms, such as those in Jones et al.
 (2023, 2025), could be useful in a variety of contexts.
-Furthermore, the all-sky search specifically for scalar boson cloud systems did not consider any grid in spin-up, which limited it to be sensitive to boson clouds whose frequency drift did not exceed one frequency bin during the entire observation time.
-Such modifications to searches, coupled with additional theoretical understanding of the GW signal itself, could increase chances of detecting a signal.
+Furthermore, the all-sky search specifically for scalar boson cloud systems did not consider the possibility that the cloud could spin-up, which, in practice, fixed a maximum spin-up to which the search was sensitive (such that the spin-up did not induce a freqeuncy shift by more than one bin $\delta f$ in $T_{\text{obs}}$, i.e.
+$\dot{f}T_{\text{obs}}<\delta f\equiv 1/T_{\text{FFT}}$).
+Such modifications to searches, coupled with additional theoretical understanding of the GW signal itself, could increase chances of detecting a signal by exploring a wider parameter space and allowing the cloud to contract over time.
 It is also difficult to say with certainty that a particular boson mass is “ruled out”, since the constraints are either population agonistic or depend on some assumptions about BH formation, spin or mass distributions.
@@
 (2024) attempts to unify the way in which constraints are set in a Bayesian way by handling the subtleties associated with population assumptions.
-The authors encourage that posteriors on mass and spin be made publicly available in order to incorporate into their future analyses.
-Furthermore, Ref.
-Guo et al.
-(2024c) shows that bosons could be transferred from a growing level to a depleting one if both BHs in a binary system have clouds around them, which may affect the GW waveform and hence would need to be considered in future GW searches.
-VII GW probes of soliton DM
-If DM has a mass of $\sim 10^{-24}-10^{-22}$ eV Hui et al.
-(2017), its Compton wavelength will be comparable to the size of the galactic halo, which actually limits the minimum mass of DM Sin (1994).
-In this mass regime DM could form so-called “solitons”, self-gravitating structures prevented from collapsing from degeneracy pressure, at the centers of galaxies Schive et al.
+The authors in Ref.
+Hoof et al.
+(2024) encourage that posteriors on mass and spin, obtained through analyses of X-ray data from BHs, be made publicly available in order to incorporate into their future work.
+VII GW probes of soliton dark matter
+Ultralight DM in the mass range $\sim[10^{-24},10^{-22}]$ eV has a Compton wavelength comparable to the size of galactic halos Hui et al.
+(2017); Sin (1994).
+In this regime, DM can form self-gravitating structures, or solitons, at the centers of galaxies, which remain stable due to quantum degeneracy pressure Schive et al.
 (2014).
-To probe this kind of DM, these masses imply that the oscillation frequencies will be of $\mathcal{O}(\text{nHz})$, making pulsar timing arrays the ideal probe Aghaie et al.
+A natural approach to probing this mass range of DM is through pulsar timing arrays, whose frequency sensitivity aligns with the nanohertz oscillation frequencies expected from DM.
+In this case, solitons can exert dynamical friction on binary systems, accelerating their inspiral and suppressing the gravitational wave strain amplitude at nanohertz frequencies Aghaie et al.
 (2024).
-In essence, DM would affect the waveform arising from the superposition of inspiraling supermassive BHs, causing the BHs to experience dynamical friction, thus weaking the signal strain in the nHz regime.
-But this is not the only way to detect this mass regime of ultralight dark matter: as we will show, LISA and next-generation ground-based GW interferometers could also be used.
-To probe this kind of DM with conventional GW sources, low-frequency information needs to be contained in relatively higher-frequency carrier GW.
-The GW frequency becomes modified during its propagation between its source and the earth as it passes through the DM soliton, i.e.
-it is “heterodyned”, or gravitationally redshifted, analogously to the Sachs-Wolfe effect that induces fluctuations in the cosmic microwave background Sachs and Wolfe (1967).
-VII.1 Impact on phase evolution of binary systems
+Alternatively, soliton DM can be framed as a low-frequency modulation of relatively higher-frequency GW.
+The GW frequency is modified during its propagation through the DM halo, undergoing a gravitational redshift or ’heterodyning’, analogous to the Sachs-Wolfe effect that imprints low-frequency modulations on high-frequency photons in the cosmic microwave background Sachs and Wolfe (1967).
+Two CW sources have been considered as high-frequency carriers for this modulation Blas et al.
+(2025): (1) inspiraling white-dwarf binaries, observable by space-based GW interferometers, and (2) non-axisymmetric rotating neutron stars that spin down over time.
+In what follows, we explore how these different probes can be used to constrain the properties of soliton.
+In Section VII.1, we demonstrate how data from pulsar timing arrays can be used to set constraints on soliton DM.
+We then explore two other approaches in Section VII.2 and Section VII.3: detecting CWs from white-dwarf binaries and isolated neutron stars, which could reveal nanohertz modulations induced by soliton DM.
+VII.1 Constraints from pulsar timing arrays
+The evidence for a GW background from NANOGrav results can be used to constrain soliton DM Aghaie et al.
+(2024).
+Essentially, ultralight dark matter would induce dynamical friction in the inspiral of supermassive BHs, speeding up the orbit and suppressing the GW strain at nanohertz frequencies.
+Using the most accurate limit at 3.92 nHz on the GW strain from supermassive BHs in the right-hand panel of Fig.
+1 of Afzal et al.
+(2023), Ref.
+Aghaie et al.
+(2024) derives constraints on the soliton DM mass-to-BH mass ratio parameter space, as shown in Fig.
+22.
+These constraints are obtained using Eq.
+17 from Aghaie et al.
+(2024), which models energy loss due to GW emission in the presence of ultralight dark matter.
+This equation incorporates astrophysical inputs such as galaxy-stellar mass functions, galaxy merger rates, and galaxy pair fractions Chen et al.
+(2019).
+The constraints are obtained by integrating Eq.
+17 over these distributions, taking into account variations in the assumptions about the ultralight dark-matter mass and the population of supermassive BHs.
+However, these constraints are sensitive to assumptions regarding astrophysical parameters.
+For example, the specific values chosen for galaxy merger rates, stellar mass functions, and the fraction of galaxies with supermassive BHs affect the final upper limits.
+Figure 22: Taken from Aghaie et al.
+(2024).
+Upper limits on the DM mass and soliton-to-supermassive black hole mass ratio parameter space, derived from NANOGrav data.
+These constraints are based on the notion that the GW background from the inspiral of supermassive BHs, for which NANOGrav has found evidence, would be modulated by the dynamical friction exerted by soliton dark matter.
+Specifically, the presence of soliton DM alters the supermassive BH inspiral dynamics, resulting in a suppression of the GW strain at nanohertz frequencies.
+The magenta lines represent two distinct sets of scaling relations derived from numerical simulations: solid lines correspond to DM-only simulations, where the soliton DM is the sole interacting component, while dashed lines incorporate the influence of supermassive BHs, suggesting a modification of the scaling relations due to the presence of these massive objects.
+VII.2 Dynamical friction on white dwarf binaries
 Space-based GW detectors will observe the quasi-infinite inspiral of galactic white dwarf binaries and extreme-mass ratio inspirals of ordinary compact objects with much heavier BHs Bartolo et al.
@@
 (2025).
-Ref Brax et al.
+Ref.
+Brax et al.
 (2024) has found that the impact of dynamical friction on the phase evolution of GW signals from white dwarf binaries exceeds that of the low-frequency DM-induced modulation for masses above $10^{-21}$ eV, essentially providing two orders of magnitude of mass that could be probed in the future.
@@
 However, the mass regime $\sim[2\times 10^{-22},3\times 10^{-21}]$ eV could be constrained if DM has stronger, quadratic couplings to ordinary matter.
-VII.2 Imprint of DM on CW signal from neutron star
-In addition to modulating the millihertz frequency of GWs from binaries in space-based GW detectors, soliton ultralight dark matter could induce changes in the CWs arising from deformed neutron stars, thus allowing the possibility of detecting this effect in future ground-based detectors Blas et al.
+Again, the possibility that DM couples to the standard model in some way is able to enhance the signal strength, as found in Sections II and IV for different types of DM.
+VII.3 Nanohertz modulations of continuous gravitational waves from neutron stars
+In addition to modulating the millihertz frequency of GWs from binaries in space-based GW detectors, soliton ultralight dark matter could induce modulations in the rotational frequency of neutron stars in a way distinct from ordinary spinning down, deformed neutron stars, thus affecting CW signals and allowing the possibility of detecting this effect in future ground-based detectors Blas et al.
 (2025).
 These GWs are sourced from the rotational power of the neutron stars, thus spinning down the neutron stars as GWs are emitted Riles (2017).
-In ET and Cosmic Explorer, we expect to detect $\mathcal{O}(100)$ or more neutron stars spinning down due to the emission of CWs Pagliaro et al.
-(2023), all of which could be modulated by ultralight dark matter solitons.
+In ET and Cosmic Explorer, we expect to detect $\mathcal{O}(100)$ or more canonical neutron stars spinning down due to the emission of CWs Pagliaro et al.
+(2023), which could be modulated by ultralight dark matter solitons.
 Using CWs from deformed neutron stars, the soliton mass regime $\sim[2\times 10^{-22},3\times 10^{-21}]$ eV becomes accessible in ET for both linear and quadratic couplings, and improves with the number of neutron stars that are detected.
-The constraints will even outperform existing pulsar timing array ones, since the signal-to-noise ratio scales linearly with the carrier GW frequency.
-This is shown in LABEL:fig:lambda12res for LISA, ET and Cosmic Explorer, and demonstrates the powerful probe that CWs from neutron stars will be of this particular DM model.
-VII.3 Pulsar-timing constraints on soliton ultralight dark matter
-The evidence for a GW background from NANOGrav results can be used to put constraints on soliton DM Aghaie et al.
+The constraints will even outperform existing pulsar timing array ones, since the SNR scales linearly with the carrier GW frequency.
+This is shown in Fig.
+23 for LISA, ET and Cosmic Explorer, and demonstrates the powerful probe that CWs from neutron stars will be of this particular DM model.
+Figure 23: Taken from Blas et al.
 (2025).
-Essentially, ultralight dark matter would induce dynamical friction in the inspiral of supermassive BHs, speeding up the orbit, which suppresses the GW strain at nanohertz frequencies.
-The supermassive BHs are too massive to be prevented from merging, as would be the case for lighter BHs Hui et al.
-(2017); Bar-Or et al.
-(2019).
-Using the most accurate limit at 3.92 nHz on the GW strain from supermassive BHs in the right-hand panel of Fig.
-1 of Afzal et al.
-(2023), Ref.
-Aghaie et al.
-(2024) derives constraints on the soltion DM mass-to-BH mass ratio parameter space, as shown in Fig.
-24.
-These bounds are derived by using Eq.
-17 of Aghaie et al.
-(2024), which depends on how much energy is lost via GW emission normally and in the presence of the ultralight dark matter soliton, and on the differential density populations of supermassive BHs.
-The latter can be written analytically in terms of the the galaxy-stellar mass functions, galaxy merger rates and galaxy pair fractions Chen et al.
-(2019).
-It is then a matter of integrating Eq.
-17 in Aghaie et al.
-(2024) over these different distributions (which may depend on decently uncertain astrophysics) for different choices of ultralight dark matter mass to derive the constraint on mass ratio of soliton DM to supermassive BH mass, and determine if the integrated strain exceeds the error bar on the GW strain obtained by NANOGrav.
-Figure 24: Taken from Aghaie et al.
-(2024).
-Upper limits from NANOGrav data on the DM mass and solition-BH mass ratio parameter space, obtained by noting that inspiraling supermassive BH background for which NANOGrav found evidence would be modulated by the dynamical friction induced by soliton DM.
-Magenta lines show the scaling relations derived through DM-only numerical simulations (solid), and ones in which supermassive BHs dominate scaling relations suggested by DM-only numerical simulations and SMBH dominance (dashed).
+Projected constraints on the linear (left) and quartic coupling parameters (right) for both LISA and ET from low-frequency soliton DM modulating (“heterodyning”) high-frequency carrier GW signals via changes to the moment of inertia of neutron stars, which alters CWs from deformed neutron stars (ET) and the GWs from inspiraling binary black holes (LISA).
+The curves for ET assume one or 200 detections of CWs from isolated neutron stars.
+The LISA curve is determined by the impact of soliton DM on GW frequencies from binary systems.
+The olive regions are excluded by the pulsar timing array analysis in Smarra et al.
+(2024), while blue and red are constrained away via Cassini tests of general relativity (violet) and on a stochastic GW background (red).
+We note that in this case, DM leaves a particular imprint on the GW signal itself, arising from deformed, isolated neutron stars.
+This differs from the discussion in Section II, where DM interacts directly with the components of the interferometer, causing differential strains that are distinct from the expected signal from GWs.
+In contrast, here we are looking for a deviation in a canonical GW source due to DM, rather than an interaction with the interferometer itself.
+Furthermore, though the effect described in Section II.3.2 of dilatons interacting with neutron stars is similar to that discussed here, it arises from a different physical mechanism.
+While dilatons alter the fundamental constants via a new coupling that leads to a oscillatory moment of inertia of neutron stars, solitons interact directly with the internal structure of the neutron star.
+This difference leads to solitons modulating the GW frequency, while dilatons would alter the pulse time of arrival in pulsar timing array experiments.
+In summary, solitons could significantly impact the GW signals arising from supermassive black hole binaries, white-dwarf binaries, and asymmetrically rotating neutron stars.
+These signals, which are observed using pulsar timing arrays, space-based interferometers, and ground-based interferometers, respectively, can be modulated by the presence of soliton dark matter.
+For supermassive BH binaries, solitons can induce dynamical friction on the orbit, leading to detectable changes in the nanohertz frequency range.
+White-dwarf binaries, could show low-frequency modulations in their GW emission, potentially revealing the imprint of soliton DM on the millihertz carrier signals.
+Similarly, asymmetrically rotating neutron stars, whose CWs are monitored by ground-based detectors, could exhibit phase shifts by passing through soliton DM.
+These potential signatures across multiple observational platforms make soliton DM a promising candidate for future studies in GW astronomy, and provide interesting probes of the DM landscape.
 VIII GW probes of WIMP dark matter
+While much of this review has focused on probing ultralight dark matter with GW interferometers, with the notable exception of macroscopic DM discussed in Section IV, it is important to also highlight the potential for using GWs to probe more conventional forms of dark matter, such as WIMPs.
+Although WIMPs are typically considered on different scales and their interactions are more strongly coupled than those of ultralight dark matter, they can still leave distinctive imprints on GW signals.
+In this section, we outline two possible indirect ways to search for WIMP dark matter using GWs.
+First, in Section VIII.1, we explore the possibility that WIMPs could induce the collapse of stellar-mass objects, both compact and not, into BHs leading to unique signatures in the GW emission from these newly formed BHs.
+Second, in Section VIII.2, we discuss how the non-detection of CWs from the Galactic Center could help constrain the millisecond pulsar hypothesis for the observed GeV excess, thereby reinforcing the alternative scenario in which annihilating WIMP DM is the underlying cause of this astrophysical signal.
+By examining these two avenues, we can see how GW observations can provide crucial insights into the nature of WIMP DM and contribute to the broader effort to identify and characterize a diverse set of DM candidates.
 VIII.1 Transmuted black holes in binaries
@@
 (2021), eventually becoming gravitationally bound to them over time.
-After repeated interactions between the WIMPs and ordinary nucleons, the WIMPs become thermalized, forming a compact core with a higher density than that of the celestial object.
-If the density of the particles is sufficient, gravitational collapse (in the case of bosons) or Chandrasekhar collapse (in the case of fermions) of the WIMP core will occur, thus resulting in a newborn BH.
-After formation, the celestial object could accrete surrounding DM on a given timescale, and if enough matter becomes bound to the object quickly enough compared to the Hawking evaporation timescale, the entire object could collapse into what is known as a transmuted black hole (TBH) Dasgupta et al.
+After WIMPs repeatedly interact with ordinary nucleons, they become thermalized, forming a compact core with a higher density than that of the celestial object.
+If the density of the particles becomes sufficient inside the celestial object, gravitational collapse (in the case of bosons) or Chandrasekhar collapse (in the case of fermions) of the WIMP core will occur, thus resulting in a newborn BH that is extremely tiny with respect to the celestial object.
+After formation, the tiny BH could accrete surrounding DM, and if enough DM becomes bound to the object quickly enough compared to the Hawking evaporation timescale, the entire object be “swallowed” by the growing BH and become “transmuted” to what is known as a transmuted black hole (TBH) Dasgupta et al.
 (2021); Ray (2023); Bhattacharya et al.
 (2023, 2024); Bhattacharya (2024).
-If the sum of the time to accrete new particles and the time to collapse is smaller than the age of the universe, TBHs should form.
+Quantitatively, if the sum of the time to accrete new particles and the time for the WIMP core to become gravitationally unstable is smaller than the age of the universe, TBHs should form.
 TBHs forming from inspiraling neutron stars were first considered in Dasgupta et al.
 (2021).
-Since DM would induce neutron stars to transmute into BHs, from the observational perspective, there is no difference between electromagnetically silent neutron stars and low-mass BHs.
+Since DM would induce neutron stars to transmute into BHs, from the observational perspective, there is no clear difference between electromagnetically silent neutron stars and low-mass BHs.
 Therefore, the non-detection of low-mass binary black holes by LIGO, Virgo and KAGRA can be used to constrain weak interaction cross-sections, since the progenitor neutron stars provide an extremely strong gravitational field to accrete and retain DM particles Bhattacharya et al.
 (2023); Bhattacharya (2024).
-In LABEL:fig:wimp-cbc-cons, we show the excluded bosonic DM mass/cross-section parameter space based on null search results from the LIGO, Virgo and KAGRA collaborations for sub-solar mass compact objects Abbott et al.
+In Fig.
+24, we show the excluded bosonic DM mass/cross-section parameter space based on null search results from the LIGO, Virgo and KAGRA collaborations for sub-solar mass compact objects Abbott et al.
 (2023).
-Additionally, a projection for this constraint is shown for 50x more exposure time, which may be possible for the future GW network [^6] https://observing.docs.ligo.org/plan/.While analyses of LIGO, Virgo and KAGRA data cannot yet constrain new parameter space for this model of DM, future searches should be able to surpass the pulsar constraint, which comes from the fact that the nearby Gyr-old pulsar PSR 0437--4715 has not actually collapsed into a BH in its lifetime.
+Additionally, a projection for this constraint is shown for a 50x increase in the probable spacetime volume, which is reasonable for the future GW network using ET or Cosmic Explorer.
+While analyses of LIGO, Virgo and KAGRA data cannot yet constrain new parameter space for this model of DM, future searches should be able to surpass the pulsar constraint, which comes from the fact that the nearby Gyr-old pulsar PSR 0437-4715 has not actually collapsed into a BH in its lifetime McDermott et al.
+(2012); Garani et al.
+(2019).
 In contrast, stars, being less compact than neutron stars, could sustain DM cores with stronger interaction cross-sections, because of the weaker gravitational field around the star Bhattacharya et al.
 (2024).
-In LABEL:fig:lisa-wimp-cons, we show the potential excluded parameter space for null detection of TBHs arising from slowly inspiraling non-compact binaries that could be visible in space-based GW detectors.
+In Fig.
+24, we show the potential excluded parameter space for null detection of TBHs arising from slowly inspiraling non-compact binaries that could be visible in space-based GW detectors.
 These binaries would be inspiraling for durations of order of the age of the universe, and thus exhibit CW signals in the millihertz regime.
 The results are parameterized in terms of $\alpha$, which denotes the fraction of stellar-mass binaries that form close enough to exist within the millihertz band.
-Essentially, $\alpha$ parametrizes our ignorance about how closely the population of sun-like stars form to each other.
-Comparing LABEL:fig:wimp-cbc-cons and LABEL:fig:lisa-wimp-cons, we note the vastly different interaction cross-section ranges from these two complementary GW probes of WIMP DM.
+Essentially, $\alpha$ parametrizes our ignorance about how closely the population of sun-like stars form close to each other.
+Comparing Fig.
+24 and Fig.
+24, we note the vastly different interaction cross-section ranges from these two complementary GW probes of WIMP DM.
 For both low-mass BHs and sun-like stars, fermionic DM has also been constrained in similar ways as bosonic DM has, see Bhattacharya et al.
 (2023, 2024) for more details.
-In contrast to TBHs, we note that if DM collects inside neutron stars, a different signature would be present.
-Depending on what equation-of-state assumed for baryonic or fermionic DM, neutron stars could actually be DM admixed neutron stars, which would have a particular equation-of-state that differs from the (currently unknown) canonical one.
-Evidence for such an effect could be found in changes to the tidal deformability parameter that could be measured by GW detectors in the future Collier et al.
-(2022).
-VIII.2 GWs to probe the GeV excess
+[…]
+VIII.2 GW probes of the galactic-center GeV excess
 Annihilating WIMP DM and a collection of unresolved millisecond pulsars are two leading explanations for the unexpected excess of GeV gamma rays observed by the Fermi-LAT experiment coming from the Galactic Center Goodenough and Hooper (2009); Hooper and Goodenough (2011).
@@
 GWs can be used to weigh in on this debate.
-If the millisecond pulsars are deformed, they will emit CWs as they rotate, and this radiation could be detectable by the LIGO, Virgo and KAGRA if the degree of deformation, called the ellipticity (defined in terms of the principle moments of inertia as $\epsilon=\frac{I_{\rm xx}-I_{\rm yy}}{I_{\rm zz}}$), is large enough Haskell et al.
+If the millisecond pulsars are deformed, they will emit CWs as they rotate.
+The amplitude of these CWs is determined by both the GW frequency and the degree of deformation, called the ellipticity, which is defined as $\epsilon\equiv\frac{I_{\rm xx}-I_{\rm yy}}{I_{\rm zz}}$ Haskell et al.
+(2015), where $I_{\rm xx},I_{\rm yy}$ and $I_{\rm zz}$ are the principle moment of inertias along each axis.
+One approach to constraining the millisecond pulsar hypothesis is to consider the bulk emission of a thousands or more sources of CWs that are superimposed, creating a stochastic GW background.
+Ref.
+Calore et al.
+(2019) modelled this stochastic GW background and projected constraints on the average ellipticity of this population as a function of the total number of millisecond pulsars in the Galactic Center.
+The authors showed that $\mathcal{O}(10^{4})$ millisecond pulsars with ellipticities of $\sim 10^{-7}$ could be detectable by current detectors and be consistent with the diffuse electromagnetic radiation coming from the Galactic Center, as shown in Fig.
+25.
+However, they did not go so far as to link GW observations with Fermi-LAT ones.
+Ref.
+Miller and Zhao (2023) took a different approach by analyzing the continuous gravitational wave (CGW) signals from individual neutron stars (NSs), specifically focusing on millisecond pulsars (MSPs).
+They explored two main scenarios for the deformation of these pulsars:
+  1. 1.
+    Magnetic Field-Induced Deformations: In the first scenario, they assumed that the deformations of the millisecond pulsars are sustained by the internal magnetic fields of the neutron stars, as proposed in previous work Lander et al.
+(2012); Mastrano and Melatos (2012); Lander (2014).
+  2. 2.
+    Model-Agnostic Deformation: In the second, more general scenario, they did not assume a specific cause for the deformation.
+Instead, they modeled the ellipticity as 1% of the total electromagnetic radiation emitted by the star, making it independent of the underlying physical mechanism.
+Using these two models, Ref.
+Miller and Zhao (2023) then assumed that the ellipticity and frequency distributions of unresolved population of millisecond pulsars can be calcualted with knonw pulsars’ parameters Manchester et al.
 (2005).
-First, a stochastic GW background of GWs from millisecond pulsars in the Galactic Center was considered, and projected constraints on the average ellipticity of this population were obtained as a function of the total number of millisecond pulsars in the Galactic Center Calore et al.
-(2019).
-The authors showed that $\mathcal{O}(10^{4})$ millisecond pulsars with ellipticities of $\sim 10^{-7}$ could be detectable by current detectors and be consistent with the diffuse electromagnetic radiation coming from the Galactic Center, as shown in LABEL:fig:Ntoteps.
-However, they did not go so far as to link GW observations with Fermi-LAT ones.
-On the other hand, Ref.
-Miller and Zhao (2023) considered different models for how the deformations on the millisecond pulsars were sustained, either through the internal magnetic field Lander et al.
-(2012); Mastrano and Melatos (2012); Lander (2014) or simply as a fraction of the total electromagnetic radiation from the star.
-It is then assumed that the unresolved population of millisecond pulsars follows the same spin and ellipticity distribution of the known pulsars Manchester et al.
-(2005), and, from that, the probability of observing a CW in the search be computed, given upper limits on ellipticities from searches for isolated neutron stars or those in binary systems Abbott et al.
-(2022b); Miller and Zhao (2023).
-Once this quantity is computed, we can determine the number of millisecond pulsars in the Galactic Center by fitting parameters of luminosity functions to match the observed Galactic Center GeV excess Hooper and Linden (2016); Ploeg et al.
+By integrating over these distributions with upper limits on ellipticities from an O3 LIGO search for isolated neutron stars Abbott et al.
+(2022b), they calculated the probability of detecting a CW from this population for both models of ellipticity noted above
+In parallel, the number of millisecond pulsars in the Galactic Center can be inferred by modeling the electromagnetic emission associated with the GeV excess.
+This is done by fitting luminosity functions, which describe the expected number of millisecond pulsars needed to explain the GeV excess as a function of various model parameters Hooper and Linden (2016); Ploeg et al.
 (2020); Dinsmore and Slatyer (2022).
-By multiplying the predicted number of millisecond pulsars with the probability of detecting them, we can rule out luminosity function parameters that would have led to at least one millisecond pulsar being detected via GWs.
-LABEL:fig:gev-cons shows the parameter space of a simplistic luminosity function, and the constraints that can be obtained through null results from CW searches, assuming particular ellipticity and frequency distributions for the millisecond pulsars.
+By combining the predicted number of millisecond pulsars, derived from the luminosity functions, with the probability of detecting them via CWs, the authors were able to rule out certain luminosity function parameters.
+Specifically, they excluded models in which the number of millisecond pulsars would have been large enough to produce at least one detectable CW, thereby exploring the possible origins of the GeV excess.
+Fig.
+25 shows the parameter space of a luminosity function that models the observed luminosity of the Galactic Center GeV excess as following a log-normal distribution.
+The color show the number of millisecond pulsars predicted by each choice of the two parameters of the luminosity function ($L_{0},\sigma_{L})$.
+Plotted as a light blue region is an exclusion area, which signifies that CW searches would have seen at least one millisecond pulsar, assuming the aforementioned ellipticity and frequency distributions for the millisecond pulsars in the galaxy.
 We can see that GW searches probe a complementary portion of the parameter space to what Fermi-LAT does: if there are more millisecond pulsars that are dimmer electromagnetically, this means more rotational power may go into GWs, and more of them are necessary to explain the GeV excess.
@@
 Likewise, Fermi performs better when less millisecond pulsars are more luminous.
-Most recently, Ref.
-Bartel and Profumo (2024) simulated a population of $\sim 40000$ millisecond pulsars in the Galactic Center that could explain the Galactic Center GeV excess, and found that the GW signal could lie between [200,1400] Hz, and would have strain amplitudes of $\mathcal{O}(10^{-30}-10^{-28})$ based on the population parameters chosen (e.g.
-moment of inertia around $\sim 10^{38}$ $\cdot$ m2, ellipticities mostly smaller than $10^{-9}$, and Boxy vs.
-spherical spatial distributions of the millisecond pulsars).
-It is argued that such signals are too weak to be detected by current and future GW observatories; however, this conclusion depends heavily on the assumed ellipticity distribution and the number of millisecond pulsars chosen to be in the population.
-If millisecond pulsars appear in larger numbers than assumed here, they would each be dimmer, since the luminosity of the GeV excess is fixed.
-In fact, the constraints from Miller and Zhao (2023) are strongest for larger populations than assumed in Bartel and Profumo (2024).
-Additionally, the ellipticity distribution employed in Bartel and Profumo (2024) only reaches a maximum of $\sim 2\times 10^{-9}$, which is consistent with the minimum ellipticity of neutron stars as predicted by Woan et al.
-(2018), but may be too conservative.
-Upper limits on ellipticities at the Galactic Center from CW searches have not yet constrained ellipticities to be less than $\sim 10^{-6}$ Abbott et al.
-(2022e, b); Steltner et al.
-(2023), which is consistent with the maximum ellipticity employed in Miller and Zhao (2023) up to $10^{-6}$.
-If GWs can rule out more and more of the luminosity function parameter space to explain the GeV excess, then this will give significant weight to the annihilating WIMP DM hypothesis.
-IX GW probes of DM spikes
-IX.1 Background
-It was first recognized in Gondolo and Silk (1999) that BHs in the DM halo accreting DM adiabatically (slowly) could allow a high-density region of DM to form, called a “spike”, which could enhance the DM annihilation rate inside the spike.
-Subsequent studies showed that if such spikes formed around supermassive BHs at the center of galaxies, they would would likely have been disrupted during host galaxy mergers Merritt et al.
+[…]
+[…]
+In summary, this section highlights the potential of using GWs to probe WIMP DM, focusing on two distinct approaches.
+First, the formation of TBHs from WIMP interactions with celestial objects can lead to unique GW signatures, such as those from inspiraling neutron stars and BHs.
+These signals, especially in the millihertz band, offer constraints on the DM interaction cross-section and can be used to probe the properties of both bosonic and fermionic WIMPs.
+Second, GW non-observations of CWs from deformed millisecond pulsars in the Galactic Center provide a complementary method to weaken the millisecond pulsar hypothesis for the GeV excess observed by the Fermi-LAT experiment, which lends strength to the annihilating WIMP hypothesis for this excess.
+By analyzing the number and ellipticity of unresolved millisecond pulsars, GW searches can rule out certain luminosity function parameters, helping to distinguish between WIMP and millisecond pulsars explanations for the observed excess.
+These two complementary methods underscore the crucial role GWs can play in advancing our understanding of WIMP DM.
+IX GW probes of DM spikes around inspiraling black holes
+It was first recognized in Gondolo and Silk (1999) that BHs embedded in the DM halo of a galaxy could accrete DM adiabatically, potentially forming a high-density region around the BH known as a “spike.” Subsequent studies demonstrated that such DM spikes around supermassive BHs at centers of galaxies would have been disrupted by galactic mergers Merritt et al.
 (2002), if the BH formed off-center Ullio et al.
-(2001), or if DM particles scattered off of nearby stars Merritt (2004); Bertone and Merritt (2005).
-Though, recently, upper bounds on the DM densities for supermassive BHs have been forecast with LISA Daniel et al.
+(2001), or through interactions with stars in the surrounding stellar population Merritt (2004); Bertone and Merritt (2005).
+However, realistic profiles incorporating relativistic corrections have shown that these spikes can cause more DM to accumulate closer to the BHs in both Schwarzchild and Kerr metrics Sadeghian et al.
+(2013); Ferrer et al.
+(2017); Bertone (2024), which begs the question of whether these objects could survive galactic dynamics.
+Though it has been shown that heavy, particle-like DM would be dispersed in equal-mass merger events through N-body simulations Bertone et al.
+(2005), it is much less clear-cut for ultralight dark matter: multiple works Boudon et al.
+(2023); Zhong et al.
+(2023); Cardoso et al.
+(2022a, b); Vicente and Cardoso (2022); Annulli et al.
+(2020a, b); Brax et al.
+(2020); Ficarra (2021); Zhang et al.
+(2023b); Choudhary et al.
+(2021); Yang et al.
+(2018) suggest that such DM spikes could survive, even up to ten orbits before merger Aurrekoetxea et al.
 (2024).
-However, intermediate-mass BHs (IMBHs) could sustain such spikes Zhao and Silk (2005); Kim et al.
-(2023), since their spikes may not have been torn apart through mergers over cosmic time.
-Furthermore, recently it has been shown that astrophysical BHs and PBHs could also host DM spikes Kohri et al.
+The extent to which DM spikes survive directly affects the amount of dynamical friction the binary will experience as it inspirals, thus altering the GWs from the inspiral.
+In light of these recent developments regarding the plausibility of forming DM spikes around supermassive BHs, Ref.
+Daniel et al.
+(2025) has forecast upper bounds on the DM densities around supermassive BHs using future LISA observations of these systems.
+These bounds provide new constraints on the DM spike formation and the nature of DM in the central regions of galaxies, offering a unique opportunity to test the high-density DM environment near supermassive BH binaries.
+Furthermore, Ref.
+Chan and Lee (2024) determines that the observed decay rate of the orbit can be completely explained by the existence of a DM spike around the supermassive BH in OJ 287, providing plausible evidence of the existence of DM spikes.
+Additionally, IMBHs can sustain DM spikes because their smaller masses and relatively longer timescales for mergers (with respect to those of supermassive BHs) prevent the spikes from being disrupted Zhao and Silk (2005); Kim et al.
+(2023).
+Similarly, recent studies suggest that both astrophysical BHs and PBHs could also host DM spikes if there is an absence of significant stellar interactions that could scatter DM particles, and if the gravitational influence of the IMBH persists over time, especially in environments where galactic mergers or other disruptive events have not destroyed the spike Kohri et al.
 (2014); Eroshenko (2016); Boucenna et al.
@@
 (2025); Bertone (2024); Dosopoulou and Silk (2025).
-The density of the DM spike is characterized by the following equation Eda et al. (2013):
-$\rho(r)$ $=\rho_{\rm sp}\left(\frac{r_{\rm sp}}{r}\right)^{\alpha}\,\,\,(r_{\rm min}\leq r\leq r_{\rm sp}),$ (108)
+These findings expand the range of environments where DM spikes might form, suggesting that IMBHs could significantly influence DM distributions.
+For simplicity, despite the more realistic DM spike profiles recently discussed, the density of the DM spike is often approximated by the following equation Eda et al. (2013); Cole et al. (2023a, b):
+$\rho(r)$ $=\rho_{\rm sp}\left(\frac{r_{\rm sp}}{r}\right)^{\gamma_{s}}\,\,\,(r_{\rm min}\leq r\leq r_{\rm sp}),$ (111)
 where $\rho_{\rm sp}$ is a normalization constant for the DM density of the spike, $r_{\rm min}$ is the minimum separation between the two objects, often taken to the radius at the innermost stable circular orbit, and $r_{\rm sp}$ is the spike radius.
 Here, $\rho_{\rm sp}=379\ M_{\odot}/\text{pc}^{3}$ and $r_{\rm sp}=0.33\text{pc}$ if the mass of the IMBH is $M_{\rm BH}=10^{3}M_{\odot}$ and the DM spike has a mass of $M_{\rm halo}=10^{6}M_{\odot}$.
-The power-law index $\alpha$ denotes how the concentration of DM particles falls off as one moves away from the host BH.
-Astrophysically speaking, measurements of $\alpha$ would allow us to glean important information regarding whether the host BH evolved from the DM halo and/or experienced mergers in the past Eda et al.
+The power-law index $\gamma_{s}$ denotes how the concentration of DM particles falls off as one moves away from the host BH.
+Astrophysically speaking, measurements of $\gamma_{s}$ would allow us to glean important information regarding whether the host BH evolved from the DM halo and/or experienced mergers in the past Eda et al.
 (2013).
-IX.2 Dephasing due to dynamical friction of IMRI systems
+We consider two scenarios where dynamical friction can influence the evolution of binary black holes: in Section IX.1, we examine how dynamical friction leads to gravitational-wave dephasing in IMRI systems; in Section IX.2, we explore how DM spikes can form around PBHs and cause dephasing in binaries with primary masses of $\mathcal{O}(1\text{--}100)M_{\odot}$ and mass ratios $\leq 10^{-3}$.
+Lastly, in Section IX.3, we describe how observations of binary black holes can be used to place limits on DM spikes, under the assumption that their orbits remain largely unaffected by dynamical friction.
+IX.1 Dephasing due to dynamical friction of intermediate mass ratio inspirals
 Ref.
@@
 This occurs because the less massive object sweeps through the DM spike as it orbits around the more massive one, losing energy via dynamical friction (in addition to via GWs) and speeding up the orbit Speeney et al.
+(2022); Mitra et al.
 (2025).
@@
 (2024); Ghodla (2025), since IMBHs inspiral and merge at much lower frequencies than stellar-mass BHs Miller (2009).
-In the future, if matched filtering analyses are performed that do not account for the DM spike, dephasing with respect to the vacuum templates may occur.
-In Fig.
-26, we show how the signal-to-noise ratio degrades as a function of the observation time, the initial frequency $f_{\rm ini}$, and the power-law index $\alpha$ of the DM spike).
-As described in Eda et al.
-(2013), dynamical friction affects the early stages of the inspiral, and as the system progresses toward, the DM spike becomes disrupted, reducing its impact on the GW signal.
-That is why in Fig.
-26, lower initial frequencies start to lose signal-to-noise ratio at smaller values of $\alpha$.
-Furthermore, it is clear that above $\alpha\simeq 2$, templates that include dynamical friction will be required to successfully recover IMRI signals in space-based GW interferometers.
 This conclusion was significantly expanded in Eda et al.
-(2015), in which the authors demonstrated how accurately the power-law index $\alpha$ could be measured from a detection of an IMRI signal in LISA.
-This accuracy is quantified in terms of the power-law index itself and the masses of the two objects that are inspiraling.
+(2015), in which the authors demonstrated how accurately the power-law index can be measured from a detection of an IMRI signal in LISA.
+This accuracy is quantified in terms of the power-law index and the masses of the two objects that are inspiraling.
 If the IMRI system consists of smaller primary and secondary masses, and a larger power-law index, the power-law index can be measured more accurately.
-This is because higher values of $\alpha$ induce a larger phase difference, and smaller masses in the IMBH system ensures that it spends more cycles in the LISA band before merging.
-Figure 26: Taken from Eda et al.
+This is because higher values of induce a larger phase difference, and smaller masses in the IMBH system ensures that it spends more cycles in the LISA band before merging.
+Recently, it has also been shown that these dephasing effects are unique enough to allow distinguishing between three BH environments: accretion disks, DM spikes, and ultralight boson clouds Cole et al.
+(2023b), as well as distinguishing between environmental effects and modified theories of gravity Yuan et al.
 (2025).
-The number of cycles $N_{\text{cycle}}$ spent in the GW band $f\in[f_{\text{min}},f_{\text{max}}]$ for a primary BH with mass $M_{\rm BH}=1000M_{\odot}$ and a companion of $1M_{\odot}$.
-At large $\alpha,$ the DM solition causes large dephasing between the injected waveform and the one that would be used to recover the signal.
-Thus, the number of cycles that are captured in-band drops significantly.
-PBHs could also host DM spikes, and must do so, since, at most masses, the fraction of DM that they could compose is less than one Green and Kavanagh (2021); Carr et al.
+As an illustration, we show in Fig.
+26 the impact of an environmental effect, the DM spike, on the SNR obtained when match-filtering with a vacuum template to analyze data in which a binary black hole system has a DM spike around it.
+Here, $m_{1}=10^{5}M_{\odot}$, which was chosen to ensure that a plausible formation mechanism exists to form this kind of BH Gondolo and Silk (1999).
+Ref.
+Cole et al.
+(2023b) argues that a percentage SNR loss of more than 30% would compromise the prospects of detecting a (weak) signal, and shows that even smaller SNR losses would bias parameter estimation.
+However, it is important to note that in LISA, we will be in a regime in which signals of this kind have enormous SNR, and that parameter estimation and distinguishing among different GW signals will be a key challenge.
+Thus, it is more likely that parameter estimates will be biased rather than signals being completely missed.
+Moreover, the dephasing caused by DM spikes underscores the need to adapt CW methods, as discussed in Section III.1 and Section VI.2, to effectively search for these systems Sethi et al.
+(2025).
+In principle, CW methods could provide quick point estimates for key binary black hole parameters, which would enable the development of a reduced set of templates that incorporate DM spikes.
+These templates could then be used for follow-up matched filtering searches.
+Figure 26: Taken from Cole et al.
+(2023b).
+Percentage SNR loss from matched filtering with a vacuum template when the injected GW signal originates from a binary black hole system surrounded by DM spikes (the injected waveform includes the effect of dynamical friction).
+The $x$ and $y$ axes correspond to the parameters of the DM spike profile defined in Eq.
+111.
+The assumed observation time is one year, with component masses $m_{1}=10^{5}M_{\odot}$ and $m_{2}=10M_{\odot}$.
+While Ref.
+Cole et al.
+(2023b) demonstrated that distinguishing between different BH environments, including the presence of DM spikes, is feasible, Ref.
+Zwick et al.
+(2024) presents a different perspective, arguing that the dephasing effects caused by these environmental factors could result from a variety of unmodeled sources.
+Moreover, the authors argue about the practicality of measuring such dephasing effects, noting that the early inspiral is where the effects would be most pronounced, but this is also where the GW emission is weakest; that a constant phase offset could mimic dephasing effects ; and that matched filtering over such long timescales would have practical limitations.
+For thee reasons.
+However, Ref.
+Wilcox et al.
+(2024) provides a counterpoint, suggesting that such dephasing could still be tracked within a parameterized post-Einsteinian framework, offering a potential solution to the challenges raised in Zwick et al.
+(2024).
+Additionally, they propose more model-agnostic approaches, such as periodic forces, that could leave characteristic signatures on the GW signal and be applied to study the evolution of the binary system.
+These differing viewpoints are crucial when considering the overall approach to detecting and interpreting GWs from systems with DM spikes.
+While Zwick et al.
+(2024) concerns highlight the potential difficulties in applying matched filtering to these cases, the model-agnostic approach in Wilcox et al.
+(2024) offer a promising alternative for tackling the problem.
+In light of these challenges, it is clear that both traditional and alternative methods, such as the CW approaches discussed in Section III.1 and Section VI.2, should be explored in tandem to fully address the complexities involved in detecting signals from binary black holes influenced by DM spikes Sethi et al.
+(2025).
+As mentioned in Section VII.2, CW methods could provide ways of addressing the three concerns laid out in Ref.
+Zwick et al.
+(2024) by providing point estimates of binary parameters that can be followed up by a sensitive and restricted matched filtering search.
+This approach can also be applied in lieu of traditional matched filtering for PBHs in ET and Cosmic Explorer.
+IX.2 Dephasing due to dynamical friction of inspiraling primordial black holes
+PBHs could also host DM spikes, since at most masses, the fraction of DM that they could compose is less than one Green and Kavanagh (2021); Carr et al.
 (2024).
@@
 Cole et al.
-(2023a) showed that ET and Cosmic Explorer could probe the existence DM spikes around binary PBH systems, where the primary has a mass of $[1,10]M_{\odot}$ and the system has a mass ratio of $10^{-3}$.
-However, just as in astrophysical binary black hole systems, these exotic mergers would be missed if the DM spike is not taken into account in matched-filtering searches in ET and Cosmic Explorer.
-Multi-band observation of inspirals/mergers constraining DM spikes are also possible with future deci-hertz and ground-based GW interferometers, which would enhance the constraints on the power-law index and spike-density factors by orders of magnitude relative to individual constraints from each of these detectors Tahelyani et al.
-(2024); Chen et al.
-(2024b).
-Recently, it has also been shown that these dephasing effects are unique enough to allow distinguishing between three BH environments: accretion disks, DM spikes, and ultralight boson clouds Cole et al.
-(2023b), as well as distinguishing between environmental effects and modified theories of gravity Yuan et al.
-(2024).
-However, Ref.
-Zwick et al.
-(2024) argues that practically speaking, dephasing could result from a variety of (also unmodelled) sources, and would be difficult (computationally and algorithmically speaking) to track with long-duration templates in matched filtering.
-Though, Ref.
-Wilcox et al.
-(2024) has indicated this could be possible within a parameterized post-Einsteinian framework.
-They thus propose more model-agnostic ways (i.e.
-periodic forces) that may influence the evolution of the binary and leave characteristic signatures on the GW signal.
-Such problems with matched filter, could, however, be handled by the use of alternative, non-matched filtering CW methods meant to handle intermediate-duration signals, e.g.
-Isi et al.
-(2019); Sun and Melatos (2019); Miller et al.
-(2018); Oliver et al.
-(2019); Banagiri et al.
-(2019); Miller et al.
-(2019); Guo and Miller (2022); Andrés-Carcasona et al.
-(2023); Alestas et al.
-(2024); Miller et al.
-(2024a); Andrés-Carcasona et al.
-(2025), and those discussed already in this review.
-Perhaps more surprising, it has been proposed that equal-mass systems could sustain DM spikes even in the last ten orbits of binary black hole merger Aurrekoetxea et al.
-(2024b).
-Initially, it had been shown through N-body simulations that heavy, particle-like DM environments would be disrupted well before equal-mass binary black hole systems merge Bertone and Merritt (2005); Kavanagh et al.
-(2018).
-But, for wave-like, ultralight dark matter, though there have been many works that have modeled the motion of binary black holes through DM environments Koo et al.
-(2024); Boudon et al.
-(2023); Zhong et al.
-(2023); Cardoso et al.
-(2022a, b); Vicente and Cardoso (2022); Annulli et al.
-(2020a, b); Brax et al.
-(2020); Bromley et al.
-(2024) or with boson clouds Cao and Tang (2023); Traykova et al.
-(2023); Chia et al.
-(2023); Baumann et al.
-(2019a, 2020); Ikeda et al.
-(2021); Cardoso et al.
-(2020); Hannuksela et al.
-(2019); Baumann et al.
-(2022b); Zhang and Yang (2020); Kumar Poddar et al.
-(2020); Wong (2019, 2020); Kavic et al.
-(2020); East (2018); Siemonsen and East (2020); Baumann et al.
-(2022a); Siemonsen et al.
-(2023b); Tomaselli et al.
-(2023); Poddar et al.
-(2024); Cao et al.
-(2024); Guo et al.
-(2024d), the point during the inspiral at which the environment becomes disrupted is not clear, motivating further studies into this Bamber et al.
-(2023).
-Ref.
-Aurrekoetxea et al.
-(2024b) found that amount of dephasing that exists within the last ten orbits of a binary black hole merger depends strongly the scalar DM mass and the masses of the BHs:
-$\mu\approx 5\times 10^{-17}\,\left(M/10^{6}M_{\odot}\right)^{-1}\,\mathrm{eV},$ (109)
-which gives the optimal DM mass to cause the most dephasing during the last ten orbits of the binary, and determines the DM density as
-$\rho_{0}\approx 10^{20}\,\left(\frac{M}{10^{6}M_{\odot}}\right)^{-2}\,\frac{M_{\odot}}{\mathrm{pc}^{3}}\,.$ (110)
-which gives the DM density necessary to induce dephasing.
-Though the DM density presented here is larger than the average, it results from a measured dephasing over a short period of time.
-If the effect of the DM spike on the binary black hole system turns on at larger separations, as may be expected from previous works Cole et al.
-(2023a, b), then smaller densities could cause significant dephasing over the lifetime of the binary.
-IX.3 Upper limits on BH environments
+(2023a) showed that ET and Cosmic Explorer could probe the existence of DM spikes around binary PBH systems, where the primary has a mass of $1,10]M_{\odot}$ and the system has a mass ratio of $10^{-3}$.
+However, similarly to IMRI systems discussed in the previous subsection, these exotic mergers could be missed if the DM spike is not taken into account in matched-filtering searches in ET and Cosmic Explorer.
+As mentioned in Sections IX.1 and VII.2, one possible approach is to adapt CW methods to search for DM spikes around binary black holes.
+This could provide a quick, though less sensitive, way to obtain point estimates for important binary black hole parameters, such as the chirp mass.
+These estimates could then be used to reduce the number of templates needed to account for DM spikes, enabling a follow-up matched filtering search.
+IX.3 Upper limits on black-hole environments
 Because vacuum phase-evolution waveforms confirmed the detection of many binary black hole mergers by LIGO, Virgo and KAGRA, any deviations due to environmental effects, such as a DM spike, must be small enough to be inconsequential to the detected GW signals.
@@
 Caneva Santoro et al.
-(2024) modeled dynamical friction and other possible environmental effects as additional (negative) PN corrections to the waveform, and placed upper limits on the density of environment around BHs by requiring the induced dephasing to be small with respect to the phase evolution of the binary black hole systems in vacuum.
+(2024) modeled dynamical friction and other possible environmental effects as additional (negative) PN corrections to the waveform, and placed upper limits on the density of environment around BHs by requiring the induced dephasing to be small enough to not alter the phase evolution of the binary black hole systems in vacuum.
 While the results were used to eliminate dynamical fragmentation Loeb (2016) as a possible binary formation channel, they are generic enough to apply to any BH environment.
-Unfortunately, the upper limits on environments affected by dynamical friction, which would be relevant for the DM spike scenario, are, at best, about seven orders of magnitude larger than the maximum density of the spike ($10^{-6}$ g/cm3 Gondolo and Silk (1999); Sadeghian et al.
+Unfortunately, the upper limits on environments affected by dynamical friction, which would be relevant for the DM spike scenario, are, at best, about seven orders of magnitude larger than the maximum density of the spike ($10^{-6}$ g/cm3) Gondolo and Silk (1999); Sadeghian et al.
 (2013).
@@
 X GW probes of atomic DM
-X.1 Background
 It is possible that the dark sector is actually much more complicated than a single new particle with a single coupling to gravity or the standard model.
-In fact, a atomic DM models exist Feng et al.
+In fact, atomic DM models exist Feng et al.
 (2009); Kaplan et al.
-(2010) composed of dark protons and dark electrons that interact via a massless dark photon with a particular coupling.
-With such a model, DM can form bounds states analogous to the Hydrogen atom, and dissipate energy via Bremsstrahlung, recombination, and collisional excitation processes Buckley and DiFranzo (2018).
-If DM collects in small-enough regions of space, it will collapse and form “dark BHs Shandera et al.
+(2010) in which dark protons and dark electrons interact via a massless dark photon [^12] This dark photon is different from both the dark photon that couples to baryons, discussed in Section II.4, and the one that kinetically mixes with the ordinary photon, discussed in Section VI.3.3.
+with a particular coupling.
+In such a model, DM can form bounds states analogous to the Hydrogen atom, and dissipate energy via Bremsstrahlung, recombination, and collisions Buckley and DiFranzo (2018); Rosenberg and Fan (2017).
+If this kind of DM collects in small-enough regions of space, it will collapse and form “dark BHs” Shandera et al.
 (2018).
@@
 (2022f, 2023).
-X.2 Constraints from observations of binary black holes
+Here, the observable GW signature of dark BHs would come from the inspiral, merger and ringdown of binary black hole systems.
+However, this signature is not unique: there does not yet exist a way of distinguishing between astrophysical, dark BHs and PBHs.
+Thus, the constraints that we present in Section X.1 and Section X.2 on BHs with $M_{\rm BH}=[1,100]M_{\odot}$ and $M_{\rm BH}=[0.1,1]M_{\odot}$, respectively, assume that the BHs that are searched for in LIGO, Virgo and KAGRA data are dark BHs.
+X.1 Constraints from observations of binary black holes
 Ref.
@@
 If the BH merger of GW190425 Abbott et al.
-(2020a) is interpreted as the merger of two dark BHs, the Chandrasekhar mass Chandrasekhar (1931) can be constrained to be below $1.4M_{\odot}$ at $>99.9\%$, bounding the lower-most value mass of the dark proton as 0.95 GeV and the spacing of energy levels in a dark atom to be near $10^{-3}$ eV Singh et al.
+(2020a) is interpreted as the merger of two dark BHs, the Chandrasekhar mass Chandrasekhar (1931) can be constrained to be below $1.4M_{\odot}$ at $99.9\%$, bounding the lower-most value mass of the dark proton as 0.95 GeV and the spacing of energy levels in a dark atom to be near $10^{-3}$ eV Singh et al.
 (2023).
-These constraints on the mass of the dark proton and the Chandrasekhar limit of the mass of dark BHs are derived using Chandrasekhar (1931); Shandera et al. (2018); Singh et al. (2021):
+In order to obtain these constraints, we note that LIGO, Virgo and KAGRA observations of binary black holes provide estimates of the chirp mass and the component masses of the system.
+Ref.
+Singh et al.
+(2021, 2023) jointly constrain, in a Bayesian way, the fraction of DM that dark BHs could compose and the minimum mass $M_{\rm min}$ of the distribution of dark BHs assumed, while marginalizing over other hyperparmeters, such as the slope and range of the mass distribution.
+These posteriors are conditioned on the output of a GW search, which is essentially predicted merger rate density that depends both on what binary black holes have been observed and power spectral density of the interferometers.
+Quantitatively, these constraints on the mass of the dark proton and the Chandrasekhar limit of the mass of dark BHs are derived using Chandrasekhar (1931); Shandera et al.
+(2018); Singh et al.
+(2021):
 $M_{\rm DC}\simeq 1.4M_{\odot}\left(\frac{m_{p}}{m_{\chi}}\right)^{2}.$ (112)
+where $m_{p}$ is the mass of the proton.
 The range of the dark proton mass $m_{\chi}$ can be obtained by setting $M_{\rm DC}=M_{\rm min}$ for the values of $M_{\rm min}$ for which the fraction of DM that dark BHs could compose is less than 1:
-$\sqrt{\frac{1.4M_{\odot}}{M_{\rm min}^{\rm max}}}m_{p}<m_{\chi}<\sqrt{\frac{1.4M_{\odot}}{M_{\rm max}^{\rm max}}}m_{p}$ (112)
-$M_{\rm min}$ is the minimum mass in a distribution of dark BHs masses that is assumed, as described below.
-Further constraints on the fraction of DM composed of dark BHs are presented in three cases, assuming that (1) none of the detected BHs by LIGO, Virgo and KAGRA are dark, (2) half are dark, and (3) GW190425 and/or GW190814 Abbott et al.
-(2020b) are dark.
+$\sqrt{\frac{1.4M_{\odot}}{M_{\rm min}^{\rm max}}}m_{p}<m_{\chi}<\sqrt{\frac{1.4M_{\odot}}{M_{\rm min}^{\rm min}}}m_{p}$ (113)
+In Singh et al.
+(2021), $M_{\rm min}=[0.054,1.50]M_{\odot}$, which implies $0.91<m_{\chi}/\,\mathrm{GeV}<4.76$ when considering both GW190425 and GW190814 Abbott et al.
+(2020b) as dark binary black hole mergers.
+Beyond a range of implied $M_{\rm min}$ and $m_{\chi}$, constraints can be placed on the fraction of DM taht PBHs compose as a function of the average mass of the binary black hole system.
+These constraints are shown in Fig.
+27, assuming that (1) none of the detected BHs by LIGO, Virgo and KAGRA are dark, (2) half are dark, and (3) GW190425 and/or GW190814 are dark.
 Here, we limit ourselves to case (1) and case (3).
-These constraints rely on an assumed mass distribution of dark BHs based on Population III star formation works Bromm and Larson (2004), which is chosen to be $P(m)\propto m^{-b}$ along with a mass range of $[M_{\rm min},rM_{\rm min}]$.
-The parameters $b=[-1,2],r=[2,100]$ determine the mass distribution of dark BHs.
-With these assumptions, we show in Fig.
-27 the constraint on the fraction of DM that dark BHs could compose as a function of the average mass of the BHs considered, in relation to other limits, in case (1) and in case (3).
+These constraints, along with those on the dark proton mass, rely on an assumed mass distribution of dark BHs based on Population III star formation works Bromm and Larson (2004), which is chosen to be $P(m)\propto m^{-b}$ along with a mass range of $M_{\rm min},rM_{\rm min}]$.
+The parameters b=[1,2],r=[2,100]b=[-1,2],r=[2,100] determine the mass distribution of dark BHs.
 Assuming all BHs are not dark, the constraint beats existing ones by several orders of magnitude; assuming that GW190425 and GW190814 are dark BHs, a region in this parameter space is carved out of possible average mass/ DM fractions.
@@
 These plots assume an atomic model of DM that forms dark BHs.
+The constraints are obtained in three cases: (1) assuming none of the detected BHs are dark; (2) assuming only GW190425 is a dark BH; and (3) assuming both GW190425 and GW190814 are dark BHs.
+The latter two cases essentially arise from a joint posterior distribution conditioned on the rate densities inferred from the search results, which also constrain the minimum allowed mass in the assumed population distribution (see text for details).
 Even in the absence of any dark BHs, the constraints obtained in Singh et al.
@@
 SB stands for stochastic background and refers to analyses of the first and second observing runs of LIGO, Virgo and KAGRA data for stochastic GW backgrounds from generic systems.
-X.3 Constraints from sub-solar mass binary black hole searches
+X.2 Constraints from sub-solar mass binary black hole searches
 The lack of smoking-gun evidence for dark BHs in current LIGO, Virgo and KAGRA observations motivates the need to consider sub-solar mass BHs, whose origins, if detected, cannot be explained by astrophysical means.
-Searches for GWs from sub-solar mass compact objects in binary systems have grown in popularity in recent years, covering mass ranges of roughly $[10^{-11},1]M_{\odot}$ Nitz and Wang (2021); Phukon et al.
+Searches for GWs from sub-solar mass compact objects in binary systems have grown in popularity in recent years, covering mass ranges of roughly $10^{-11},1]M_{\odot}$ Nitz and Wang (2021); Phukon et al.
 (2021); Abbott et al.
 (2022f); Miller et al.
-(2022b); Nitz and Wang (2022); Abbott et al.
+(2022a); Nitz and Wang (2022); Abbott et al.
 (2023, 2022b); Miller et al.
-(2024b), and, traditionally, have put constraints on the fraction of DM that PBHs could compose.
+(2024), and, traditionally, have put constraints on the fraction of DM that PBHs could compose.
 However, as discussed in Singh et al.
@@
 28, using different matched filtering methods, again assuming a power-law dependence for the mass distribution of dark BHs and as a function of the minimum mass of this distribution.
-Nuisance parameters in the mass distribution are marginalized over.
-These limits do not go below $\sim 10^{-2}M_{\odot}$ because the maximum mass probed is between $[2,1000]M_{\odot}$, and so the distribution tends to disfavor smaller masses.
+These limits do not go below $\sim 10^{-2}M_{\odot}$ because the maximum mass probed is between $2,1000]M_{\odot}$, and so the distribution tends to disfavor smaller masses.
 Moreover, the limits do not go above a solar mass because the search is only for sub-solar mass objects.
@@
 (2021) of [0.91,4.76] GeV since there are no assumed detections of dark BHs.
-Additionally, the energy spacing between dark atomic energy levels can also be constrained, assuming values for the mass of the dark fermions, the dark fine structure constant, and the Jeans mass, and using Eq.
-1 of Singh et al.
-(2021) and Eq.
-4 of Shandera et al.
-(2018).
+Additionally, constraints can be placed on the energy spacing between dark atomic levels.
+In atomic DM models, the collapse of dark gas into compact stars (and hence into dark BHs) requires efficient radiative cooling, which is governed by the energy level spacing
+$\Delta E\sim\tfrac{1}{2}\,\alpha_{\chi}^{2}m_{e_{\chi}},$ (114)
+where $\alpha_{\chi}$ is the dark fine-structure constant and $m_{e_{\chi}}$ the mass of the dark electron.
+By combining the Chandrasekhar mass relation with assumptions about the dark fine-structure constant, the dark fermion masses, and the Jeans mass for collapse, one can use Eq.
+(1) of Singh et al.
+(2021) together with Eq.
+(4) of Shandera et al.
+(2018) to infer the allowed range of $\Delta E$.
+For the GW190425 interpretation, this yields a characteristic value of order $\sim 10^{-3}\,\mathrm{eV}$.
 Figure 28: Taken from Abbott et al.
@@
 (2022f).
+This plot differs from Fig.
+27 primarily in the mass range that is considered in the sub-solar mass search versus the ones in, e.g.
+Abbott et al.
+(2021c).
+At lower masses, the constraints on $f_{\rm DBH}$ weaken with respect to those at higher masses.
+Although sub-solar mass searches for dark BHs have placed promising constraints, they are typically limited to masses above $\sim 10^{-2}M_{\odot}$, largely due to the prohibitive computational cost of conducting matched filtering searches for lighter-mass binaries.
+In contrast, CW methods—which search for time-frequency tracks rather than relying on matched filtering —can access lower-mass regimes Miller et al.
+(2021b); Abbott et al.
+(2022b); Miller (2024b); Alestas et al.
+(2024).
+While originally developed for PBH searches, these techniques are equally applicable to testing the dark BH hypothesis.
+As such, CW approaches offer a valuable path toward exploring a broader parameter space than is feasible with traditional matched filtering methods.
 XI Conclusions
 There is no shortage of observable signatures that DM could leave on GW sources when the particle, mass and couplings of DM the standard model are varied.
-We do not assume to have covered every single DM model in this work, but we hope to have provided an overview for the interested reader that wishes to pursue more details about different ways that GWs can be used to probe the existence of particle DM.
-In this work, we have explored (1) how ultralight dark matter could couple to GW interferometers, (2) how macroscopic DM could transit through the interferometers, (3) how boson cloud systems formed through a superradiance process around rotating BHs could emit GWs by annihilation, (4) how soliton DM could induce changes in GWs arising from binary systems and isolated neutron stars, (5) how WIMP DM could collect around celestial objects and induce them to collapse into BHs detectable via GW emission, (6) how environmental effects of DM spikes around BHs could affect GW signals in future detectors, and (7) how GW observations and non-observations can be used to probe atomic models of DM.
+We do not assume to have covered every single DM model in this work, but we hope to have provided an overview for the interested reader that wishes to pursue more details about different ways that GWs can be used to probe particle DM.
+In this work, we have explored (1) how ultralight dark matter could couple to GW interferometers, (2) how macroscopic DM can transit through the interferometers, (3) how boson cloud systems formed through a superradiance process around rotating BHs can emit GWs by annihilation, (4) how soliton DM can induce changes in GWs arising from binary systems and isolated neutron stars, (5) how WIMP DM can collect around celestial objects and induce them to collapse into BHs detectable via GW emission, (6) how environmental effects of DM spikes around BHs can affect GW signals in future detectors, and (7) how GW observations and non-observations of binary black holes can be used to probe atomic models of DM.
 Some of these types for DM have been extensively constrained through analysis of terrestrial GW detector data; others will require the advent of space-based and next-generation ground-based GW interferometers to obtain insight into these types of DM.
@@
 Moreover, it has been assumed in this work, and almost all those cited, that DM is the way to solve the puzzling cosmological observations on a variety of scales.
-However, modified theories of gravity are real contenders to solve this problem as well, though their community is a lot smaller and not quite as popular as the DM one.
+However, modified theories of gravity are real contenders to solve this problem as well.
 These two communities – DM and modified gravity – act entirely independently, choosing to break, respectively, one of the two assumptions of “classical” physics: (1) the matter in the universe is mostly baryonic and luminous, and (2) general relativity and, in the limit of slow-moving objects, Newtonian theory, are correct Martens and Lehmkuhl (2020).
 It may be that these assumptions need to be broken simultaneously, or that these two distinct fields have more in common than it appears, as argued in Martens and Lehmkuhl (2020).
-We only briefly mentioned PBHs in this review, which have been reviewed extensively elsewhere Green and Kavanagh (2021); Miller (2024).
+We only briefly mentioned PBHs in this review, which have been reviewed extensively elsewhere Green and Kavanagh (2021); Miller (2024a).
 PBHs could also comprise a fraction of or the totality of DM, and recent observations of binary black holes by the LIGO, Virgo and KAGRA with low spins whose merging rates are consistent with PBHs have reignited interested in this hypothesis Bird et al.
@@
 in the atomic DM models, or that there is a whole dark sector.
-Nevertheless, GWs provide a new window to probe the existence of DM in a variety of ways.
-Most of the actual analyses performed on GW data do not assume particular models for DM, and only constrain them later when one interprets null (or actual) results in terms of a DM model.
+Nevertheless, GWs provide a new window to probe the origins of DM in a variety of ways.
+Most of the actual analyses performed on GW data do not assume particular models for DM, and only constrain them later when one interprets results in terms of a DM model.
 GW searches are therefore powerful, relatively model-agnostic probes of DM that permit the detection of a variety of DM signatures.
@@
 We thank William East, Alexandre Gottel, Dana Jones, Jun’ya Kume, Sachiko Kuroyanagi, Ling Sun, and Yue Zhao for reviewing this manuscript.
+We also thank the three anonymous referees that provided valuable feedback on this review article.
 This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation

### v5

@@
 I Introduction
-Our knowledge of physics has been incomplete for decades, if not for over a century Bertone and Hooper (2018).
+Our knowledge of physics has been incomplete for decades, if not for over a century Bertone:2016nfn.
 Dark energy and dark matter (DM) comprise 69% and 26% of the universe, respectively, and yet we have not been able to directly detect either.
-The experimental evidence for the existence of DM is overwhelming: stars orbiting the center of the spiral galaxies would have flown out of their orbits if DM did not hold them in place Freeman (1970); Rogstad and Shostak (1972); Whitehurst and Roberts (1972); Roberts and Rots (1973); anisotropies in the cosmic microwave background can result from DM potential wells in the early universe Peebles (1982); Burles and Tytler (1998); DM can seed the formation of large-scale structures visible today, which would have been impossible with ordinary matter alone Pagels and Primack (1982); Davis et al.
-(1982); Blumenthal et al.
-(1984); Davis et al.
-(1985); Navarro et al.
-(1996); and gravitational lensing of light by galaxies in between us and the light source can be explained by DM inside the galaxy Paczynski (1986); Alcock et al.
-(1993); Aubourg et al.
-(1993); Alcock et al.
-(2000); Lasserre (2000); Tisserand et al.
-(2007).
+The experimental evidence for the existence of DM is overwhelming: stars orbiting the center of the spiral galaxies would have flown out of their orbits if DM did not hold them in place Freeman:1970mx; rogstad1972gross; whitehurst1972high; roberts1973comparison; anisotropies in the cosmic microwave background can result from DM potential wells in the early universe Peebles:1982ff; Burles:1997fa; DM can seed the formation of large-scale structures visible today, which would have been impossible with ordinary matter alone Pagels:1981ke; davis1982survey; Blumenthal:1984bp; Davis:1985rj; Navarro:1995iw; and gravitational lensing of light by galaxies in between us and the light source can be explained by DM inside the galaxy Paczynski:1985jf; SupernovaCosmologyProject:1993faz; Aubourg:1993wb; MACHO:2000qbb; Lasserre:2000xw; EROS-2:2006ryy.
 While the observational evidence shows that DM exists, it does not explain what it is made of.
-Can it be a single new particle, such as the axion Peccei and Quinn (1977a, b); Weinberg (1978) or dark photon Galison and Manohar (1984); Holdom (1986); Pierce et al.
-(2018); Filippi and De Napoli (2020), or macroscopic, in the form of PBHs formed within $\sim 1$ second of the Big Bang Green and Kavanagh (2021)?
-Or, can particle DM exist alongside PBHs, which may comprise a portion or the totality of DM Hawking (1971); Carr et al.
-(2010); Dai et al.
-(2009); Clesse and García-Bellido (2015); Carr et al.
-(2016); Clesse and García-Bellido (2017); Clesse and Garcia-Bellido (2022); Miller (2024a)?
-The numerous hypotheses regarding the origins of DM suggest that the mass of its constituents can vary across hundreds of orders of magnitude Belenchia et al.
-(2022).
-Significant time and resources, however, have been spent on searching a somewhat narrow mass range comprising so-called Weakly Interacting Massive Particles (WIMPs) Jungman et al.
-(1996).
-WIMPs have been probed through collider experiments Goodman et al.
-(2010) and DM/ nucleus interactions by experiments such as CDMS Akerib et al.
-(2003, 2006); Ahmed et al.
-(2009), SuperCDMS Agnese et al.
-(2019), XENON Angle et al.
-(2008, 2011); Aprile et al.
-(2012, 2016, 2024), and LZ (LUX-Zeppelin) Malling et al.
-(2011); Akerib et al.
-(2013, 2016); Mount et al.
-(2017); Akerib et al.
-(2020).
+Can it be a single new particle, such as the axion Peccei:1977hh; Peccei:1977ur; Weinberg:1977ma or dark photon Galison:1983pa; Holdom:1985ag; Pierce:2018xmy; Filippi:2020kii, or macroscopic, in the form of PBHs formed within $\sim 1$ second of the Big Bang Green:2020jor?
+Or, can particle DM exist alongside PBHs, which may comprise a portion or the totality of DM Hawking:1971ei; Carr:2009jm; Dai:2009hx; Clesse:2015wea; Carr:2016drx; Clesse:2016vqa; Clesse:2020ghq; Miller:2024rca?
+The numerous hypotheses regarding the origins of DM suggest that the mass of its constituents can vary across hundreds of orders of magnitude Belenchia:2021rfb.
+Significant time and resources, however, have been spent on searching a somewhat narrow mass range comprising so-called Weakly Interacting Massive Particles (WIMPs) Jungman:1995df.
+WIMPs have been probed through collider experiments Goodman:2010ku and DM/ nucleus interactions by experiments such as CDMS CDMS:2003yjm; CDMS:2005rss; CDMS:2008uih, SuperCDMS SuperCDMS:2018gro, XENON XENON:2007uwm; XENON10:2011prx; XENON100:2011cza; XENON:2016jmt; XENON:2024wpa, and LZ (LUX-Zeppelin) LZ:2011rhn; LUX:2012kmp; LUX:2016sci; Mount:2017qzi; LZ:2019sgr.
 Other experiments have tried to detect annihilating DM indirectly through its gamma-ray emission in the Galactic Center, galaxy clusters, or other areas, e.g.
-MAGIC Aleksic et al.
-(2012); Aleksić et al.
-(2016), VERITAS Holder et al.
-(2009), HESS Aharonian et al.
-(2006) and Fermi Hooper and Goodenough (2011), or through cosmic rays, e.g.
-AMS Aguilar et al.
-(2007, 2013); Bergstrom et al.
-(2013) or Fermi-LAT Abdo et al.
-(2009); Ackermann et al.
-(2012).
-Despite these efforts, however, WIMPS remain undetected [^1] DAMA claims to see a periodic oscillation of a DM signal, though this result has not yet been reproduced Petriello and Zurek (2008), and so it is worthwhile to ask whether other models can better explain DM.
-Thus, the experimental development from the particle physics community has widened to include probing ultralight ($\ll 1$ eV) DM through (1) torsion balance experiments, such as Eöt-Wash Su et al.
-(1994); Schlamminger et al.
-(2008) and the MICROSCOPE satellite Touboul et al.
-(2012); Bergé et al.
-(2018), whose results can be interpreted as constraints on DM, (2) resonant cavity experiments, such as ADMX Du et al.
-(2018), (3) the Event Horizon telescope Doeleman et al.
-(2009); Akiyama et al.
-(2019), which can measure photons lensed by boson clouds and thus affect BH images Chen et al.
-(2023), and which can measure changes in the polarization of the electric field when the light passes through the axion background Chen et al.
-(2022, 2022), and (4) astrometry experiments, such as Gaia Prusti et al.
-(2016) and the Roman Space Telescope Akeson et al.
-(2019), that can measure stochastic metric perturbations sourced by ultralight dark matter Kim (2024).
-In light of the diversity of experiments mentioned, it is also worth asking whether high-precision GW interferometers, such as LIGO, Virgo, KAGRA and GEO600 Affeldt et al.
-(2014); Dooley et al.
-(2016); Aasi et al.
-(2015); Acernese et al.
-(2015); Aso et al.
-(2013), can be used to search for DM, as previously reviewed in Bertone et al.
-(2020); Bertone (2024).
-While it has been demonstrated that constraints on axions can be set using current observations of binary neutron star inspirals (GW170817) Hook and Huang (2018); Croon et al.
-(2018); Zhang et al.
-(2021) and the Hulse-Taylor binary Kumar Poddar et al.
-(2020), in this review, we would like to explore how continuous gravitational-wave (CWs) can be used as a probe of DM.
+MAGIC MAGIC:2011vay; Aleksic:2014poa, VERITAS Holder:2008ux, HESS HESS:2006fka and Fermi Hooper:2010mq, or through cosmic rays, e.g.
+AMS AMS01:2007rrn; AMS:2013fma; Bergstrom:2013jra or Fermi-LAT Fermi-LAT:2009yfs; Fermi-LAT:2011baq.
+Despite these efforts, however, WIMPS remain undetected [^1] DAMA claims to see a periodic oscillation of a DM signal, though this result has not yet been reproduced Petriello:2008jj, and so it is worthwhile to ask whether other models can better explain DM.
+Thus, the experimental development from the particle physics community has widened to include probing ultralight ($\ll 1$ eV) DM through (1) torsion balance experiments, such as Eöt-Wash Su:1994gu; Schlamminger:2007ht and the MICROSCOPE satellite Touboul:2012ui; Berge:2017ovy, whose results can be interpreted as constraints on DM, (2) resonant cavity experiments, such as ADMX Du:2018uak, (3) the Event Horizon telescope Doeleman:2009te; EventHorizonTelescope:2019uob, which can measure photons lensed by boson clouds and thus affect BH images Chen:2022kzv, and which can measure changes in the polarization of the electric field when the light passes through the axion background Chen:2021lvo; Chen:2021lvo, and (4) astrometry experiments, such as Gaia Gaia:2016zol and the Roman Space Telescope Akeson:2019biv, that can measure stochastic metric perturbations sourced by ultralight dark matter Kim:2024xcr.
+In light of the diversity of experiments mentioned, it is also worth asking whether high-precision GW interferometers, such as LIGO, Virgo, KAGRA and GEO600 Affeldt:2014rza; Dooley:2015fpa; 2015CQGra..32g4001L; 2015CQGra..32b4001A; Aso:2013eba, can be used to search for DM, as previously reviewed in Bertone:2019irm; Bertone:2024rxe.
+While it has been demonstrated that constraints on axions can be set using current observations of binary neutron star inspirals (GW170817) Hook:2017psm; Croon:2017zcu; Zhang:2021mks and the Hulse-Taylor binary KumarPoddar:2019jxe, in this review, we would like to explore how continuous gravitational-wave (CWs) can be used as a probe of DM.
 CWs are quasi-monochromatic, quasi-infinite-duration signals that canonically arise from asymmetrically rotating, lumpy neutron stars.
-Over the last few decades, much research has focused on developing ways to probe such neutron stars, both those that are known electromagnetically and those that may only be emitting GWs (“gravitars”) Sieniawska and Bejger (2019); Tenorio et al.
-(2021); Riles (2023); Piccinni (2022); Miller (2023); Wette (2023).
+Over the last few decades, much research has focused on developing ways to probe such neutron stars, both those that are known electromagnetically and those that may only be emitting GWs (“gravitars”) Sieniawska:2019hmd; Tenorio:2021wmz; Riles:2022wwz; Piccinni:2022vsd; Miller:2023qyw; Wette:2023dom.
 Though the signal model is simple, the unknown sky positions and the uncertain neutron-star physics (e.g.
-the unknown equation of state, “spin wandering” Akmal et al.
-(1998); Cutler (2002), changes in the magnetosphere Lyne et al.
-(2010), etc.) complicate these searches, and thus methods that are robust not only to non-Gaussian noise disturbances but also to theoretical uncertainties had to be developed to look for such systems Krishnan et al.
-(2004); Dhurandhar et al.
-(2008); Astone et al.
-(2014); Suvorova et al.
-(2016); Bayley et al.
-(2019).
-Until recently, such searches have only yielded constraints on the maximum possible deformation of the size of known Abbott et al.
-(2021a, 2022a); Ashok et al.
-(2021, 2024); Abac et al.
-(2025a) and unknown neutron stars Abbott et al.
-(2022b); Steltner et al.
-(2023).
+the unknown equation of state, “spin wandering” Akmal:1998cf; Cutler:2002, changes in the magnetosphere Lyne:2010ad, etc.) complicate these searches, and thus methods that are robust not only to non-Gaussian noise disturbances but also to theoretical uncertainties had to be developed to look for such systems Krishnan:2004sv; Dhurandhar:2007vb; Astone:2014esa; Suvorova:2016rdc; Bayley:2019bcb.
+Until recently, such searches have only yielded constraints on the maximum possible deformation of the size of known LIGOScientific:2020lkw; LIGOScientific:2021hvc; Ashok:2021fnj; Ashok:2024fts; LIGOScientific:2025kei and unknown neutron stars KAGRA:2022dwb; Steltner:2023cfk.
 In this review, we show how CW methods have been generalized to search for the following diverse set of DM models:
   1. 1.
-    Ultralight bosons (scalar, vector, or tensor): Oscillating fields can couple to standard-model particles, leading to periodic modulations in the interferometer arm lengths, refractive indices, neutron-star moment of inertias, or timings of atomic clocks Khmelnitsky and Rubakov (2014); Pierce et al.
-(2018); Guo et al.
-(2019); Grote and Stadnik (2019); Ismail et al.
-(2023).
+    Ultralight bosons (scalar, vector, or tensor): Oscillating fields can couple to standard-model particles, leading to periodic modulations in the interferometer arm lengths, refractive indices, neutron-star moment of inertias, or timings of atomic clocks Khmelnitsky:2013lxt; Pierce:2018xmy; Guo:2019ker; grote2019novel; Ismail:2022ukp.
 CW methods are sensitive to these narrowband, quasi-monochromatic signals and can be applied to searches for this kind of DM in ground-based interferometers, pulsar timing arrays, or future space-based antennas.
@@
   2. 2.
-    Macroscopic dark matter objects: Compact dark objects of masses of $\mathcal{O}(1-10^{9})$ kg passing through or near GW interferometers could induce transient but quasi-periodic disturbances in the test masses Hall et al.
-(2018); Du et al.
-(2023).
+    Macroscopic dark matter objects: Compact dark objects of masses of $\mathcal{O}(1-10^{9})$ kg passing through or near GW interferometers could induce transient but quasi-periodic disturbances in the test masses Hall:2016usm; Du:2023dhk.
 Searches using matched filtering can place constraints on Yukawa-type interactions or on the rate of such transient events.
@@
   3. 3.
-    Boson clouds around rotating black holes: Ultralight bosons can grow around black holes via superradiance, accelerated when the Compton wavelength of the boson matches the size of the BH, forming clouds that extract energy and momentum from the BH and emit nearly monochromatic GWs through annihilation or level transitions Arvanitaki and Dubovsky (2011); Arvanitaki et al.
-(2015a); Brito et al.
-(2015, 2017).
+    Boson clouds around rotating black holes: Ultralight bosons can grow around black holes via superradiance, accelerated when the Compton wavelength of the boson matches the size of the BH, forming clouds that extract energy and momentum from the BH and emit nearly monochromatic GWs through annihilation or level transitions Arvanitaki:2010sy; Arvanitaki:2014wva; Brito:2015oca; Brito:2017zvb.
 CW searches can target these long-lived signals and test the existence of bosons in the $10^{-13},10^{-11}]$ eV mass range in ground-based GW interferometers, and between $10^{-17},10^{-15}]$ eV in future space-based GW antennas.
   4. 4.
-    Soliton dark matter: Gravitational interactions can lead to the formation of solitonic cores of ultralight dark matter at the centers of galaxies Schive et al.
-(2014); Hui et al.
-(2017).
+    Soliton dark matter: Gravitational interactions can lead to the formation of solitonic cores of ultralight dark matter at the centers of galaxies Schive:2014dra; Hui:2016ltb.
 Compact binaries or isolated neutron stars moving through solitons experience additional gravitational potentials that can modify the phase evolution of the GWs as binary systems inspiral, and as neutron stars spin down, respectively.
-For binaries, this manifests as a long-lived, cumulative dephasing effect during the inspiral Aghaie et al.
-(2024), while for isolated nonaxisymmetric neutron stars, frequency modulations in the emitted CWs Blas et al.
-(2025) are induced.
+For binaries, this manifests as a long-lived, cumulative dephasing effect during the inspiral Aghaie:2023lan, while for isolated nonaxisymmetric neutron stars, frequency modulations in the emitted CWs Blas:2024duy are induced.
 In both cases, the effect is coherent over long timescales, making CW search methods particularly well suited to test these scenarios.
@@
   5. 5.
-    WIMP dark matter: Even though WIMPs are not inherently wave-like, their capture in stars can induce their collapse into BHs Singh et al.
-(2021); Bhattacharya et al.
-(2024).
+    WIMP dark matter: Even though WIMPs are not inherently wave-like, their capture in stars can induce their collapse into BHs Singh:2020wiq; Bhattacharya:2024pmp.
 CW searches can probe this scenario indirectly, by looking for the long-lived inspirals of BHs space-based GW interferometers.
-Additionally, null results from CW searches from millisecond pulsars in the Galactic Center can indirectly preference WIMP DM as responsible for the gamma-ray GeV excess emanating from the center of the Milky Way Calore et al.
-(2019); Miller and Zhao (2023).
+Additionally, null results from CW searches from millisecond pulsars in the Galactic Center can indirectly preference WIMP DM as responsible for the gamma-ray GeV excess emanating from the center of the Milky Way Calore:2018sbp; Miller:2023qph.
   6. 6.
-    Dark matter spikes around compact objects: Dense DM profiles near BHs inspiraling can generate dynamical friction as the secondary object flows through the DM spike of the first, leading to small but measurable dephasing in these long-lived inspirals Gondolo and Silk (1999); Eda et al.
-(2013, 2015); Kavanagh et al.
-(2020); Coogan et al.
-(2022); Cole et al.
-(2023a, b); Bertone (2024).
+    Dark matter spikes around compact objects: Dense DM profiles near BHs inspiraling can generate dynamical friction as the secondary object flows through the DM spike of the first, leading to small but measurable dephasing in these long-lived inspirals Gondolo:1999ef; Eda:2013gg; Eda:2014kra; Kavanagh:2020cfn; Coogan:2021uqv; Cole:2022ucw; Cole:2022yzw; Bertone:2024rxe.
 While originally such effects were considered around extreme-mass ratio inspirals (EMRIs), intermediate-mass ratio inspirals (IMRIs) and ordinary binaries have been shown to allow the presence of DM spikes, and such dephasing could be measurable in ET, Cosmic Explorer or space-based GW interferometers, when inspiraling binaries will spend hours, days or years in-band.
@@
   7. 7.
-    Atomic dark matter: If DM possesses a composite structure with a dark sector analogue of electromagnetism, it could form bound states or “dark atoms” Feng et al.
-(2009); Kaplan et al.
-(2010); Buckley and DiFranzo (2018); Shandera et al.
-(2018).
+    Atomic dark matter: If DM possesses a composite structure with a dark sector analogue of electromagnetism, it could form bound states or “dark atoms” Feng:2009mn; Kaplan:2009de; Buckley:2017ttd; Shandera:2018xkn.
 Such atomic DM can exhibit dissipative dynamics, allowing it to cool and collapse differently from standard cold DM.
 This can lead to the formation of compact dark objects, modify binary black hole merger populations, or even produce sub-solar mass binaries that would be invisible electromagnetically.
-While GW searches have constrained only the presence of such atomic DM using matched filtering searches, planetary- or asteroid-mass systems (masses $0.1M_{\odot}$) could also have been formed from atomic DM Fan et al.
-(2013a, b); Buckley and DiFranzo (2018), and would emit longer-duration GWs.
-Thus, CW searches can be applied to look for such systems Miller et al.
-(2022a); Abbott et al.
-(2022b); Miller et al.
-(2024).
+While GW searches have constrained only the presence of such atomic DM using matched filtering searches, planetary- or asteroid-mass systems (masses $0.1M_{\odot}$) could also have been formed from atomic DM Fan:2013tia; Fan:2013yva; Buckley:2017ttd, and would emit longer-duration GWs.
+Thus, CW searches can be applied to look for such systems Miller:2021knj; KAGRA:2022dwb; Miller:2024fpo.
 In each of aforementioned scenarios, considering DM within the framework of CWs allows us to adapt methods originally designed to probe neutron stars to search for DM.
@@
 $\approx$ $1.69\times 10^{54}\left(\frac{10^{-12}\text{ eV}}{m_{\rm DM}}\right)^{4},$
-where $v_{0}\simeq 220$ km/s is the virial velocity Smith et al.
-(2007), $m_{\rm DM}$ is the mass of the ultralight dark matter particle, $\lambda_{\rm C}$ is the Compton wavelength of DM, and $\rho_{\text{DM}}=0.3\,\mathrm{GeV}/$cm3 is the DM energy density de Salas and Widmark (2021).
+where $v_{0}\simeq 220$ km/s is the virial velocity smith2007rave, $m_{\rm DM}$ is the mass of the ultralight dark matter particle, $\lambda_{\rm C}$ is the Compton wavelength of DM, and $\rho_{\text{DM}}=0.3\,\mathrm{GeV}/$cm3 is the DM energy density deSalas:2020hbh.
 The large occupation number $N_{0}$ implies that ultralight DM can be approximated as a plane wave that oscillates at a fixed frequency $f_{0}$, which is given by the DM mass:
 $f_{0}=\frac{m_{\rm DM}}{2\pi}\simeq 241\text{ Hz}\left(\frac{m_{\rm DM}}{10^{-12}\text{ eV}}\right),$ (2)
-In other words, the observable effect is that the interferometer components will forever oscillate at a fixed frequency since they always sit in the DM field Pierce et al. (2018).
+In other words, the observable effect is that the interferometer components will forever oscillate at a fixed frequency since they always sit in the DM field Pierce:2018xmy.
 Despite the fixed frequency implied by Eq.
@@
 This spread in velocities causes a small but finite spread in the oscillation frequency, which in turn limits the coherence of the field.
-As a result, the DM wave can be treated as coherent only over a characteristic time scale $T_{\text{coh}}$, after which the accumulated phase dispersion between different velocity components causes the signal to lose coherence Carney et al.
-(2021).
-The wave coherent time $T_{\text{coh}}$ is given by Carney et al.
-(2021):
+As a result, the DM wave can be treated as coherent only over a characteristic time scale $T_{\text{coh}}$, after which the accumulated phase dispersion between different velocity components causes the signal to lose coherence Carney:2019cio.
+The wave coherent time $T_{\text{coh}}$ is given by Carney:2019cio:
 $T_{\text{coh}}=\frac{4\pi}{m_{\rm DM}v_{0}^{2}}=1.4\times 10^{4}\text{ s}\left(\frac{10^{-12}\text{ eV}}{m_{\rm DM}}\right),$ (3)
@@
 Here, simulated vector DM particles are interacting with one of the LIGO detectors, which produces a stochastic, narrow-band signal in the frequency domain.
-Figure 3: Taken from Miller et al.
-(2021a).
+Figure 3: Taken from Miller:2020vsl.
 We show the strain time series $h(t)$ (left) and the Fourier transform of it (right) of a simulated dark photon DM signal.
@@
 We will assume that we are working within $T_{\text{coh}}$, in order to avoid technical difficulties from working with particles of slightly different frequencies.
-However, we note that in practice, this consideration is necessary when assessing the sensitivity of GW interferometers to DM/standard-model interactions Pierce et al.
-(2018); Miller et al.
-(2021a); Vermeulen et al.
-(2021); Göttel et al.
-(2024).
+However, we note that in practice, this consideration is necessary when assessing the sensitivity of GW interferometers to DM/standard-model interactions Pierce:2018xmy; Miller:2020vsl; Vermeulen:2021epa; Gottel:2024cfj.
 II.2 Axions
-Axions were originally proposed as a solution to the strong CP problem in QCD Peccei and Quinn (1977a, b), and more generally, axion-like particles arise in many extensions of the standard model.
+Axions were originally proposed as a solution to the strong CP problem in QCD Peccei:1977hh; Peccei:1977ur, and more generally, axion-like particles arise in many extensions of the standard model.
 If sufficiently light and weakly interacting, they can also constitute a viable component of DM.
 A key feature of axions is their ability to couple to standard-model fields, in particular photons, through a Chern–Simons interaction.
-This coupling is described by the Lagrangian Chern and Simons (1974):
+This coupling is described by the Lagrangian Chern:1974ft:
 $\mathcal{L}_{I}$ $=\frac{g_{\text{a}\gamma}}{4}a(t)F_{\mu\nu}\tilde{F}^{\mu\nu}$
@@
 […]
-the equations of motion for $A_{\text{L/R}}(t,\bm{k})$ can be derived Nagano et al. (2019):
+the equations of motion for $A_{\text{L/R}}(t,\bm{k})$ can be derived Nagano:2019rbw:
 $\ddot{A}_{\text{L/R}}+\omega^{2}_{\text{L/R}}A_{\text{L/R}}$ $=0$ (8)
@@
 11 implies that in the presence of an axion field, the phase velocities of circularly polarized photons in interferometers are modulated.
-In order to detect this effect in LIGO, Virgo and KAGRA instruments, additional but simplistic optical components are required in order to measure the optical path difference between two orthogonal polarizations of light ($p$- and $s$-polarizations) Nagano et al.
-(2021).
+In order to detect this effect in LIGO, Virgo and KAGRA instruments, additional but simplistic optical components are required in order to measure the optical path difference between two orthogonal polarizations of light ($p$- and $s$-polarizations) Nagano:2021kwx.
 In practice, GW interferometers use linearly polarized light ($p$-polarized), which becomes partially $s$-polarized due to axion-induced modulation.
-Similar deviations in polarization can also be probed with pulsar timing arrays Xue et al.
-(2024); Porayko et al.
-(2025), and experimental constraints on axions from them will be discussed in Section III.3.5.
+Similar deviations in polarization can also be probed with pulsar timing arrays Xue:2024zjq; EPTA:2024gxu, and experimental constraints on axions from them will be discussed in Section III.3.5.
 Notably, ground-based interferometers and pulsar timing arrays are sensitive to very different axion mass ranges, roughly $10^{-13},10^{-11}]$ eV and $10^{-23},10^{-21}]$ eV, respectively, but can achieve comparable absolute constraints on the axion-photon coupling $g_{\text{a}\gamma}$, which vary strongly with the axion mass.
@@
 II.3 Dilaton scalar dark matter
-Another class of ultralight dark matter candidates are dilatons, scalar fields that can couple to standard-model particles and induce time-dependent variations in fundamental constants Preskill et al.
-(1983); Abbott and Sikivie (1983); Dine and Fischler (1983); Cho and Keum (1998); Cho and Kim (2009); Arvanitaki et al.
-(2015b); Stadnik and Flambaum (2015a, b, 2016).
+Another class of ultralight dark matter candidates are dilatons, scalar fields that can couple to standard-model particles and induce time-dependent variations in fundamental constants Preskill:1982cy; Abbott:1982af; Dine:1982ah; Cho:1998aa; Cho:2007cy; Arvanitaki:2014faa; Stadnik2015a; Stadnik2015b; Stadnik2016.
 In particular, a dilaton field can cause the electron mass and other physical constants to oscillate at the dilaton frequency.
-Similarly to the axion, the scalar ultralight dark-matter field $\phi$ can be written as Arvanitaki et al.
-(2015b); Derevianko (2018); Vermeulen et al.
-(2021):
+Similarly to the axion, the scalar ultralight dark-matter field $\phi$ can be written as Arvanitaki:2014faa; Derevianko:2016vpm; Vermeulen:2021epa:
 $\phi(t,\vec{x})=\left(\frac{\sqrt{2\rho_{\rm{DM}}}}{m_{\rm DM}}\right)\cos\left(m_{\rm DM}t-\vec{k}\cdot\vec{x}+\Psi\right),$ (13)
-where $\vec{x}$ is a position vector, $\vec{k}=m_{\rm DM}\vec{v}_{0}$ is the wave vector, and $\Psi$ is a phase factor, and has a corresponding Lagrangian $\mathcal{L}_{\mathrm{int}}$ Vermeulen et al.
-(2021):
+where $\vec{x}$ is a position vector, $\vec{k}=m_{\rm DM}\vec{v}_{0}$ is the wave vector, and $\Psi$ is a phase factor, and has a corresponding Lagrangian $\mathcal{L}_{\mathrm{int}}$ Vermeulen:2021epa:
 $\mathcal{L}_{\mathrm{int}}\supset\frac{\phi}{\Lambda_{\gamma}}\frac{F_{\mu\nu}F^{\mu\nu}}{4}-\frac{\phi}{\Lambda_{e}}m_{e}\bar{\psi}_{e}\psi_{e},$ (14)
@@
 II.3.1 Changes to size and indices of refraction of solids
-The oscillations in fundamental constants induced by a dilaton field lead to modulations of the Bohr radii of atoms in interferometer components Grote and Stadnik (2019), leading to measurable changes in the size and refractive index of the beam splitter Vermeulen et al.
-(2021), and the arm mirrors Göttel et al.
-(2024).
+The oscillations in fundamental constants induced by a dilaton field lead to modulations of the Bohr radii of atoms in interferometer components grote2019novel, leading to measurable changes in the size and refractive index of the beam splitter Vermeulen:2021epa, and the arm mirrors Gottel:2024cfj.
 Oscillations in the beam splitter size cause photons traveling down each arm to traverse slightly different paths along its surface, producing a differential phase shift.
-Initially, it was assumed that the beam-splitter effect was the dominant contribution Vermeulen et al.
-(2021), but more recent work has shown that small differences in the thicknesses of the mirrors in each arm also lead to non-negligible path-length variations in the interferometer Göttel et al.
-(2024).
+Initially, it was assumed that the beam-splitter effect was the dominant contribution Vermeulen:2021epa, but more recent work has shown that small differences in the thicknesses of the mirrors in each arm also lead to non-negligible path-length variations in the interferometer Gottel:2024cfj.
 We will now calculate the impact of dilatons on GW interferometers.
@@
 In the context of GW interferometers, this effect translates into a differential displacement of the beam splitter along the two arms.
-For the $x$-arm, assuming light incidence at 45 degrees and 50% mirror reflectivity, the induced change in arm length, $\delta L_{x}$, can be computed as in Grote and Stadnik (2019):
+For the $x$-arm, assuming light incidence at 45 degrees and 50% mirror reflectivity, the induced change in arm length, $\delta L_{x}$, can be computed as in grote2019novel:
 $\delta L_{x}\approx\delta[\sqrt{2}nl-l/(2\sqrt{2})-w/2],$ (23)
-where $w$ is the thickness of the mirrors, $n$ is the index of the beamsplitter, and $l$ is the length of the beamsplitter. Along the $y$-arm, the change in arm length is Vermeulen et al. (2021):
+where $w$ is the thickness of the mirrors, $n$ is the index of the beamsplitter, and $l$ is the length of the beamsplitter. Along the $y$-arm, the change in arm length is Vermeulen:2021epa:
 $\delta L_{y}=-\delta l/(2\sqrt{2})-\delta w/2,$ (24)
@@
 $\approx\left(\frac{1}{\Lambda_{\gamma}}+\frac{1}{\Lambda_{e}}\right)\left(\frac{n\,l\,\sqrt{2\,\rho_{\mathrm{DM}}}}{m_{\rm DM}\,}\right)\cos\left(m_{\rm DM}t\right),$ (26)
-We have only considered the contribution arising from the change in length of the beamsplitter, not from the change in index of refraction, because the former is much larger than the latter Grote and Stadnik (2019).
+We have only considered the contribution arising from the change in length of the beamsplitter, not from the change in index of refraction, because the former is much larger than the latter grote2019novel.
 In addition to dilatons interacting with atoms in the beamsplitter, the reference cavity of the detector can provide a region in which dilatons interact appreciably with light, allowing for potential detection.
-Laser light locked to an optical cavity made out of a solid material will have its frequency modulated due to the presence of a dilaton DM field, compared to light locked to a free-space suspended cavity, whose frequency would remain stable Derevianko (2018); Geraci et al.
-(2019); Hall and Aggarwal (2022).
+Laser light locked to an optical cavity made out of a solid material will have its frequency modulated due to the presence of a dilaton DM field, compared to light locked to a free-space suspended cavity, whose frequency would remain stable Derevianko:2016vpm; Geraci:2018fax; Hall:2022zvi.
 The strain amplitude of such a signal, which would be distinct from that in Eq.
@@
 Currently, however, only the spatial strain affecting mirrors and the beamsplitter has motivated dedicated search methods for dilaton dark matter (see Section III.1) and has been constrained using data from GEO600 and LIGO (see Section III.2).
-A search strategy targeting the temporal modulation of the laser frequency is still under development Hall and Aggarwal (2022).
+A search strategy targeting the temporal modulation of the laser frequency is still under development Hall:2022zvi.
 II.3.2 Changes to neutron-star moment of inertia
 Beyond laboratory-based detectors, dilaton-induced fluctuations in fundamental constants could also have observable consequences in astrophysical systems.
-For example, they could alter moment of inertia of neutron stars Kaplan et al.
-(2022), which would produce deviations in the arrival time of pulses, to be discussed in Section III.3.
+For example, they could alter moment of inertia of neutron stars Kaplan:2022lmz, which would produce deviations in the arrival time of pulses, to be discussed in Section III.3.
 In order to quantify the extent to which the neutron star moment of inertia will change, we must know how protons and neutrons will change in response to dilatons.
-This motivates us to consider the following interactions in QCD of quarks in the presence of dilatons Damour and Donoghue (2010); Kaplan et al.
-(2022):
+This motivates us to consider the following interactions in QCD of quarks in the presence of dilatons Damour:2010rp; Kaplan:2022lmz:
 $\mathcal{L}_{\phi,\rm QCD}\supset\frac{\phi}{\Lambda}\left(\frac{d_{g}\beta_{3}}{2g_{3}}G_{\mu\nu}^{A}G_{A}^{\mu\nu}-\sum_{q=u,d}(d_{q}+\gamma_{q}d_{g})m_{q}\bar{q}q\right)$ (29)
@@
 The shift in nucleon masses can then be obtained by noting that nucleons are composite objects whose masses arise mostly from QCD binding energy, with a smaller contribution from the quark masses.
-The dilaton couples to both the quark masses and the gluon field strength, so the fractional change in a nucleon mass is a weighted sum of these effects Damour and Donoghue (2010); Kaplan et al.
-(2022):
+The dilaton couples to both the quark masses and the gluon field strength, so the fractional change in a nucleon mass is a weighted sum of these effects Damour:2010rp; Kaplan:2022lmz:
 $\frac{\delta m_{p,n}}{m_{p,n}}\simeq\frac{1}{\Lambda}\Big(d_{g}+C_{n}\,d_{\hat{m}}\Big)\phi\,$, (32)
@@
 where $I_{0}$ is the unperturbed moment of inertia and $\omega_{0}$ is the unperturbed angular frequency of the neutron star.
-Assuming a simplistic model of a non-rotating, spherically symmetric neutron star composed only of neutrons with a polytropic equation-of-state, the radius and moment of inertia can be written as Shapiro and Teukolsky (1983):
+Assuming a simplistic model of a non-rotating, spherically symmetric neutron star composed only of neutrons with a polytropic equation-of-state, the radius and moment of inertia can be written as Shapiro:1983du:
 $R$ $\propto M_{0}^{-1/3}m_{n}^{-8/3}$ (35)
@@
 The sum is over the different types of particles present in the pulsar: electrons, muons, protons and neutrons, $Y_{f}\equiv N_{f}/(N_{n}+N_{p})$ is a weighting function that gives the number of $f$ particles relative to the sum of protons $N_{p}$ and neutrons $N_{n}$ (which dominate the pulsar’s mass).
-For neutron stars observed by pulsar timing arrays, the following is assumed for $Y_{f}$: $Y_{n}=0.9$, $Y_{p}=Y_{e}=0.1$ and $Y_{\mu}=0.05$ Bell et al.
-(2019); Kaplan et al.
-(2022).
+For neutron stars observed by pulsar timing arrays, the following is assumed for $Y_{f}$: $Y_{n}=0.9$, $Y_{p}=Y_{e}=0.1$ and $Y_{\mu}=0.05$ Bell:2019pyc; Kaplan:2022lmz.
 By plugging in the relations in Eq.
@@
 $\begin{split}\{y_{g},\,y_{\hat{m}},\,y_{\mu},\,y_{e}\}=&\,\eta\left\{1,\,C_{n},\,6\times 10^{-3},\,5\times 10^{-5}\right\}\\ +&\delta\eta\left\{1,\,C_{n},\,0,\,0\right\}\,.\end{split}$ (41)
-Using the above formalism, dilatons have been constrained with recent pulsar timing array experiments Porayko et al.
-(2018); Smarra et al.
-(2023); Afzal et al.
-(2023), which will be extensively discussed in Section III.3.
+Using the above formalism, dilatons have been constrained with recent pulsar timing array experiments Porayko:2018sfa; EuropeanPulsarTimingArray:2023egv; NANOGrav:2023hvm, which will be extensively discussed in Section III.3.
 II.3.3 Changes to atomic clock times
 Changes in fundamental constants would also affect the way in which atomic clocks measure time.
-Since pulsar timing array experiments reference their measurements of the ToAs of pulses to terrestrial time (TT) using atomic clocks composed primarily of cesium, any oscillation of the fundamental constants will affect the atomic clocks, and thus the times measured Flambaum et al.
-(2004).
+Since pulsar timing array experiments reference their measurements of the ToAs of pulses to terrestrial time (TT) using atomic clocks composed primarily of cesium, any oscillation of the fundamental constants will affect the atomic clocks, and thus the times measured Flambaum:2004tm.
 Likewise, a shift in the frequency of the atomic clock will give rise to an apparent shift in the measured pulsar spin frequency, given by Eq.
 38.
-From Flambaum et al. (2004); Kaplan et al. (2022), the frequency of an atomic clock is:
+From Flambaum:2004tm; Kaplan:2022lmz, the frequency of an atomic clock is:
 $f\propto\Big(m_{e}\alpha^{2}\Big)\Big[\alpha^{2}F_{\rm rel}(Z\alpha)\Big]\left(\mu\frac{m_{e}}{m_{p}}\right)^{\zeta}$ (42)
@@
 The fluctuations in fundamental constants will cause modulations in the frequencies of atomic clocks composed of atoms $A$.
-Using previously derived expressions for variations in the electron mass, fine-structure constant, and quark and nucleon masses, the fractional frequency shift can be obtained Flambaum et al.
-(2004); Kaplan et al.
-(2022):
+Using previously derived expressions for variations in the electron mass, fine-structure constant, and quark and nucleon masses, the fractional frequency shift can be obtained Flambaum:2004tm; Kaplan:2022lmz:
 […]
 where $\delta F_{\rm rel}/F_{\rm rel}=K_{A}$, and the derivative of the nuclear magnetic moment is $\delta\mu/\mu=C_{A}\,\delta m_{q}/m_{q}$.
-For Cesium, $K_{A}=0.83$ depends on $Z$ and $\alpha$, and $C_{A}=0.110$ Flambaum et al.
-(2004).
+For Cesium, $K_{A}=0.83$ depends on $Z$ and $\alpha$, and $C_{A}=0.110$ Flambaum:2004tm.
 The index $q$ runs over up and down quarks.
@@
 where $\xi_{A}\equiv 4+K_{A}$, and $\hat{C}_{A}=C_{A}(m_{u}+m_{d})^{2}/2m_{u}m_{d}$.
-This type of DM interaction would leave imprints on pulsar timing array experiments Porayko et al.
-(2018); Smarra et al.
-(2023); Afzal et al.
-(2023), which will be discussed in Section III.3.
+This type of DM interaction would leave imprints on pulsar timing array experiments Porayko:2018sfa; EuropeanPulsarTimingArray:2023egv; NANOGrav:2023hvm, which will be discussed in Section III.3.
 We would like to draw a parallel between Eqs.
@@
 vectors, which we refer to as “dark photons”.
-Like dilatons, dark photons could explain the entirety of the relic abundance of DM, which could arise from the misalignment mechanism Nelson and Scholtz (2011); Arias et al.
-(2012); Graham et al.
-(2016), parametric resonance or the tachyonic instability of a scalar field Agrawal et al.
-(2020); Co et al.
-(2019); Bastero-Gil et al.
-(2019); Dror et al.
-(2019), or from cosmic string network decays Long and Wang (2019).
+Like dilatons, dark photons could explain the entirety of the relic abundance of DM, which could arise from the misalignment mechanism nelson2011dark; arias2012wispy; graham2016vector, parametric resonance or the tachyonic instability of a scalar field agrawal2020relic; Co:2018lka; Bastero-Gil:2018uel; Dror:2018pdh, or from cosmic string network decays long2019dark.
 They could couple to standard-model particles, either to baryon or baryon-lepton number.
 In the interferometers, these interactions would occur everywhere, but would be most pronounced in the four interferometer test masses in the main resonant cavities of the instrument.
-The dark photons [^4] These are different dark photons than those that kinetically mix with the ordinary photon, which will be discussed later in the context of luminous superradiance Siemonsen et al.
-(2023a).
-would cause a “dark” force on the mirrors, causing quasi-sinusoidal oscillations Pierce et al.
-(2018); Miller et al.
-(2021a), as will be shown below.
+The dark photons [^4] These are different dark photons than those that kinetically mix with the ordinary photon, which will be discussed later in the context of luminous superradiance Siemonsen:2022ivj.
+would cause a “dark” force on the mirrors, causing quasi-sinusoidal oscillations Pierce:2018xmy; Miller:2020vsl, as will be shown below.
 To derive the impact of dark photons on GW interferometers, we will begin by describing dark photons analogously to ordinary photons, i.e. with a four-vector potential:
@@
 The acceleration of a given mirror can thus be derived using Eqs.
-49 and 48 Guo et al.
-(2019); Pierce et al.
-(2018):
+49 and 48 Guo:2019ker; Pierce:2018xmy:
 $\vec{a}_{j}(t,\vec{x}_{j})$ $=\simeq\epsilon e\frac{Q_{D,j}}{M_{j}}m_{\rm DM}|\vec{A}_{0}|\hat{A}\cos(m_{\rm DM}t-\vec{k}\cdot\vec{x}_{j}+\Upsilon),$
@@
 Thus, the interferometer experiences a measurable differential acceleration, and thus a differential strain.
-Assuming that all mirrors have the same charge-to-mass ratio, integrating LABEL:eqn:accel twice over time, and averaging over random polarization and propagation directions, the strain on the interferometers caused by a dark photon DM signal can be computed Pierce et al.
-(2018):
+Assuming that all mirrors have the same charge-to-mass ratio, integrating LABEL:eqn:accel twice over time, and averaging over random polarization and propagation directions, the strain on the interferometers caused by a dark photon DM signal can be computed Pierce:2018xmy:
 $\sqrt{\langle h^{2}_{D}\rangle}$ $=C\frac{Q}{M}e\sqrt{2\rho_{\text{DM}}}v_{0}\frac{\epsilon}{f_{0}},$
 $\simeq 6.56\times 10^{-26}\left(\frac{\epsilon}{10^{-22}}\right)\left(\frac{100\text{ Hz}}{f_{0}}\right),$ (54)
-where $C=\sqrt{2}/3$ (for an interferometer with two perpendicular arms) is a geometrical factor obtained by averaging over all possible dark photon propagation and polarization directions, and different detector geometries (the calculation for $C$ is shown in the appendix of Pierce et al.
-(2018) and in Morisaki et al.
-(2021)).
+where $C=\sqrt{2}/3$ (for an interferometer with two perpendicular arms) is a geometrical factor obtained by averaging over all possible dark photon propagation and polarization directions, and different detector geometries (the calculation for $C$ is shown in the appendix of Pierce:2018xmy and in Morisaki:2020gui).
 A second strain results due to the so-called “finite light travel time effect, in which the mirrors in the interferometers have moved in the time that it takes the light to reach them from the beamsplitter.
 This strain can actually be larger than that given in Eq.
-54 and can be derived Morisaki et al.
-(2021):
+54 and can be derived Morisaki:2020gui:
 $\sqrt{\langle h_{C}^{2}\rangle}$ $=\frac{\sqrt{3}}{2}\sqrt{\langle h_{D}^{2}\rangle}\frac{2\pi f_{0}L}{v_{0}},$
@@
 Modifications to gravity due to an additional spin-2 particle could act as DM.
-Specifically, bimetric gravity Hassan and Rosen (2012); Schmidt-May and von Strauss (2016), a theory in which a massless and massive spin-2 field interact, provides a plausible candidate for ultralight dark matter Marzola et al.
-(2018).
-Here, we consider a massive spin-2 field $M_{\mu\nu}$ described by the Fierz-Pauli Lagrangian density Fierz and Pauli (1939); Armaleo et al. (2020):
+Specifically, bimetric gravity Hassan:2011zd; Schmidt-May:2015vnx, a theory in which a massless and massive spin-2 field interact, provides a plausible candidate for ultralight dark matter Marzola:2017lbt.
+Here, we consider a massive spin-2 field $M_{\mu\nu}$ described by the Fierz-Pauli Lagrangian density Fierz:1939ix; Armaleo:2020yml:
 ${\mathcal{L}}$ $\,:=\frac{1}{2}M_{\mu\nu}\mathcal{E}^{\mu\nu\rho\sigma}M_{\rho\sigma}-\frac{1}{4}m_{\rm DM}^{2}\left(M_{\mu\nu}M^{\mu\nu}-M^{2}\right)\,$, (56)
-where $M:=g^{\mu\nu}M_{\mu\nu}$ and $\mathcal{E}^{\mu\nu\rho\sigma}$ is the Lichnerowicz operator, defined in equation 2.2 of Armaleo et al. (2020).
-For the Friedman-Lemaitre-Robertson-Walker (FLRW) background metric, the equations of motion for this field, assuming it is ultralight, can be derived at late times as in  Marzola et al.
-(2018); Aoki and Maeda (2018); Kun et al.
-(2018).
+where $M:=g^{\mu\nu}M_{\mu\nu}$ and $\mathcal{E}^{\mu\nu\rho\sigma}$ is the Lichnerowicz operator, defined in equation 2.2 of Armaleo:2020yml.
+For the Friedman-Lemaitre-Robertson-Walker (FLRW) background metric, the equations of motion for this field, assuming it is ultralight, can be derived at late times as in  Marzola:2017lbt; Aoki:2017cnz; Kun:2018ino.
 The field spatial component can be written in a similar way as for dark photons and scalar bosons at given position:
@@
 TB is tensor boson.
-Armaleo et al.
-(2020); thus, the perturbation $h_{ij}$ is:
+Armaleo:2020yml; thus, the perturbation $h_{ij}$ is:
 $h_{ij}(t)$ $=\frac{\alpha_{\rm TB}}{M_{\text{Pl}}}M_{ij}(t)=\frac{\alpha_{\rm TB}\sqrt{2\rho_{\text{DM}}}}{m_{\rm DM}M_{\text{Pl}}}\cos{\left(m_{\rm DM}t+\Upsilon\right)}\varepsilon_{ij}(\vec{x})\,$, (58)
 where $\alpha_{\rm TB}$ is a dimensionless constant that quantifies the difference between the strengths of each of the spin-2 fields, and $M_{\text{Pl}}$ is the Planck mass.
-To obtain the strain on GW interferometers, we note that the detector response function can be parameterized in terms of a tensor, often written as $D^{ij}$ Romano and Cornish (2017).
-By contracting $h_{ij}$ with $D^{ij}$, the resulting strain can be derived Armaleo et al.
-(2021).
+To obtain the strain on GW interferometers, we note that the detector response function can be parameterized in terms of a tensor, often written as $D^{ij}$ Romano:2016dpx.
+By contracting $h_{ij}$ with $D^{ij}$, the resulting strain can be derived Armaleo:2020efr.
 $h(t)=D^{ij}h_{ij}=\frac{\alpha_{\rm TB}\sqrt{\rho_{\text{DM}}}}{\sqrt{2}mM_{\text{Pl}}}\cos{\left(m_{\rm DM}t+\Upsilon\right)}\Delta\varepsilon,$ (59)
@@
 We note that an interesting discrepancy arises when comparing the results from Refs.
-Aoki and Mukohyama (2016); Manita et al.
-(2024) with those from Refs.
-Marzola et al.
-(2018); Armaleo et al.
-(2021) regarding the strain induced by tensor dark matter.
+Aoki:2016zgp; Manita:2023mnc with those from Refs.
+Marzola:2017lbt; Armaleo:2020efr regarding the strain induced by tensor dark matter.
 The two sets of works differ cause a factor of 2 difference in the strain, which can be traced back to a factor of eight difference in their respective Lagrangians.
@@
 Here, ultralight dark matter can source time-varying gravitational potentials, leading to observable effects such as oscillations in pulsar timing residuals, which will be discussed in Section III.3.
-A commonly studied example of a minimally coupled scalar DM field $\phi$ is described by the Lagrangian Khmelnitsky and Rubakov (2014):
+A commonly studied example of a minimally coupled scalar DM field $\phi$ is described by the Lagrangian Khmelnitsky:2013lxt:
 $\mathcal{L}=\sqrt{-g}\left[\frac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{2}m_{\rm DM}^{2}\phi^{2}\right]\,.$ (60)
@@
 As before, the scalar field density is normalized by $\rho_{\text{DM}}$.
-Scalar DM can cause the arrival times of pulses from millisecond pulsars on earth to change, in an analogous way that GWs do Khmelnitsky and Rubakov (2014), i.e.
-introduce stochastic fluctuations of the metric Kim and Mitridate (2024); Kim (2023).
-Such a difference in the arrival times can be written as Khmelnitsky and Rubakov (2014); Porayko et al.
-(2018):
+Scalar DM can cause the arrival times of pulses from millisecond pulsars on earth to change, in an analogous way that GWs do Khmelnitsky:2013lxt, i.e.
+introduce stochastic fluctuations of the metric Kim:2023kyy; Kim:2023pkx.
+Such a difference in the arrival times can be written as Khmelnitsky:2013lxt; Porayko:2018sfa:
 $\delta t_{\text{DM}}=\frac{\Psi_{\text{c}}(\vec{x})}{2m_{\rm DM}}[\hat{\phi}^{2}_{\text{E}}\sin{(2m_{\rm DM}+\gamma_{\text{E}})}-\hat{\phi}^{2}_{\text{P}}\sin{(2m_{\rm DM}+\gamma_{\text{P}})}],$ (62)
@@
 This kind of DM typically involve scalar fields that couple to the trace of the energy-momentum tensor.
-These models are closely related to the dilaton-type scalar DM discussed in Section II.3, but with specific couplings motivated by scalar–tensor theories of gravity, such as Fierz–Jordan–Brans–Dicke (FJBD) and Damour–Esposito–Farese (DEF) theories Fierz and Pauli (1939); Fierz (1956); Dicke (1962); Jordan (1959); Damour and Esposito-Farese (1992, 1993).
+These models are closely related to the dilaton-type scalar DM discussed in Section II.3, but with specific couplings motivated by scalar–tensor theories of gravity, such as Fierz–Jordan–Brans–Dicke (FJBD) and Damour–Esposito–Farese (DEF) theories Fierz:1939ix; Fierz:1956zz; Dicke:1961gz; Jordan:1959eg; Damour:1992we; Damour:1993hw.
 In such models, DM couples universally to the gravitational sector through a conformal rescaling of the metric.
-This coupling can induce a gravity-mediated force between neutron stars and the ultralight dark matter field, altering the moment of inertia of rotating pulsars – a manifestation of the Nordtvedt effect Nordtvedt (1968).
+This coupling can induce a gravity-mediated force between neutron stars and the ultralight dark matter field, altering the moment of inertia of rotating pulsars – a manifestation of the Nordtvedt effect Nordtvedt:1968qs.
 These variations lead to deterministic changes in pulsar ToA and constitute a violation of the equivalence principle.
-This kind of conformal scalar DM has been extensively studied and constrained using binary pulsar timing measurements Damour and Taylor (1992); Kramer et al.
-(2006), which will be discussed inSection III.3.
+This kind of conformal scalar DM has been extensively studied and constrained using binary pulsar timing measurements Damour:1991rd; Kramer:2006nb, which will be discussed inSection III.3.
 In the case of a linear FJBD coupling, the pulsar timing residuals take the form:
 $\Delta t(t)$ $=2\alpha s_{I}\left.\frac{\sqrt{\rho_{\text{DM}}}}{M_{\text{Pl}}m_{\rm DM}^{2}}\hat{\phi}(\vec{x})\sin(m_{\rm DM}t+\theta(\vec{x}))\right|_{t_{\text{start }}-d}^{t_{\text{end }}-d},$ (64)
-where $s_{I}$ is the angular moment sensitivity parameter computed in Kuntz and Barausse (2024) and is of $\mathcal{O}(1)$ in FJBD theory, $\alpha$ is the strength of the coupling, $\theta(\vec{x})$ is a random phase, and $\hat{\phi}$ is a stochastic parameter whose value depends on whether the correlated, uncorrelated or pulsar correlated scenarios are considered (see Section III.3 for a description of these terms).
+where $s_{I}$ is the angular moment sensitivity parameter computed in Kuntz:2024jxo and is of $\mathcal{O}(1)$ in FJBD theory, $\alpha$ is the strength of the coupling, $\theta(\vec{x})$ is a random phase, and $\hat{\phi}$ is a stochastic parameter whose value depends on whether the correlated, uncorrelated or pulsar correlated scenarios are considered (see Section III.3 for a description of these terms).
 In summary, both dilaton-type and conformal scalar DM predict couplings that can modify either the properties of matter or the effective gravitational interaction, leading to observable signatures.
@@
 III.1 Methods
-As explained in Section II.1, the expected signal frequency is fixed by the DM mass, but has some stochastic variations of $\mathcal{O}(10^{-6})$ Centers et al.
-(2021).
+As explained in Section II.1, the expected signal frequency is fixed by the DM mass, but has some stochastic variations of $\mathcal{O}(10^{-6})$ Centers:2019dyn.
 Because the interferometers always exist in the DM field, the signal is always present.
@@
 III.1.1 Cross-correlation
-Conceptually, cross-correlation Pierce et al.
-(2018); Guo et al.
-(2019) requires at least two separate time-domain datasets that are Fourier transformed and multiplied together to compute the cross-power in each frequency bin.
+Conceptually, cross-correlation Pierce:2018xmy; Guo:2019ker requires at least two separate time-domain datasets that are Fourier transformed and multiplied together to compute the cross-power in each frequency bin.
 The cross power is then divided by the auto-power (the power spectral density) of each detector, and then summed over all the Fourier transforms to arrive at a measure of power at each frequency.
 It is then divided by the standard deviation of the noise to compute the signal-to-noise ratio (SNR).
-Mathematically, the cross-correlated signal strength for detector pair $IJ$ is Pierce et al. (2018); Guo et al. (2019):
+Mathematically, the cross-correlated signal strength for detector pair $IJ$ is Pierce:2018xmy; Guo:2019ker:
 $S_{IJ,j}=\frac{1}{N_{\rm FFT}}\sum_{i=1}^{N_{\rm FFT}}\frac{z_{I,ij}z_{J,ij}^{*}}{P_{I,ij}P_{J,ij}},$ (65)
@@
 Typically, $T_{\text{FFT}}=1800$ s, independently of the DM mass, which ensures a relatively straightforward analysis but implies the sensitivity towards DM interactions is not optimal at each mass analyzed.
-In addition to the signal strength, we have to obtain an estimate of the variance of the noise data. This can be computed from the individual power spectral densities in each FFT as Guo et al. (2019):
+In addition to the signal strength, we have to obtain an estimate of the variance of the noise data. This can be computed from the individual power spectral densities in each FFT as Guo:2019ker:
 $\sigma_{IJ,j}^{2}=\frac{1}{N_{\rm FFT}}\left\langle\frac{1}{2P_{I,ij}P_{J,ij}}\right\rangle_{N_{\rm FFT}},$ (66)
@@
 If the SNR exceeds a certain threshold, which is set both theoretically assuming Gaussian noise and by the trials factor (accounting for the size of the parameter space), then a particular frequency is classified as being “significant”.
-Depending on the type of ultralight dark matter being searched for, the detectability of a signal in cross-correlation searches between interferometers depends on the spin of the DM: spin-0 (scalar), spin-1 (vector), or spin-2 (tensor) Manita et al.
-(2024).
+Depending on the type of ultralight dark matter being searched for, the detectability of a signal in cross-correlation searches between interferometers depends on the spin of the DM: spin-0 (scalar), spin-1 (vector), or spin-2 (tensor) Manita:2023mnc.
 The SNR for a given detector pair is determined by the DM energy density fraction in that spin state, the strength of its couplings to standard-model particles, the observation time, and the relative orientation, location and geometry of the detectors.
 The latter is characterized by an “overlap reduction function” (ORF), which determines the sensitivity loss when applying cross-correlation to non-collocated, non-aligned detector pairs.
-For these models of DM, two kinds of ORFs arise: one from the spatial displacement of mirrors, and another from the finite-light time travel effects of the signal on the mirrors, a previously explained in Section II.4 Morisaki et al.
-(2021).
+For these models of DM, two kinds of ORFs arise: one from the spatial displacement of mirrors, and another from the finite-light time travel effects of the signal on the mirrors, a previously explained in Section II.4 Morisaki:2020gui.
 In particular, cross-correlation is not nearly as sensitive to the finite-light travel time effect as to the spatial effect when it was applied to search for spin-1 DM, which will be shown in Section III.2.
@@
 68, it can be translated directly from SNR to coupling strengths using Eq.
-120-122 in Manita et al.
-(2024)).
+120-122 in Manita:2023mnc).
 It is also useful to calculate a minimum detectable strain amplitude, which can also be used to place constraints on the DM coupling constants for scalar, vector, or tensor ultralight dark matter.
 Given the SNR, the corresponding strain amplitude can be estimated as:
-$h_{0,j}=\sqrt{\frac{2\,\mathrm{SNR}\,P_{I,j}\,P_{J,j}}{\gamma^{2}\,T_{\text{obs}}\,T_{\text{FFT}}}},$ (68)
-where $\gamma$ is the ORF that depends on the way that different kinds of DM couple to the interferometers Manita et al.
-(2024).
+$h_{0,j}=\left(\frac{2\,\mathrm{SNR}}{\gamma}\right)^{1/2}\left(\frac{\,P_{I,j}\,P_{J,j}}{T_{\text{obs}}\,T_{\text{FFT}}}\right)^{1/4},$ (68)
+where $\gamma$ is the ORF that depends on the way that different kinds of DM couple to the interferometers Manita:2023mnc.
 By inserting the SNR returned from a search (Eq.
-67), we can obtain an estimate of the minimum detectable strain amplitude for a potential DM signal, and then map those amplitude values to constraints on DM couplings in spin-0, spin-1 or spin-2 using the strain amplitude equations derived in Section II, analogously to translating the SNR directly to coupling strengths in Manita et al.
-(2024).
-III.1.2 Excess power method
+67), we can obtain an estimate of the minimum detectable strain amplitude for a potential DM signal, and then map those amplitude values to constraints on DM couplings in spin-0, spin-1 or spin-2 using the strain amplitude equations derived in Section II, analogously to translating the SNR directly to coupling strengths in Manita:2023mnc.
+III.1.2 BSD Excess power method
 Cross-correlation is able to capture the phase information in each FFT; however, at the moment the software has not yet been implemented to change $T_{\text{FFT}}$ to match $T_{\text{coh}}$, which would result in optimal sensitivity to each DM mass.
-Thus, another method Miller et al.
-(2021a) was developed that can vary $T_{\text{FFT}}$ as a function of frequency, and that is more robust against noise disturbances.
+Thus, another method Miller:2020vsl called “BSD excess power” was developed that can vary $T_{\text{FFT}}$ as a function of frequency that employs Band-Sampled Data (BSD) structures that allow easy changes of $T_{\text{FFT}}$ piccinibsd, and that is more robust against noise disturbances.
 It relies on creating time-frequency spectrograms in which frequencies in each FFT are only kept if their equalized powers are (1) above a given threshold $\theta_{\rm thr}=2.5$ and (2) local maxima.
@@
 The $CR$, like the SNR, follows a normal distribution.
-Using the $CR$ and assuming Gaussian noise, the minimum strain amplitude of a sinusoidal signal can be derived that would, in a frequentest interpretation, produce a detectable signal in a fraction $\geq\Gamma$ of a repeated number of experiments Astone et al.
-(2014); Miller et al.
-(2021a):
+Using the $CR$ and assuming Gaussian noise, the minimum strain amplitude of a sinusoidal signal can be derived that would, in a frequentest interpretation, produce a detectable signal in a fraction $\geq\Gamma$ of a repeated number of experiments Astone:2014esa; Miller:2020vsl:
 $h_{\text{0,min}}$ $\approx\frac{\mathcal{G}}{T_{\text{obs}}^{1/4}T_{\text{FFT}}^{1/4}}\sqrt{\frac{P_{I}(f)}{\theta_{\text{thr}}}}$
@@
 The $\mathcal{O}(1)$ differences arise from averages over the spin-1 or spin-2 polarizations, and are valid for ground-based L-shaped GW interferometers (for triangular shapes, the averages will be slightly different).
-The $\frac{\sqrt{2}}{3}$ factor comes from an average over the DM Maxwell-Boltzmann distributed velocities as discussed in Section II.4, while the $\frac{2\pi}{2.4308}$ factor comes from convolving a sinusoid with a rectangular window function when computing the sensitivity of GW interferometers to a sinusoidal signal Astone et al.
-(2014).
+The $\frac{\sqrt{2}}{3}$ factor comes from an average over the DM Maxwell-Boltzmann distributed velocities as discussed in Section II.4, while the $\frac{2\pi}{2.4308}$ factor comes from convolving a sinusoid with a rectangular window function when computing the sensitivity of GW interferometers to a sinusoidal signal Astone:2014esa.
 III.1.3 Logarithmic power spectral density (LPSD) method
 To maximize sensitivity to ultralight dark-matter signals, it is important to let $T_{\text{FFT}}$ match $T_{\text{coh}}$ as the DM mass varies.
-The logarithmic power spectral density (LPSD) method Tröbs and Heinzel (2006); Vermeulen et al.
-(2021); Göttel et al.
-(2024) implements this idea by constructing spectra that are logarithmically spaced in frequency, in contrast to the excess power method in Section III.1.2 that use linearly spaced spectra every 1 Hz.
+The logarithmic power spectral density (LPSD) method trobs2006improved; Vermeulen:2021epa; Gottel:2024cfj implements this idea by constructing spectra that are logarithmically spaced in frequency, in contrast to the BSD excess power method in Section III.1.2 that use linearly spaced spectra every 1 Hz.
 While this approach is computationally intensive due to the need for a separate implementation from the standard discrete FFT, it captures subtle features in the data that could otherwise be missed.
@@
 74 to the logarithmically estimated spectra, one can extract an estimate of the strain $h(\omega)$ from the data.
-This estimate can then be mapped to constraints on the coupling constants $\Lambda_{\gamma}$ and $\Lambda_{e}$, providing a direct connection between the interferometer measurement and the underlying ultralight dark matter model Göttel et al.
-(2024); Cowan et al.
-(2011).
+This estimate can then be mapped to constraints on the coupling constants $\Lambda_{\gamma}$ and $\Lambda_{e}$, providing a direct connection between the interferometer measurement and the underlying ultralight dark matter model Gottel:2024cfj; Cowan:2010js.
 In this way, the LPSD method offers a way to detect logarithmically-spaced frequency modulations that might be missed by linear spectral analyses.
 III.1.4 Stochastic summing method
-The stochastic summing method Nakatsuka et al.
-(2023) exploits the fact that when $T_{\text{FFT}}>T_{\text{coh}}$, the signal power from ultralight dark matter is spread stochastically over a range of frequencies rather than concentrated at a single frequency.
+The stochastic summing method Nakatsuka:2022gaf exploits the fact that when $T_{\text{FFT}}>T_{\text{coh}}$, the signal power from ultralight dark matter is spread stochastically over a range of frequencies rather than concentrated at a single frequency.
 Instead of searching for a narrow peak, this method sums the power across all frequency bins where the signal is expected, producing a detection statistic that captures the total available signal power.
@@
 The outer sum is over all chunks of length $T_{\text{FFT}}$; the inner sum is over the frequency spread of the signal $\delta f_{v}$.
-The sensitivity of the stochastic summing method has been studied for both axions and dark photons Nakatsuka et al.
-(2023), using a combination of analytical estimates and numerical simulations.
+The sensitivity of the stochastic summing method has been studied for both axions and dark photons Nakatsuka:2022gaf, using a combination of analytical estimates and numerical simulations.
 In the absence of a signal and assuming Gaussian noise, a detection threshold can be established based on a chosen false-alarm probability.
@@
 The exact sensitivity depends on the spectral shape of the signal, which in turn is influenced by the relative motion of the DM with respect to the interferometer.
-Conservative and optimistic assumptions about the direction of the DM velocity lead to $\mathcal{O}(1)$ variations in the sensitivity, which are accounted for in this method Nakatsuka et al.
-(2023).
+Conservative and optimistic assumptions about the direction of the DM velocity lead to $\mathcal{O}(1)$ variations in the sensitivity, which are accounted for in this method Nakatsuka:2022gaf.
 Once the minimum detectable strain amplitude is determined, it can be translated into upper limits on the relevant coupling constants for each type of dark matter interaction, similarly to what has been discussed in Sections III.1.1 and III.1.2.
@@
 In practice, because data are typically divided into segments with $T_{\text{FFT}}\sim T_{\text{coh}}$, the analytic likelihood approach can become unstable.
-Consequently, numerical simulations are used to estimate the sensitivity and derive robust upper limits, as done in a search on O3 KAGRA data Abac et al.
-(2024), which will be discussed in Section III.2.
+Consequently, numerical simulations are used to estimate the sensitivity and derive robust upper limits, as done in a search on O3 KAGRA data KAGRA:2024ipf, which will be discussed in Section III.2.
 III.1.5 Distinguishing amongst dark-matter models
-Only a few works Miller et al.
-(2022b); Manita et al.
-(2024) have explored ways to determine whether a particular type of DM interacted with GW detectors.
-In Miller et al.
-(2022b), the authors employ the Wiener filter Wiener et al.
-(1964) to compute the cross power across different detector pairs during follow-up stages, using different $T_{\text{FFT}}$ to exploit the fact that different types of DM couple differently to different detector baselines.
+Only a few works Miller:2022wxu; Manita:2023mnc have explored ways to determine whether a particular type of DM interacted with GW detectors.
+In Miller:2022wxu, the authors employ the Wiener filter wiener1964extrapolation to compute the cross power across different detector pairs during follow-up stages, using different $T_{\text{FFT}}$ to exploit the fact that different types of DM couple differently to different detector baselines.
 In Wiener filtering, the goal is to find a waveform that, when subtracted from the data, leaves only noise.
@@
 Low residuals indicate that the candidate waveform closely matches the data, whereas high residuals indicate poor agreement.
-Miller et al.
-(2022b) showed that the Wiener filter produces small residuals when the data contain a signal consistent with the model and large residuals when the wrong signal is applied.
+Miller:2022wxu showed that the Wiener filter produces small residuals when the data contain a signal consistent with the model and large residuals when the wrong signal is applied.
 This property allowed the method to confidently veto spurious candidates in real O3 data.
 Conversely, this means that the method has the potential to confirm the detection of a certain type of DM interaction signal.
-This is in contrast to other methods and previous searches Guo et al.
-(2019); Miller et al.
-(2021a); Abbott et al.
-(2022c), which focused mostly on rejecting outliers that were shown to be compatabile with noise.
-After the work of Miller et al.
-(2022b), Ref.
-Manita et al.
-(2024) uses standard cross correlation to argue that the overlap reduction function can be used as way to distinguish amongst different spins of DM particles, since the cross power will be different.
+This is in contrast to other methods and previous searches Guo:2019ker; Miller:2020vsl; LIGOScientific:2021odm, which focused mostly on rejecting outliers that were shown to be compatabile with noise.
+After the work of Miller:2022wxu, Ref.
+Manita:2023mnc uses standard cross correlation to argue that the overlap reduction function can be used as way to distinguish amongst different spins of DM particles, since the cross power will be different.
 In particular, Ref.
-Manita et al.
-(2024) how the cross-correlation SNR would change as a function of the type of DM and which baseline was considered.
+Manita:2023mnc how the cross-correlation SNR would change as a function of the type of DM and which baseline was considered.
 In both cases, cross power is important, since the individual power spectra from one detector are indistinguishable for all types of ultralight dark matter, which arises because the phase evolutions for the considered types of ultralight dark matter here are the same.
 III.2 Ground- and space-based interferometers
-Searches for DM particles interacting with the LIGO, Virgo, KAGRA and GEO600 interferometers have been performed using data from the most recent observing run, O3, resulting in constraints on dilatons and dark photons Abbott et al.
-(2022c); Vermeulen et al.
-(2021); Göttel et al.
-(2024); Abac et al.
-(2024).
-Furthermore, LISA Pathfinder data were also used to set constraints on dark photon DM Miller and Mendes (2023); Frerick et al.
-(2024).
+Searches for DM particles interacting with the LIGO, Virgo, KAGRA and GEO600 interferometers have been performed using data from the most recent observing run, O3, resulting in constraints on dilatons and dark photons LIGOScientific:2021odm; Vermeulen:2021epa; Gottel:2024cfj; KAGRA:2024ipf.
+Furthermore, LISA Pathfinder data were also used to set constraints on dark photon DM Miller:2023kkd; Frerick:2023xnf.
 The theoretical framework for each particle is described in Section II.3 and Section II.4, respectively.
@@
 In Ref.
-Abbott et al.
-(2022c), two complementary methods were employed to search for spin-1 ultralight dark matter that can couple to baryons in the test masses of LIGO in O3 data – cross-correlation and excess power, both of which are described in Section III.1.1 and Section III.1.2, respectively.
-Cross-correlation benefits from utilizing the phase information of the signal, something that is lost when using the time-frequency excess power method.
-On the other hand, the excess power method matches its coherence time $T_{\text{FFT}}$ to the signal coherence time $T_{\text{coh}}$ in every one-Hz band analyzed, allowing an improved sensitivity across the frequency range compared to cross-corelation, which fixed $T_{\text{FFT}}=1800$ s.
-Both searches returned no significant DM candidate signals; thus, 95% confidence-levle upper limits were set on the coupling strength $\epsilon^{2}$ of dark photons to baryons for both methods, as shown in Fig.
+LIGOScientific:2021odm, two complementary methods were employed to search for spin-1 ultralight dark matter that can couple to baryons in the test masses of LIGO in O3 data – cross-correlation and BSD excess power, both of which are described in Section III.1.1 and Section III.1.2, respectively.
+Cross-correlation benefits from utilizing the phase information of the signal, something that is lost when using the time-frequency BSD excess power method.
+On the other hand, the BSD excess power method matches its coherence time $T_{\text{FFT}}$ to the signal coherence time $T_{\text{coh}}$ in every one-Hz band analyzed, allowing an improved sensitivity across the frequency range compared to cross-corelation, which fixed $T_{\text{FFT}}=1800$ s.
+Both searches returned no significant DM candidate signals; thus, 95% confidence-level upper limits were set on the coupling strength $\epsilon^{2}$ of dark photons to baryons for both methods, as shown in Fig.
 4..
 We can see, across a broad frequency range, that results from GW interferometers surpass by orders of magnitude constraints from Eöt-Wash and the MICROSCOPE experiments.
-These upper limits were derived for cross-correlation and excess power following the procedure outlined in Section III.1.1 and Section III.1.2, respectively.
-Both methods employed the Feldman-Cousins approach Feldman and Cousins (1998) to set upper limits, which is robust against non-Gaussian noise and has been shown to produce conservative upper limits with respect to those that would have been obtained through simulations Miller et al.
-(2021a).
-Figure 4: Adapted from Abbott et al.
-(2022c).
+These upper limits were derived for cross-correlation and BSD excess power following the procedure outlined in Section III.1.1 and Section III.1.2, respectively.
+Both methods employed the Feldman-Cousins approach Feldman:1997qc to set upper limits, which is robust against non-Gaussian noise and has been shown to produce conservative upper limits with respect to those that would have been obtained through simulations Miller:2020vsl.
+Figure 4: Adapted from LIGOScientific:2021odm.
 Upper limits obtained from analyzing LIGO O3 data on the square of the coupling of dark photons to baryons $U(1)_{\rm B}$ in the LIGO mirrors.
 The physics behind this form of DM has been discussed in Section II.4.
-The limits derived from each of the methods discussed in Sections III.1.1 and III.1.2 are show in red (cross-correlation) and cyan (excess power), respectively.
-MICROSCOPE Bergé et al.
-(2018) and Eöt-Wash torsion balance upper limits are plotted as a comparison to the results here Schlamminger et al.
-(2008).
+The limits derived from each of the methods discussed in Sections III.1.1 and III.1.2 are show in red (cross-correlation) and cyan (BSD excess power), respectively.
+MICROSCOPE Berge:2017ovy and Eöt-Wash torsion balance upper limits are plotted as a comparison to the results here Schlamminger:2007ht.
 To produce limits on the square of the dark photon/baryon-lepton coupling, $U(1)_{\rm B-L}$ in the LIGO mirrors, these limits should be multiplied by four.
-While cross-correlation and excess power analyses were used for dark photon searches, the LPSD technique (described in Section III.1.3) was employed in Ref.
-Göttel et al.
-(2024) to search for scalar dilaton dark matter in the LIGO O3 dataset.
+While cross-correlation and BSD excess power analyses were used for dark photon searches, the LPSD technique (described in Section III.1.3) was employed in Ref.
+Gottel:2024cfj to search for scalar dilaton dark matter in the LIGO O3 dataset.
 This analysis placed strong limits on the coupling of scalar dark matter to electrons and photons.
 In Fig.
-5, we can see, in blue, the constraints that have been derived, which supersede, at least at low masses, those that were derived in a GEO600 search for the same kind of DM Vermeulen et al.
-(2021).
+5, we can see, in blue, the constraints that have been derived, which supersede, at least at low masses, those that were derived in a GEO600 search for the same kind of DM Vermeulen:2021epa.
 The enhancement comes from the improved low-frequency sensitivity of LIGO relative to GEO600, as well as incorporating the interaction of scalar DM with the mirrors, as well as the beam splitter, into the search.
@@
 These upper limits were derived by following the procedure outlined in Section III.1.3 and Section III.1.2.
-Figure 5: Taken from Göttel et al.
-(2024).
+Figure 5: Taken from Gottel:2024cfj.
 95% confidence-level upper limits on $\Lambda_{i}^{-1}$, the coupling of dilaton DM to electrons or photons, as a function of mass and frequency from LIGO O3 data, which would have cause time-dependent oscillations of the sizes and indices of refraction of the beamsplitter and the LIGO mirrors.
 The physics behind this form of DM has been discussed in Section II.3.1.a) and b) show these constraints compared to existing ones on $\Lambda_{e}$ and $\Lambda_{\gamma}$, respectively.
-The results from the LIGO O3 search are shown by the thick blue line, constraints from direct experimental searches for DM Vermeulen et al.
-(2021); Aiello et al.
-(2022); Aharony et al.
-(2021); Savalle et al.
-(2021); Kennedy et al.
-(2020); Antypas et al.
-(2019, 2021); Tretiak et al.
-(2022); Oswald et al.
-(2022); Campbell et al.
-(2021); Beloy et al.
-(2021); Zhang et al.
-(2023a); Fukusumi et al.
-(2023); Sherrill et al.
-(2023) are shown in thin grey, and constraints from searches for “fifth forces” Bergé et al.
-(2022); Hees et al.
-(2018) are depicted by the dashed red lines.
+The results from the LIGO O3 search are shown by the thick blue line, constraints from direct experimental searches for DM Vermeulen:2021epa; Aiello:2021wlp; Aharony:2019iad; Savalle:2020vgz; Kennedy:2020bac; Antypas:2019qji; Antypas:2020rtg; Tretiak:2022ndx; Oswald:2021vtc; Campbell:2020fvq; BACON:2020ubh; Zhang:2022ewz; Fukusumi:2023kqd; Sherrill:2023zah are shown in thin grey, and constraints from searches for “fifth forces” Berge:2021yye; Hees:2018fpg are depicted by the dashed red lines.
 III.2.2 KAGRA
 KAGRA offers unique opportunities to search for DM.
-In particular, the different materials used in KAGRA’s mirrors result in significantly different charge-to-mass ratios for baryon-lepton number, enhancing the sensitivity to vector DM coupling to this quantum number Michimura et al.
-(2020).
+In particular, the different materials used in KAGRA’s mirrors result in significantly different charge-to-mass ratios for baryon-lepton number, enhancing the sensitivity to vector DM coupling to this quantum number Michimura:2020vxn.
 While LIGO and Virgo currently set the strongest limits on dark photon coupling to baryon-lepton number – obtainable from the results in Fig.
-4 by multiplying the limits on baryon number coupling by four Abbott et al.
-(2022c) – this difference in mirror materials means that once KAGRA reaches comparable sensitivity levels, it will outperform LIGO in constraining such couplings Michimura et al.
-(2020).
+4 by multiplying the limits on baryon number coupling by four LIGOScientific:2021odm – this difference in mirror materials means that once KAGRA reaches comparable sensitivity levels, it will outperform LIGO in constraining such couplings Michimura:2020vxn.
 This enhancement is particularly apparent in KAGRA’s auxiliary length control channels, PRCL and MICH, which record displacements of mirrors made from both fused silica (beam splitter and power recycling mirrors) and cryogenic sapphire (test masses).
@@
 54.
-The raw outputs of PRCL and MICH correspond to time-varying displacements of the form Abac et al. (2024):
+The raw outputs of PRCL and MICH correspond to time-varying displacements of the form KAGRA:2024ipf:
 $\delta L_{\rm MICH}$ $=\delta(l_{\rm x}-l_{\rm y}),$ (76)
@@
 See Fig.
-1 of Abac et al.
-(2024) for a schematic of the interferometer.
-No significant signals were found in this analysis of KAGRA O3 data Abac et al.
-(2024), and upper limits on the coupling strength were derived following the procedure outlined in Section III.1.4.
+1 of KAGRA:2024ipf for a schematic of the interferometer.
+No significant signals were found in this analysis of KAGRA O3 data KAGRA:2024ipf, and upper limits on the coupling strength were derived following the procedure outlined in Section III.1.4.
 These limits, shown in Fig.
 6 for each channel, exhibit many narrow-band noise artifacts that limit sensitivity at low frequencies.
-Although not yet competitive with LIGO constraints, future KAGRA searches are expected to provide powerful limits on this type of DM due to the pronounced difference in baryon-lepton charge-to-mass ratios among its mirrors Michimura et al.
-(2020).
-Figure 6: Taken from Abac et al.
-(2024).
+Although not yet competitive with LIGO constraints, future KAGRA searches are expected to provide powerful limits on this type of DM due to the pronounced difference in baryon-lepton charge-to-mass ratios among its mirrors Michimura:2020vxn.
+Figure 6: Taken from KAGRA:2024ipf.
 95% confidence-level upper limit on the strength of the coupling between dark photons and baryon-lepton number using KAGRA O3 data and the method described in Section III.1.4, derived from two particular interferometer channels: MICH (blue line) and PRCL (orange line).
@@
 […]
-Such a search was preformed using LISA Pathfinder data Miller and Mendes (2023) using the excess power method described in Section III.1.2 for ultralight dark photons coupling to baryons, however, no physical constraints on the coupling strength were obtained, since the noise level was too high, the arm length was too short (only $40$ cm compared to the expected LISA arm lengths of $\mathcal{O}(10^{6})$ km), and only one arm existed.
+Such a search was preformed using LISA Pathfinder data Miller:2023kkd using the BSD excess power method described in Section III.1.2 for ultralight dark photons coupling to baryons, however, no physical constraints on the coupling strength were obtained, since the noise level was too high, the arm length was too short (only $40$ cm compared to the expected LISA arm lengths of $\mathcal{O}(10^{6})$ km), and only one arm existed.
 However, the search procedure designed will serve as a roadmap for future analyses of spaced-based detector data, since many problems, including limited sampling at low frequency, gaps, noise non-stationarities and glitches had to be handled in this analysis of data from a space-based detector.
 After this search was performed, however, it was realized that data from a different channel would produce tighter constraints, particularly in the case of baryon-lepton coupling.
-Here, the relative acceleration of the spacecraft and one of the test masses would produce a much stronger signal than that which was searched for previously because the charge-to-mass ratios of each are different Frerick et al.
-(2024).
+Here, the relative acceleration of the spacecraft and one of the test masses would produce a much stronger signal than that which was searched for previously because the charge-to-mass ratios of each are different Frerick:2023xnf.
 This echoes the logic presented in Section III.2.2 to use KAGRA to search for dark photons coupling to baryon-lepton number.
@@
 7, by using a reference acceleration power spectral density of the spacecraft with respect to one of the test masses.
-We can see that the upper limits from the relative acceleration of space-craft/ test mass are much better than that from Miller and Mendes (2023), labeled (decoherence).
+We can see that the upper limits from the relative acceleration of space-craft/ test mass are much better than that from Miller:2023kkd, labeled (decoherence).
 The constraint on B-L coupling is stronger with respect to that of baryon coupling because the test mass and spacecraft are primarily made of gold, and gold/carbon, respectively.
-Figure 7: Taken from Frerick et al.
-(2024).
+Figure 7: Taken from Frerick:2023xnf.
 Upper limits on the coupling of dark photons to baryon-lepton number considering the relative acceleration of the spacecraft and one of the test masses in LISA Pathfinder (dark blue region).
@@
 Constraints from other experiments (Eöt-Wash and MICROSCOPE) and future projections for LISA, MAGIS-100 and torision balance are shown as well.
-Details on the curves coming from STE-QUEST and Asteroids can be found in O’HARE (2020).
-Only B-L constraints are shown from Frerick et al.
-(2024) because those arising from coupling to baryons are not yet strong enough to probe a tighter constraint than that provided by MICROSCOPE.
-However, many conservative assumptions were made in Frerick et al.
-(2024), including the composition of the spacecraft and the use of a reference acceleration power spectral density, both of which could, if studied in more detail, lead to more stringent constraints for both the baryon and B-L coupling scenarios.
+Details on the curves coming from STE-QUEST and Asteroids can be found in ciaran_o_hare_2020_3932430.
+Only B-L constraints are shown from Frerick:2023xnf because those arising from coupling to baryons are not yet strong enough to probe a tighter constraint than that provided by MICROSCOPE.
+However, many conservative assumptions were made in Frerick:2023xnf, including the composition of the spacecraft and the use of a reference acceleration power spectral density, both of which could, if studied in more detail, lead to more stringent constraints for both the baryon and B-L coupling scenarios.
 III.3 Pulsar timing arrays
@@
 Nevertheless, these effects can be mitigated by cross-correlating data from many pulsars in the sky to form a pulsar timing array.
-As described in Section II.3, ultralight scalar DM can also induce variations in pulse ToAs Khmelnitsky and Rubakov (2014), though the relevant mass range lies many orders of magnitude below that probed by ground-based interferometers.
-Such scalar fields can cause coherent oscillations in the gravitational potential or effective constants of nature, and may also generate stochastic perturbations in the metric Kim and Mitridate (2024), which could lead to further timing variations over long baselines Kim (2023).
+As described in Section II.3, ultralight scalar DM can also induce variations in pulse ToAs Khmelnitsky:2013lxt, though the relevant mass range lies many orders of magnitude below that probed by ground-based interferometers.
+Such scalar fields can cause coherent oscillations in the gravitational potential or effective constants of nature, and may also generate stochastic perturbations in the metric Kim:2023kyy, which could lead to further timing variations over long baselines Kim:2023pkx.
 Throughout this section, we will see that, independently of the physical mechanism of coupling of DM to gravity or the standard model, the shapes of the constraints on the coupling constant appear similar.
@@
 In both correlated cases, $\hat{\phi}_{P}^{2}=\hat{\phi}_{E}^{2}$.
-Figure 8: Taken from Smarra et al.
-(2023).
+Figure 8: Taken from EuropeanPulsarTimingArray:2023egv.
 95% confidence-level upper limits from analyses of European pulsar timing array data on ultralight dark matter interacting only gravitationally and changing the ToAs of pulses in an analogous way that GWs do.
@@
 $\rho_{\text{DM}}=0.4\text{GeV}/\text{cm}^{3}$.
-For comparison, upper limits on these quantities are shown from previous searches (PPTA Bayes 2018, NANOGrav 2018) Porayko and Postnov (2014); Porayko et al.
-(2018).
+For comparison, upper limits on these quantities are shown from previous searches (PPTA Bayes 2018, NANOGrav 2018) Porayko:2014rfa; Porayko:2018sfa.
 Furthermore, the dashed line indicates the signal amplitude assuming that this form of DM comprises all of DM; thus, curves that lie above this dashed line do not represent physical constraints.
@@
 The right panel shows that scalar ultralight dark-matter particles with mass $24.0<\text{log}_{10}\text{~}(m_{\rm DM}/\text{eV})<-23.7$ can comprise at most $30-40\%$ of $\rho_{\text{DM}}$, while particles with masses $23.7<\text{log}_{10}\text{~}(m_{\rm DM}/\text{eV})<-23.3$ could comprise up to $\sim 70$ % of $\rho_{\text{DM}}$.
-Additionally, NANOGrav data has been be used to constrain this kind of ultralight dark matter Afzal et al.
-(2023), though the constraints are weaker than from the EPTA.
+Additionally, NANOGrav data has been be used to constrain this kind of ultralight dark matter NANOGrav:2023hvm, though the constraints are weaker than from the EPTA.
 In particular, NANOGrav cannot yet probe a physical constraint for the fraction of DM energy density that this model of DM could compose, likely because their dataset only spans 15 years, while the European pulsar timing array collaboration analyzed almost 25 years of data.
@@
 This model of DM can also be constrained with pulsar timing array data, and in particular Ref.
-Smarra et al.
-(2024) uses European pulsar timing array data to constrain both linear (FJBD-type) and quadratic (DEF-type) conformal couplings.
+Smarra:2024kvv uses European pulsar timing array data to constrain both linear (FJBD-type) and quadratic (DEF-type) conformal couplings.
 In Fig.
@@
 However, despite showing upper limits (that assume no signal), Ref.
-Smarra et al.
-(2024) notes that there is some evidence for additional signal power at two masses: $m\sim 10^{-22.7}$ eV and $m\sim 10^{-21.4}$ eV, though other physical processes could be responsible for these excesses.
+Smarra:2024kvv notes that there is some evidence for additional signal power at two masses: $m\sim 10^{-22.7}$ eV and $m\sim 10^{-21.4}$ eV, though other physical processes could be responsible for these excesses.
 It is worth noting that the effective change in the neutron-star moment of inertia in this scenario arises from the universal rescaling of the metric due to the conformal coupling, and is therefore distinct from the composition-dependent moment of inertia variations discussed in Section II.3, where scalar fields couple directly to quarks and nucleons.
 The constraints arising from the interaction of dilatons with standard-model particles in neutron stars will be discussed in the next subsection.
-Figure 9: Taken from Smarra et al.
-(2024).
-Constraints on the coupling $\log_{10}\alpha$ of conformal DM to gravity at the 95% confidence-level (solid line) for the correlated case compared to the analytical estimate obtained by using the upper limits from the European pulsar timing array search in Smarra et al.
-(2024) to estimate $\alpha$ (brown dashed lines).
+Figure 9: Taken from Smarra:2024kvv.
+Constraints on the coupling $\log_{10}\alpha$ of conformal DM to gravity at the 95% confidence-level (solid line) for the correlated case compared to the analytical estimate obtained by using the upper limits from the European pulsar timing array search in Smarra:2024kvv to estimate $\alpha$ (brown dashed lines).
 The physics behind this form of DM has been discussed in Section II.6.
@@
 Here, this model of DM would change the moment of inertia of neutron stars used in pulsar timing arrays, resulting in changes of the ToA of the pulses.
-If conformal DM does not comprise all of DM, the limits will weaken, which is shown by the dotted line, for values of $\rho$ given in Smarra et al.
-(2023).
-AP4 indicates that these limits are derived assuming the AP4 equation-of-state of neutron stars Akmal et al.
-(1998).
+If conformal DM does not comprise all of DM, the limits will weaken, which is shown by the dotted line, for values of $\rho$ given in EuropeanPulsarTimingArray:2023egv.
+AP4 indicates that these limits are derived assuming the AP4 equation-of-state of neutron stars Akmal:1998cf.
 III.3.3 Variation of fundamental constants constraints
@@
 The strongest constraints from pulsar timing arrays are on the couplings to the electron and muons.
-This is in contrast to constraints that come from atomic clock experiments, which also aim to find ultralight dark matter but are insensitive to the electron coupling, since changing the electron mass does not affect the spacing between energy levels Kaplan et al.
-(2022).
+This is in contrast to constraints that come from atomic clock experiments, which also aim to find ultralight dark matter but are insensitive to the electron coupling, since changing the electron mass does not affect the spacing between energy levels Kaplan:2022lmz.
 In terms of (the lack of) other constraints on muons, laboratory experiments do not have enough muons to harness on earth to study the coupling of DM to them, while, in contrast, neutron stars have an abundance of muons.
@@
 In other words, these constraints on ultralight dark matter coupling to standard-model particles are valid when the coupling of DM to gravity can be neglected, i.e.
-when Kaplan et al.
-(2022):
+when Kaplan:2022lmz:
 $d_{i}\gtrsim\frac{4.5\times 10^{-9}}{y_{i}}\left(\frac{10^{-23}\,\text{eV}}{m_{\rm DM}}\right)\,$, (78)
@@
 Furthermore, these constraints assume that scalar particles constitute all of DM, but if they only comprise a fraction $f$ of DM, then the constraints will be weakened by $\sqrt{f}$, because of the $\sqrt{\rho_{\text{DM}}}$ term in the amplitude of these interaction signals.
-Figure 10: Taken from Afzal et al.
-(2023).
+Figure 10: Taken from NANOGrav:2023hvm.
 Upper limits from an analysis of NANOGrav data on the coupling of ultralight dark matter to various standard-model particles: quarks, electrons, muons, photons and gluons, shown in red for the correlated case, and red dashed for the uncorrelated case.
@@
 The black dashed line, and the gray shaded region, indicate where the amplitude of the signal would be less than the amplitude of the signal arising from DM coupling only to gravity.
-Current constraints “Rb/Cs atomic clocks” (purple) are from Hees et al.
-(2016), “Al/Hg atomic clocks” (turquoise) are from Beloy et al.
-(2021), “MICROSCOPE” (teal) are from Bergé et al.
-(2018), “H/Si clock shift” (orange) are from Beloy et al.
-(2021), and “NS binary system” are from Kumar Poddar et al.
-(2019) and Dror et al.
-(2020).
+Current constraints “Rb/Cs atomic clocks” (purple) are from Hees:2016gop, “Al/Hg atomic clocks” (turquoise) are from BACON:2020ubh, “MICROSCOPE” (teal) are from Berge:2017ovy, “H/Si clock shift” (orange) are from BACON:2020ubh, and “NS binary system” are from KumarPoddar:2019ceq and Dror:2019uea.
 The shape of these constraints is similar to those in Fig.
@@
 Furthermore, these constraints assume that vector bosons constitute all of DM, but if they only comprise a fraction $f$ of DM, then the constraints will be weakened by $\sqrt{f}$.
-Figure 11: Taken from Afzal et al.
-(2023).
+Figure 11: Taken from NANOGrav:2023hvm.
 Upper limits from analyses of NANOGrav pulsar timing array data on the coupling of DM to baryon and baryon-lepton number, shown for the correlated (solid lines) and uncorrelated (dashed lines) cases.
@@
 The black dashed line and gray shaded region indicate where the amplitude of the signal would be less than that from gravitational coupling alone.
-Constraints from equivalence principle tests (“MICROSCOPE”) are shown in teal (Bergé et al., 2018), and previous limits from the PPTA Collaboration are shown in yellow (Xue et al., 2022).
+Constraints from equivalence principle tests (“MICROSCOPE”) are shown in teal (Berge:2017ovy), and previous limits from the PPTA Collaboration are shown in yellow (PPTA:2021uzb).
 The similarity in shape between the curves in this figure and those in Fig.
@@
 In particular, Ref.
-Chowdhury et al.
-(2024) showed that a vector ultralight dark matter model coupled to muons ($L_{\mu}-L_{\tau}$) can provide an even better fit to NANOGrav’s observations than a conventional stochastic GW background.
+Chowdhury:2023xvy showed that a vector ultralight dark matter model coupled to muons ($L_{\mu}-L_{\tau}$) can provide an even better fit to NANOGrav’s observations than a conventional stochastic GW background.
 In this scenario, pulsars undergo oscillations due to their muon content, leading to a spin-dependent frequency shift in the pulses.
@@
 As described in Section II.2, axion-like dark matter induces an oscillatory modulation of the polarization plane of linearly polarized light—cosmic birefringence—via its coupling to photons.
-Recent analyses of European pulsar timing array data, particularly from the Parkes Pulsar Polarization Array (PPPA), have performed the first dedicated searches for these axion-induced polarization oscillations Xue et al.
-(2024); Porayko et al.
-(2025).
+Recent analyses of European pulsar timing array data, particularly from the Parkes Pulsar Polarization Array (PPPA), have performed the first dedicated searches for these axion-induced polarization oscillations Xue:2024zjq; EPTA:2024gxu.
 By tracking polarization angles over long time baselines, these searches place upper limits on the axion-photon coupling constant, $g_{a\gamma}$, for axion masses in the ultra-light regime.
@@
 Together, they cover a broad mass range, making multi-messenger searches a powerful strategy for testing axion dark matter scenarios.
-Figure 12: Taken from Xue et al.
-(2024).
+Figure 12: Taken from Xue:2024zjq.
 Upper limits from analyses of European pulsar timing array data on the coupling of axion-like DM to photons.
@@
 Key timescales are shown with vertical dotted lines, included $T_{\text{obs}}=18$ years.
-Smoothed constraints are shown as dashed lines in cyan, yellow, green, and purple, representing limits from polarization angle (PA) variations observed in the Crab Nebula by POLARBEAR Adachi et al.
-(2024), analyses of PPTA DR1 and Crab Nebula data from QUIJOTE Castillo et al.
-(2022), measurements of CMB polarization by SPT-3G Ferguson et al.
-(2022), and studies of axion-like DM-induced washout effects during recombination from the Planck mission Fedderke et al.
-(2019), respectively.
-Additional constraints from the CAST experiment Anastassopoulos et al.
-(2017), as well as observations of SN1987A Payez et al.
-(2015) and X-ray spectral distortions in quasar H1821+643 Reynés et al.
-(2021), are shown as black dashed lines.
+Smoothed constraints are shown as dashed lines in cyan, yellow, green, and purple, representing limits from polarization angle (PA) variations observed in the Crab Nebula by POLARBEAR POLARBEAR:2024vel, analyses of PPTA DR1 and Crab Nebula data from QUIJOTE Castillo:2022zfl, measurements of CMB polarization by SPT-3G SPT-3G:2022ods, and studies of axion-like DM-induced washout effects during recombination from the Planck mission Fedderke:2019ajk, respectively.
+Additional constraints from the CAST experiment CAST:2017uph, as well as observations of SN1987A Payez:2014xsa, updated in Fiorillo:2025gnd, and X-ray spectral distortions in quasar H1821+643 Reynes:2021bpe, are shown as black dashed lines.
 The lower panel shows the Bayes factors (BFs, essentially, the preference for DM to be in the data versus noise) for each of the scenarios.
@@
 Of course, ultralight dark matter could take on any mass a priori, and ground-based GW detectors only allow us to probe a small region of this parameter space, from $\sim 10^{-14}-10^{-12}$ eV.
-Space-based GW detectors will permit being sensitive to masses a few orders of magnitude smaller, while high-frequency gravitational-wave (HFGW) detectors could see DM particles a few orders of magnitude larger Aggarwal et al.
-(2021).
+Space-based GW detectors will permit being sensitive to masses a few orders of magnitude smaller, while high-frequency gravitational-wave (HFGW) detectors could see DM particles a few orders of magnitude larger Aggarwal:2020olq.
 However, it should be noted that when placing constraints on ultralight dark matter, the stochasticity of the signal could affect the limits: if we happen to “get unlucky” and observe the signal when the phase offset reduces the signal amplitude to zero, as can be seen at the zero points in Fig.
@@
 Refs.
-Nakatsuka et al.
-(2023); Boddy et al.
-(2025) have shown that for ground- and space-based interferometers, that this effect has an $\mathcal{O}(1)$ impact on upper limits, but needs to be mitigated for pulsar timing arrays by averaging over the signals coming from many pulsars.
+Nakatsuka:2022gaf; Boddy:2025oxn have shown that for ground- and space-based interferometers, that this effect has an $\mathcal{O}(1)$ impact on upper limits, but needs to be mitigated for pulsar timing arrays by averaging over the signals coming from many pulsars.
 The prospects for future ground-based GW interferometers, such as ET and Cosmic Explorer, to observe ultralight dark matter interactions are bright.
-Within a decade or so of observation, such detectors could actually be competitive at masses between $10^{-15},10^{-14}]$ eV, thus permitting an extremely strong probe of new physics Pierce et al.
-(2018); Morisaki et al.
-(2021).
+Within a decade or so of observation, such detectors could actually be competitive at masses between $10^{-15},10^{-14}]$ eV, thus permitting an extremely strong probe of new physics Pierce:2018xmy; Morisaki:2020gui.
 Though GW interferometers can probe extremely small DM/standard model couplings, we do not have an estimate of a minimum strength of these interactions.
@@
 As mentioned in Section I, the mass of DM could span many orders of magnitude.
-In particular, macroscopic dark matter, with masses in the range $M_{\rm DM}\sim[1,10^{9}]$ kg, has been proposed as a viable candidate Hall et al.
-(2018).
-These candidates could include composite objects such as nuclear-density nuggets Itoh (1970); Witten (1984) or Q-balls Coleman (1985),
+In particular, macroscopic dark matter, with masses in the range $M_{\rm DM}\sim[1,10^{9}]$ kg, has been proposed as a viable candidate Hall:2016usm.
+These candidates could include composite objects such as nuclear-density nuggets Itoh:1970uw; Witten:1984rs or Q-balls Coleman:1985ki,
 If such objects were to transit through or near a GW interferometer, they could exert measurable gravitational forces on the test masses.
@@
 DM would have an impact parameter of $\sim 10$ km, which is of the same order as the length of the interferometer arms (4 km).
-However, gravitationally interacting DM would not be able to produce such an event rate Hall et al.
-(2018); Du et al.
-(2023), so we will need to consider interactions of DM with the standard model to enhance the event rate.
+However, gravitationally interacting DM would not be able to produce such an event rate Hall:2016usm; Du:2023dhk, so we will need to consider interactions of DM with the standard model to enhance the event rate.
 In Section IV.1, we will discuss the physics of Yukawa-like interactions that macroscopic DM can have with the standard model, which will enhance the strength of the transiting DM object.
@@
 $r$ is a length scale, $\lambda$ is the screening length, $s$ is the spin of the interaction mediator, and $M_{i,j}$ are the two masses.
-By noting that the physical size of the detector and range of the interaction force are much larger than $r_{\rm DM}$, but much smaller than the separation between DM particles $L_{\rm DM}$, the interaction can be modeled in terms of an effective geometric cross section and derived using the Born approximation Born (1926):
+By noting that the physical size of the detector and range of the interaction force are much larger than $r_{\rm DM}$, but much smaller than the separation between DM particles $L_{\rm DM}$, the interaction can be modeled in terms of an effective geometric cross section and derived using the Born approximation Born:1926oue:
 $\frac{d\sigma}{d\Omega}=\frac{M_{\rm DM}^{2}}{4\pi^{2}}\left|\int V(\vec{r})e^{i(\vec{k}_{i}-\vec{k}_{f})\cdot\vec{r}}d^{3}\vec{r}\right|^{2}=\frac{M_{\rm DM}^{2}}{4\pi^{2}}\left|\tilde{V}(\vec{k}_{i}-\vec{k}_{f})\right|^{2}$ (82)
@@
 $\sigma_{\text{DM-DM}}=16\,\pi\times\frac{G^{2}\,M_{\text{DM}}^{2}\,\delta_{\text{DM}}^{4}}{v_{0}^{4}}\times\log\left[\frac{\lambda}{r_{\text{DM}}}\right],$ (84)
-While $|\delta_{\rm SM}|<5\times 10^{-4}$ has been strongly constrained by equivalence principle experiments Adelberger et al.
-(2009), the DM coupling constant could be much larger, as the only constraints come from how DM self-interaction would have influenced structure formation Spergel and Steinhardt (2000) and collisions between galaxies Harvey et al.
-(2015).
-In particular, the lack of observed deceleration of DM in the bullet cluster collision constrained the cross-section of DM to be less than $\sim 1$ cm2/g Randall et al.
-(2008), which provides limits on the coupling constant of self-interacting DM to be:
+While $|\delta_{\rm SM}|<5\times 10^{-4}$ has been strongly constrained by equivalence principle experiments Adelberger:2009zz, the DM coupling constant could be much larger, as the only constraints come from how DM self-interaction would have influenced structure formation Spergel:1999mh and collisions between galaxies Harvey:2015hha.
+In particular, the lack of observed deceleration of DM in the bullet cluster collision constrained the cross-section of DM to be less than $\sim 1$ cm2/g Randall:2008ppe, which provides limits on the coupling constant of self-interacting DM to be:
 $|\delta_{\text{DM}}|\lesssim 5\times 10^{9}\times\left(\frac{1~\text{kg}}{M_{\text{DM}}}\right)^{1/4},$ (85)
@@
 Based on the current upper limits on $\delta_{\rm SM}$ and $\delta_{\rm DM}$, could allow an observation of a transit event in GW interferometers.
-Furthermore, $\delta_{\rm DM}\gg$ 1 would actually alleviate some problems with the cold DM scenario, including overly dense areas of dwarf galaxies found in simulations, and a prediction that the Milky Way produces more stars than are observed Boylan-Kolchin et al.
-(2011).
+Furthermore, $\delta_{\rm DM}\gg$ 1 would actually alleviate some problems with the cold DM scenario, including overly dense areas of dwarf galaxies found in simulations, and a prediction that the Milky Way produces more stars than are observed Boylan-Kolchin:2011qkt.
 IV.2 Projected constraints on Yukawa-like interactions
-In order to search for transiting macroscopic DM, it has been proposed in Hall et al. (2018) to use matched filtering. To do so, firstly we must calculate the matched filtering SNR $\rho$:
+In order to search for transiting macroscopic DM, it has been proposed in Hall:2016usm to use matched filtering. To do so, firstly we must calculate the matched filtering SNR $\rho$:
 $\rho^{2}=4\int\frac{|a(f)|^{2}}{S_{n}(f)}\,df;$ (86)
 Here, $a(f)$ is the Fourier transform of the differential acceleration $a(t)$ between the test masses, and $S_{n}(f)$ is the power spectral density of the detector.
-The differential acceleration $a(t)$ arises from the DM transiting through the interferometer test masses, and is calculated with the difference in accelerations between the input mirror and end mirror Saulson (2017) in each arm.
+The differential acceleration $a(t)$ arises from the DM transiting through the interferometer test masses, and is calculated with the difference in accelerations between the input mirror and end mirror Saulson:2017jlf in each arm.
 The form of $a(t)$ can be derived by taking the gradient of Eq.
@@
 In searches with GW data, a threshold must be set that delineates candidates that are most likely noise with those that could be of astrophysical origin.
-To estimate the number of detectable transiting DM events, an SNR threshold of $\rho>8$ is applied, which is standard in GW searches Maggiore (2008), and the detection rate as a function of DM mass, and coupling $g\equiv\delta_{\rm SM}\delta_{\rm DM}$ is computed:
+To estimate the number of detectable transiting DM events, an SNR threshold of $\rho>8$ is applied, which is standard in GW searches maggiore2008gravitational, and the detection rate as a function of DM mass, and coupling $g\equiv\delta_{\rm SM}\delta_{\rm DM}$ is computed:
 $\mathcal{R}\sim\Phi\sigma_{\rm DM-SM}A$ (95)
@@
 These results demonstrate that macroscopic DM can be detectable, depending on the parameters of the interaction.
-Figure 13: Taken from Hall et al.
-(2018).
+Figure 13: Taken from Hall:2016usm.
 Expected number of macroscopic DM transit events per year detectable by advanced LIGO at design sensitivity, as a function of DM mass, screening length $\lambda$, and coupling $g=\delta_{\rm SM}\delta_{\rm DM}$.
@@
 The threshold fixes the minimum detectable coupling accessible to a matched filtering search.
-Further work in Du et al.
-(2023) has expanded on this analysis, incorporating additional effects of DM transits, such as the Doppler effect, Shapiro delay, and Einstein delay, each of which contributes to the strain observed in the interferometer.
+Further work in Du:2023dhk has expanded on this analysis, incorporating additional effects of DM transits, such as the Doppler effect, Shapiro delay, and Einstein delay, each of which contributes to the strain observed in the interferometer.
 These effects arise from the interaction of the DM object with the interferometer’s test masses, leading to small deviations in the expected strain signal.
@@
 To make the effects measurable by GW interferometers, Ref.
-Du et al.
-(2023) confirmed that DM must couple with standard-model particles, agreeeing with the findings of Hall et al.
-(2018).
-The strength of this coupling is given in Du et al.
-(2023) by the effective parameter $\tilde{\alpha}\sim\delta_{\rm SM}\delta_{\rm DM}$, where $\tilde{\alpha}=g$ in Hall et al.
-(2018).
+Du:2023dhk confirmed that DM must couple with standard-model particles, agreeeing with the findings of Hall:2016usm.
+The strength of this coupling is given in Du:2023dhk by the effective parameter $\tilde{\alpha}\sim\delta_{\rm SM}\delta_{\rm DM}$, where $\tilde{\alpha}=g$ in Hall:2016usm.
 Thus, detecting these weak gravitational signals requires a sufficiently large value of $\tilde{\alpha}$, which can be constrained by observations of DM transits.
@@
 14 shows the projected 90% sensitivity on $\tilde{\alpha}$, which can be probed by GW detectors for macroscopic DM.
-The results, shown for two different screening lengths ($\lambda=1\,\text{m}$ and $10^{6}\,\text{m}$), are compared to existing constraints from the Bullet Cluster Spergel and Steinhardt (2000), MICROSCOPE Bergé et al.
-(2018), and neutron-star kinetic heating experiments Gresham et al.
-(2023).
+The results, shown for two different screening lengths ($\lambda=1\,\text{m}$ and $10^{6}\,\text{m}$), are compared to existing constraints from the Bullet Cluster Spergel:1999mh, MICROSCOPE Berge:2017ovy, and neutron-star kinetic heating experiments Gresham:2022biw.
 While GW detectors may not surpass these existing constraints, they provide complementary information, especially if part of the DM population interacts with the standard model via a new fifth force.
-Figure 14: Taken from Du et al.
-(2023).
+Figure 14: Taken from Du:2023dhk.
 Projected sensitivity at 90% confidence to the standard-model/DM interaction parameter $\tilde{\alpha}\sim\delta_{\rm SM}\delta_{\rm DM}$ from macroscopic DM transits through several GW detectors.
@@
 Here, $T_{\text{obs}}=1$, and $\lambda=1$ m and $\lambda=10^{6}$ m, and an SNR threshold of 2 are assumed.
-The gray line represents combined limits from Bullet Cluster observations Spergel and Steinhardt (2000); Kahlhoefer et al.
-(2014) and the MICROSCOPE experiment Bergé et al.
-(2018); Fayet (2019).
-Purple and red lines indicate bounds from neutron star kinetic heating Gresham et al.
-(2023), with and without additional short-range DM–baryon interactions, respectively.
+The gray line represents combined limits from Bullet Cluster observations Spergel:1999mh; Kahlhoefer:2013dca and the MICROSCOPE experiment Berge:2017ovy; Fayet:2018cjy.
+Purple and red lines indicate bounds from neutron star kinetic heating Gresham:2022biw, with and without additional short-range DM–baryon interactions, respectively.
 While deterministic signals might be challenging to detect due to the weak gravitational interactions, there is also potential for detecting a stochastic background arising from numerous transiting DM events.
-This stochastic background could be visible if the Yukawa interaction is sufficiently strong, as explored in Hall et al.
-(2018); Du et al.
-(2023).
+This stochastic background could be visible if the Yukawa interaction is sufficiently strong, as explored in Hall:2016usm; Du:2023dhk.
 Future work will focus on whether a stochastic signal could be distinguished from the background noise in upcoming GW observations.
@@
 Some bosons will fall in; others will scatter off, and the extent to which either of these scenarios happens depends on the mass and spin of both the black hole and the new boson.
-If the following so-called “superradiance condition”, which relates the BH angular frequency $\Omega$ to the boson angular frequency $\omega$ via the magnetic quantum number $m$, is met Brito et al.
-(2015):
+If the following so-called “superradiance condition”, which relates the BH angular frequency $\Omega$ to the boson angular frequency $\omega$ via the magnetic quantum number $m$, is met Brito:2015oca:
 $\omega<m\Omega,$ (96)
@@
 More quantitatively, solving the Klein-Gordon equation for a Kerr BH hole in the presence of a massive scalar field results in a Schrondinger-like equation that exhibits a $1/r$ potential, resulting from the gravitational interaction between the bosons and the BH.
-The solution of this equation shows that the energy states of the boson cloud are quantized analogously to those of the hydrogen atom, allowing us to describe the boson cloud as a “gravitational atom in the sky” Baumann et al.
-(2019a).
+The solution of this equation shows that the energy states of the boson cloud are quantized analogously to those of the hydrogen atom, allowing us to describe the boson cloud as a “gravitational atom in the sky” Baumann:2018vus.
 Qualitatively, superradiance will occur regardless of the spin of the boson; however, the timescales to build the cloud, as well as to deplete it, will differ.
@@
 96 is no longer satisfied.
-In particular, the timescales $\tau_{\rm inst}^{\rm(s)}$ and $\tau_{\rm inst}^{\rm(v)}$ for which this occurs scalar and vector clouds, respectively, are Isi et al.
-(2019) [^9] Note that we do not consider tensor boson clouds in this work:
+In particular, the timescales $\tau_{\rm inst}^{\rm(s)}$ and $\tau_{\rm inst}^{\rm(v)}$ for which this occurs scalar and vector clouds, respectively, are Isi:2018pzk [^9] Note that we do not consider tensor boson clouds in this work:
 $\tau_{\rm inst}^{\rm(s)}\approx 27\,{\rm days}\left(\frac{M_{\rm BH}}{10\,M_{\odot}}\right)\left(\frac{0.1}{\alpha}\right)^{9}\frac{1}{\chi_{i}}\,$, (97)
@@
 where $m_{b}$ is the mass of the boson and $M_{\rm BH}$ is the mass of the BH.
-This new boson could have a spin of 0, 1 or 2, where spin-0 could correspond to the pseudo-scalar QCD axion (or axion-like) particle Brito et al.
-(2015, 2020), spin-1 could indicate dark photons arising through kinetic mixing with the ordinary photon Siemonsen et al.
-(2023a) [^10] Note that this dark photon is not the dark photon described in Section II.4., or spin-2 tensor bosons that could arise from modified theories of gravity Brito et al.
-(2020).
+This new boson could have a spin of 0, 1 or 2, where spin-0 could correspond to the pseudo-scalar QCD axion (or axion-like) particle Brito:2015oca; Brito:2020lup, spin-1 could indicate dark photons arising through kinetic mixing with the ordinary photon Siemonsen:2022ivj [^10] Note that this dark photon is not the dark photon described in Section II.4., or spin-2 tensor bosons that could arise from modified theories of gravity Brito:2020lup.
 After the cloud builds up in an isolated BH system, GWs can be emitted in three different ways: (1) through boson-boson annihilation, (2) through boson energy-level transitions, and (3) after a “bosenova”, all of which will be discussed in Sections V.1, V.2 and V.4.
-Furthermore, the GW frequency is fixed by the boson mass, which could in principle take on any ultralight mass; thus, it is important to note that GW interferometers, from pulsar timing arrays to HFGW instruments, will allow an extensive mass range to be probed, from $10^{-22},10^{-9}]\,\mathrm{eV}$ Aggarwal et al.
-(2021); Brzeminski et al.
-(2025); Arvanitaki et al.
-(2015a); Brito et al.
-(2017); Ng et al.
-(2020).
-It has been shown, under a wide range of assumptions of BH populations, spins, redshift distributions, etc., that such systems could be detectable by CW searches in advanced LIGO, Virgo and KAGRA data and/or in future detectors, such as ET and Cosmic Explorer Arvanitaki et al.
-(2015a); Zhu et al.
-(2020); and if such sources are not seen, particular boson masses can be disfavored – this will be the subject of Section VI.
-Furthermore, both galactic BHs Palomba et al.
-(2019) and remnants of mergers of BHs outside the galaxy could also be detected, the latter of which, in the scalar case, only being visible in ET and Cosmic Explorer Arvanitaki et al.
-(2010, 2015a); Isi et al.
-(2019), which will be discussed more in Section VI.1.
-Annihilating vector boson cloud systems, however, emit stronger, though shorter, GW signals Brito et al.
-(2015); Baryakhtar et al.
-(2017); Siemonsen et al.
-(2023b); Jones et al.
-(2023, 2025), and potentially may contain electromagnetic counterparts if the vector boson kinetically mixes with the ordinary photon Siemonsen et al.
-(2023a).
+Furthermore, the GW frequency is fixed by the boson mass, which could in principle take on any ultralight mass; thus, it is important to note that GW interferometers, from pulsar timing arrays to HFGW instruments, will allow an extensive mass range to be probed, from $10^{-22},10^{-9}]\,\mathrm{eV}$ Aggarwal:2020olq; Brzeminski:2024drp; Arvanitaki:2014wva; Brito:2017zvb; Ng:2020jqd.
+It has been shown, under a wide range of assumptions of BH populations, spins, redshift distributions, etc., that such systems could be detectable by CW searches in advanced LIGO, Virgo and KAGRA data and/or in future detectors, such as ET and Cosmic Explorer Arvanitaki:2014wva; Zhu:2020tht; and if such sources are not seen, particular boson masses can be disfavored – this will be the subject of Section VI.
+Furthermore, both galactic BHs Palomba:2019vxe and remnants of mergers of BHs outside the galaxy could also be detected, the latter of which, in the scalar case, only being visible in ET and Cosmic Explorer Arvanitaki:2009fg; Arvanitaki:2014wva; Isi:2018pzk, which will be discussed more in Section VI.1.
+Annihilating vector boson cloud systems, however, emit stronger, though shorter, GW signals Brito:2015oca; Baryakhtar:2017ngi; Siemonsen:2022yyf; Jones:2023fzz; Jones:2024fpg, and potentially may contain electromagnetic counterparts if the vector boson kinetically mixes with the ordinary photon Siemonsen:2022ivj.
 In this section, we will discuss the three main ways in which GWs can be emitted from isolated BH/ boson cloud systems: annihilation, transitions and bosenovae (Sections V.1, V.2 and V.4).
@@
 Ultralight bosons are their own anti-particles, which means that they continuously annihilate (even as the cloud is growing), producing GW radiation at a frequency fixed by the boson mass and slightly redshifted by the BH mass (in the non-relativistic limit $\alpha\ll 1$, which predominately applies to the equations shown in this review).
-In the scalar boson case Abbott et al.
-(2022d):
+In the scalar boson case LIGOScientific:2021jlr:
 $f_{\text{gw}}$ $\simeq 483\,{\rm Hz}\left(\frac{m_{\text{b}}}{10^{-12}~\text{eV}}\right)$
@@
 Qualitatively, the GWs emitted are monochromatic because the mass of the boson is fixed, and all the bosons are in the same energy state.
-As alluded to before, the timescales for GW emission are much longer than the time it takes to build the cloud Isi et al.
-(2019):
+As alluded to before, the timescales for GW emission are much longer than the time it takes to build the cloud Isi:2018pzk:
 $\tau_{\rm GW}^{\rm(s)}\approx 2\times 6.5\times 10^{4}\,{\rm yr}\left(\frac{M_{\rm BH}}{10\,M_{\odot}}\right)\left(\frac{0.1}{\alpha}\right)^{15}\hskip-2.0pt\frac{1}{\chi_{i}}.$ (102)
 $\tau_{\rm GW}^{\rm(v)}\approx 2\times 1\,{\rm day}\left(\frac{M_{\rm BH}}{10\,M_{\odot}}\right)\left(\frac{0.1}{\alpha}\right)^{11}\hskip-2.0pt\frac{1}{\chi_{i}}.$ (103)
-where we have explicitly noted a factor of 2 that is missing from the original reference Isi et al. (2019), as pointed out in Siemonsen et al. (2023b).
+where we have explicitly noted a factor of 2 that is missing from the original reference Isi:2018pzk, as pointed out in Siemonsen:2022yyf.
 As the number of bosons starts to decrease, there will also be a classical depletion of the cloud over time.
-This reduction in bosons causes the cloud to contract, resulting in the following positive drifts in frequency over time Isi et al.
-(2019) [^11] We note that relativistic computations of the frequency drift have also been performed Siemonsen et al.
-(2023b); May et al.
-(2025), which are useful for real searches of GWs arising from boson clouds but not less important for our discussions here.:
+This reduction in bosons causes the cloud to contract, resulting in the following positive drifts in frequency over time Isi:2018pzk [^11] We note that relativistic computations of the frequency drift have also been performed Siemonsen:2022yyf; May:2024npn, which are useful for real searches of GWs arising from boson clouds but not less important for our discussions here.:
 $\dot{f}_{\rm gw}^{\rm(s)}\approx 3\times 10^{-14}\,{\rm Hz/s}\left(\frac{10M_{\odot}}{M_{\rm BH}}\right)^{2}\left(\frac{\alpha}{0.1}\right)^{19}\chi_{i}^{2}\,$, (104)
 $\dot{f}_{\rm gw}^{\rm(v)}\approx 1\times 10^{-6}\,{\rm Hz/s}\left(\frac{10M_{\odot}}{M_{\rm BH}}\right)^{2}\left(\frac{\alpha}{0.1}\right)^{15}\chi_{i}^{2}\,.$ (105)
-In the non-relativistic limit $\alpha\ll 1$, the signals will have the corresponding amplitudes of Isi et al. (2019):
+In the non-relativistic limit $\alpha\ll 1$, the signals will have the corresponding amplitudes of Isi:2018pzk:
 $h_{0}^{\rm(s)}\approx 8\times 10^{-28}\left(\frac{M_{\rm BH}}{10M_{\odot}}\right)\left(\frac{\alpha}{0.1}\right)^{7}\left(\frac{\rm Mpc}{r}\right)\left(\frac{\chi-\chi_{f}}{0.1}\right)$ (106)
@@
 These transitions between specific energy levels are expected to generate detectable gravitational wave signals that could last for durations much longer than the typical observation times of ground-based gravitational wave interferometers.
-The frequency of these signals is given by Arvanitaki et al.
-(2015a):
+The frequency of these signals is given by Arvanitaki:2014wva:
 $f_{\rm gw}\sim 15\,\text{Hz}\left(\frac{m_{b}}{10^{-11}\,\text{eV}}\right)$ (109)
-The frequency of the signal would drift over time according to the duration of the signal and the axion decay constant $f_{a}$ Arvanitaki et al. (2015a):
+The frequency of the signal would drift over time according to the duration of the signal and the axion decay constant $f_{a}$ Arvanitaki:2014wva:
 $\frac{df}{dt}\simeq 10^{-11}\,\frac{\text{Hz}}{\text{s}}\left(\frac{f_{\rm gw}}{90\,\text{Hz}}\right)\left(\frac{M_{\rm BH}}{10\,M_{\odot}}\right)\left(\frac{10^{17}\,\text{GeV}}{f_{a}}\right)^{2}\left(\frac{5\,\text{yr}}{T_{\rm obs}}\right)^{2}$ (110)
@@
 104 and 105, self-interactions were assumed to be zero.
-While the expected event rate of these transitions in ground-based detectors is unlikely to exceed one event in around 3.5 years of observation with advanced LIGO Arvanitaki et al.
-(2015a), the presence of self-interactions could potentially enhance the transition rates, making these signals detectable Collaviti et al.
-(2025).
-In the vector boson case, transitions can also occur, and interestingly, overtone modes may grow faster than the fundamental modes, particularly for small angular numbers, fine-structure constants, and black hole spin values Siemonsen and East (2020).
+While the expected event rate of these transitions in ground-based detectors is unlikely to exceed one event in around 3.5 years of observation with advanced LIGO Arvanitaki:2014wva, the presence of self-interactions could potentially enhance the transition rates, making these signals detectable Collaviti:2024mvh.
+In the vector boson case, transitions can also occur, and interestingly, overtone modes may grow faster than the fundamental modes, particularly for small angular numbers, fine-structure constants, and black hole spin values Siemonsen:2019ebd.
 These higher overtones could saturate more quickly than the fundamental modes, allowing transitions to lower states and increasing the overall energy of the cloud, potentially leading to detectable gravitational wave emission.
@@
 This could result in stronger GW emission once the cloud reaches its maximum energy state.
-The timescale for superradiance saturation is sensitive to the coupling strength between bosons, and self-interactions may allow the cloud to reach a critical size faster, enhancing the CW signal emitted during this phase Yoshino and Kodama (2012, 2015a, 2015b).
+The timescale for superradiance saturation is sensitive to the coupling strength between bosons, and self-interactions may allow the cloud to reach a critical size faster, enhancing the CW signal emitted during this phase Yoshino:2012kn; Yoshino:2014wwa; Yoshino:2015nsa.
   2. 2.
@@
 This decay modifies the timescale of the inspiral phase and the subsequent merger, altering the frequency evolution of the emitted GWs.
-A faster decay due to self-interactions could lead to a sharper rise in the frequency of the GW signal Boudon et al.
-(2024); Aurrekoetxea et al.
-(2024); Takahashi et al.
-(2024).
+A faster decay due to self-interactions could lead to a sharper rise in the frequency of the GW signal Boudon:2023vzl; Aurrekoetxea:2024cqd; Takahashi:2024fyq.
   3. 3.
@@
 As a result, the rate of boson annihilation can be significantly enhanced.
-This increased annihilation rate leads to stronger and potentially more detectable GW signals from the cloud Baryakhtar et al.
-(2021).
+This increased annihilation rate leads to stronger and potentially more detectable GW signals from the cloud Baryakhtar:2020gao.
   4. 4.
@@
 Self-interactions can facilitate the transition of bosons between higher energy states more rapidly, increasing the amplitude of the emitted gravitational waves.
-This could make energy-level transitions a more significant source of detectable GW signals, especially if the cloud is in a high-energy state Baryakhtar et al.
-(2021).
-These effects become particularly relevant when considering the potential detection of boson clouds in future GW observatories, such as Cosmic Explorer and ET Collaviti et al.
-(2025).
+This could make energy-level transitions a more significant source of detectable GW signals, especially if the cloud is in a high-energy state Baryakhtar:2020gao.
+These effects become particularly relevant when considering the potential detection of boson clouds in future GW observatories, such as Cosmic Explorer and ET Collaviti:2024mvh.
 Recent searches for these signals have not accounted for these interactions, and their inclusion may shift current detection prospects by enhancing the signal strength or changing the emission characteristics.
@@
 Note that while spin-0 and spin-1 boson clouds are relatively well understood theoretically, spin-2 clouds are much more complicated.
-The superradiant instability of spin-2 fields occurs much faster than the others East and Siemonsen (2023), and since the backreaction of the instability has not been computed due to difficulties arising from a nonlinear coupling between the spin-2 particle and gravity, it is not yet known whether the observational signatures are similar to the vector or scalar cases, and whether there is a dependency on the specific nonlinear theory chosen.
-Such ultralight bosons could also be used to probe the quantum nature of BHs Mitra et al.
-(2024).
+The superradiant instability of spin-2 fields occurs much faster than the others East:2023nsk, and since the backreaction of the instability has not been computed due to difficulties arising from a nonlinear coupling between the spin-2 particle and gravity, it is not yet known whether the observational signatures are similar to the vector or scalar cases, and whether there is a dependency on the specific nonlinear theory chosen.
+Such ultralight bosons could also be used to probe the quantum nature of BHs Mitra:2023sny.
 V.4 Bosenova signal
-If the self-interactions between bosons become stronger than the gravitational binding energy of the cloud, the cloud begins to destabilize and cannot be described analogously to the hydrogen atom Fukuda and Nakayama (2020); Collaviti et al.
-(2025); Omiya et al.
-(2024); Jones et al.
-(2025).
+If the self-interactions between bosons become stronger than the gravitational binding energy of the cloud, the cloud begins to destabilize and cannot be described analogously to the hydrogen atom Fukuda:2019ewf; Collaviti:2024mvh; Omiya:2024xlz; Jones:2024fpg.
 The cloud may undergo a significant reconfiguration, leading to a dramatic release of energy, a process referred to as a “bosenova.” This occurs when the potential energy of the boson cloud becomes comparable to or exceeds the self-interaction energy, leading to the rapid rearrangement of the bosonic matter.
-The phenomenon of a bosenova has been observed in condensed matter systems, in which a Bose-Einstein condensate undergoes a similar process due to a change in interactions induced by an external magnetic field Donley et al.
-(2001).
-However, the occurrence of a bosenova in the context of ultralight bosons around black holes remains highly uncertain, with studies suggesting that self-interactions may limit bosenovas to a restricted region of the allowed parameter space, particularly for certain combinations of black hole and boson masses Baryakhtar et al.
-(2021).
-In particular, in the axion case, originally proposed in Arvanitaki and Dubovsky (2011), bosenovas would likely not occur.
-While early papers assumed that the boson cloud reached large enough amplitudes to produce GW emission Yoshino and Kodama (2012, 2015a, 2015b), subsequent studies in the non-relativistic Baryakhtar et al.
-(2021) and relativistic Omiya et al.
-(2023) regimes confirmed that dissipative nonlinear effects would prevent large-enough cloud amplitudes from occurring to cause a bosenova.
-In the vector case, bosenovas could actually occur if strings are formed the Higgs-Abelian mode East (2022); East and Huang (2022); Brzeminski et al.
-(2025).
+The phenomenon of a bosenova has been observed in condensed matter systems, in which a Bose-Einstein condensate undergoes a similar process due to a change in interactions induced by an external magnetic field donley2001dynamics.
+However, the occurrence of a bosenova in the context of ultralight bosons around black holes remains highly uncertain, with studies suggesting that self-interactions may limit bosenovas to a restricted region of the allowed parameter space, particularly for certain combinations of black hole and boson masses Baryakhtar:2020gao.
+In particular, in the axion case, originally proposed in Arvanitaki:2010sy, bosenovas would likely not occur.
+While early papers assumed that the boson cloud reached large enough amplitudes to produce GW emission Yoshino:2012kn; Yoshino:2014wwa; Yoshino:2015nsa, subsequent studies in the non-relativistic Baryakhtar:2020gao and relativistic Omiya:2022gwu regimes confirmed that dissipative nonlinear effects would prevent large-enough cloud amplitudes from occurring to cause a bosenova.
+In the vector case, bosenovas could actually occur if strings are formed the Higgs-Abelian mode East:2022ppo; East:2022rsi; Brzeminski:2024drp.
 Though GWs from boson clouds described above may seem independent, they could in principle be happening at the same time.
-In particular, if ultralight scalar or vector bosons have some self-interactions, and If these clouds are composed of string axions (as opposed to the QCD axion), a non-linear self-interaction may cause a bosenova if, during superradiance, the value of the scalar field approaches the axion decay constant $f_{a}$, which could be below the grand unification theory (GUT) scale Arvanitaki et al.
-(2010); Yoshino and Kodama (2012, 2015a, 2015b).
+In particular, if ultralight scalar or vector bosons have some self-interactions, and If these clouds are composed of string axions (as opposed to the QCD axion), a non-linear self-interaction may cause a bosenova if, during superradiance, the value of the scalar field approaches the axion decay constant $f_{a}$, which could be below the grand unification theory (GUT) scale Arvanitaki:2009fg; Yoshino:2012kn; Yoshino:2014wwa; Yoshino:2015nsa.
 This bosenova would halt the superradiance process, as about 5% of the boson cloud mass would fall back into the BH, but superradiance would immediately start again until the next bosenova is triggered.
 Such periodic bosenovas could sustain superradiance for much longer durations than those in the conventional scenario, e.g.
-in Arvanitaki and Dubovsky (2011), permitting annihilating clouds even around extremely old BHs, and even affect GWs from binary black hole systems Xie and Huang (2024) .
+in Arvanitaki:2010sy, permitting annihilating clouds even around extremely old BHs, and even affect GWs from binary black hole systems Xie:2022uvp .
 V.5 Clouds in binary black hole systems
-If boson clouds form around binary black hole systems, they should imprint some signature in the GWs arising from their inspiral Baumann et al.
-(2019a).
-In particular, the GW signal will be modified due to tidal disruptions induced by the companion, the multiple moments of the cloud Baumann et al.
-(2019a, b, 2020) or through a dynamical friction drag force induced by the boson clouds.
+If boson clouds form around binary black hole systems, they should imprint some signature in the GWs arising from their inspiral Baumann:2018vus.
+In particular, the GW signal will be modified due to tidal disruptions induced by the companion, the multiple moments of the cloud Baumann:2018vus; Baumann:2019eav; Baumann:2019ztm or through a dynamical friction drag force induced by the boson clouds.
 These kinds of signals could be visible in future space-based detectors, since the deviations primarily occur in the early inspiral stages of the system, i.e.
-at frequencies much lower than currently accessible by ground-based GW interferometers Guo et al.
-(2024a).
-In both cases, observations of the inspiral of binary systems, and in particular EMRI systems Zhang and Yang (2019); Brito and Shah (2023), could indicate the presence of boson clouds around one or both of the objects if there is some dephasing with respect to the waveform in vacuum.
+at frequencies much lower than currently accessible by ground-based GW interferometers Guo:2023lbv.
+In both cases, observations of the inspiral of binary systems, and in particular EMRI systems Zhang:2018kib; Brito:2023pyl, could indicate the presence of boson clouds around one or both of the objects if there is some dephasing with respect to the waveform in vacuum.
 Additionally, once the binary reaches a certain stage of its evolution, the impact of both the spin-induced quadrupole moment and tidal forces will lessen, since the clouds will become disrupted by their companions.
-If we can observe variations in both the spin-induced quadrupole moment and the tidal deformability over the course of the lifetime of the binary, it would indicate the presence of a (disrupted) boson cloud Baumann et al.
-(2019a).
-More recently, it has been shown that as the orbital separation in a binary black hole system decreases and becomes comparable to the size of the boson cloud, the boson cloud could be “ionized” Takahashi et al.
-(2022); Baumann et al.
-(2022a, b).
+If we can observe variations in both the spin-induced quadrupole moment and the tidal deformability over the course of the lifetime of the binary, it would indicate the presence of a (disrupted) boson cloud Baumann:2018vus.
+More recently, it has been shown that as the orbital separation in a binary black hole system decreases and becomes comparable to the size of the boson cloud, the boson cloud could be “ionized” Takahashi:2021yhy; Baumann:2021fkf; Baumann:2022pkl.
 This ionization occurs when the energy from the binary black hole system becomes large enough to unbind the cloud from its host BH.
@@
 Ref.
-Krishnendu et al.
-(2017) and Ref.
-Rahman and Bhattacharyya (2023) showed that $\kappa$ can be used as a distinguishing parameter for different kinds of objects, and also a test of the no-hair theorem: as an example, $\kappa=1$ for BHs Hansen (1974), while $\kappa\sim[1.4,8]$ for neutron stars depending on the equation-of-state Laarakkers and Poisson (1999); Pappas and Apostolatos (2012).
+Krishnendu:2017shb and Ref.
+Rahman:2021eay showed that $\kappa$ can be used as a distinguishing parameter for different kinds of objects, and also a test of the no-hair theorem: as an example, $\kappa=1$ for BHs Hansen:1974zz, while $\kappa\sim[1.4,8]$ for neutron stars depending on the equation-of-state Laarakkers:1997hb; Pappas:2012ns.
 In the case of boson cloud systems, $\kappa$ will take on vastly different values, of $\mathcal{O}(1000)$, depending on the boson cloud and BH masses, and the extent to which the Kerr metric is altered in the presence of a cloud with enough mass compared to the BH.
 V.5.2 Tidal disruptions
-Because the boson cloud is much less compact than the BH, we expect that tidal disruptions from the companion could occur when the two objects in the binary black hole are close enough to each other Ng et al.
-(2020).
+Because the boson cloud is much less compact than the BH, we expect that tidal disruptions from the companion could occur when the two objects in the binary black hole are close enough to each other Ng:2020jqd.
 To characterize the effects of tides in a binary system, the tidal deformability parameter $\Lambda$ is used, which depends on the mass and radius of the compact objects.
-For BHs in vacuum, this quantity is 0, while for neutron stars, it can take on values of $\mathcal{O}(100-1000)$ Chatziioannou (2020).
-For boson clouds, $\Lambda\sim 10^{7}$, which is a remarkable departure from that expected from conventional astrophysical binaries Baumann et al.
-(2019b).
-In Payne et al.
-(2022), the authors considered the impact that superradiance would have on hierarchical BH mergers, and showed that the reduction of spin in BHs implies smaller recoil velocities of binary black hole mergers.
+For BHs in vacuum, this quantity is 0, while for neutron stars, it can take on values of $\mathcal{O}(100-1000)$ Chatziioannou:2020pqz.
+For boson clouds, $\Lambda\sim 10^{7}$, which is a remarkable departure from that expected from conventional astrophysical binaries Baumann:2019eav.
+In Payne:2021ahy, the authors considered the impact that superradiance would have on hierarchical BH mergers, and showed that the reduction of spin in BHs implies smaller recoil velocities of binary black hole mergers.
 Smaller recoil velocities imply that remnants could remain inside clusters, leading to higher chances of hierarchical formation and merger of binary black holes.
@@
 This condition is analogous to a resonance in atomic physics, where the energy difference between two states is exactly matched by an external driving frequency.
-In this case, the system stays in equilibrium, and energy is transferred between energy levels that are in resonance Baumann et al.
-(2019a).
+In this case, the system stays in equilibrium, and energy is transferred between energy levels that are in resonance Baumann:2018vus.
   2. 2.
     Bohr Resonance: In addition to hyperfine resonance, the growing mode (the mode that is being excited by the inspiral) mixes with decaying modes of the cloud.
-This interaction allows energy to be transferred between different states, influencing the GW signal Baumann et al.
-(2019a).
+This interaction allows energy to be transferred between different states, influencing the GW signal Baumann:2018vus.
   3. 3.
@@
 This transition process, characterized by the Landau-Zener mechanism, significantly alters the frequency and amplitude evolution of the GW signal.
-These non-adiabatic transitions can provide observable changes in the signal, offering additional insights into the properties of the boson cloud Baumann et al.
-(2020).
-These effects, all of which could be present simultaneously, would provide “smoking-gun” signatures of the presence of boson clouds around rotating black holes in a binary system Baumann et al.
-(2019a).
+These non-adiabatic transitions can provide observable changes in the signal, offering additional insights into the properties of the boson cloud Baumann:2019ztm.
+These effects, all of which could be present simultaneously, would provide “smoking-gun” signatures of the presence of boson clouds around rotating black holes in a binary system Baumann:2018vus.
 V.5.4 Ionization
-When the binary system separation becomes comparable to the size of the cloud, an “ionization” of the boson cloud occurs as the cloud becomes unbounded from its host BH Takahashi et al.
-(2022); Baumann et al.
-(2022a, b).
+When the binary system separation becomes comparable to the size of the cloud, an “ionization” of the boson cloud occurs as the cloud becomes unbounded from its host BH Takahashi:2021yhy; Baumann:2021fkf; Baumann:2022pkl.
 The energy from the binary drives this effect, and is significantly larger than that emitted via GWs; thus, the inspiral of the binary becomes dominated by the boson cloud, instead of simply perturbed by it.
-Ionization will also tend to circularize the binary if it formed through dynamical capture, while leaving any orbital inclinations unaffected Tomaselli et al.
-(2023).
+Ionization will also tend to circularize the binary if it formed through dynamical capture, while leaving any orbital inclinations unaffected Tomaselli:2023ysb.
 It has been recently found, however, that as the resonances in the boson cloud evolve during the orbit, the boson cloud is unperturbed and ordinarily visible in LISA if the cloud and binary are counter rotating; otherwise, the cloud is destroyed because of the resonances, but a distinct mark is left on the binary.
-That is, the binary is forced to co-rotate with the boson cloud, and its eccentricity is driven to a particular value, which allows the possibility of doing statistics with detected binary black hole systems in LISA to probe the boson cloud hypothesis Tomaselli et al.
-(2024).
+That is, the binary is forced to co-rotate with the boson cloud, and its eccentricity is driven to a particular value, which allows the possibility of doing statistics with detected binary black hole systems in LISA to probe the boson cloud hypothesis Tomaselli:2024dbw.
 Taken together, the various phenomena discussed above—from level transitions and resonances to the dramatic ionization of the cloud—highlight the rich interplay between boson clouds and binary black holes.
 Each mechanism offers a distinct pathway by which the presence of a boson cloud can alter binary evolution and leave characteristic imprints on the emitted GWs.
-These theoretical insights provide the foundation for future searches with space-based GW data Bertone et al.
-(2020); Bertone (2024).
+These theoretical insights provide the foundation for future searches with space-based GW data Bertone:2019irm; Bertone:2024rxe.
 In the following section, however, we turn to current ground-based GW interferometer data, which so far constrain only annihilation signals from boson clouds, as discussed in Section V.1.
@@
 Different GW interferometers probe vastly different mass regimes of ultralight dark matter boson clouds.
-While ground-based detectors are sensitive to annihilation signals from bosons of masses between $\sim 10^{-14}-10^{-12}$ eV Arvanitaki et al.
-(2015a), space-based detectors can probe a few orders of magnitude lower than that, $\sim 10^{-20}-10^{-16}$ eV Baumann et al.
-(2019a), and HFGW detectors could see systems with masses around $\sim 10^{-9}$ eV Aggarwal et al.
-(2021).
+While ground-based detectors are sensitive to annihilation signals from bosons of masses between $\sim 10^{-14}-10^{-12}$ eV Arvanitaki:2014wva, space-based detectors can probe a few orders of magnitude lower than that, $\sim 10^{-20}-10^{-16}$ eV Baumann:2018vus, and HFGW detectors could see systems with masses around $\sim 10^{-9}$ eV Aggarwal:2020olq.
 At the moment, only ground-based GW interferometers exist, and so we focus on methods and search results from the most recent LIGO, Virgo and KAGRA observing runs.
@@
   1. 1.
-    Known galactic BHs (e.g. Cygnus X-1 Fabian et al. (1989))
+    Known galactic BHs (e.g. Cygnus X-1 Fabian:1989ej)
   2. 2.
-    Known remnants of binary black hole mergers (e.g. GW150914 Abbott et al. (2016))
+    Known remnants of binary black hole mergers (e.g. GW150914 Abbott:2016blz)
   3. 3.
-    Known binary black hole inspirals and mergers Abac et al. (2025b)
+    Known binary black hole inspirals and mergers LIGOScientific:2025hdt
   4. 4.
@@
 On the contrary, remnants of BH mergers are extremely young: their clouds build up quickly and they could begin emitting days or weeks after formation, as described in Section V.
-Their age is thus precisely measured, as well as the final mass and spin (within $10\%$ in the case of GW150914 Abbott et al.
-(2016)).
+Their age is thus precisely measured, as well as the final mass and spin (within $10\%$ in the case of GW150914 Abbott:2016blz).
 These estimates can be even more precise than those that arise from some x-ray binaries.
@@
 Thus, the scaling of GW amplitude with distance implies that we could not detect remnants of mergers with annihilating scalar boson clouds around them.
-We note, however, that rapidly annihilating vector boson clouds could be detected out to Gpc Jones et al.
-(2023, 2025) in current-generation GW interferometers, making merger remnants a viable target even now.
+We note, however, that rapidly annihilating vector boson clouds could be detected out to Gpc Jones:2023fzz; Jones:2024fpg in current-generation GW interferometers, making merger remnants a viable target even now.
 On the other hand, with known binary black hole inspirals, boson clouds could alter the GW signal in a variety of ways, as detailed in Section V.5.
 Thus, using the spins and masses inferred from LIGO, Virgo and KAGRA measurements, the allowed boson mass can be constrained: if certain BH masses are detected with low spins, this could be evidence that boson clouds have spun down BHs.
-On the contrary, if BHs with high spins are detected, this could rule out the presence of boson clouds existing in a particular mass range Ng et al.
-(2021a).
+On the contrary, if BHs with high spins are detected, this could rule out the presence of boson clouds existing in a particular mass range Ng:2020ruv.
 Such an analysis allows us to obtain some evidence for or rule out the presence of boson clouds around rotating binary black holes, but does not necessarily permit us to claim confidently that we have detected a boson cloud.
 Finally, unknown galactic BHs could take on a wide range of masses, ages and spins depending on when and how they formed.
-The number of expected BHs in the galaxy is $\mathcal{O}(10^{9})$, which comes from considering the age of the galaxy ($\sim 10^{11}$ years) and the supernova rate ($\sim 1$ per century) Madau and Dickinson (2014).
+The number of expected BHs in the galaxy is $\mathcal{O}(10^{9})$, which comes from considering the age of the galaxy ($\sim 10^{11}$ years) and the supernova rate ($\sim 1$ per century) Madau:2014bja.
 Thus, the BHs we observe are such a small portion of the total number that exist in the local universe, and may not be representative of the full population.
 However, searches for such systems are entirely blind: we do not know any BH mass or spin parameters, nor do we know the location of such systems.
-While this may seem to be a daunting task, computationally efficient methods from the CW community have been leveraged to do these kinds of searches D’Antonio et al.
-(2018); Isi et al.
-(2019); Jones et al.
-(2023, 2025).
+While this may seem to be a daunting task, computationally efficient methods from the CW community have been leveraged to do these kinds of searches DAntonio:2018sff; Isi:2018pzk; Jones:2023fzz; Jones:2024fpg.
 We will now list the methods that have been developed to search for these sources, and then delve into the constraints that have been set using real LIGO, Virgo and KAGRA data.
@@
 Thus, a multi-pronged approach has been taken within the community to probe both known and unknown sources of GWs.
-VI.2.1 Excess power method
+VI.2.1 SD Excess power method
 One approach to searching for boson clouds around rotating BHs is the excess-power method, which models the expected signal as approximately monochromatic over the entire observation time.
 Because the source could lie anywhere on the sky, the signal is modulated by the Earth’s motion, leading to a Doppler-shifted frequency.
-To recover such a signal, the data are transformed into time–frequency maps and corrected for different sky positions, so that a truly monochromatic signal would appear as a narrow line D’Antonio et al.
-(2018).
+To recover such a signal, the data are transformed into time–frequency maps and corrected for different sky positions, so that a truly monochromatic signal would appear as a narrow line DAntonio:2018sff.
 This strategy is similar to that described in Section III.1.2, but here the modulation arises from the Doppler effect rather than from the random velocity distribution of individual DM particles.
@@
 Another method that is more robust against stochastic frequency variations relies on Hidden Markov Models (HMMs) to find tracks in the time-frequency plane.
-The main idea is to model a time-frequency GW signal probabilistically as a Markov chain of transitions between “hidden” (unobservable) frequency states and use a detection statistic to relate “observed” frequency states with the hidden ones Quinn and Hannan (2001).
+The main idea is to model a time-frequency GW signal probabilistically as a Markov chain of transitions between “hidden” (unobservable) frequency states and use a detection statistic to relate “observed” frequency states with the hidden ones quinn2001estimation.
 In the framework for this method, the probability to jump to a new state at a given time only depends upon the previous hidden state.
@@
 In Fig.
-16, taken from Isi et al.
-(2019), we show the optimal path found by the Viterbi method for injected boson cloud signals in Gaussian noise that has an associated random walk of the frequency, which is meant to simulate unknown theoretical features of the signal.
+16, taken from Isi:2018pzk, we show the optimal path found by the Viterbi method for injected boson cloud signals in Gaussian noise that has an associated random walk of the frequency, which is meant to simulate unknown theoretical features of the signal.
 We can see that in both cases, with less and more variability of the signal, the Viterbi algorithm can find the appropriate track.
-Recently, the Viterbi algorithm has also been adapted to search for GWs from annihilating vector boson clouds, which have timescales significantly shorter than in the scalar case Jones et al.
-(2023, 2025), and for which numerical waveforms must be used to gauge sensitivity Siemonsen and East (2020); Siemonsen et al.
-(2023b).
+Recently, the Viterbi algorithm has also been adapted to search for GWs from annihilating vector boson clouds, which have timescales significantly shorter than in the scalar case Jones:2023fzz; Jones:2024fpg, and for which numerical waveforms must be used to gauge sensitivity Siemonsen:2019ebd; Siemonsen:2022yyf.
 (a) Less variability
 (b) More variability
-Figure 16: Taken from Isi et al.
-(2019).
+Figure 16: Taken from Isi:2018pzk.
 This figure illustrates how the Viterbi HMM algorithm described in Section VI.2.2 works to identify tracks in the time-frequency plane that would arise from GW emission from annihilating boson clouds around rotating BHs.
@@
 We have already discussed in Section III.1.1 cross correlating detector data to search for ultralight dark matter that could interact directly with GW interferometers; however, it is important to note that such a method could also be used to search for boson clouds around rotating BHs, both deterministically and for a stochastic GW background composed of the superposition of cosmological or astrophysical GW signals.
-Furthermore, virtually any method in the CW community Riles (2023) could be tuned to look for such boson cloud systems because such methods already analyze almost monochromatic signals focusing on spin-downs but could also extend their parameter spaces to include spin-ups.
+Furthermore, virtually any method in the CW community Riles:2022wwz could be tuned to look for such boson cloud systems because such methods already analyze almost monochromatic signals focusing on spin-downs but could also extend their parameter spaces to include spin-ups.
 VI.3 Search results for scalar boson clouds
@@
 VI.3.1 All-sky search constraints
-The excess power method described in Section VI.2.1 has been used to analyze LIGO O3 data to search for the presence of annihilating scalar boson clouds around isolated BHs within our galaxy Abbott et al.
-(2022d, e).
+The BSD excess power method described in Section VI.2.1 has been used to analyze LIGO O3 data to search for the presence of annihilating scalar boson clouds around isolated BHs within our galaxy LIGOScientific:2021jlr; KAGRA:2022osp.
 This search was very computationally expensive, since each sky position had to be analyzed individually.
-No evidence for scalar boson cloud systems was found in either case, and thus upper limits were placed on the presence of boson clouds with different ages, spins, and distances away from earth in two different ways Palomba et al.
-(2019); Dergachev and Papa (2019); Abbott et al.
-(2022d):
+No evidence for scalar boson cloud systems was found in either case, and thus upper limits were placed on the presence of boson clouds with different ages, spins, and distances away from earth in two different ways Palomba:2019vxe; Dergachev:2019wqa; LIGOScientific:2021jlr:
   1. 1.
@@
 The search on LIGO O3 data can thus exclude the existence of BH/boson cloud systems that have these particular combinations of parameters at certain distances away from us.
-Figure 17: Taken from Abbott et al.
-(2022d).
+Figure 17: Taken from LIGOScientific:2021jlr.
 Using upper limits from the O3 LIGO search for GWs from annihilating boson cloud systems, exclusion region of particular boson mass /BH mass pairs are derived assuming different ages and distances (left: 1kpc; right: 15 kpc) for BHs with initial spins of $\chi_{i}=0.9$.
@@
 Constraints from the second way are shown in Fig.
-18, in which a Kroupa BH mass distribution $f(m)\propto M_{\rm BH}^{-2.3}$ Elbert et al.
-(2018) between $5,~50]M_{\odot}$ and $5,~100]M_{\odot}$ is assumed, as well as a uniform initial BH spin distribution between [0.2,0.9].
+18, in which a Kroupa BH mass distribution $f(m)\propto M_{\rm BH}^{-2.3}$ Elbert:2017sbr between $5,~50]M_{\odot}$ and $5,~100]M_{\odot}$ is assumed, as well as a uniform initial BH spin distribution between [0.2,0.9].
 For a fixed BH age, the maximum distance reach as a function of boson mass indicates that at least 5% of GWs from BH/boson cloud systems would have been detected by the search, i.e.
@@
 As expected, on average when $50M_{\odot}$ is the maximum BH mass considered, the distance that can be reached is also smaller, since the GW amplitudes scales strongly with the BH mass.
-Figure 18: Taken from Abbott et al.
-(2022d).
-Using upper limits from the O3 LIGO search for GWs from annihilating boson cloud systems, the maximum distance that the search in Abbott et al.
-(2022d) could reach at which at least 5% of a simulated BH population would have been detectable, that is, with a strain greater than the value of the upper limit, is shown here.
+Figure 18: Taken from LIGOScientific:2021jlr.
+Using upper limits from the O3 LIGO search for GWs from annihilating boson cloud systems, the maximum distance that the search in LIGOScientific:2021jlr could reach at which at least 5% of a simulated BH population would have been detectable, that is, with a strain greater than the value of the upper limit, is shown here.
 The physics behind this form of DM has been discussed in Section V.1.
 To create this plot, Ref.
-Abbott et al.
-(2022d) assumed a uniform spin distribution between [0.2,0.9], different ages, and a Kroupa mass distribution with maximum mass $50M_{\odot}$ (left) and $100M_{\odot}$ (right).
+LIGOScientific:2021jlr assumed a uniform spin distribution between [0.2,0.9], different ages, and a Kroupa mass distribution with maximum mass $50M_{\odot}$ (left) and $100M_{\odot}$ (right).
 These limits depend on the distributions assumed and are thus less agnostic than those presented in Fig.
 17.
-We note that the search constraints present in, for example, Steltner et al.
-(2023); Ming et al.
-(2024); Covas et al.
-(2025), could also be cast in terms of all-sky or directed search constraints on scalar or vector boson clouds, but with particular caveats that would have be worked out depending on the actual search parameter space.
+We note that the search constraints present in, for example, Steltner:2023cfk; Ming:2024dug; Covas:2024nzs, could also be cast in terms of all-sky or directed search constraints on scalar or vector boson clouds, but with particular caveats that would have be worked out depending on the actual search parameter space.
 VI.3.2 Cygnus X-1 constraints
 Cygnus X-1 is an x-ray binary that is close by, has a large spin, $\sim 0.9$, and has relatively well-measured orbital parameters, which make it a reasonable BH to target.
-Since the sky position is known, searches based on the Viterbi algorithm discussed in Section VI.2.2 can use much longer coherence times than those used in Abbott et al.
-(2022d), thus permitting exquisite sensitivity towards almost monochromatic signals, while also being robust against theoretical uncertainties in the GW emission from Cygnus X-1, e.g.
+Since the sky position is known, searches based on the Viterbi algorithm discussed in Section VI.2.2 can use much longer coherence times than those used in LIGOScientific:2021jlr, thus permitting exquisite sensitivity towards almost monochromatic signals, while also being robust against theoretical uncertainties in the GW emission from Cygnus X-1, e.g.
 due to uncertainties on the measured semi-major axis, etc., due to the presence of an accretion disk, etc.
@@
 The white line indicates the axion decay constant that is fixed by the axion mass.
-Constraints on self-interacting DM using the search results on O2 data in Sun et al.
-(2020) are shown only in the left-hand plot, since that search was designed to analyze 211 annihilation from Cygnus X-1.
+Constraints on self-interacting DM using the search results on O2 data in Sun:2019mqb are shown only in the left-hand plot, since that search was designed to analyze 211 annihilation from Cygnus X-1.
 Furthermore, current and future ground-based interferometers will be able to place constraints on the axion decay constant at frequencies of $\mathcal{O}(100)$ Hz (left plot), while space-based interferometers, such as DECIGO and MAGIS-Space are required to probe transition signals at frequencies below 1 Hz (right panel).
-The figure also highlights the possibility of multi-band GW astronomy: it would be possible to detect a transition from the 322 to 211 energy levels in space-based observatories while simultaneously observing the CW arising from annihilation of bosons in the 211 energy level in ground-based interferometers Collaviti et al.
-(2025).
+The figure also highlights the possibility of multi-band GW astronomy: it would be possible to detect a transition from the 322 to 211 energy levels in space-based observatories while simultaneously observing the CW arising from annihilation of bosons in the 211 energy level in ground-based interferometers Collaviti:2024mvh.
 Annihilation signals from the 322 level (middle plot) would be too weak to be detected by any of the considered current or future GW interferometers.
-Figure 19: Taken from Sun et al.
-(2020).
+Figure 19: Taken from Sun:2019mqb.
 Upper limits at 95% confidence on the strain amplitude of annihilating scalar boson clouds coming from the BH in Cygnus X-1.
@@
 The parameters of Cygnux X-1 used to compute the amplitude curves are: $M=14.8M_{\odot}$, $\chi_{i}=0.99$, and $d=1.86$ kpc.
-Figure 20: Taken from Collaviti et al.
-(2025).
+Figure 20: Taken from Collaviti:2024mvh.
 Parameter space reachable by current and next-generation GW interferometers for (left) 211 annihilation, (middle) 322 annihilation, and (right) 322 $\rightarrow$ 211 transition signals from a scalar boson cloud around Cygnus X-1, assuming an age of $10^{5}$ yr and the parameters for Cygnus X-1 given in Fig.
@@
 VI.3.3 Constraints on luminous dark photon clouds
-As alluded to in Section V, if dark photon DM kinetically mixes with the ordinary photon and forms clouds around rotating BHs, it would produce both electromagnetic and GW signals, which, if detected, could enable a major multi-messenger discovery Siemonsen et al.
-(2023a).
+As alluded to in Section V, if dark photon DM kinetically mixes with the ordinary photon and forms clouds around rotating BHs, it would produce both electromagnetic and GW signals, which, if detected, could enable a major multi-messenger discovery Siemonsen:2022ivj.
 This additional coupling implies that soon after exponential growth of the boson cloud begins, electromagnetic fields act on charged particles in the vicinity of the BH, producing electron/positron pairs essentially out of vacuum, i.e.
-through the photon-assisted Schwinger mechanism Sauter (1931); Heisenberg and Euler (1936); Schwinger (1951).
-At this point, a pair-production cascade ensues Dunne et al.
-(2009); Monin and Voloshin (2010), sourced by the electromagnetic fields, creating a plasma around the BH in a state of turbulent quasi-equilibrium.
+through the photon-assisted Schwinger mechanism Sauter:1931zz; Heisenberg:1936nmg; Schwinger:1951nm.
+At this point, a pair-production cascade ensues Dunne:2009gi; Monin:2010qj, sourced by the electromagnetic fields, creating a plasma around the BH in a state of turbulent quasi-equilibrium.
 This plasma emits copious amounts of electromagnetic radiation, sourced by energy injections through dissipative processes in the cloud (e.g.
 at magnetic reconnection sites, Landau damping, turbulence).
-The luminosity of these sources depends on the strength of the kinetic mixing parameter, but could be at the level of supernovae or known pulsars, and and also periodic Siemonsen et al.
-(2023a).
-If the dark photon has a mass around $\sim 10^{-12}$ eV, it will induce the formation of clouds around stellar-mass BHs, and give off radio emission that would look remarkably similar to that from known pulsars in the ATNF catalog Manchester et al.
-(2005).
+The luminosity of these sources depends on the strength of the kinetic mixing parameter, but could be at the level of supernovae or known pulsars, and and also periodic Siemonsen:2022ivj.
+If the dark photon has a mass around $\sim 10^{-12}$ eV, it will induce the formation of clouds around stellar-mass BHs, and give off radio emission that would look remarkably similar to that from known pulsars in the ATNF catalog Manchester:2004bp.
 The absence of observed CWs from pulsars, combined with the possibility that dark-photon clouds could act as “pulsar mimickers,” motivates searches for such clouds around sources in the ATNF catalog that might be misidentified as pulsars.
 Ref.
-Mirasola et al.
-(2025) explores this scenario by analyzing a subset of spinning-up “pulsars” in the catalog that could instead be stellar-mass BHs surrounded by luminous dark-photon clouds.
+Mirasola:2025car explores this scenario by analyzing a subset of spinning-up “pulsars” in the catalog that could instead be stellar-mass BHs surrounded by luminous dark-photon clouds.
 Three analysis strategies are considered: one assuming perfect phase coherence between the electromagnetic emission and any associated GW signal; one allowing for small frequency deviations around the measured pulsar frequency; and one that is most robust to theoretical uncertainties but least sensitive to GW emission.
@@
 Even under these assumptions, the analysis requires only $\sim 10$ detectable “events” (i.e., BHs with GW amplitudes exceeding the strain upper limit), which is conservative compared to the $\mathcal{O}(10^{3})$ events expected at certain coupling–mass values (see Fig.
-3 of Mirasola et al.
-(2025)).
+3 of Mirasola:2025car).
 Ultimately, the study excludes dark-photon masses in the range $\sim[10^{-13},10^{-12}]$ eV with kinetic mixing larger than roughly $10^{-9},10^{-7}]$.
-Figure 21: Taken from Mirasola et al.
-(2025).
-For each of the radio pulsar mimickers targeted in Mirasola et al.
-(2025), the upper limit at 95% confidence on the kinetic mixing parameter $\epsilon$ is plotted as a function of the boson mass for which at least 10 systems would have been detected with signal amplitudes above the strain upper limits.
+Figure 21: Taken from Mirasola:2025car.
+For each of the radio pulsar mimickers targeted in Mirasola:2025car, the upper limit at 95% confidence on the kinetic mixing parameter $\epsilon$ is plotted as a function of the boson mass for which at least 10 systems would have been detected with signal amplitudes above the strain upper limits.
 Each vertical line corresponds to one radio pulsar, whose frequency fixes the mass of the dark photon.
@@
 using a larger maximum mass and spin for the assumed BH population with respect to the “pessimistic” population also constrained.
-The bounds reproduced from other experiments can be found in O’Hare (2020), specifically from the conversion of CMB photons into dark photons, which would cause distortions in the black body spectrum measured by COBE/FIRAS Mirizzi et al.
-(2009); Caputo et al.
-(2020) (dark gray), and from measurements by Planck and unWISE galaxies of cross correlations between temperature anisotropies in the CMB McCarthy et al.
-(2024) (light gray).
+The bounds reproduced from other experiments can be found in AxionLimits, specifically from the conversion of CMB photons into dark photons, which would cause distortions in the black body spectrum measured by COBE/FIRAS Mirizzi:2009iz; Caputo:2020bdy (dark gray), and from measurements by Planck and unWISE galaxies of cross correlations between temperature anisotropies in the CMB McCarthy:2024ozh (light gray).
 The upper limits arising from this search surpass those from existing experiments, and indicate that between $10$ and $10^{4}$ events could have been detected at couplings of between $\sim[10^{-10},10^{-7}]$.
@@
 The superposition of GWs arising from annihilating scalar boson clouds could form a stochastic GW background detectable by current-generation GW interferometers.
-In Tsukada et al.
-(2019), the authors designed two models of the source of this stochastic GW background: (1) from remnants of core collapse supernovae, and (2) from remnants of mergers of BHs or neutron stars.
-The authors also performed search on LIGO O1 data using cross-correlation to look for evidence of a stochastic GW background from scalar boson clouds Tsukada et al.
-(2019).
+In Tsukada:2018mbp, the authors designed two models of the source of this stochastic GW background: (1) from remnants of core collapse supernovae, and (2) from remnants of mergers of BHs or neutron stars.
+The authors also performed search on LIGO O1 data using cross-correlation to look for evidence of a stochastic GW background from scalar boson clouds Tsukada:2018mbp.
 No signal was found, and no range of boson masses could be robustly excluded unfortunately, except under optimistic assumptions about BH formation rates and spin distributions, which would disfavor the regime $2\times 10^{-13}\leq m_{b}\leq 3.8\times 10^{-13}$ eV at 95% confidence.
@@
 This limitation highlighted the degeneracy between boson mass and poorly constrained astrophysical parameters, particularly BH spins.
-The search was repeated in the second observing run as well, finding no evidence for any stochastic GW background of vector boson clouds around rotating BHs Tsukada et al.
-(2021), in particular disfavoring vector boson cloud masses between $0.8,6]\times 10^{-13}$ eV at 95% confidence.
+The search was repeated in the second observing run as well, finding no evidence for any stochastic GW background of vector boson clouds around rotating BHs Tsukada:2020lgt, in particular disfavoring vector boson cloud masses between $0.8,6]\times 10^{-13}$ eV at 95% confidence.
 Ref.
-Yuan et al.
-(2022) has also disfavored bosons with masses between $1.5,17]\times 10^{-13}$ eV depending upon what is assumed for the uniform spin distribution, with wider constraints present when higher-spinning BHs dominate the priors.
+Yuan:2022bem has also disfavored bosons with masses between $1.5,17]\times 10^{-13}$ eV depending upon what is assumed for the uniform spin distribution, with wider constraints present when higher-spinning BHs dominate the priors.
 Beyond scalar and vector cases, Ref.
-Guo et al.
-(2024b) investigated tensor bosons, computing the corresponding stochastic GW background and performing a dedicated search.
+Guo:2023gfc investigated tensor bosons, computing the corresponding stochastic GW background and performing a dedicated search.
 Their analysis disfavored boson masses between $10^{-13.4}$ and $10^{-11.7}$ eV, again with the precise range depending on the assumed range of uniform spin distribution .
@@
 VI.3.5 Constraints from the spins of detected binary black holes
-We note that constraints on boson clouds can be inferred from the spin measurements of binary black hole systems Arvanitaki et al.
-(2015a, 2017); Brito et al.
-(2017); Baryakhtar et al.
-(2017); Cardoso et al.
-(2018); Ghosh and Sachdeva (2021).
-Recently, a Bayesian method has been developed by combining the information from each system Ng et al.
-(2021b), assuming a distribution for the initial BH spins and requiring that the BHs have had enough time to undergo superradiance before merging.
-Such a study has been applied to the second GW transient catalog (GWTC-2) Abbott et al.
-(2021b), and has resulted in essentially a joint posterior on the scalar boson mass and the spin distribution of BHs at their formation Ng et al.
-(2021a).
+We note that constraints on boson clouds can be inferred from the spin measurements of binary black hole systems Arvanitaki:2014wva; Arvanitaki:2016qwi; Brito:2017zvb; Baryakhtar:2017ngi; Cardoso:2018tly; Ghosh:2021zuf.
+Recently, a Bayesian method has been developed by combining the information from each system Ng:2019jsx, assuming a distribution for the initial BH spins and requiring that the BHs have had enough time to undergo superradiance before merging.
+Such a study has been applied to the second GW transient catalog (GWTC-2) LIGOScientific:2020ibl, and has resulted in essentially a joint posterior on the scalar boson mass and the spin distribution of BHs at their formation Ng:2020ruv.
 The latter is required because BHs born with high spins that have been spun down by superradiance is partially degenerate with BHs simply born with low spins in a universe without superradiance.
-The study presented in Ng et al.
-(2021a) tentatively rules out scalar bosons in the mass range $1.3,2.7]\times 10^{-13}$ eV.
-These constraints from measured BH spins can disfavor particular masses or even confirm the presence of an ultralight boson; however, they rely on assumptions about the spin distribution at formation, on whether the boson clouds have time to form (which could be impacted by eccentricity at formation Peters (1964); Wen and Gair (2005)), and are a function of the (strongest) BH mergers that we see with GWs.
+The study presented in Ng:2020ruv tentatively rules out scalar bosons in the mass range $1.3,2.7]\times 10^{-13}$ eV.
+These constraints from measured BH spins can disfavor particular masses or even confirm the presence of an ultralight boson; however, they rely on assumptions about the spin distribution at formation, on whether the boson clouds have time to form (which could be impacted by eccentricity at formation Peters:1964zz; Wen:2005xn), and are a function of the (strongest) BH mergers that we see with GWs.
 Thus, these kinds of analyses complement those previously that look for CWs around isolated, rotating BHs.
 Both have their advantages: CW analyses described above allow us to probe unknown sources and known, nearby sources whose signals we expect to be stronger than those arising from remnants of mergers.
-Likewise, the spins and masses of BHs formed in binary system are well measured, and the constraints obtained in Ng et al.
-(2021a) will only improve as we detect more and more BHs.
+Likewise, the spins and masses of BHs formed in binary system are well measured, and the constraints obtained in Ng:2020ruv will only improve as we detect more and more BHs.
 VI.4 Prospects for GW probes of boson clouds
@@
 Moreover, scalar boson clouds are by far easier to work with theoretically than vector or tensors, which has restricted the search parameter space primarily to what can be probed by CW searches.
-However, as the theoretical calculations of vector boson cloud waveforms improve Siemonsen and East (2020); Siemonsen et al.
-(2023b), new methods are being developed to search for annihilating vector boson clouds as well Jones et al.
-(2023, 2025).
+However, as the theoretical calculations of vector boson cloud waveforms improve Siemonsen:2019ebd; Siemonsen:2022yyf, new methods are being developed to search for annihilating vector boson clouds as well Jones:2023fzz; Jones:2024fpg.
 Additionally, there are also significant astrophysical uncertainties, namely: what are the distributions of BH spins, masses, distances and ages in the galaxy and beyond.
@@
 The chosen targets of GW searches should also be evaluated.
-While having accurate measurements of the mass, spin and age of remnants of binary black hole mergers would allow us to ignore astrophysical population models Arvanitaki et al.
-(2015b), these systems are currently too far to be detected if the clouds are composed of scalar bosons Isi et al.
-(2019), though could be seen up to Gpc now if they are composed of vector bosons Jones et al.
-(2023).
-On the flipside, GWs from vector boson clouds suffer from even more theoretical uncertainties than scalars Siemonsen and East (2020), and so the robustness of search results by targeting remnants would have to be evaluated.
+While having accurate measurements of the mass, spin and age of remnants of binary black hole mergers would allow us to ignore astrophysical population models Arvanitaki:2014faa, these systems are currently too far to be detected if the clouds are composed of scalar bosons Isi:2018pzk, though could be seen up to Gpc now if they are composed of vector bosons Jones:2023fzz.
+On the flipside, GWs from vector boson clouds suffer from even more theoretical uncertainties than scalars Siemonsen:2019ebd, and so the robustness of search results by targeting remnants would have to be evaluated.
 All-sky searches for such boson cloud systems could also be enhanced, especially in light of boson self-interactions that could alter the GW signal significantly as discussed in Section V.3.
-Potentially, new methods to track quickly-evolving waveforms, such as those in Jones et al.
-(2023, 2025), could be useful in a variety of contexts.
+Potentially, new methods to track quickly-evolving waveforms, such as those in Jones:2023fzz; Jones:2024fpg, could be useful in a variety of contexts.
 Furthermore, the all-sky search specifically for scalar boson cloud systems did not consider the possibility that the cloud could spin-up, which, in practice, fixed a maximum spin-up to which the search was sensitive (such that the spin-up did not induce a freqeuncy shift by more than one bin $\delta f$ in $T_{\text{obs}}$, i.e.
@@
 Ref.
-Hoof et al.
-(2024) attempts to unify the way in which constraints are set in a Bayesian way by handling the subtleties associated with population assumptions.
+Hoof:2024quk attempts to unify the way in which constraints are set in a Bayesian way by handling the subtleties associated with population assumptions.
 The authors in Ref.
-Hoof et al.
-(2024) encourage that posteriors on mass and spin, obtained through analyses of X-ray data from BHs, be made publicly available in order to incorporate into their future work.
+Hoof:2024quk encourage that posteriors on mass and spin, obtained through analyses of X-ray data from BHs, be made publicly available in order to incorporate into their future work.
 VII GW probes of soliton dark matter
-Ultralight DM in the mass range $\sim[10^{-24},10^{-22}]$ eV has a Compton wavelength comparable to the size of galactic halos Hui et al.
-(2017); Sin (1994).
-In this regime, DM can form self-gravitating structures, or solitons, at the centers of galaxies, which remain stable due to quantum degeneracy pressure Schive et al.
-(2014).
+Ultralight DM in the mass range $\sim[10^{-24},10^{-22}]$ eV has a Compton wavelength comparable to the size of galactic halos Hui:2016ltb; Sin:1992bg.
+In this regime, DM can form self-gravitating structures, or solitons, at the centers of galaxies, which remain stable due to quantum degeneracy pressure Schive:2014dra.
 A natural approach to probing this mass range of DM is through pulsar timing arrays, whose frequency sensitivity aligns with the nanohertz oscillation frequencies expected from DM.
-In this case, solitons can exert dynamical friction on binary systems, accelerating their inspiral and suppressing the gravitational wave strain amplitude at nanohertz frequencies Aghaie et al.
-(2024).
+In this case, solitons can exert dynamical friction on binary systems, accelerating their inspiral and suppressing the gravitational wave strain amplitude at nanohertz frequencies Aghaie:2023lan.
 Alternatively, soliton DM can be framed as a low-frequency modulation of relatively higher-frequency GW.
-The GW frequency is modified during its propagation through the DM halo, undergoing a gravitational redshift or ’heterodyning’, analogous to the Sachs-Wolfe effect that imprints low-frequency modulations on high-frequency photons in the cosmic microwave background Sachs and Wolfe (1967).
-Two CW sources have been considered as high-frequency carriers for this modulation Blas et al.
-(2025): (1) inspiraling white-dwarf binaries, observable by space-based GW interferometers, and (2) non-axisymmetric rotating neutron stars that spin down over time.
+The GW frequency is modified during its propagation through the DM halo, undergoing a gravitational redshift or ’heterodyning’, analogous to the Sachs-Wolfe effect that imprints low-frequency modulations on high-frequency photons in the cosmic microwave background Sachs:1967er.
+Two CW sources have been considered as high-frequency carriers for this modulation Blas:2024duy: (1) inspiraling white-dwarf binaries, observable by space-based GW interferometers, and (2) non-axisymmetric rotating neutron stars that spin down over time.
 In what follows, we explore how these different probes can be used to constrain the properties of soliton.
@@
 VII.1 Constraints from pulsar timing arrays
-The evidence for a GW background from NANOGrav results can be used to constrain soliton DM Aghaie et al.
-(2024).
+The evidence for a GW background from NANOGrav results can be used to constrain soliton DM Aghaie:2023lan.
 Essentially, ultralight dark matter would induce dynamical friction in the inspiral of supermassive BHs, speeding up the orbit and suppressing the GW strain at nanohertz frequencies.
 Using the most accurate limit at 3.92 nHz on the GW strain from supermassive BHs in the right-hand panel of Fig.
-1 of Afzal et al.
-(2023), Ref.
-Aghaie et al.
-(2024) derives constraints on the soliton DM mass-to-BH mass ratio parameter space, as shown in Fig.
+1 of NANOGrav:2023hvm, Ref.
+Aghaie:2023lan derives constraints on the soliton DM mass-to-BH mass ratio parameter space, as shown in Fig.
 22.
 These constraints are obtained using Eq.
-17 from Aghaie et al.
-(2024), which models energy loss due to GW emission in the presence of ultralight dark matter.
-This equation incorporates astrophysical inputs such as galaxy-stellar mass functions, galaxy merger rates, and galaxy pair fractions Chen et al.
-(2019).
+17 from Aghaie:2023lan, which models energy loss due to GW emission in the presence of ultralight dark matter.
+This equation incorporates astrophysical inputs such as galaxy-stellar mass functions, galaxy merger rates, and galaxy pair fractions Chen:2018znx.
 The constraints are obtained by integrating Eq.
@@
 For example, the specific values chosen for galaxy merger rates, stellar mass functions, and the fraction of galaxies with supermassive BHs affect the final upper limits.
-Figure 22: Taken from Aghaie et al.
-(2024).
+Figure 22: Taken from Aghaie:2023lan.
 Upper limits on the DM mass and soliton-to-supermassive black hole mass ratio parameter space, derived from NANOGrav data.
@@
 VII.2 Dynamical friction on white dwarf binaries
-Space-based GW detectors will observe the quasi-infinite inspiral of galactic white dwarf binaries and extreme-mass ratio inspirals of ordinary compact objects with much heavier BHs Bartolo et al.
-(2016); Babak et al.
-(2017).
-The long-durations of such signals permit the possibility of detecting a low-frequency modulation when passing through the ultralight dark matter soliton Wang and Zhong (2023); Brax et al.
-(2024); Blas et al.
-(2025).
+Space-based GW detectors will observe the quasi-infinite inspiral of galactic white dwarf binaries and extreme-mass ratio inspirals of ordinary compact objects with much heavier BHs Bartolo:2016ami; Babak:2017tow.
+The long-durations of such signals permit the possibility of detecting a low-frequency modulation when passing through the ultralight dark matter soliton Wang:2023phr; Brax:2024yqh; Blas:2024duy.
 Ref.
-Brax et al.
-(2024) has found that the impact of dynamical friction on the phase evolution of GW signals from white dwarf binaries exceeds that of the low-frequency DM-induced modulation for masses above $10^{-21}$ eV, essentially providing two orders of magnitude of mass that could be probed in the future.
+Brax:2024yqh has found that the impact of dynamical friction on the phase evolution of GW signals from white dwarf binaries exceeds that of the low-frequency DM-induced modulation for masses above $10^{-21}$ eV, essentially providing two orders of magnitude of mass that could be probed in the future.
 Unfortunately, even at these masses, the density of DM would have to be at least $10^{4}$ times higher than in the solar system in order to observe any changes to the phase evolution GW.
-This is broadly consistent with previous work by Wang and Zhong (2023), which assumed an enhancement of the DM density by eight orders of magnitude relative to that on earth, and showed that DM with a mass of $\sim 10^{-23}$ eV could be detected with LISA.
+This is broadly consistent with previous work by Wang:2023phr, which assumed an enhancement of the DM density by eight orders of magnitude relative to that on earth, and showed that DM with a mass of $\sim 10^{-23}$ eV could be detected with LISA.
 Furthermore, Ref.
-Blas et al.
-(2025) assumes a representative population of white dwarf binaries, a fraction of which are detectable in LISA, and concludes that DM masses above $\sim 10^{-23}$ eV will remain unconstrained by LISA.
+Blas:2024duy assumes a representative population of white dwarf binaries, a fraction of which are detectable in LISA, and concludes that DM masses above $\sim 10^{-23}$ eV will remain unconstrained by LISA.
 However, the mass regime $\sim[2\times 10^{-22},3\times 10^{-21}]$ eV could be constrained if DM has stronger, quadratic couplings to ordinary matter.
@@
 VII.3 Nanohertz modulations of continuous gravitational waves from neutron stars
-In addition to modulating the millihertz frequency of GWs from binaries in space-based GW detectors, soliton ultralight dark matter could induce modulations in the rotational frequency of neutron stars in a way distinct from ordinary spinning down, deformed neutron stars, thus affecting CW signals and allowing the possibility of detecting this effect in future ground-based detectors Blas et al.
-(2025).
-These GWs are sourced from the rotational power of the neutron stars, thus spinning down the neutron stars as GWs are emitted Riles (2017).
-In ET and Cosmic Explorer, we expect to detect $\mathcal{O}(100)$ or more canonical neutron stars spinning down due to the emission of CWs Pagliaro et al.
-(2023), which could be modulated by ultralight dark matter solitons.
+In addition to modulating the millihertz frequency of GWs from binaries in space-based GW detectors, soliton ultralight dark matter could induce modulations in the rotational frequency of neutron stars in a way distinct from ordinary spinning down, deformed neutron stars, thus affecting CW signals and allowing the possibility of detecting this effect in future ground-based detectors Blas:2024duy.
+These GWs are sourced from the rotational power of the neutron stars, thus spinning down the neutron stars as GWs are emitted Riles:2017evm.
+In ET and Cosmic Explorer, we expect to detect $\mathcal{O}(100)$ or more canonical neutron stars spinning down due to the emission of CWs Pagliaro:2023bvi, which could be modulated by ultralight dark matter solitons.
 Using CWs from deformed neutron stars, the soliton mass regime $\sim[2\times 10^{-22},3\times 10^{-21}]$ eV becomes accessible in ET for both linear and quadratic couplings, and improves with the number of neutron stars that are detected.
@@
 23 for LISA, ET and Cosmic Explorer, and demonstrates the powerful probe that CWs from neutron stars will be of this particular DM model.
-Figure 23: Taken from Blas et al.
-(2025).
+Figure 23: Taken from Blas:2024duy.
 Projected constraints on the linear (left) and quartic coupling parameters (right) for both LISA and ET from low-frequency soliton DM modulating (“heterodyning”) high-frequency carrier GW signals via changes to the moment of inertia of neutron stars, which alters CWs from deformed neutron stars (ET) and the GWs from inspiraling binary black holes (LISA).
@@
 The LISA curve is determined by the impact of soliton DM on GW frequencies from binary systems.
-The olive regions are excluded by the pulsar timing array analysis in Smarra et al.
-(2024), while blue and red are constrained away via Cassini tests of general relativity (violet) and on a stochastic GW background (red).
+The olive regions are excluded by the pulsar timing array analysis in Smarra:2024kvv, while blue and red are constrained away via Cassini tests of general relativity (violet) and on a stochastic GW background (red).
 We note that in this case, DM leaves a particular imprint on the GW signal itself, arising from deformed, isolated neutron stars.
@@
 VIII.1 Transmuted black holes in binaries
-Non-annihilating WIMPs that interact with nucleons could lose energy through single or multiple scatterings off of celestial objects Bramante et al.
-(2017); Bell et al.
-(2021), eventually becoming gravitationally bound to them over time.
+Non-annihilating WIMPs that interact with nucleons could lose energy through single or multiple scatterings off of celestial objects Bramante:2017xlb; Bell:2020lmm, eventually becoming gravitationally bound to them over time.
 After WIMPs repeatedly interact with ordinary nucleons, they become thermalized, forming a compact core with a higher density than that of the celestial object.
 If the density of the particles becomes sufficient inside the celestial object, gravitational collapse (in the case of bosons) or Chandrasekhar collapse (in the case of fermions) of the WIMP core will occur, thus resulting in a newborn BH that is extremely tiny with respect to the celestial object.
-After formation, the tiny BH could accrete surrounding DM, and if enough DM becomes bound to the object quickly enough compared to the Hawking evaporation timescale, the entire object be “swallowed” by the growing BH and become “transmuted” to what is known as a transmuted black hole (TBH) Dasgupta et al.
-(2021); Ray (2023); Bhattacharya et al.
-(2023, 2024); Bhattacharya (2024).
+After formation, the tiny BH could accrete surrounding DM, and if enough DM becomes bound to the object quickly enough compared to the Hawking evaporation timescale, the entire object be “swallowed” by the growing BH and become “transmuted” to what is known as a transmuted black hole (TBH) Dasgupta:2020mqg; Ray:2023auh; Bhattacharya:2023stq; Bhattacharya:2024pmp; Bhattacharya:2024cpm.
 Quantitatively, if the sum of the time to accrete new particles and the time for the WIMP core to become gravitationally unstable is smaller than the age of the universe, TBHs should form.
-TBHs forming from inspiraling neutron stars were first considered in Dasgupta et al.
-(2021).
+TBHs forming from inspiraling neutron stars were first considered in Dasgupta:2020mqg.
 Since DM would induce neutron stars to transmute into BHs, from the observational perspective, there is no clear difference between electromagnetically silent neutron stars and low-mass BHs.
-Therefore, the non-detection of low-mass binary black holes by LIGO, Virgo and KAGRA can be used to constrain weak interaction cross-sections, since the progenitor neutron stars provide an extremely strong gravitational field to accrete and retain DM particles Bhattacharya et al.
-(2023); Bhattacharya (2024).
+Therefore, the non-detection of low-mass binary black holes by LIGO, Virgo and KAGRA can be used to constrain weak interaction cross-sections, since the progenitor neutron stars provide an extremely strong gravitational field to accrete and retain DM particles Bhattacharya:2023stq; Bhattacharya:2024cpm.
 In Fig.
-24, we show the excluded bosonic DM mass/cross-section parameter space based on null search results from the LIGO, Virgo and KAGRA collaborations for sub-solar mass compact objects Abbott et al.
-(2023).
+24, we show the excluded bosonic DM mass/cross-section parameter space based on null search results from the LIGO, Virgo and KAGRA collaborations for sub-solar mass compact objects LIGOScientific:2022hai.
 Additionally, a projection for this constraint is shown for a 50x increase in the probable spacetime volume, which is reasonable for the future GW network using ET or Cosmic Explorer.
-While analyses of LIGO, Virgo and KAGRA data cannot yet constrain new parameter space for this model of DM, future searches should be able to surpass the pulsar constraint, which comes from the fact that the nearby Gyr-old pulsar PSR 0437-4715 has not actually collapsed into a BH in its lifetime McDermott et al.
-(2012); Garani et al.
-(2019).
-In contrast, stars, being less compact than neutron stars, could sustain DM cores with stronger interaction cross-sections, because of the weaker gravitational field around the star Bhattacharya et al.
-(2024).
+While analyses of LIGO, Virgo and KAGRA data cannot yet constrain new parameter space for this model of DM, future searches should be able to surpass the pulsar constraint, which comes from the fact that the nearby Gyr-old pulsar PSR 0437-4715 has not actually collapsed into a BH in its lifetime McDermott:2011jp; Garani:2018kkd.
+In contrast, stars, being less compact than neutron stars, could sustain DM cores with stronger interaction cross-sections, because of the weaker gravitational field around the star Bhattacharya:2024pmp.
 In Fig.
@@
 24, we note the vastly different interaction cross-section ranges from these two complementary GW probes of WIMP DM.
-For both low-mass BHs and sun-like stars, fermionic DM has also been constrained in similar ways as bosonic DM has, see Bhattacharya et al.
-(2023, 2024) for more details.
+For both low-mass BHs and sun-like stars, fermionic DM has also been constrained in similar ways as bosonic DM has, see Bhattacharya:2023stq; Bhattacharya:2024pmp for more details.
 […]
 VIII.2 GW probes of the galactic-center GeV excess
-Annihilating WIMP DM and a collection of unresolved millisecond pulsars are two leading explanations for the unexpected excess of GeV gamma rays observed by the Fermi-LAT experiment coming from the Galactic Center Goodenough and Hooper (2009); Hooper and Goodenough (2011).
-At first, the GeV excess appeared spherically symmetric and well-fit by annihilating DM models Hooper and Linden (2011); Gordon and Macias (2013); Daylan et al.
-(2016); Calore et al.
-(2015a); Abazajian et al.
-(2014).
-But still, an astrophysical explanation of millisecond pulsars cannot yet be ruled out Abazajian (2011); Calore et al.
-(2015b); Yuan and Zhang (2014); Petrović et al.
-(2015); Ye and Fragione (2022).
+Annihilating WIMP DM and a collection of unresolved millisecond pulsars are two leading explanations for the unexpected excess of GeV gamma rays observed by the Fermi-LAT experiment coming from the Galactic Center Goodenough:2009gk; Hooper:2010mq.
+At first, the GeV excess appeared spherically symmetric and well-fit by annihilating DM models Hooper:2011ti; Gordon:2013vta; Daylan:2014rsa; Calore:2014nla; Abazajian:2014fta.
+But still, an astrophysical explanation of millisecond pulsars cannot yet be ruled out Abazajian:2010zy; Calore:2014xka; Yuan:2014rca; Petrovic:2014xra; Ye:2022yxt.
 GWs can be used to weigh in on this debate.
 If the millisecond pulsars are deformed, they will emit CWs as they rotate.
-The amplitude of these CWs is determined by both the GW frequency and the degree of deformation, called the ellipticity, which is defined as $\epsilon\equiv\frac{I_{\rm xx}-I_{\rm yy}}{I_{\rm zz}}$ Haskell et al.
-(2015), where $I_{\rm xx},I_{\rm yy}$ and $I_{\rm zz}$ are the principle moment of inertias along each axis.
+The amplitude of these CWs is determined by both the GW frequency and the degree of deformation, called the ellipticity, which is defined as $\epsilon\equiv\frac{I_{\rm xx}-I_{\rm yy}}{I_{\rm zz}}$ Haskell:2015psa, where $I_{\rm xx},I_{\rm yy}$ and $I_{\rm zz}$ are the principle moment of inertias along each axis.
 One approach to constraining the millisecond pulsar hypothesis is to consider the bulk emission of a thousands or more sources of CWs that are superimposed, creating a stochastic GW background.
 Ref.
-Calore et al.
-(2019) modelled this stochastic GW background and projected constraints on the average ellipticity of this population as a function of the total number of millisecond pulsars in the Galactic Center.
+Calore:2018sbp modelled this stochastic GW background and projected constraints on the average ellipticity of this population as a function of the total number of millisecond pulsars in the Galactic Center.
 The authors showed that $\mathcal{O}(10^{4})$ millisecond pulsars with ellipticities of $\sim 10^{-7}$ could be detectable by current detectors and be consistent with the diffuse electromagnetic radiation coming from the Galactic Center, as shown in Fig.
@@
 Ref.
-Miller and Zhao (2023) took a different approach by analyzing the continuous gravitational wave (CGW) signals from individual neutron stars (NSs), specifically focusing on millisecond pulsars (MSPs).
+Miller:2023qph took a different approach by analyzing the continuous gravitational wave (CGW) signals from individual neutron stars (NSs), specifically focusing on millisecond pulsars (MSPs).
 They explored two main scenarios for the deformation of these pulsars:
   1. 1.
-    Magnetic Field-Induced Deformations: In the first scenario, they assumed that the deformations of the millisecond pulsars are sustained by the internal magnetic fields of the neutron stars, as proposed in previous work Lander et al.
-(2012); Mastrano and Melatos (2012); Lander (2014).
+    Magnetic Field-Induced Deformations: In the first scenario, they assumed that the deformations of the millisecond pulsars are sustained by the internal magnetic fields of the neutron stars, as proposed in previous work Lander:2011yr; Mastrano:2011aa; Lander:2013oea.
   2. 2.
@@
 Using these two models, Ref.
-Miller and Zhao (2023) then assumed that the ellipticity and frequency distributions of unresolved population of millisecond pulsars can be calcualted with knonw pulsars’ parameters Manchester et al.
-(2005).
-By integrating over these distributions with upper limits on ellipticities from an O3 LIGO search for isolated neutron stars Abbott et al.
-(2022b), they calculated the probability of detecting a CW from this population for both models of ellipticity noted above
+Miller:2023qph then assumed that the ellipticity and frequency distributions of unresolved population of millisecond pulsars can be calcualted with knonw pulsars’ parameters Manchester:2004bp.
+By integrating over these distributions with upper limits on ellipticities from an O3 LIGO search for isolated neutron stars KAGRA:2022dwb, they calculated the probability of detecting a CW from this population for both models of ellipticity noted above
 In parallel, the number of millisecond pulsars in the Galactic Center can be inferred by modeling the electromagnetic emission associated with the GeV excess.
-This is done by fitting luminosity functions, which describe the expected number of millisecond pulsars needed to explain the GeV excess as a function of various model parameters Hooper and Linden (2016); Ploeg et al.
-(2020); Dinsmore and Slatyer (2022).
+This is done by fitting luminosity functions, which describe the expected number of millisecond pulsars needed to explain the GeV excess as a function of various model parameters Hooper:2016rap; Ploeg:2020jeh; Dinsmore:2021nip.
 By combining the predicted number of millisecond pulsars, derived from the luminosity functions, with the probability of detecting them via CWs, the authors were able to rule out certain luminosity function parameters.
@@
 IX GW probes of DM spikes around inspiraling black holes
-It was first recognized in Gondolo and Silk (1999) that BHs embedded in the DM halo of a galaxy could accrete DM adiabatically, potentially forming a high-density region around the BH known as a “spike.” Subsequent studies demonstrated that such DM spikes around supermassive BHs at centers of galaxies would have been disrupted by galactic mergers Merritt et al.
-(2002), if the BH formed off-center Ullio et al.
-(2001), or through interactions with stars in the surrounding stellar population Merritt (2004); Bertone and Merritt (2005).
-However, realistic profiles incorporating relativistic corrections have shown that these spikes can cause more DM to accumulate closer to the BHs in both Schwarzchild and Kerr metrics Sadeghian et al.
-(2013); Ferrer et al.
-(2017); Bertone (2024), which begs the question of whether these objects could survive galactic dynamics.
-Though it has been shown that heavy, particle-like DM would be dispersed in equal-mass merger events through N-body simulations Bertone et al.
-(2005), it is much less clear-cut for ultralight dark matter: multiple works Boudon et al.
-(2023); Zhong et al.
-(2023); Cardoso et al.
-(2022a, b); Vicente and Cardoso (2022); Annulli et al.
-(2020a, b); Brax et al.
-(2020); Ficarra (2021); Zhang et al.
-(2023b); Choudhary et al.
-(2021); Yang et al.
-(2018) suggest that such DM spikes could survive, even up to ten orbits before merger Aurrekoetxea et al.
-(2024).
+It was first recognized in Gondolo:1999ef that BHs embedded in the DM halo of a galaxy could accrete DM adiabatically, potentially forming a high-density region around the BH known as a “spike.” Subsequent studies demonstrated that such DM spikes around supermassive BHs at centers of galaxies would have been disrupted by galactic mergers Merritt:2002vj, if the BH formed off-center Ullio:2001fb, or through interactions with stars in the surrounding stellar population Merritt:2003qk; Bertone:2005hw.
+However, realistic profiles incorporating relativistic corrections have shown that these spikes can cause more DM to accumulate closer to the BHs in both Schwarzchild and Kerr metrics Sadeghian:2013laa; Ferrer:2017xwm; Bertone:2024rxe, which begs the question of whether these objects could survive galactic dynamics.
+Though it has been shown that heavy, particle-like DM would be dispersed in equal-mass merger events through N-body simulations Bertone:2004pz, it is much less clear-cut for ultralight dark matter: multiple works Boudon:2023qbu; Zhong:2023xll; Cardoso:2022nzc; Cardoso:2022vpj; Vicente:2022ivh; Annulli:2020lyc; Annulli:2020ilw; Brax:2019npi; Ficarra:2021qeh; Zhang:2022rex; Choudhary:2020pxy; Yang:2017lpm suggest that such DM spikes could survive, even up to ten orbits before merger Aurrekoetxea:2024cqd.
 The extent to which DM spikes survive directly affects the amount of dynamical friction the binary will experience as it inspirals, thus altering the GWs from the inspiral.
 In light of these recent developments regarding the plausibility of forming DM spikes around supermassive BHs, Ref.
-Daniel et al.
-(2025) has forecast upper bounds on the DM densities around supermassive BHs using future LISA observations of these systems.
+Daniel:2025mna has forecast upper bounds on the DM densities around supermassive BHs using future LISA observations of these systems.
 These bounds provide new constraints on the DM spike formation and the nature of DM in the central regions of galaxies, offering a unique opportunity to test the high-density DM environment near supermassive BH binaries.
 Furthermore, Ref.
-Chan and Lee (2024) determines that the observed decay rate of the orbit can be completely explained by the existence of a DM spike around the supermassive BH in OJ 287, providing plausible evidence of the existence of DM spikes.
-Additionally, IMBHs can sustain DM spikes because their smaller masses and relatively longer timescales for mergers (with respect to those of supermassive BHs) prevent the spikes from being disrupted Zhao and Silk (2005); Kim et al.
-(2023).
-Similarly, recent studies suggest that both astrophysical BHs and PBHs could also host DM spikes if there is an absence of significant stellar interactions that could scatter DM particles, and if the gravitational influence of the IMBH persists over time, especially in environments where galactic mergers or other disruptive events have not destroyed the spike Kohri et al.
-(2014); Eroshenko (2016); Boucenna et al.
-(2018); Hertzberg et al.
-(2020); Bertone et al.
-(2025); Bertone (2024); Dosopoulou and Silk (2025).
+Chan:2024yht determines that the observed decay rate of the orbit can be completely explained by the existence of a DM spike around the supermassive BH in OJ 287, providing plausible evidence of the existence of DM spikes.
+Additionally, IMBHs can sustain DM spikes because their smaller masses and relatively longer timescales for mergers (with respect to those of supermassive BHs) prevent the spikes from being disrupted Zhao:2005zr; Kim:2022mdj.
+Similarly, recent studies suggest that both astrophysical BHs and PBHs could also host DM spikes if there is an absence of significant stellar interactions that could scatter DM particles, and if the gravitational influence of the IMBH persists over time, especially in environments where galactic mergers or other disruptive events have not destroyed the spike Kohri:2014lza; Eroshenko:2016yve; Boucenna:2017ghj; Hertzberg:2019exb; Bertone:2024wbn; Bertone:2024rxe; Dosopoulou:2025jth.
 These findings expand the range of environments where DM spikes might form, suggesting that IMBHs could significantly influence DM distributions.
-For simplicity, despite the more realistic DM spike profiles recently discussed, the density of the DM spike is often approximated by the following equation Eda et al. (2013); Cole et al. (2023a, b):
+For simplicity, despite the more realistic DM spike profiles recently discussed, the density of the DM spike is often approximated by the following equation Eda:2013gg; Cole:2022ucw; Cole:2022yzw:
 $\rho(r)$ $=\rho_{\rm sp}\left(\frac{r_{\rm sp}}{r}\right)^{\gamma_{s}}\,\,\,(r_{\rm min}\leq r\leq r_{\rm sp}),$ (111)
@@
 The power-law index $\gamma_{s}$ denotes how the concentration of DM particles falls off as one moves away from the host BH.
-Astrophysically speaking, measurements of $\gamma_{s}$ would allow us to glean important information regarding whether the host BH evolved from the DM halo and/or experienced mergers in the past Eda et al.
-(2013).
+Astrophysically speaking, measurements of $\gamma_{s}$ would allow us to glean important information regarding whether the host BH evolved from the DM halo and/or experienced mergers in the past Eda:2013gg.
 We consider two scenarios where dynamical friction can influence the evolution of binary black holes: in Section IX.1, we examine how dynamical friction leads to gravitational-wave dephasing in IMRI systems; in Section IX.2, we explore how DM spikes can form around PBHs and cause dephasing in binaries with primary masses of $\mathcal{O}(1\text{--}100)M_{\odot}$ and mass ratios $\leq 10^{-3}$.
@@
 Ref.
-Eda et al.
-(2013) first showed that the phase evolution of an intermediate-mass ratio inspiral (a binary composed to a $\mathcal{O}(1000)M_{\odot}$ and an $\mathcal{O}(1)M_{\odot}$ object) would be significantly modified at milli-hertz frequencies due to the presence of a DM spike around the intermediate-mass BH.
-This occurs because the less massive object sweeps through the DM spike as it orbits around the more massive one, losing energy via dynamical friction (in addition to via GWs) and speeding up the orbit Speeney et al.
-(2022); Mitra et al.
-(2025).
-Such effects will primarily be relevant in space-based GW detectors Duque et al.
-(2024); Rahman et al.
-(2024); Chowdhuri et al.
-(2024); Ghodla (2025), since IMBHs inspiral and merge at much lower frequencies than stellar-mass BHs Miller (2009).
-This conclusion was significantly expanded in Eda et al.
-(2015), in which the authors demonstrated how accurately the power-law index can be measured from a detection of an IMRI signal in LISA.
+Eda:2013gg first showed that the phase evolution of an intermediate-mass ratio inspiral (a binary composed to a $\mathcal{O}(1000)M_{\odot}$ and an $\mathcal{O}(1)M_{\odot}$ object) would be significantly modified at milli-hertz frequencies due to the presence of a DM spike around the intermediate-mass BH.
+This occurs because the less massive object sweeps through the DM spike as it orbits around the more massive one, losing energy via dynamical friction (in addition to via GWs) and speeding up the orbit Speeney:2022ryg; Mitra:2025tag.
+Such effects will primarily be relevant in space-based GW detectors Duque:2023seg; Rahman:2023sof; AbhishekChowdhuri:2023rfv; Ghodla:2024fit, since IMBHs inspiral and merge at much lower frequencies than stellar-mass BHs Miller:2008fi.
+This conclusion was significantly expanded in Eda:2014kra, in which the authors demonstrated how accurately the power-law index can be measured from a detection of an IMRI signal in LISA.
 This accuracy is quantified in terms of the power-law index and the masses of the two objects that are inspiraling.
@@
 This is because higher values of induce a larger phase difference, and smaller masses in the IMBH system ensures that it spends more cycles in the LISA band before merging.
-Recently, it has also been shown that these dephasing effects are unique enough to allow distinguishing between three BH environments: accretion disks, DM spikes, and ultralight boson clouds Cole et al.
-(2023b), as well as distinguishing between environmental effects and modified theories of gravity Yuan et al.
-(2025).
+Recently, it has also been shown that these dephasing effects are unique enough to allow distinguishing between three BH environments: accretion disks, DM spikes, and ultralight boson clouds Cole:2022yzw, as well as distinguishing between environmental effects and modified theories of gravity Yuan:2024duo.
 As an illustration, we show in Fig.
 26 the impact of an environmental effect, the DM spike, on the SNR obtained when match-filtering with a vacuum template to analyze data in which a binary black hole system has a DM spike around it.
-Here, $m_{1}=10^{5}M_{\odot}$, which was chosen to ensure that a plausible formation mechanism exists to form this kind of BH Gondolo and Silk (1999).
+Here, $m_{1}=10^{5}M_{\odot}$, which was chosen to ensure that a plausible formation mechanism exists to form this kind of BH Gondolo:1999ef.
 Ref.
-Cole et al.
-(2023b) argues that a percentage SNR loss of more than 30% would compromise the prospects of detecting a (weak) signal, and shows that even smaller SNR losses would bias parameter estimation.
+Cole:2022yzw argues that a percentage SNR loss of more than 30% would compromise the prospects of detecting a (weak) signal, and shows that even smaller SNR losses would bias parameter estimation.
 However, it is important to note that in LISA, we will be in a regime in which signals of this kind have enormous SNR, and that parameter estimation and distinguishing among different GW signals will be a key challenge.
 Thus, it is more likely that parameter estimates will be biased rather than signals being completely missed.
-Moreover, the dephasing caused by DM spikes underscores the need to adapt CW methods, as discussed in Section III.1 and Section VI.2, to effectively search for these systems Sethi et al.
-(2025).
+Moreover, the dephasing caused by DM spikes underscores the need to adapt CW methods, as discussed in Section III.1 and Section VI.2, to effectively search for these systems Sethi:2025ixx.
 In principle, CW methods could provide quick point estimates for key binary black hole parameters, which would enable the development of a reduced set of templates that incorporate DM spikes.
 These templates could then be used for follow-up matched filtering searches.
-Figure 26: Taken from Cole et al.
-(2023b).
+Figure 26: Taken from Cole:2022yzw.
 Percentage SNR loss from matched filtering with a vacuum template when the injected GW signal originates from a binary black hole system surrounded by DM spikes (the injected waveform includes the effect of dynamical friction).
@@
 While Ref.
-Cole et al.
-(2023b) demonstrated that distinguishing between different BH environments, including the presence of DM spikes, is feasible, Ref.
-Zwick et al.
-(2024) presents a different perspective, arguing that the dephasing effects caused by these environmental factors could result from a variety of unmodeled sources.
+Cole:2022yzw demonstrated that distinguishing between different BH environments, including the presence of DM spikes, is feasible, Ref.
+Zwick:2024yzh presents a different perspective, arguing that the dephasing effects caused by these environmental factors could result from a variety of unmodeled sources.
 Moreover, the authors argue about the practicality of measuring such dephasing effects, noting that the early inspiral is where the effects would be most pronounced, but this is also where the GW emission is weakest; that a constant phase offset could mimic dephasing effects ; and that matched filtering over such long timescales would have practical limitations.
@@
 However, Ref.
-Wilcox et al.
-(2024) provides a counterpoint, suggesting that such dephasing could still be tracked within a parameterized post-Einsteinian framework, offering a potential solution to the challenges raised in Zwick et al.
-(2024).
+Wilcox:2024sqs provides a counterpoint, suggesting that such dephasing could still be tracked within a parameterized post-Einsteinian framework, offering a potential solution to the challenges raised in Zwick:2024yzh.
 Additionally, they propose more model-agnostic approaches, such as periodic forces, that could leave characteristic signatures on the GW signal and be applied to study the evolution of the binary system.
 These differing viewpoints are crucial when considering the overall approach to detecting and interpreting GWs from systems with DM spikes.
-While Zwick et al.
-(2024) concerns highlight the potential difficulties in applying matched filtering to these cases, the model-agnostic approach in Wilcox et al.
-(2024) offer a promising alternative for tackling the problem.
-In light of these challenges, it is clear that both traditional and alternative methods, such as the CW approaches discussed in Section III.1 and Section VI.2, should be explored in tandem to fully address the complexities involved in detecting signals from binary black holes influenced by DM spikes Sethi et al.
-(2025).
+While Zwick:2024yzh concerns highlight the potential difficulties in applying matched filtering to these cases, the model-agnostic approach in Wilcox:2024sqs offer a promising alternative for tackling the problem.
+In light of these challenges, it is clear that both traditional and alternative methods, such as the CW approaches discussed in Section III.1 and Section VI.2, should be explored in tandem to fully address the complexities involved in detecting signals from binary black holes influenced by DM spikes Sethi:2025ixx.
 As mentioned in Section VII.2, CW methods could provide ways of addressing the three concerns laid out in Ref.
-Zwick et al.
-(2024) by providing point estimates of binary parameters that can be followed up by a sensitive and restricted matched filtering search.
+Zwick:2024yzh by providing point estimates of binary parameters that can be followed up by a sensitive and restricted matched filtering search.
 This approach can also be applied in lieu of traditional matched filtering for PBHs in ET and Cosmic Explorer.
 IX.2 Dephasing due to dynamical friction of inspiraling primordial black holes
-PBHs could also host DM spikes, since at most masses, the fraction of DM that they could compose is less than one Green and Kavanagh (2021); Carr et al.
-(2024).
+PBHs could also host DM spikes, since at most masses, the fraction of DM that they could compose is less than one Green:2020jor; Carr:2023tpt.
 Thus, Ref.
-Cole et al.
-(2023a) showed that ET and Cosmic Explorer could probe the existence of DM spikes around binary PBH systems, where the primary has a mass of $1,10]M_{\odot}$ and the system has a mass ratio of $10^{-3}$.
+Cole:2022ucw showed that ET and Cosmic Explorer could probe the existence of DM spikes around binary PBH systems, where the primary has a mass of $1,10]M_{\odot}$ and the system has a mass ratio of $10^{-3}$.
 However, similarly to IMRI systems discussed in the previous subsection, these exotic mergers could be missed if the DM spike is not taken into account in matched-filtering searches in ET and Cosmic Explorer.
@@
 Knowing this, Ref.
-Caneva Santoro et al.
-(2024) modeled dynamical friction and other possible environmental effects as additional (negative) PN corrections to the waveform, and placed upper limits on the density of environment around BHs by requiring the induced dephasing to be small enough to not alter the phase evolution of the binary black hole systems in vacuum.
-While the results were used to eliminate dynamical fragmentation Loeb (2016) as a possible binary formation channel, they are generic enough to apply to any BH environment.
-Unfortunately, the upper limits on environments affected by dynamical friction, which would be relevant for the DM spike scenario, are, at best, about seven orders of magnitude larger than the maximum density of the spike ($10^{-6}$ g/cm3) Gondolo and Silk (1999); Sadeghian et al.
-(2013).
+CanevaSantoro:2023aol modeled dynamical friction and other possible environmental effects as additional (negative) PN corrections to the waveform, and placed upper limits on the density of environment around BHs by requiring the induced dephasing to be small enough to not alter the phase evolution of the binary black hole systems in vacuum.
+While the results were used to eliminate dynamical fragmentation Loeb:2016fzn as a possible binary formation channel, they are generic enough to apply to any BH environment.
+Unfortunately, the upper limits on environments affected by dynamical friction, which would be relevant for the DM spike scenario, are, at best, about seven orders of magnitude larger than the maximum density of the spike ($10^{-6}$ g/cm3) Gondolo:1999ef; Sadeghian:2013laa.
 The upper limits are shown to drastically improve in DECIGO because the dynamical friction (and other environments) affect the early inspiral of the binary black hole system much more than at the time of merger.
-These findings are consistent with those found in N-Body simulations about how the dynamics of the DM spike affect the inspiral, and the detectability of these effects by LISA Cardoso and Maselli (2020); Kavanagh et al.
-(2020).
+These findings are consistent with those found in N-Body simulations about how the dynamics of the DM spike affect the inspiral, and the detectability of these effects by LISA Cardoso:2019rou; Kavanagh:2020cfn.
 X GW probes of atomic DM
 It is possible that the dark sector is actually much more complicated than a single new particle with a single coupling to gravity or the standard model.
-In fact, atomic DM models exist Feng et al.
-(2009); Kaplan et al.
-(2010) in which dark protons and dark electrons interact via a massless dark photon [^12] This dark photon is different from both the dark photon that couples to baryons, discussed in Section II.4, and the one that kinetically mixes with the ordinary photon, discussed in Section VI.3.3.
+In fact, atomic DM models exist Feng:2009mn; Kaplan:2009de in which dark protons and dark electrons interact via a massless dark photon [^12] This dark photon is different from both the dark photon that couples to baryons, discussed in Section II.4, and the one that kinetically mixes with the ordinary photon, discussed in Section VI.3.3.
 with a particular coupling.
-In such a model, DM can form bounds states analogous to the Hydrogen atom, and dissipate energy via Bremsstrahlung, recombination, and collisions Buckley and DiFranzo (2018); Rosenberg and Fan (2017).
-If this kind of DM collects in small-enough regions of space, it will collapse and form “dark BHs” Shandera et al.
-(2018).
-Such dark BHs could form in binary systems, as astrophysical BHs do; thus, LIGO, Virgo and KAGRA observations of binary black hole mergers can constrain the existence of these objects, just as they do for PBHs Singh et al.
-(2021); Abbott et al.
-(2022f, 2023).
+In such a model, DM can form bounds states analogous to the Hydrogen atom, and dissipate energy via Bremsstrahlung, recombination, and collisions Buckley:2017ttd; Rosenberg:2017qia.
+If this kind of DM collects in small-enough regions of space, it will collapse and form “dark BHs” Shandera:2018xkn.
+Such dark BHs could form in binary systems, as astrophysical BHs do; thus, LIGO, Virgo and KAGRA observations of binary black hole mergers can constrain the existence of these objects, just as they do for PBHs Singh:2020wiq; LIGOScientific:2021job; LIGOScientific:2022hai.
 Here, the observable GW signature of dark BHs would come from the inspiral, merger and ringdown of binary black hole systems.
@@
 Ref.
-Shandera et al.
-(2018) first considered the prospects of detecting dark BH mergers with advanced LIGO and ET, finding that these interferometers could set stringent constraints on the fraction of DM that dark BHs could compose.
-If the BH merger of GW190425 Abbott et al.
-(2020a) is interpreted as the merger of two dark BHs, the Chandrasekhar mass Chandrasekhar (1931) can be constrained to be below $1.4M_{\odot}$ at $99.9\%$, bounding the lower-most value mass of the dark proton as 0.95 GeV and the spacing of energy levels in a dark atom to be near $10^{-3}$ eV Singh et al.
-(2023).
+Shandera:2018xkn first considered the prospects of detecting dark BH mergers with advanced LIGO and ET, finding that these interferometers could set stringent constraints on the fraction of DM that dark BHs could compose.
+If the BH merger of GW190425 Abbott:2020uma is interpreted as the merger of two dark BHs, the Chandrasekhar mass Chandrasekhar:1931ih can be constrained to be below $1.4M_{\odot}$ at $99.9\%$, bounding the lower-most value mass of the dark proton as 0.95 GeV and the spacing of energy levels in a dark atom to be near $10^{-3}$ eV Singh:2022wvw.
 In order to obtain these constraints, we note that LIGO, Virgo and KAGRA observations of binary black holes provide estimates of the chirp mass and the component masses of the system.
 Ref.
-Singh et al.
-(2021, 2023) jointly constrain, in a Bayesian way, the fraction of DM that dark BHs could compose and the minimum mass $M_{\rm min}$ of the distribution of dark BHs assumed, while marginalizing over other hyperparmeters, such as the slope and range of the mass distribution.
+Singh:2020wiq; Singh:2022wvw jointly constrain, in a Bayesian way, the fraction of DM that dark BHs could compose and the minimum mass $M_{\rm min}$ of the distribution of dark BHs assumed, while marginalizing over other hyperparmeters, such as the slope and range of the mass distribution.
 These posteriors are conditioned on the output of a GW search, which is essentially predicted merger rate density that depends both on what binary black holes have been observed and power spectral density of the interferometers.
-Quantitatively, these constraints on the mass of the dark proton and the Chandrasekhar limit of the mass of dark BHs are derived using Chandrasekhar (1931); Shandera et al.
-(2018); Singh et al.
-(2021):
+Quantitatively, these constraints on the mass of the dark proton and the Chandrasekhar limit of the mass of dark BHs are derived using Chandrasekhar:1931ih; Shandera:2018xkn; Singh:2020wiq:
 $M_{\rm DC}\simeq 1.4M_{\odot}\left(\frac{m_{p}}{m_{\chi}}\right)^{2}.$ (112)
@@
 $\sqrt{\frac{1.4M_{\odot}}{M_{\rm min}^{\rm max}}}m_{p}<m_{\chi}<\sqrt{\frac{1.4M_{\odot}}{M_{\rm min}^{\rm min}}}m_{p}$ (113)
-In Singh et al.
-(2021), $M_{\rm min}=[0.054,1.50]M_{\odot}$, which implies $0.91<m_{\chi}/\,\mathrm{GeV}<4.76$ when considering both GW190425 and GW190814 Abbott et al.
-(2020b) as dark binary black hole mergers.
+In Singh:2020wiq, $M_{\rm min}=[0.054,1.50]M_{\odot}$, which implies $0.91<m_{\chi}/\,\mathrm{GeV}<4.76$ when considering both GW190425 and GW190814 Abbott:2020khf as dark binary black hole mergers.
 Beyond a range of implied $M_{\rm min}$ and $m_{\chi}$, constraints can be placed on the fraction of DM taht PBHs compose as a function of the average mass of the binary black hole system.
@@
 Here, we limit ourselves to case (1) and case (3).
-These constraints, along with those on the dark proton mass, rely on an assumed mass distribution of dark BHs based on Population III star formation works Bromm and Larson (2004), which is chosen to be $P(m)\propto m^{-b}$ along with a mass range of $M_{\rm min},rM_{\rm min}]$.
+These constraints, along with those on the dark proton mass, rely on an assumed mass distribution of dark BHs based on Population III star formation works Bromm:2003vv, which is chosen to be $P(m)\propto m^{-b}$ along with a mass range of $M_{\rm min},rM_{\rm min}]$.
 The parameters b=[1,2],r=[2,100]b=[-1,2],r=[2,100] determine the mass distribution of dark BHs.
@@
 However, it must be emphasized that while GW190425 and GW190814 being dark BHs are not yet excluded by other experiments, this work does not provide a way of distinguishing astrophysical BHs from dark BHs.
-Figure 27: Taken from Singh et al.
-(2021).
+Figure 27: Taken from Singh:2020wiq.
 Constraints on the fraction of DM that dark BHs could compose relative to other experimental upper limits using all binary black holes detected in the first three observing runs of LIGO, Virgo and KAGRA.
@@
 The latter two cases essentially arise from a joint posterior distribution conditioned on the rate densities inferred from the search results, which also constrain the minimum allowed mass in the assumed population distribution (see text for details).
-Even in the absence of any dark BHs, the constraints obtained in Singh et al.
-(2021) significantly outperform existing microlensing experiments Allsman et al.
-(2001); Tisserand et al.
-(2007); Niikura et al.
-(2019) and those from dynamics of dwarf galaxies Brandt (2016); Koushiappas and Loeb (2017).
-Additional constraints arise from PBH searches Abbott et al.
-(2019); Gaggero et al.
-(2017); Raidal et al.
-(2017, 2019); Ali-Haimoud et al.
-(2017).
+Even in the absence of any dark BHs, the constraints obtained in Singh:2020wiq significantly outperform existing microlensing experiments Macho:2000nvd; EROS-2:2006ryy; Niikura:2019kqi and those from dynamics of dwarf galaxies Brandt:2016aco; Koushiappas:2017chw.
+Additional constraints arise from PBH searches LIGOScientific:2019kan; Gaggero:2016dpq; Raidal:2017mfl; Raidal:2018bbj; Ali-Haimoud:2017rtz.
 SB stands for stochastic background and refers to analyses of the first and second observing runs of LIGO, Virgo and KAGRA data for stochastic GW backgrounds from generic systems.
@@
 The lack of smoking-gun evidence for dark BHs in current LIGO, Virgo and KAGRA observations motivates the need to consider sub-solar mass BHs, whose origins, if detected, cannot be explained by astrophysical means.
-Searches for GWs from sub-solar mass compact objects in binary systems have grown in popularity in recent years, covering mass ranges of roughly $10^{-11},1]M_{\odot}$ Nitz and Wang (2021); Phukon et al.
-(2021); Abbott et al.
-(2022f); Miller et al.
-(2022a); Nitz and Wang (2022); Abbott et al.
-(2023, 2022b); Miller et al.
-(2024), and, traditionally, have put constraints on the fraction of DM that PBHs could compose.
-However, as discussed in Singh et al.
-(2021), these searches do not assume a particular formation mechanism for these compact objects, and the binaries can thus be interpreted as arising from not only PBHs, but also dark BHs.
+Searches for GWs from sub-solar mass compact objects in binary systems have grown in popularity in recent years, covering mass ranges of roughly $10^{-11},1]M_{\odot}$ Nitz:2021mzz; Phukon:2021cus; LIGOScientific:2021job; Miller:2021knj; Nitz:2022ltl; LIGOScientific:2022hai; KAGRA:2022dwb; Miller:2024fpo, and, traditionally, have put constraints on the fraction of DM that PBHs could compose.
+However, as discussed in Singh:2020wiq, these searches do not assume a particular formation mechanism for these compact objects, and the binaries can thus be interpreted as arising from not only PBHs, but also dark BHs.
 In the most recent analysis of LIGO, Virgo and KAGRA data from the second half of the third observing run, constraints on dark BHs with masses below $1M_{\odot}$ were placed, as shown in Fig.
@@
 Moreover, the limits do not go above a solar mass because the search is only for sub-solar mass objects.
-From this work Abbott et al.
-(2023), the mass of the dark proton has been constrained to be [0.66,8.8] GeV, which is actually broader than the constraint in Singh et al.
-(2021) of [0.91,4.76] GeV since there are no assumed detections of dark BHs.
+From this work LIGOScientific:2022hai, the mass of the dark proton has been constrained to be [0.66,8.8] GeV, which is actually broader than the constraint in Singh:2020wiq of [0.91,4.76] GeV since there are no assumed detections of dark BHs.
 Additionally, constraints can be placed on the energy spacing between dark atomic levels.
@@
 By combining the Chandrasekhar mass relation with assumptions about the dark fine-structure constant, the dark fermion masses, and the Jeans mass for collapse, one can use Eq.
-(1) of Singh et al.
-(2021) together with Eq.
-(4) of Shandera et al.
-(2018) to infer the allowed range of $\Delta E$.
+(1) of Singh:2020wiq together with Eq.
+(4) of Shandera:2018xkn to infer the allowed range of $\Delta E$.
 For the GW190425 interpretation, this yields a characteristic value of order $\sim 10^{-3}\,\mathrm{eV}$.
-Figure 28: Taken from Abbott et al.
-(2023).
+Figure 28: Taken from LIGOScientific:2022hai.
 Constraints on the fraction of DM that dark BHs could compose in the atomic DM scenario.
 A mass distribution for dark BHs is assumed with the shown minimum mass on the $x$-axis, and the constraints are shown for three different matched filtering algorithms: GstLAL (dotted), MBTA (dashed) and PyCBC (solid).
-Constraints from a previous search for sub-solar mass compact objects in the first half of the third observing run of LIGO, Virgo and KAGRA are also shown  Abbott et al.
-(2022f).
+Constraints from a previous search for sub-solar mass compact objects in the first half of the third observing run of LIGO, Virgo and KAGRA are also shown  LIGOScientific:2021job.
 This plot differs from Fig.
 27 primarily in the mass range that is considered in the sub-solar mass search versus the ones in, e.g.
-Abbott et al.
-(2021c).
+LIGOScientific:2021djp.
 At lower masses, the constraints on $f_{\rm DBH}$ weaken with respect to those at higher masses.
 Although sub-solar mass searches for dark BHs have placed promising constraints, they are typically limited to masses above $\sim 10^{-2}M_{\odot}$, largely due to the prohibitive computational cost of conducting matched filtering searches for lighter-mass binaries.
-In contrast, CW methods—which search for time-frequency tracks rather than relying on matched filtering —can access lower-mass regimes Miller et al.
-(2021b); Abbott et al.
-(2022b); Miller (2024b); Alestas et al.
-(2024).
+In contrast, CW methods—which search for time-frequency tracks rather than relying on matched filtering —can access lower-mass regimes Miller:2020kmv; KAGRA:2022dwb; Miller:2024khl; Alestas:2024ubs.
 While originally developed for PBH searches, these techniques are equally applicable to testing the dark BH hypothesis.
@@
 However, modified theories of gravity are real contenders to solve this problem as well.
-These two communities – DM and modified gravity – act entirely independently, choosing to break, respectively, one of the two assumptions of “classical” physics: (1) the matter in the universe is mostly baryonic and luminous, and (2) general relativity and, in the limit of slow-moving objects, Newtonian theory, are correct Martens and Lehmkuhl (2020).
-It may be that these assumptions need to be broken simultaneously, or that these two distinct fields have more in common than it appears, as argued in Martens and Lehmkuhl (2020).
-We only briefly mentioned PBHs in this review, which have been reviewed extensively elsewhere Green and Kavanagh (2021); Miller (2024a).
-PBHs could also comprise a fraction of or the totality of DM, and recent observations of binary black holes by the LIGO, Virgo and KAGRA with low spins whose merging rates are consistent with PBHs have reignited interested in this hypothesis Bird et al.
-(2016); Clesse and García-Bellido (2017); Sasaki et al.
-(2016); Croon et al.
-(2023).
+These two communities – DM and modified gravity – act entirely independently, choosing to break, respectively, one of the two assumptions of “classical” physics: (1) the matter in the universe is mostly baryonic and luminous, and (2) general relativity and, in the limit of slow-moving objects, Newtonian theory, are correct Martens:2020lto.
+It may be that these assumptions need to be broken simultaneously, or that these two distinct fields have more in common than it appears, as argued in Martens:2020lto.
+We only briefly mentioned PBHs in this review, which have been reviewed extensively elsewhere Green:2020jor; Miller:2024rca.
+PBHs could also comprise a fraction of or the totality of DM, and recent observations of binary black holes by the LIGO, Virgo and KAGRA with low spins whose merging rates are consistent with PBHs have reignited interested in this hypothesis Bird:2016dcv; Clesse:2016vqa; Sasaki:2016jop; Croon:2022tmr.
 While analyses of PBHs typically receive criticism for using monochromatic mass functions to place constraints, all of the results presented here implicitly assume that DM is only composed of particles that are being constrained [^13] The exception to this statement is superradiance probes of BHs– these bosons need not be DM..